### v2

# Trees of Graphs as Boundaries of Hyperbolic Groups

Abstract.
We characterize those 1-ended word hyperbolic groups whose Gromov boundaries are homeomorphic to trees of graphs (i.e. to inverse limits of graphs that have particularly simple finitary descriptions). These are groups with the simplest connected Gromov boundaries of topological dimension 1. The characterization is expressed in terms of algebraic properties of the Bowditch JSJ splitting of the corresponding groups (i.e. the canonical JSJ splitting over 2-ended subgroups).
Key words and phrases:
hyperbolic group, Gromov boundary, trees of graph, JSJ splitting
2010 Mathematics Subject Classification:
20F65, 20F67
The first named author was partially supported by the ERC grant GroIsRan and by an NSERC Postdoctoral Fellowship
The second author was partially supported by (Polish) Narodowe Centrum Nauki, grant no UMO-2017/25/B/ST1/01335.
1. Introduction
In this paper we address the following classical and widely open problems:
  1. (1)
    to describe explicitly the class of topological spaces which appear as Gromov boundaries of word hyperbolic groups, and
  2. (2)
    to characterize those word hyperbolic groups whose Gromov boundary is homeomorphic to a given topological space.
Concerning the above problem (1), the paper [Ś16] provides its reduction to the case when the boundary is connected, i.e., the corresponding group is 1-ended. Some further partial insight into problem (1) is provided in [Paw15], by showing that Gromov boundaries of hyperbolic groups belong to the class of topological spaces called Markov compacta. These are, roughly speaking, the spaces which can be realized as inverse limits of sequences of polyhedra built recursively out of an initial finite polyhedron according to some finite set of modification (or substitution) rules. In dimension 1, the easiest to describe reasonably rich class of Markov compacta is that of so called regular trees of graphs, as introduced in Section 6 of this paper. The main result of this paper is the complete characterization of those 1-ended hyperbolic groups $G$ whose Gromov boundaries $\partial G$ are trees of graphs (see Theorem A below).
In the statement of our main result, we refer to the concept of a regular tree of graphs (as mentioned above), as well as to the more general concept of a tree of graphs, as introduced in Section 4. The statement refers also to some features of the Bowditch JSJ splitting of a group $G$, as introduced in [Bow98], and recalled in detail in Section 3. This splitting is a canonical splitting of a word hyperbolic group $G$ over 2-ended subgroups. It is defined when $G$ is not cocompact Fuchsian. ($G$ is cocompact Fuchsian if it acts properly and cocompactly on the hyperbolic plane, a condition equivalent to $\partial G\cong S^{1}$.) Here we mention just a few details concerning the Bowditch JSJ splitting, which are necessary for the statement of Theorem A.
Technically, the Bowditch JSJ splitting of $G$ is the action of $G$ on the associated Bass-Serre tree ${T}_{G}$, called the Bowditch JSJ tree for $G$. This tree is naturally bipartitioned into black and white vertices. The black vertices are either rigid or flexible, and their stabilizers are the essential factors of the splitting (which are also called rigid and flexible, respectively). The white vertex stabilizers are maximal 2-ended subgroups of $G$. The $G$-action on ${T}_{G}$ preserves all the above mentioned types of vertices. We introduce the notion of a rigid cluster in $T_{G}$, as follows. We say that two rigid vertices in $T_{G}$ are adjacent if their combinatorial distance in $T_{G}$ is equal to 2. A rigid cluster is a subtree of $T_{G}$ spanned on any equivalence class of the equivalence relation on the set of rigid vertices of $T_{G}$ induced by adjacency. A rigid cluster factor of $G$ is the stabilizing subgroup of any rigid cluster, i.e., the subgroup consisting of all elements which map this cluster to itself. In particular, any rigid factor of $G$ is contained in some rigid cluster factor, but in general rigid cluster factors are larger than rigid factors, and they are in some sense “combined” from collections of rigid factors. See Section 3.2, and in particular Remark 3.6, for more detailed explanations, and for the description of a related splitting that we call the reduced Bowditch JSJ splitting of $G$.
Recall finally that a graph is 2-connected if it is finite, connected, nonempty, not equal to a single vertex, and cutpoint-free.
Theorem A.
Let $G$ be a 1-ended hyperbolic group that is not cocompact Fuchsian. Then the following conditions are equivalent:
  1. (1)
    $\partial G$ is homeomorphic to a tree of graphs;
  2. (2)
    $\partial G$ is homeomorphic to a regular tree of 2-connected graphs;
  3. (3)
    each rigid cluster factor of $G$ is virtually free.
As a complement to Theorem A, we present also in the paper the description of the structure of a regular tree of graphs appearing in condition (2) of this result, given explicitly in terms of the features of the corresponding Bowditch JSJ splitting of $G$ (see Theorem 10.1, and the description of the corresponding graphical connecting system $\mathcal{R}_{G}$ given in Section 10). In fact, as we show in Remark 15.1, an explicit description of $\partial G$ as a regular tree of graphs can be similarly obtained from a splitting of $G$ coming from a much more general class of splittings over $2$-ended subgroups with maximal $2$-ended and virtually free vertex groups.
One can view Theorem A as confirming that regular trees of graphs are indeed the simplest topological spaces in the complexity hierarchy among connected 1-dimensional Gromov boundaries of hyperbolic groups. Theorem A shows also that the zoo of connected 1-dimensional Gromov boundaries is quite complicated, and there is a big gap in it between trees of graphs and the boundaries of those groups which contain as their rigid factors groups whose boundary is either the Sierpinski curve or the Menger curve (see Remark 15.6 for more detailed discussion of this issue). The closer understanding of the species inside this gap seems to be an interesting challenge for future investigations.
As a byproduct of the methods developed for the proof of Theorem A, we present also another result, Theorem B below. In the light of the comments in the previous paragraph, it can be viewed as characterizing those 1-ended hyperbolic groups $G$ whose boundaries $\partial G$ belong to some subclass in the class of trees of graphs, which consists of the simplest spaces among them, namely the so called trees of $\theta$-graphs. A $\theta$-graph is any connected graph having precisely 2 vertices, whose any edge has both these vertices as its endpoints (obviously, $\theta$-graphs typically contain multiple edges). A $\theta$-graph is thick if it has at least three edges. A tree of thick $\theta$-graphs is a tree of graphs whose all consituent graphs are thick $\theta$-graphs (compare Section 4).
Theorem B.
Let $G$ be a 1-ended hyperbolic group that is not cocompact Fuchsian. Then the following conditions are equivelent:
  1. (1)
    $\partial G$ is homeomorphic to a tree of thick $\theta$-graphs;
  2. (2)
    $\partial G$ is homeomorphic to a regular tree of thick $\theta$-graphs;
  3. (3)
    $G$ has no rigid factor in its Bowditch JSJ splitting.
1.1. Examples.
We now illustrate our results with few examples (see Figures 1–4). The details concerning these examples, and other examples of a different nature, will be discussed and presented in Section 14. In each of the examples below, the 1-ended hyperbolic group $G_{i}$ is the fundamental group of the corresponding “graph of surfaces” $M_{i}$. Recall also that the Bowditch JSJ splitting of a group $G$ (as well as its reduced Bowditch JSJ splitting) can be expressed by means of the related graph of groups decomposition $\mathcal{G}$ of $G$, and this is the approach that we follow below. In all four discussed examples $G_{i}$, $i\in\{1,2,3,4\}$, the edge groups of the corresponding graph of groups decompositions $\mathcal{G}_{i}$ coincide with the corresponding incident white vertex groups, so we skip the indication of the form of these edge groups in the figures.
Figure 1. “Graph of surfaces” $M_{1}$ and JSJ splitting $\mathcal{G}_{1}$ of its fundamental group $G_{1}=\pi_{1}M_{1}$.
Consider the group $G_{1}=\pi_{1}M_{1}$, presented in Figure 1. The Bowditch JSJ graph of groups decomposition of $G_{1}$, denoted $\mathcal{G}_{1}$, is presented in the right part of this figure. Since all factors of this decomposition are easily seen to be flexible (because they are the conjugates of the surface subgroups $\pi_{1}F_{i}$ equipped with peripheral subgroups corresponding to the boundary curves), condition (3) of Theorem B holds for $G_{1}$, and hence the boundary $\partial G_{1}$ is homeomorphic to some regular tree of thick $\theta$-graphs.
Figure 2 shows an example in which the corresponding group $G_{2}=\pi_{1}M_{2}$ contains, up to conjugacy, precisely one rigid factor in its Bowditch JSJ splitting. This factor corresponds to the subgroup $\pi_{1}H_{0}<\pi_{1}M_{2}$, and it is obviously isomorphic to the free group of rank 2. Moreover, it is not hard to observe that the rigid clusters in the Bowditch JSJ tree of $G_{2}$ coincide with the rigid vertices, and hence the rigid cluster factors coincide with the rigid factors (actually, in this case the reduced Bowditch JSJ splitting coincides with the ordinary Bowditch JSJ splitting). It follows that the group $G_{2}$ satisfies condition (3) of Theorem A, and hence its Gromov boundary $\partial G_{2}$ is homeomorphic to some regular tree of 2-connected graphs. In Section 14 we will give an explicit description of this regular tree of graphs, in terms of the so called graphical connecting system $\mathcal{R}_{G_{2}}$ associated to $G_{2}$ in the way presented in Section 10.
Figure 2. “Graph of surfaces” $M_{2}$ and JSJ splitting $\mathcal{G}_{2}$ of its fundamental group $G_{2}=\pi_{1}M_{2}$.
The group $G_{3}=\pi_{1}M_{3}$ presented in Figure 3 contains, up to conjugacy, precisely two rigid factors in its Bowditch JSJ splitting, and they correspond to the subgroups $\pi_{1}H_{i}<\pi_{1}M_{3}$, $i=1,2$. It can be deduced from the corresponding graph of groups decomposition $\mathcal{G}_{3}$ of $G_{3}$ that the rigid clusters are in this case nontrivial, and that the rigid cluster factors of $G_{3}$ are the conjugates of the subgroup $\pi_{1}(H_{1}\cup H_{2})$. (The corresponding reduced Bowditch JSJ graph of groups decomposition of $G_{3}$ is presented as $\mathcal{G}^{r}_{3}$ in the same figure.) Since obviously these rigid cluster factors are free, condition (3) of Theorem A is again satisfied, and thus the boundary $\partial G_{3}$ is also homeomorphic to some regular tree of 2-connected graphs.
Figure 3. “Graph of surfaces” $M_{3}$ and splittings $\mathcal{G}_{3}$ and $\mathcal{G}_{3}^{r}$ of its fundamental group $G_{3}=\pi_{1}M_{3}$.
Our final example, presented in Figure 4, shows some 1-ended hyperbolic group $G_{4}=\pi_{1}M_{4}$ whose boundary $\partial G_{4}$ is not a tree of graphs, even though its all rigid factors are free. The rigid factors of $G_{4}$ correspond to the subgroups $\pi_{1}H_{i}<\pi_{1}M_{4}$, for $i=1,2$, and the rigid cluster factors correspond to the subgroup $\pi_{1}(H_{1}\cup H_{2})$. Since that latter group is the fundamental group of a closed surface, it is not virtually free, and hence condition (3) of Theorem A is not fulfilled. As a consequence, the boundary $\partial G_{4}$ is not homeomorphic to any tree of graphs.
Figure 4. “Graph of surfaces” $M_{4}$ and splittings $\mathcal{G}_{4}$ and $\mathcal{G}_{4}^{r}$ of its fundamental group $G_{4}=\pi_{1}M_{4}$.
1.2. Organization of the paper.
The paper is organized as follows. The first part of the paper (Sections 2–5) focuses on the proofs of the implications (1$\Rightarrow$(3) in Theorem A, and (1$\Rightarrow$(3) in Theorem B. The second part of the paper (Sections 6–13) deals with the proofs of the implications (3$\Rightarrow$(2) in Theorem A, and (3$\Rightarrow$(2) in Theorem B. The remaining implications ((2$\Rightarrow$(1) in Theorem A, and (2$\Rightarrow$(1) in Theorem B) necessary to get the full proofs of Theorems A and B are obvious.
Our exposition in both parts of the paper consists of many preparations for the proofs, including the complete descriptions and explanations of the terms appearing in the statements of our two main results. More precisely, in the first part, the organization is as follows. In Section 2 we recall the construction called the limit of a tree system of metric compacta. It is used later, in Section 4, to describe trees of graphs (and trees of $\theta$-graphs), and it is also used throughout the paper, as the main technical ingredient in the arguments. In Section 3.1 we recall the concept of the canonical (Bowditch) JSJ splitting of a 1-ended hyperbolic group, and we describe related concepts of the reduced Bowditch JSJ splitting, and of rigid cluster factors of this reduced splitting. In Section 5 we study some configurations of local cutpoints in trees of graphs (and trees of $\theta$-graphs), and we deduce the proofs of the implications (1$\Rightarrow$(3) in Theorem A, and (1$\Rightarrow$(3) in Theorem B.
In the second part of the paper, the organization is as follows. In Section 6 we introduce regular trees of graphs. Sections 7–9 provide various technicalities and terminology which allow to state in Section 10 (as Theorem 10.1) the result which describes explicitly the form of the regular trees of graphs which appear in our main results. In particular, we discuss in Section 9 some graphs (the so called Whitehead graphs, and their extended variants) which are naturally associated to the virtually free rigid factors (or rather rigid cluster factors) of the JSJ splittings. Sections 10–12 contain some further technical preparations, and in Section 13 we present the proofs of those parts of Theorems A and B, which identify the Gromov boundaries of the corresponding groups as explicitly described regular trees of graphs.
In Section 14, we discuss the examples presented above in this introduction, as well as some further examples, of a different sort, namely some right-angled Coxeter groups.
In Section 15, we make several final comments, including explaining how the description of $\partial G$ as a regular tree of graphs in terms of the Bowditch JSJ splitting can be generalized to a much more general class of splittings.
1.3. Acknowledgments.
The second author expresses his thanks to Jason Manning for directing his attention to Whitehead graphs as objects which could potentially play a role in solving the problem studied in this paper (which indeed turned out to be the case).
2. Tree systems and their limits
In this preparatory section we present (or rather recall) the construction called limit of a tree system of spaces (or, less formally, a tree of spaces). This construction is a slight extension of that presented in Section 1 of [Ś20] (see also [PZ16]). We will use this construction in the next sections to describe and study trees of graphs, the class of spaces on which we focus in this paper.
2.1. Some terminology and notation concerning trees
We will mainly deal with trees $T$ where the vertex set $V_{T}$ is endowed with a bipartition $V_{T}=V_{T}^{b}\sqcup V_{T}^{w}$ into black and white vertices such that every edge of $T$ has endpoints of different colours. We denote the set of all (unoriented) edges of $T$ by $E_{T}$, and for each $e\in E_{T}$ we denote by $b(e)$ and $w(e)$ the black and the white endpoint of $e$, respectively. We will usually denote black vertices with letters $t,s$ and white ones with $u,w$. We use also other letters, such as $v,z,x,y$, to denote vertices whose colour is not essential.
We denote by $[v,z]$ the combinatorial (embedded) path in $T$ connecting a vertex $v$ to a vertex $z$. (If $v,z$ are adjacent, $[v,z]$ denotes also the edge with the endpoints $v,z$.) A ray in $T$ is an infinite (embedded) combinatorial path having an initial vertex. We usually denote rays by $\rho$, and we denote by $\rho(0)$ the initial vertex and by $e_{1}(\rho)$ the initial (unoriented) edge of a ray $\rho$.
For any $v\in V_{T}$, we denote by $N^{T}_{v}$, or simply by $N_{v}$, the set of all vertices of $T$ adjacent to $v$. We denote by $d(v)=d_{T}(v)$ the degree of a vertex $v$ in $T$, i.e., the cardinality of the set $N^{T}_{v}$.
We denote by $\partial T$ the set of ends of $T$, i.e., the set of equivalence classes of rays in $T$ with respect to the relation of coincidence except possibly at some finite initial parts. We denote the end determined by a ray $\rho$ as $[\rho]$.
We consider subtrees of $T$ of some special form, that we call b-subtrees. A subtree $S$ of $T$ is a b-subtree if for each white vertex $u$ of $S$ the set $N^{T}_{u}$ is contained in the vertex set of $S$. We apply to b-subtrees $S$ the earlier introduced notation $V_{S}$, $V_{S}^{b}$, $V_{S}^{w}$, $E_{S}$, $N^{S}_{v}$, $d_{S}(v)$, etc. For a b-subtree $S$ of $T$ we define
$N_{S}=\{u\in V_{T}:\text{$u\notin V_{S}$ and $u$ is adjacent to some $t\in V_{S}$}\}.$
Note that $N_{S}$ consists solely of white vertices of $T$. Note also that, in case when $S$ is equal to a single black vertex $t$, this notation agrees with the earlier introduced notation for the set $N_{t}$.
2.2. Spaces with peripherals and tree systems
Recall that a family of subsets of a compact metric space is null if for each $\varepsilon>0$ all but finitely many of these subsets have diameters less than $\varepsilon$. Note that nullness of a family of subspaces does not depend on the choice of a metric compatible with the topology of the corresponding space.
Definition 2.1.
A space with peripherals is a pair $(K,\Omega)$ such that $K$ is a compact metric space and $\Omega$ is a countable null family of pairwise disjoint closed subsets of $K$. We call the sets from $\Omega$ the peripheral subspaces of $K$, or simply the peripherals of $K$.
Example 2.2.
Say that a graph $\Gamma$ is essential if it is nonempty and contains no isolated vertex. (We allow that $\Gamma$ contains loop edges and multiple edges.) Given a finite essential graph $\Gamma$, we describe the associated space with peripherals called the punctured graph $\Gamma$, denoted $\Gamma^{\circ}$, as follows. Denote by $|\Gamma|$ the realization of $\Gamma$, i.e., the underlying topological space of $\Gamma$, which is a compact metric space. For each point $x\in|\Gamma|$ consider a normal neighbourhood of $x$ in $\Gamma$, i.e., a connected closed neighbourhood $B(x)$ of $x$ whose intersection with the vertex set of $\Gamma$ either coincides with $\{x\}$ (when $x$ is a vertex of $\Gamma$), or is empty. Denote by $\operatorname{int}B(x)$ and $\operatorname{bd}B(x)$ the interior and the boundary (in $|\Gamma|$) of a normal neighbourhood $B(x)$, respectively. A standard dense family of normal neighbourhoods in $\Gamma$ is any family of pairwise disjoint normal neighbourhoods in $\Gamma$ which contains normal neighbourhoods of all vertices of $\Gamma$, and whose union is dense in $|\Gamma|$. Note that each such family is automatically countable and null. Moreover, such family is unique up to a homeomorphism of the ambient space $|\Gamma|$ which is identity on the vertex set of $\Gamma$ and which maps all edges of $\Gamma$ to themselves.
Let ${\mathcal{N}}_{\Gamma}$ be a standard dense family of normal neighbourhoods in $\Gamma$. Put
$\Gamma^{\circ}:=|\Gamma|\setminus\bigcup\{\operatorname{int}B:B\in{\mathcal{N}}_{\Gamma}\}$
and
$\Omega=\{\operatorname{bd}B:B\in{\mathcal{N}}_{\Gamma}\}.$
Note that, by what was said above, $\Gamma^{\circ}$ is a compact metric space, and $\Omega$ is a countable and null family of closed subsets of $\Gamma^{\circ}$. The punctured graph $\Gamma$ is the space with peripherals $(\Gamma^{\circ},\Omega)$.
Note that the peripheral sets $\operatorname{bd}B(x)$ of $\Gamma^{\circ}$ naturally correspond to vertices and edges of $\Gamma$, according to whether $x$ is a vertex or an interior point of an edge. This correspondence describes what we call a type of a peripheral set. The space with peripherals $(\Gamma^{\circ},\Omega)$ is unique up to a homeomorphism which respects types of peripheral sets.
If the underlying space $|\Gamma|$ of a graph $\Gamma$ is homeomorphic to the circle $S^{1}$, it is not hard to realize that the space with peripherals $(\Gamma^{\circ},\Omega)$ as above does not depend, up to homeomorphism, on the number of edges in $\Gamma$. This justifies the notation $(S^{1})^{\circ}$ for this space with peripherals, and the term punctured circle, with which we will be referring to it later.
Punctured graphs will be used in the next section, to describe trees of graphs.
Definition 2.3.
A tree system of spaces is a tuple $\Theta=(T,\{K_{t}\},\{\Sigma_{u}\},\{\varphi_{e}\})$ such that:
  1. (TS1)
    $T$ is a countable tree (called the underlying tree of $\Theta$), equipped with a proper colouring of its vertices in black and white (i.e. each edge has one of its vertices colored black and the other colored white), such that the degree $d(u)$ of each white vertex $u$ is finite;
  2. (TS2)
    to each black vertex $t\in V_{T}^{b}$ there is associated a nonempty compact metric space $K_{t}$, called a constituent space of $\Theta$;
  3. (TS3)
    to each white vertex $u\in V_{T}^{w}$ there is associated a nonempty compact metric space $\Sigma_{u}$, called a peripheral space of $\Theta$;
  4. (TS4)
    to each edge $e\in E_{T}$ there is associated a topological embedding
    $\varphi_{e}:\Sigma_{w(e)}\to K_{b(e)},$
    called a connecting map of $\Theta$;
  5. (TS5)
    for each black vertex $t\in V_{T}^{b}$ the pair
    $(K_{t},\{\varphi_{[t,u]}(\Sigma_{u}):u\in N_{t}\})$
    is a space with peripherals, i.e., the family $\Omega_{t}:=\{\varphi_{[t,u]}(\Sigma_{u}):u\in N_{t}\}$ is null and consists of pairwise disjoint sets; pairs $(K_{t},\Omega_{t})$ are called the constituent spaces with peripherals of $\Theta$.
Note that, by condition (TS5), a black vertex of the underlying tree $T$ of a system $\Theta$ typically has infinite (countable) degree.
2.3. Limit of a tree system
We now describe, following the exposition in Section 1 of [Ś20], an object called the limit of a tree system $\Theta=(T,\{K_{t}\},\{\Sigma_{u}\},\{\varphi_{e}\})$, denoted $\lim\Theta$, starting with its description as a set. Denote by #$\#\Theta$ the quotient
$(\bigsqcup_{t\in V_{T}^{b}}K_{t})/\sim$
of the disjoint union of the sets $K_{t}$ by the equivalence relation $\sim$ induced by the equivalences $\varphi_{[u,t]}(x)\sim\varphi_{[u,s]}(x)$ for any $u\in V_{T}^{w}$, any $x\in\Sigma_{u}$, and any vertices $t,s\in N_{u}$. This set may be viewed as obtained from the family $K_{t}:t\in V$ as a result of gluings provided by the maps
$\varphi_{[u,s]}\varphi_{[u,t]}^{-1}:\varphi_{[u,t]}(\Sigma_{u})\to\varphi_{[u,s]}(\Sigma_{u}).$
Observe that, by the fact that the peripheral subspaces in any $K_{t}$ are pairwise disjoint, the equivalence classes of the relation $\sim$ are easy to describe, and they are all finite (some of them have cardinality 1, and others have cardinalities equal to the degrees of the appropriate white vertices of $T$). Observe also that any set $K_{t}$ canonically injects in #$\#\Theta$. Define $\lim\Theta$ to be the disjoint union #$\#\Theta\sqcup\partial T$, where $\partial T$ is the set of ends of $T$.
To put appropriate topology on the set $\lim\Theta$ we need more terminology. Given a family $\mathcal{A}$ of subsets in a set $X$, we say that $U\subset X$ is saturated with respect to $\mathcal{A}$ (shortly, $\mathcal{A}$-saturated) if for any $A\in{\mathcal{A}}$ we have either $A\subset U$ or $A\cap U=\emptyset$.
For any finite b-subtree $F$ of $T$, denote by
$\Theta_{F}=(F,\{K_{t}:t\in V_{F}^{b}\},\{\Sigma_{u}:u\in V_{F}^{w}\},\{\varphi_{e}:e\in E_{F}\})$
the tree system $\Theta$ restricted to $F$. The set $K_{F}:=\#\Theta_{F}$, equipped with the natural quotient topology (with which it is clearly a metrizable compact space), will be called the partial union of the system $\Theta$ related to $F$. Observe that any partial union $K_{F}$ is canonically a subset in #$\#\Theta$, and thus also in $\lim\Theta$.
For any finite b-subtree $F$ and for any $u\in N_{F}$ denote by $t_{u}$ the (unique) vertex of $F$ adjacent to $u$ (which is necessarily a black vertex). Put ${\Omega}_{F}=\{\varphi_{[u,t_{u}]}(\Sigma_{u}):u\in N_{F}\}$, and view the elements of this family as subsets in the partial union $K_{F}$. Observe that the family ${\Omega}_{F}$ consists of pairwise disjoint closed subsets, and that this family is countable and null (so that $(K_{F},\Omega_{F})$ is a space with peripherals). Let $U\subset K_{F}$ be a subset which is saturated with respect to $\Omega_{F}$. Put $N_{U}:=\{u\in N_{F}:\varphi_{[u,t_{u}]}(\Sigma_{u})\subset U\}$ and
$D_{U}:=\{t\in V_{T}^{b}:\text{$[u,t]\cap F=\emptyset$ for some $u\in N_{U}$}\}$
(i.e. $D_{U}$ is the set of these vertices $t\in V_{T}^{b}\setminus V_{F}$ for which the shortest path in $T$ connecting $t$ to $F$ passes through a vertex $u\in N_{U}$).
Again, for any finite b-subtree $F$ of $T$ let $R_{F}$ be the set of all rays $\rho$ in $T$ with $\rho(0)\in N_{F}$, and with $\rho$ disjoint from $F$. Note that the map $\rho\to[\rho]$ is then a bijection from $R_{F}$ to the set $\partial T$ of all ends of $T$. Given a subset $U\subset K_{F}$ as above (i.e. saturated with respect to $\Omega_{F}$), put
$R_{U}:=\{\rho\in R_{F}:\rho(0)\in N_{U}\},$
and put also $\partial_{U}:=\{[\rho]\in\partial T:\rho\in R_{U}\}$.
Finally, for any subset $U\subset K_{F}$ as above, put
${\mathcal{G}}(U):=U\cup(\bigcup_{t\in D_{U}}K_{t})\cup\partial_{U},$
where all summands in the above union are viewed as subsets of $\lim\Theta$; clearly, ${\mathcal{G}}(U)$ is then also a subset in $\lim\Theta$. We consider the topology in the set $\lim\Theta$ given by the basis $\mathcal{B}$ consisting of all sets ${\mathcal{G}}(U)$, for all finite b-subtrees $F$ in $T$, and all open subsets $U\subset K_{F}$ saturated with respect to $\Omega_{F}$. We leave it as an easy and instructive exercise to check that the family $\mathcal{B}$ is closed under finite intersections, and hence it satisfies the axioms for a basis of a topology. Moreover, we have the following result, which is a slight extension of Proposition 1.C.1 of [Ś20], and which can be justified by the same proof as the latter proposition.
Proposition 2.4.
For any tree system $\Theta$ of metric compacta the limit $\lim\Theta$, with topology given by the above described basis $\mathcal{B}$, is a compact metrizable space.
2.4. Functoriality with respect to subtrees
A tree system can be naturally restricted to any b-subtree of the underlying tree. Namely, let $S\subset T$ be a b-subtree of $T$. The restriction of $\Theta$ to $S$ is the tree system
$\Theta_{S}=(S,\{K_{t}:t\in V_{S}^{b}\},\{\Sigma_{u}:u\in V_{S}^{w}\},\{\varphi_{e}:e\in E_{S}\}).$
Next observation follows by a direct argument (compare the proof of Claim 1 inside the proof of Theorem 3.A.1 in [Ś20]).
Fact 2.5.
The natural map #$\#\Theta_{S}\cup\partial_{S}\to\#\Theta\cup\partial_{T}$ induced by the inclusion $S\to T$, and by the identities of the spaces $K_{t}:t\in V_{S}$, is injective. Moreover, this map viewed as a map from $\lim\Theta_{S}$ to $\lim\Theta$ is a topological embedding.
In the case when $S=F$ is a finite b-subtree of $T$, it is not hard to see that the limit $\lim\Theta_{F}$ coincides with the partial union $K_{F}$. It follows then from Fact 2.5 that each of the partial unions $K_{F}$ naturally embeds (as a topological space) in $\lim\Theta$. In particular, each of the constituent spaces $K_{t}$ naturally embeds in $\lim\Theta$. A similar observation shows that each of the peripheral spaces $\Sigma_{u}$ naturally embeds in $\lim\Theta$. The following result is proved as Proposition 1.C.1 in [Ś20].
Fact 2.6.
The families $K_{t}:t\in V^{b}_{T}$ and $\Sigma_{u}:u\in V^{w}_{T}$, viewed as families of subsets of the limit space $\lim\Theta$, are null. Moreover, the family $\Sigma_{u}:u\in V^{w}_{T}$ consists of pairwise disjoint subsets.
2.5. An isomorphism of tree systems
Let $\Theta=(T,\{K_{t}\},\{\Sigma_{u}\},\{\varphi_{e}\})$ and $\Theta^{\prime}=(T^{\prime},\{K^{\prime}_{t}\},\{\Sigma^{\prime}_{u}\},\{\varphi^{\prime}_{e}\})$ be two tree systems of metric compacta. An isomorphism $F:\Theta\to\Theta^{\prime}$ is a tuple
$F=(\lambda,\{f_{t}:t\in V^{b}_{T}\},\{g_{u}:u\in V^{w}_{T}\})$
such that:
  1. (I1)
    $\lambda:T\to T^{\prime}$ is an isomorphism of trees which respects colours of vertices;
  2. (I2)
    for each $t\in V^{b}_{T}$ and each $u\in V^{w}_{T}$ the maps $f_{t}:K_{t}\to K^{\prime}_{\lambda(t)}$ and $g_{u}:\Sigma_{u}\to\Sigma^{\prime}_{\lambda(u)}$ are homeomorphisms;
  3. (I3)
    for each $e\in E_{T}$ we have $\varphi^{\prime}_{\lambda(e)}\circ g_{w(e)}=f_{b(e)}\circ\varphi_{e}$.
An easy consequence of the definition of the limit (of a tree system of metric compacta) is the following.
Lemma 2.7.
If $\Theta,\Theta^{\prime}$ are isomorphic tree systems of metric compacta then their limits $\lim\Theta$ and $\lim\Theta^{\prime}$ are homeomorphic. More precisely, an isomorphism $F:\Theta\to\Theta^{\prime}$ induces (in the obvious way) a homeomorphism $\lim\Theta\to\lim\Theta^{\prime}$.
2.6. An inverse system associated to a tree system
Let
$\Theta=(T,\{K_{t}\},\{\Sigma_{u}\},\{\varphi_{e}\})$
be a tree system. For any finite b-subtree $F\subset T$, consider the family $\Omega_{F}$ of peripheral subspaces of the partial union $K_{F}$, as described in Section 2.3, and denote by $K_{F}^{*}=K_{F}/\Omega_{F}$ the quotient of $K_{F}$ in which all the sets from $\Omega_{F}$ are shrunk to points. More precisely, let ${\mathcal{D}}_{F}$ be the decomposition of $K_{F}$ consisting of the sets of $\Omega_{F}$ and the singletons from the complement $K_{F}\setminus\bigcup\Omega_{F}$, and let $K_{F}^{*}=K_{F}/{\mathcal{D}}_{F}$. Since the family $\Omega_{F}$ is null, the decomposition ${\mathcal{D}}_{F}$ is upper semicontinuous, and hence the quotient $K_{F}^{*}$ is metrizable (see [Dav86], Proposition 3 on page 14 and Proposition 2 on page 13). Consequently, $K_{F}^{*}$ is a metric compactum, and we call it the reduced partial union of $\Theta$ related to $F$. We denote by $q_{F}:K_{F}\to K_{F}^{*}$ the quotient map resulting from the above description of $K_{F}^{*}$.
For any pair $F_{1}\subset F_{2}$ of finite b-subtrees of $T$, define a map $f_{F_{1}F_{2}}:K_{F_{2}}^{*}\to K_{F_{1}}^{*}$ as follows. For each white vertex $u\in N_{F_{1}}\cap V_{F_{2}}$ denote by ${\mathcal{V}}_{u}$ the set of all black vertices $s\in V_{F_{2}}^{b}\setminus V_{F_{1}}$ such that the shortest path in $T$ connecting $s$ with $F_{1}$ passes through $u$. Then the family ${\mathcal{V}}_{u}:u\in N_{F_{1}}\cap V_{F_{2}}$ is a partition of $V_{F_{2}}^{b}\setminus V_{F_{1}}$. For each $u\in N_{F_{1}}\cap V_{F_{2}}$, denote by $S_{u}$ the subtree of $T$ spanned on the set ${\mathcal{V}}_{u}$, and note that it is actually a subtree of $F_{2}$. Viewing each $K_{S_{u}}=\lim\Theta_{S_{u}}$ canonically as a subset in $K_{F_{2}}$, consider the corresponding subset $q_{F_{2}}(K_{S_{u}})$ in $K_{F_{2}}^{*}$. Observe that, by shrinking each of the subsets $q_{F_{2}}(K_{S_{u}}):u\in N_{F_{1}}\cap V_{F_{2}}$ to a point we get a quotient of $K_{F_{2}}^{*}$ which is canonically homeomorphic to (and which we identify with) $K_{F_{1}}^{*}$. Take the corresponding quotient map as $f_{F_{1}F_{2}}$, and observe that this map is continuous and surjective.
Given any finite b-subtrees $F_{1}\subset F_{2}\subset F_{3}$ of $T$, it is not hard to see that $f_{F_{1}F_{2}}\circ f_{F_{2}F_{3}}=f_{F_{1}F_{3}}$. Consequently, the system
${\mathcal{S}}_{\Theta}=(\{K_{F}^{*}:F\in{\mathcal{F}}_{T}^{b}\},\{f_{FF^{\prime}}:F\subset F^{\prime}\})$
is an inverse system of metric compacta over the poset ${\mathcal{F}}_{T}^{b}$ of all finite b-subtrees of $T$. We call it the standard inverse system associated to $\Theta$.
Next result is a slight extension of Proposition 1.D.1 in [Ś20], and its proof is the same as the proof of that proposition (hence we omit it).
Proposition 2.8.
Let $\Theta$ be a tree system, and let ${\mathcal{S}}_{\Theta}$ be the standard inverse system associated to $\Theta$. Then the limit $\lim\Theta$ is canonically homeomorphic to the inverse limit $\lim_{\leftarrow}{\mathcal{S}}_{\Theta}$.
2.7. Consolidation of a tree system
We describe an operation which turns one tree system of spaces into another by merging the constituent spaces of the initial system, and forming a new system out of bigger pieces (corresponding to a family of pairwise disjoint b-subtrees in the underlying tree of the initial system). As we indicate below (see Theorem 2.9), this operation does not affect the limit of a system.
Let $\Theta=(T,\{K_{t}\},\{\Sigma_{u}\},\{\varphi_{e}\})$ be a tree system of metric compacta. Let $\Pi$ be a partition of a tree $T$ into b-subtrees, i.e., a family of b-subtrees $S\subset T$ such that the black vertex sets $V_{S}^{b}:S\in{\Pi}$ are pairwise disjoint and cover all of $V_{T}^{b}$. We allow that some of the subtrees $S\in{\Pi}$ are just single black vertices of $T$.
We define a consolidation of $\Theta$ with respect to $\Pi$ to be a tree system
$\Theta^{\Pi}=(T^{\Pi},\{K_{S}:S\in\Pi\},\{\Sigma^{\Pi}_{u}:u\in V^{b}_{T^{\Pi}}\},\{\varphi^{\Pi}_{e}:e\in E_{T^{\Pi}}\})$
described as follows. As a set $V^{b}_{T^{\Pi}}$ of black vertices of $T^{{\Pi}}$ we take the family $\Pi$, and as a set $V^{w}_{T^{\Pi}}$ of white vertices we take the set of those white vertices $u\in V^{w}_{T}$ which remain outside all of the subtrees $S\in\Pi$. As the edge set $E_{T^{\Pi}}$ we take the set $\{e\in E_{T}:e\notin\cup_{S\in{\Pi}}\,E_{S}\}$, so that each edge $e$ from this set is meant to connect the white vertex $w(e)\in V^{w}_{T^{\Pi}}$ with the subtree $S\in\Pi$ (viewed as a black vertex of $T^{\Pi}$) which contains the vertex $b(e)$ of $T$. Intuitively speaking, the above described tree $T^{\Pi}$ is a “dual tree” of the partition $\Pi$.
To describe the constituent spaces $K_{S}$ of the tree system $\Theta^{\Pi}$, for any subtree $S\in{\Pi}$ denote by $\Theta_{S}$ the restriction of $\Theta$ to $S$, and put $K_{S}:=\lim\Theta_{S}$. To describe the peripheral spaces $\Sigma_{u}^{\Pi}$ and the connecting maps $\varphi_{e}^{\Pi}$, note that for any vertex $u\in V^{w}_{T^{\Pi}}$ (viewed as a white vertex of $T$) and any edge $e$ with $w(e)=u$, the map $\varphi_{e}:\Sigma_{u}\to K_{b(e)}$ may be viewed as a map to the space $K_{S}$ for this $S\in\Pi$ which contains the vertex $b(e)$. Thus, for $u$ and $e$ as above, we put $\Sigma_{u}^{\Pi}:=\Sigma_{u}$ and $\varphi_{e}^{\Pi}:=\varphi_{e}$. Now, viewing any subtree $S\in\Pi$ as a vertex of $T^{\Pi}$, for any vertex $u\in N^{T^{\Pi}}_{S}$ denote by $t^{S}_{u}$ this vertex of the subtree $S$ which is adjacent to $u$ in $T$, and note that, as a consequence of Fact 2.6, the family
$\Omega^{\Pi}_{S}:=\{\varphi^{\Pi}_{[u,t^{S}_{u}]}(\Sigma^{\Pi}_{u}):u\in N^{T^{\Pi}}_{S}\}$
of subsets of the space $K_{S}$ is null. It follows that the pair $(K_{S},\Omega^{\Pi}_{S})$ is a space with peripherals. This justifies that the just described tuple
$\Theta^{\Pi}=(T^{\Pi},\{K_{S}\},\{\Sigma^{\Pi}_{u}\},\{\varphi^{\Pi}_{e}\})$
is a tree system of metric compacta.
The following result has been proved as Theorem 3.A.1 in [Ś20].
Theorem 2.9.
For any tree system $\Theta$ of metric compacta, and any consolidation $\Theta^{\Pi}$ of $\Theta$, the limits $\lim\Theta$ and $\lim\Theta^{\Pi}$ are canonically homeomorphic.
3. The reduced Bowditch JSJ splitting and rigid cluster factors
In this section we introduce a new concept of the reduced Bowditch JSJ splitting of a 1-ended hyperbolic group, which is a variation of the classical Bowditch JSJ splitting. In particular, we describe the factors of this splitting, especially the so called rigid cluster factors which appear in the statement of our main result, Theorem A of the introduction. We start with recalling basic features of the classical Bowditch JSJ splitting.
3.1. The Bowditch JSJ splitting of a 1-ended hyperbolic group
We refer the reader to [Bow98] or to Section 5 in [KW22] for a detailed exposition of the Bowditch JSJ splitting. Here we only mention briefly that this splitting describes in some clever way the configuration of all local cutpoints in the Gromov boundary $\partial G$ of a 1-ended hyperbolic group $G$, and relates it to a certain canonical JSJ splitting of $G$ (so that the Bowditch JSJ tree appearing as part of the data of this construction is canonically identified with the Bass-Serre tree of the corresponding JSJ splitting of $G$).
More precisely, let $G$ be a 1-ended hyperbolic group which is not cocompact Fuchsian (i.e. its Gromov boundary $\partial G$ is connected and not homeomorphic to the circle $S^{1}$). Recall that, under these assumptions, $\partial G$ is a connected and locally connected compact metrizable space without global cutpoints [Bow98]. Moreover, each local cutpoint $p$ of $\partial G$ has finite degree (where degree of $p$ is, by definition, the number of ends of a locally compact space $\partial G\setminus\{p\}$). The set of all local cutpoints of $\partial G$ is split into classes, so that any two points from the same class form a cut pair, any two points in the same class have the same degree and that this degree is equal to the number of connected components in the complement of the corresponding cut pair. This splitting into classes is natural, in the sense that classes are mapped to classes by homeomorphisms of $\partial G$. Each class is either infinite (in which case it consists of points of degree 2) or it is of size 2 (i.e. it consists of precisely 2 points). Two classes $\zeta_{1}$ and $\zeta_{2}$ are separated if they lie in distinct components of the complement $\partial G\setminus\{x,y\}$ of a cut pair $x,y\in\eta$ from some other class $\eta$. An object called the Bowditch JSJ tree for $G$, which we denote by ${T}_{G}$, has vertices $v$ (accompanied with some quasi-convex subgroups $G(v)<G$) of the following four types:
  1. (v1)
    vertices $v$ related to infinite classes of local cutpoints of degree 2, the so called necklaces, for which the associated groups $G(v)$ are virtually free and coincide with the stabilizers of the corresponding necklaces under the action of $G$ on $\partial G$; the boundary $\partial G(v)\subset\partial G$ coincides then with the closure of the corresponding necklace, and it is homeomorphic to the Cantor set; such vertices $v$ have countable infinite degree in ${T}_{G}$;
  2. (v2)
    vertices $v$ related to classes of size 2, for which the associated groups $G(v)$ are equal to the $G$-stabilizers of the corresponding classes; these groups $G(v)$ are maximal 2-ended subgroups of $G$ and their boundaries $\partial G(v)$ coincide with the corresponding classes; the degree of such a vertex $v$ in the tree ${T}_{G}$ is equal to the degree of any local cutpoint in the corresponding class, which is an integer $\geq 2$;
  3. (v3)
    vertices $v$ related to infinite maximal families of pairwise not separated classes called stars, for which the associated groups $G(v)$ are equal to the $G$-stabilizers of unions of the corresponding stars; the degree of such a vertex $v$ in the tree ${T}_{G}$ is equal to the cardinality of the family of classes in the corresponding star, which is always infinite countable;
  4. (v4)
    vertices $v$ corresponding to pairs consisting of a necklace $C$ and a star $S$ containing this necklace; if we denote by $v_{1}$ and $v_{3}$ the vertices corresponding to the necklace $C$ and to the star $S$, respectively, then $G(v)=G({v_{1}})\cap G({v_{3}})$, which is a maximal 2-ended subgroup of $G$; the degree of such a vertex $v$ in the tree ${T}_{G}$ is equal to 2, and the boundary $\partial G(v)$ is contained in $C$.
The edges $e$ of ${T}_{G}$, and the associated groups $G(e)$, have the following forms:
  1. (e1)
    an edge $e$ connecting a pair of vertices $v_{1},v_{2}$ of types (v1) and (v2), respectively; such an edge appears if and only if the class of size 2 corresponding to $v_{2}$ is contained in the closure of the infinite class corresponding to $v_{1}$; moreover, in this case the class corresponding to $v_{2}$ consists of local cutpoints of degrees $\geq 3$;
  2. (e2)
    an edge $e$ connecting a pair of vertices $v_{2},v_{3}$ of types (v2) and (v3), respectively; such an edge appears if and only if the class corresponding to $v_{2}$ belongs to the star corresponding to $v_{3}$;
  3. (e3)
    given any vertex $v$ of type (v4), let $v_{1},v_{3}$ be the vertices related to the corresponding necklace and star; for each such $v$ we have in ${T}_{G}$ the edges $e=[v,v_{1}]$ and $e^{\prime}=[v,v_{3}]$, and these are the only edges of ${T}_{G}$ adjacent to $v$.
In all cases (e1)(e3) above the associated group $G(e)$ is the intersection of the two vertex groups $G(v)$ corresponding to the endpoints of $e$. For each edge $e$ the group $G(e)$ is 2-ended, and it is a finite index subgroup of the corresponding group $G(v)$, where $v$ is the (unique) vertex of type (v2) or (v4) in $e$. We have a natural partition of the vertices of ${T}_{G}$ into black ones (vertices of types (v1) and (v3)) and white ones (types (v2) and (v4)). No two vertices of the same colour are adjacent in ${T}_{G}$. The degrees in ${T}_{G}$ of all white vertices are finite, and those of the black vertices are countable infinite.
The tree ${T}_{G}$ is naturally acted upon by $G$. This action preserves colours of the vertices, and the groups $G(v)$ and $G(e)$ coincide with the stabilizers of vertices $v$ and of edges $e$, with respect to this action.
Groups $G(v)$ for vertices $v$ of type (v1) are called flexible or hanging factors of the Bowditch JSJ splitting of $G$, while those for vertices of type (v3) are called rigid factors. The corresponding vertices $v$ are also called flexible and rigid, accordingly. We will adapt this terminology in the next sections of this paper.
The following observation is well known, and will be explained in more details later in this paper (see Section 8.1, Definition 8.3 and the comments in the paragraph right before this definition; see also Example 8.8 together with Definition 8.7).
Fact 3.1.
Let $v$ be a black vertex of ${T}_{G}$, i.e., either a rigid or a flexible vertex. Put $K_{v}:=\partial G(v)$, and for any vertex $u$ of ${T}_{G}$ adjacent to $v$, $u\in N_{v}$, put $\Sigma_{u}:=\partial G(u)=\partial G([u,v])$. Then the pair $(K_{v},\{\Sigma_{u}:u\in N_{v}\})$ is a space with peripherals. Moreover, if $v$ is flexible then this space with peripherals is (homeomorphic to) the punctured circle $(S^{1})^{\circ}$.
We now turn to the description of a certain tree system canonically associated to the Bowditch JSJ splitting of $G$, and to the expression of the Gromov boundary of $G$ as the limit of this tree system. Denote by $V_{{T}_{G}}^{b}$ and $V_{{T}_{G}}^{w}$ the sets of all black and white vertices of the tree ${T}_{G}$, respectively. Consider the following tree system $\Theta=\Theta(G)$:
$\Theta(G)=({T}_{G},\{K_{t}=\partial G(t):t\in V_{{\mathcal{T}}_{G}}^{b}\},\{\Sigma_{u}=\partial G(u):u\in V_{{T}_{G}}^{w}\},\{\varphi_{e}:e\in E_{{T}_{G}}\}),$
where each $\varphi_{e}:\partial G({w(e)})\to\partial G({b(e)})$ is the natural inclusion. We will call $\Theta(G)$ the tree system associated to the Bowditch JSJ splitting of $G$, or concisely the tree system associated to $G$. The next result follows fairly directly by the arguments provided by Alexandre Martin (see Corollary 9.19 in [Mar14] or Theorem 2.25 in [CM17]). In the proof provided below we explain briefly how to apply these results in our situation, by referring to the exposition in [CM17].
Theorem 3.2.
Let $G$ be a 1-ended hyperbolic group which is not cocompact Fuchsian, and let $\Theta(G)$ be the associated tree system. Then the Gromov boundary $\partial G$ is homeomorphic to the limit of the tree system $\Theta(G)$, i.e.,
$\partial G\cong\lim\Theta(G).$
Remark 3.3.
The homeomorphism $\partial G\cong\lim\Theta(G)$ in the above statement is natural in the sense that it is induced by the inclusions $\partial G(t)\subset\partial G$ and $\partial G(u)\subset\partial G$, where $t\in V_{{\mathcal{T}}_{G}}^{b}$, $u\in V_{{T}_{G}}^{w}$.
Proof.
We sketch here an explanation of how Theorem 3.2 follows from the result of Alexandre Martin mentioned right before the statement of this theorem.
In Section 2.6.2 of [CM17] Alexandre Martin and Christopher Cashen consider an object $X$ called a proper geodesic hyperbolic tree of spaces, and the associated object $\partial X$ (which may be viewed as the associated tree of boundaries). It is not hard to see that any limit $\lim\Theta(G)$, viewed as a set, can be identied with appropriately chosen Martin’s set $\partial X$. (Namely, take as the underlying tree $T$ of $X$ the Bass-Serre tree $T_{G}$, take as the vertex spaces $X_{v}$ of $X$ the corresponding vertex groups $G(v)$, and as the edge spaces $X_{e}$ the corresponding edge stabilizers $G(e)$.) A less clear fact is that in case of $\lim\Theta(G)$ (viewed as special case of Martin’s $\partial X$) the topology on $\partial X$ described in Definitions 2.19 and 2.21 in [CM17] coincides with the topology on $\lim\Theta(G)$ provided in Section 2.3 of the present paper. We devote the remaining part of this proof to the justification of this coincidence of topologies, noting that then Theorem 3.2 follows fairly directly from Theorem 2.22 in [CM17].
Martin introduces his topology on $\partial X$ by means of describing families of neighbourhoods of points. (These neighbourhoods are not necessarily open in the resulting topology, but they contain open subneighbourhoods of the corresponding points.) In the case where $\partial X=\lim\Theta(G)$, we need to show that the identity map is a homeomorphism from $\partial X$ equipped with Martin’s topology to $\partial X$ equipped with our topology (as described in Section 2.3). However, since both these topological spaces are compact metrizable, it is sufficient to show that the identity map is continuous.
Suppose we are given a topological space $Z$ whose topology is given by a family $N(z):z\in Z$ of neighbourhood systems for points $z$, and a topological space $Z^{\prime}$ whose topology is induced by a basis $\mathcal{B}$ of open sets. Recall that then a map $f:Z\to Z^{\prime}$ is continuous iff for any open set $E\in{\mathcal{B}}$ and any $z\in f^{-1}(E)$ there is a neighbourhood $N\in N(z)$ of point $z$ such that $N\subset f^{-1}(E)$. In particular, if $Z=Z^{\prime}$ (as sets) then the identity map $\hbox{id}_{Z}:Z\to Z^{\prime}$ is continuous (with respect to the topologies mentioned above) iff for any $E\in{\mathcal{B}}$ and any $z\in E$ there is $N\in N(z)$ such that $N\subset E$.
Suppose now that $\partial X=\lim\Theta(G)$. Recall that $\lim\Theta(G)$, as a set, is the disjoint union of the following subsets: $\partial T_{G}$, $\Sigma_{u}:u\in V^{w}_{T_{G}}$ and $\dot{K}_{t}:t\in V^{b}_{T_{G}}$, where $\dot{K}_{t}=K_{t}\setminus\bigcup\{\Sigma_{u}:u\in N_{t}\}$. By a direct application of Definitions 2.19 and 2.21 of [CM17] in our context we get that Martin’s topology on $\partial X=\lim\Theta(G)$ is described by the following families of neighbourhoods of points:
  1. (1)
    for a point $\eta\in\partial T_{G}\subset\lim\Theta(G)$, and for any neighbourhood $U$ of $\eta$ in $V_{T_{G}}\cup\partial T_{G}$, we have a neighbourhood $V_{U}(\eta)$ of $\eta$ given by
    $V_{U}(\eta)=(U\cap\partial T_{G})\cup\bigcup\{\dot{K}_{t}:t\in U\}\cup\bigcup\{\Sigma_{u}:\{u\}\cup N_{u}\subset U\};$
  2. (2)
    for a point $\xi\in\dot{K}_{t}\subset\lim\Theta(G)$, and for any neighbourhood $U_{t}$ of $\xi$ in $K_{t}$, define $\hbox{Cone}_{\{U_{t}\}}(\xi)$ as the set consisting of all those $x\in V_{T_{G}}\cup\partial T_{G}\setminus\{t\}$ for which we have $\varphi_{e}(\Sigma_{w(e)})\subset U_{t}$, where $e$ is the last edge on the path from $x$ to $t$ in $T_{G}$; we then have a neighbourhood $V_{\{U_{t}\}}(\xi)$ of $\xi$ given by
    $V_{\{U_{t}\}}(\xi)=(\partial T_{G}\cap\hbox{Cone}_{\{U_{t}\}}(\xi))\cup\bigcup\{K_{s}:s\in\hbox{Cone}_{\{U_{t}\}}(\xi)\}\cup U_{t};$
  3. (3)
    for a point $\xi\in\Sigma_{u_{0}}\subset\lim\Theta(G)$, and for any family ${\mathcal{U}}=\{U_{u_{0}}\}\cup\{U_{t}:t\in N_{u_{0}}\}$ of neighbourhoods of $\xi$ in the spaces $\Sigma_{u_{0}}$ and $\{K_{t}:t\in N_{u_{0}}\}$, respectively, define $\hbox{Cone}_{\mathcal{U}}(\xi)$ as the set consisting of all those $x\in V_{T_{G}}\cup\partial T_{G}\setminus(\{u_{0}\}\cup N_{u_{0}})$ for which we have $\varphi_{e}(\Sigma_{w(e)})\subset U_{b(e)}$, where $e$ is the last edge on the shortest path in $T_{G}$ from $x$ to $N_{u_{0}}$; we then have a neighbourhood $V_{\mathcal{U}}(\xi)$ of $\xi$ given by
    $V_{\mathcal{U}}(\xi)=(\partial T_{G}\cap\hbox{Cone}_{\mathcal{U}}(\xi))\cup\bigcup\{K_{s}:s\in\hbox{Cone}_{\mathcal{U}}(\xi)\}\cup\big{[}\bigcup\{U_{t}:t\in N_{u_{0}}\}\setminus U_{u_{0}}\big{]}\cup$
    $\cup\big{[}U_{u_{0}}\cap\bigcap\{U_{t}:t\in N_{u_{0}}\}\big{]}.$
We are now ready for showing that the identity map of the set $\partial X=\lim\Theta(G)$, for the Martin’s topology in the source, and for the topology as in Section 2.3 in the target, is continuous. To do this, consider any open set $E\in{\mathcal{B}}$ (where $\mathcal{B}$ is as in Section 2.3), i.e. a set of form $E={\mathcal{G}}(W)$ for an appropriate $\Omega_{F}$-saturated open subset $W$ in a partial union $K_{F}$. We will proceed with the argument in the following three cases. First, let $\eta\in E$ be a point which belongs to $\partial T_{G}$. We need to find a neighbourhood of the form $N=V_{U}(\eta)$ as in (1) above, such that $N\subset E$. Let $t_{0}$ be this vertex of the b-subtree $F$ which is “closest” to $\eta$ in $T_{G}$, and let $e=[u_{0},t_{0}]$ be the last edge on the path in $T_{G}$ from $\eta$ to $t_{0}$. Put $U$ to be the set of all those elements $z$ in $(V_{T_{G}}\cup\partial T_{G})\setminus V_{F}$ for which the path in $T_{G}$ from $z$ to $t_{0}$ passes through $u_{0}$ (including $z=u_{0}$). $U$ is obviously a neighbourhood of $\eta$ in $V_{T_{G}}\cup\partial T_{G}$, and it is not hard to observe that the corresponding neighbourhood $V_{U}(\eta)$ is a subset of $E$, as required.
In the second case, let $\xi\in E$ be a point which belongs to $\dot{K}_{t}$, for some $t\in V^{b}_{T_{G}}$. We need to find a neighbourhood of the form $N=V_{\{U_{t}\}}(\xi)$ as in (2) above, such that $N\subset E$. Consider first the subcase when $t\notin V_{F}$, and let $e$ be the first edge on the path in $T_{G}$ connecting $t$ with the nearest vertex of $F$. Put $U_{t}=K_{t}\setminus\varphi_{e}(\Sigma_{w(e)})$. Obviously, $U_{t}$ is then a neighbourhood of $\xi$ in $K_{t}$, and the corresponding neighbourhood $V_{\{U_{t}\}}(\xi)$ is a subset of $E$, as required. In the second subcase, let $t\in V_{F}$. Observe that then $\xi\in W$. Let $e_{1},\dots,e_{k}$ be the edges of $F$ adjacent to $t$, and put $U_{t}=(K_{t}\cap W)\setminus\cup_{i=1}^{k}\varphi_{e_{i}}(\Sigma_{w(e_{i})})$. Again, it is easily seen that $U_{t}$ is a neighbourhood of $\xi$ in $K_{t}$, and that the corresponding neighbourhood $V_{\{U_{t}\}}(\xi)$ is a subset of $E$, as required.
In the final third case, let $\xi\in E$ be a point which belongs to $\Sigma_{u_{0}}$, for some $u_{0}\in V^{w}_{T_{G}}$. We need to find a neighbourhood of the form $N=V_{\mathcal{U}}(\xi)$ as in (3) above, such that $N\subset E$. We will do this by considering three subcases related to the possible position of $u_{0}$ with respect to $F$. First, consider the subcase where $u_{0}\in V_{F}$. Observe that then $\xi\in W$. Note also that, since $F$ is a b-subtree of $T_{G}$, the set $N_{u_{0}}$ of all neighbours of $u_{0}$ in $T_{G}$ is actually contained in $V_{F}$. For each $t\in N_{u_{0}}$, let ${\mathcal{E}}_{t}$ be the set of edges of $F$ adjacent to $t$ and distinct from $[u_{0},t]$. For each $t\in N_{u_{0}}$ put $U_{t}=(W\cap K_{t})\setminus\bigcup\{\varphi_{e}(\Sigma_{w(e)}):e\in{\mathcal{E}}_{t}\}$; put also $U_{u_{0}}=W\cap\Sigma_{u_{0}}$ and ${\mathcal{U}}=\{U_{u_{0}}\}\cup\{U_{t}:t\in N_{u_{0}}\}$. Note that for each $t\in N_{u_{0}}$, $U_{t}$ is a neighbourhood of $\xi$ in $K_{t}$, and similarly $U_{u_{0}}$ is a neighbourhood of $\xi$ in $\Sigma_{u_{0}}$. Moreover, it is not hard to see that the corresponding neigbourhood $V_{\mathcal{U}}(\xi)$ is then contained in $E$, as required. In the second subcase, we suppose that $u_{0}\notin V_{F}$, but $N_{u_{0}}\cap F\neq\emptyset$. For each $t\in N_{u_{0}}\setminus V_{F}$ put $U_{t}=K_{t}$. For $t\in N_{u_{0}}\cap V_{F}$, let ${\mathcal{E}}_{t}^{F}$ be the set of all edges in $F$ adjacent to $t$, and put $U_{t}=(W\cap K_{t})\setminus\bigcup\{\varphi_{e}(\Sigma_{w(e)}):e\in{\mathcal{E}}_{t}\}$. Put also $U_{u_{0}}=\Sigma_{u_{0}}$ and ${\mathcal{U}}=\{U_{u_{0}}\}\cup\{U_{t}:t\in N_{u_{0}}\}$. As before, note that for each $t\in N_{u_{0}}$, $U_{t}$ is a neighbourhood of $\xi$ in $K_{t}$, and similarly $U_{u_{0}}$ is a neighbourhood of $\xi$ in $\Sigma_{u_{0}}$. Likewise, it is not hard to see that the corresponding neigbourhood $V_{\mathcal{U}}(\xi)$ is then contained in $E$, as required. Finally, in the third subcase, we suppose that $N_{u_{0}}\cap F=\emptyset$. Denote by $t_{1}\in N_{u_{0}}$ this vertex which lies on the path from $u_{0}$ to $F$. Denote also by $u_{1}$ the white vertex adjacent to $t_{1}$ and distinct from $u_{0}$ which also lies on the path from $u_{0}$ to $F$. For any $t\in N_{u_{0}}\setminus\{t_{1}\}$, put $U_{t}=K_{t}$, and put $U_{t_{1}}=K_{t_{1}}\setminus\varphi_{[u_{1},t_{1}]}(\Sigma_{u_{1}})$. Put $U_{u_{0}}=\Sigma_{u_{0}}$ and ${\mathcal{U}}=\{U_{u_{0}}\}\cup\{U_{t}:t\in N_{u_{0}}\}$, and note again that $V_{\mathcal{U}}(\xi)$ is contained in $E$, as required.
This finishes the proof.
3.2. The reduced Bowditch JSJ splitting and its rigid cluster factors
Let $G$ be a 1-ended hyperbolic group which is not cocompact Fuchsian and let ${T}_{G}$ be the Bowditch JSJ tree of $G$. Let $\sim$ be the smallest equivalence relation on the vertex set of ${T}_{G}$ for which $u\sim v$ whenever $u$ and $v$ are rigid vertices at distance $2$. Equivalence classes of $\sim$ are either white vertex singletons, flexible vertex singletons or rigid classes (i.e. sets of rigid vertices). The convex hulls of the rigid classes are the rigid clusters of ${T}_{G}$. Each rigid cluster of ${T}_{G}$ is a subtree of $T_{G}$ (possibly a single vertex).
It is not hard to observe that embedded cycles in the quotient graph ${T}_{G}/{\sim}$ have length two, and there are no loop edges in this quotient. So, by replacing any multi-edge of ${T}_{G}/{\sim}$ with a single edge, we obtain a tree. The latter tree contains some vertices of degree 1, namely those which correspond to the white vertices in ${T}_{G}$ whose all neighbours are rigid. We delete all those vertices (and all edges adjacent to them), thus obtaining a tree denoted as ${T}^{r}_{G}$, which we call the reduced Bowditch JSJ tree for $G$. The action of $G$ descends to ${T}^{r}_{G}$ and there is a natural map ${T}_{G}\to{T}^{r}_{G}$ which is $G$-equivariant. (This $G$-equivariant map can be described as composition of the quotient map ${T}_{G}\to{T}_{G}/{\sim}$ with the map which collapses to points all those edges of the tree obtained from ${T}_{G}/{\sim}$ which are adjacent to the degree 1 vertices.) Technically speaking, the reduced Bowditch JSJ splitting for $G$ is the above induced action of $G$ on the tree ${T}^{r}_{G}$.
We now describe closer the nature of the just defined reduced Bowditch JSJ splitting, and in particular of its factors. Observe that the bipartite structure of the tree ${T}_{G}$ descends to ${T}^{r}_{G}$, so that we can speak of black and white vertices of the latter. The white vertices of ${T}^{r}_{G}$ are naturally in bijective correspondence with those white vertices of ${T}_{G}$ which are adjacent to at least one flexible vertex. Their stabilizers (for the induced action of $G$ on ${T}^{r}_{G}$) coincide with the corresponding stabilizers for the $G$-action on ${T}_{G}$, and so they are maximal 2-ended subgroups of $G$. There are two kinds of black vertices in ${T}^{r}_{G}$. The first kind of black vertices of ${T}^{r}_{G}$ are those in bijective correspondence with the flexible vertices of ${T}_{G}$, and so we still call them flexible. Their stabilizers under the action of $G$ on ${T}^{r}_{G}$ do not change, and we call them the flexible factors of the reduced Bowditch JSJ splitting of $G$. The second kind of black vertices of ${T}^{r}_{G}$ are those in bijective correspondence with the rigid clusters of ${T}_{G}$, and we call them the rigid cluster vertices of ${T}^{r}_{G}$. Their stabilizers, which coincide with the stabilizers of the corersponding rigid clusters in the $G$-action on ${T}_{G}$, are called the rigid cluster factors of the reduced Bowditch JSJ splitting of $G$.
The next observations follow directly from the above description.
Fact 3.4.
Let $v^{\prime}$ be a flexible vertex of the reduced Bowditch tree ${T}^{r}_{G}$, and let $v$ be the vertex of ${T}_{G}$ corresponding to $v^{\prime}$ (so that the vertex stabilizers $G(v^{\prime})$ and $G(v)$ coincide). Then the family of edge stabilizers of the edges adjacent to $v^{\prime}$ in ${T}^{r}_{G}$ coincides with the family of edge stabilizers of the edges adjacent to $v$ in ${T}_{G}$.
Fact 3.5.
Each rigid cluster vertex $v$ of the reduced Bowditch tree ${T}^{r}_{G}$ is isolated, i.e., there is no rigid cluster vertex in ${T}^{r}_{G}$ at distance 2 from $v$. Equivalently, all vertices of ${T}^{r}_{G}$ lying at distance 2 from a rigid cluster vertex $v$ are flexible.
Remark 3.6.
Given a rigid cluster $S\subset{T}_{G}$, and the corresponding rigid cluster factor $G(v_{S})$, consider the action of $G(v_{S})$ on $S$. The vertex stabilizers of this action are easily seen to coincide with the corresponding vertex stabilizers of the $G$-action on ${T}_{G}$. Thus, this action describes some splitting of the rigid cluster factor $G(v_{S})$, along maximal 2-ended subgroups, whose factors are some appropriate rigid factors of the Bowditch JSJ splitting of $G$. This observation is helpful for understanding the nature of rigid cluster factors, even though we will make no essential use of it.
Since the factors of any splitting of a hyperbolic group along quasi-convex subgroups are also quasi-convex (see e.g. Proposition 1.2 in [Bow98]), and since 2-ended subgroups are quasi-convex, we get the following.
Corollary 3.7.
Each rigid cluster factor of a 1-ended hyperbolic group $G$ is quasi-convex in $G$. In particular, each rigid cluster factor is a non-elementary hyperbolic group.
The reduced Bowditch JSJ splitting of a group $G$ induces in a natural way a tree system described as follows:
$\Theta^{r}(G):=({T}^{r}_{G},\{K_{t}=\partial G(t):t\in V^{b}_{{T}^{r}_{G}}\},\{\Sigma_{u}=\partial G(u):u\in V^{w}_{{T}^{r}_{G}}\},\{\varphi_{e}:e\in E_{{T}^{r}_{G}}\}),$
where each $\varphi_{e}:\partial G({w(e)})\to\partial G({b(e)})$ is the natural inclusion. The same argument as in the proof of Theorem 3.2 yields also the following result.
Theorem 3.8.
Let $G$ be a 1-ended hyperbolic group which is not cocompact Fuchsian, and let $\Theta^{r}(G)$ be the associated tree system (related to the reduced Bowditch JSJ splitting of $G$). Then the Gromov boundary $\partial G$ is naturally homeomorphic to the limit of the tree system $\Theta^{r}(G)$, i.e.,
$\partial G\cong\lim\Theta^{r}(G).$
4. General trees of graphs
In this section we describe trees of graphs—a class of topological spaces which is the main object of study in this paper. As we will see, these spaces are compact metrizable, and the most interesting subclass (that of trees of 2-connected graphs) consists of spaces that are connected, locally connected, cutpoint-free, and of topological dimension 1. Throughout all of this section we make use of the terminology and notation introduced in Section 2. We start with basic definitions.
Definition 4.1.
A tree system of punctured graphs is a tree system of spaces in which:
  1. (1)
    each constituent space with peripherals is a punctured graph (as described in Example 2.2);
  2. (2)
    each white vertex of the underlying tree has degree 2.
We will use the notation
$\Theta=(T,\{\Gamma_{t}^{\circ}\},\{\Sigma_{u}\},\{\varphi_{e}\})$
for a tree system of punctured graphs (where for each $t\in V^{b}_{T}$, $\Gamma_{t}$ is some finite essential graph, i.e., nonempty and without isolated vertices).
Definition 4.2.
A tree of graphs is any topological space which is homeomorphic to the limit of a tree system of punctured graphs.
4.1. Trees of graphs as inverse limits of graphs
We will now present an alternative description of trees of graphs, as inverse limits of some special inverse systems of graphs. (As we will see below, these inverse systems in fact coincide, up to isomorphism, with the systems ${\mathcal{S}}_{\Theta}$ as described in Section 2.6, associated to the corresponding tree systems $\Theta$ of punctured graphs.) This alternative approach to trees of graphs is related to (and generalizes) that given in Section 2 of [Ś21], where only a special case of the so called reflection trees of graphs has been treated, using a similar description.
Let $\Theta$ be a tree system of punctured graphs. For any graph $\Gamma_{t}$ whose punctured version $\Gamma_{t}^{\circ}$ appears as a constituent space in $\Theta$, denote by $D_{t}\subset|\Gamma_{t}|$ a countable dense subset of the realization $|\Gamma_{t}|$ containing all vertices of $\Gamma_{t}$. Note that such $D_{t}$ is unique up to a homeomorphism of $|\Gamma_{t}|$ that is the identity on the vertices and that preserves all the edges. Consider also the inverse system ${\mathcal{S}}_{\Theta}=[\{K_{F}^{\ast}\},\{f_{F_{1}F_{2}}\}]$ associated to $\Theta$. Observe that, viewing any black vertex $t\in V^{b}_{T}$ as a b-subtree of $T$, and denoting by $\{[\varphi_{[u,t]}(\Sigma_{u})]:u\in N_{t}\}$ the set of points of the quotient space $K_{t}^{\ast}$ obtained from $\Gamma_{t}^{\circ}$ by shrinking the peripheral sets $\varphi_{[u,t]}(\Sigma_{u}):u\in N_{t}$, there is a homeomorphism $h_{t}:K_{t}^{\ast}\to|\Gamma_{t}|$ such that:
  1. (1)
    $h_{t}$ maps the subset $\{[\varphi_{[u,t]}(\Sigma_{u})]:u\in N_{t}\}$ of $K_{t}^{*}$ onto the subset $D_{t}$ of $|\Gamma_{t}|$;
  2. (2)
    $h_{t}$ respects types of the shrunk peripheral sets $[\varphi_{[u,t]}(\Sigma_{u})]$, i.e., if $\varphi_{[u,t]}(\Sigma_{u})$ is the boundary of a normal neighbourhood of some vertex of $\Gamma$, then $h_{t}([\varphi_{[u,t]}(\Sigma_{u})])$ is equal to this vertex, and if $\varphi_{[u,t]}(\Sigma_{u})$ is the boundary of a normal neighbourhood of some interior point of some edge $a$ of $\Gamma$, then $h_{t}([\varphi_{[u,t]}(\Sigma_{u})])$ is equal to some interior point of the same edge $a$.
To better understand $h_{t}$, consider an edge $e\subset|\Gamma_{t}|$ and its punctured version $e^{\circ}\subset\Gamma_{t}^{\circ}\subset|\Gamma_{t}|$. Up to homeomorphism, we can view $e^{\circ}$ as the standard Cantor set on a closed interval $I\subset e$. Let $h\colon I\to I$ be the Cantor function with respect to $e^{\circ}$ and note that $h$ descends to a continuous map $h^{e}\colon(e^{\circ})^{*}\to I$ from the image $(e^{\circ})^{*}$ of $e^{\circ}$ in $K_{t}^{\ast}$. Then $h_{t}|_{(e^{\circ})^{*}}$ can be taken to be the composition of $h^{e}$ with an orientation preserving homeomorphism $I\to e$ which maps dyadic rationals in $I$ onto the subset $D_{t}\cap e$.
The above identifications of the spaces $K_{t}^{*}$ of the inverse system ${\mathcal{S}}_{\Theta}$ with realizations $|\Gamma_{t}|$ of the graphs $\Gamma_{t}$ induce further identifications of all other spaces $K_{F}^{*}$ of this inverse system with certain graphs $X_{F}$ which are obtained out of the graphs $\Gamma_{t}:t\in V^{b}_{F}$ by means of operations of “connected sum” for graphs. We now describe these graphs $X_{F}$, together with the description of maps between them, which correspond to the bonding maps of the inverse system ${\mathcal{S}}_{\Theta}$.
Let $X$ be the realization of a finite graph, and let $x\in X$. The blow-up of $X$ at $x$ is the space obtained from $X$ by deleting $x$, and by taking the completion of the resulting space which consists of adding as many points as the number of connected components into which $x$ splits its normal neighbourhood $B(x)$ in $X$. We denote this blow-up by $X^{\#}(x)$, and the set of points added at the completion by $P_{x}$ (we use the term blow-up divisor at $x$ for this set $P_{x}$). Similarly, if $J$ is a finite subset of $X$, the blow-up of $X$ at $J$, denoted $X^{\#}(J)$, is the space obtained from $X$ by performing blow-ups at all points of $J$ (in arbitrary order). Moreover, for any two finite subsets $J_{1}\subset J_{2}$ of $X$, denote by $\rho^{X}_{J_{2}J_{1}}:X^{\#}(J_{2})\to X^{\#}(J_{1})$ the blow-down map which maps the sets $P_{x}:x\in J_{2}\setminus J_{1}$ to the corresponding points $x$, and which is the identity on the remaining part of $X^{\#}(J_{2})$.
Now, let $F$ be a finite b-subtree of the underlying tree $T$ of the tree system $\Theta$. Recalling that we have identified graphs $|\Gamma_{t}|$ with the corresponding spaces $K_{t}^{*}$, for each $t\in V^{b}_{F}$ consider the subset $J_{t}^{F}\subset|\Gamma_{t}|$ given by $J_{t}^{F}:=\{h_{t}([\varphi_{[t,u]}(\Sigma_{u})]):u\in N^{F}_{t}\}$. Define $X_{F}$ as quotient of the disjoint union
$\bigsqcup_{t\in V^{b}_{F}}|\Gamma_{t}|^{\#}(J^{F}_{t})$
by the equivalence relation $\sim$ determined by the following gluing maps. Recall that, by condition (2) of Definition 4.1, for any white vertex $u$ of $F$ there are exactly two black vertices in $F$ adjacent to $u$. Denoting these adjacent black vertices by $t$ and $s$, consider the blow-up divisors $P_{h_{t}([\varphi_{[t,u]}(\Sigma_{u})])}\subset|\Gamma|^{\#}(J_{t}^{F})$ and $P_{h_{s}([\varphi_{[s,u]}(\Sigma_{u})])}\subset|\Gamma|^{\#}(J_{s}^{F})$. Note that there is a natural bijection
$b_{s,t}:P_{h_{s}([\varphi_{[s,u]}(\Sigma_{u})])}\to P_{h_{t}([\varphi_{[t,u]}(\Sigma_{u})])}$
induced by the map $\varphi_{[t,u]}\circ\varphi^{-1}_{[s,u]}:\varphi_{[s,u]}(\Sigma_{u})\to\varphi_{[t,u]}(\Sigma_{u})$ (under the canonical identification of each of the two divisors with the corresponding peripheral subspace), and that $b_{t,s}=b_{s,t}^{-1}$. The equivalence relation $\sim$ is then determined by all of the maps $b_{s,t}$ as above related to all white vertices $u$ of $F$. Note that the described quotient space $X_{F}$ may be viewed as obtained by appropriate iterated connected sum of the family $\Gamma_{t}:t\in V^{b}_{F}$ of graphs (see the paragraph right before Lemma 4.8 for a precise description of the concept of connected sum for graphs). Note also that $X_{F}$ is in a natural way homeomomorphic to the space $K^{*}_{F}$, via homeomorphism induced by the homeomorphisms $h_{t}^{-1}:|\Gamma_{t}|\to K^{*}_{t}$, for $t\in V^{b}_{F}$.
In order to make our alternative description of the inverse system ${\mathcal{S}}_{\Theta}$ complete, we need to describe the bonding maps between the spaces $X_{F}$ which, under the above identifications with the spaces $K_{F}^{*}$, correspond to the bonding maps $f_{FF^{\prime}}$. To do this, consider any two finite b-subtrees $F\subset F^{\prime}$ of $T$, and define the map $\pi_{FF^{\prime}}:X_{F^{\prime}}\to X_{F}$ to be a natural map from the larger connected sum to the smaller one, given as follows. For $t\in V^{b}_{F}$, the restriction of $\pi_{FF^{\prime}}$ to $|\Gamma_{t}|^{\#}(J_{t}^{F^{\prime}})$ (viewed as a subset in $X_{F^{\prime}}$) coincides with the blow-down map $\rho^{|\Gamma_{t}|}_{J_{t}^{F^{\prime}}J_{t}^{F}}$ (whose image $|\Gamma_{t}|^{\#}(J_{t}^{F})$ is viewed as a subset in $X_{F}$). For $t\in V^{b}_{F^{\prime}}\setminus V^{b}_{F}$, the subset $|\Gamma_{t}|^{\#}(J_{t}^{F^{\prime}})\subset X_{F^{\prime}}$ is mapped by $\pi_{FF^{\prime}}$ to the point $h_{s}([\varphi_{[u,s]}(\Sigma_{u})])\in|\Gamma_{s}|^{\#}(J_{s}^{F})\subset X_{F}$, where $s\in V^{b}_{F}$ is the vertex of the subtree $F$ that is closest to $t$, and where $u\in V^{w}_{F^{\prime}}$ is the vertex adjacent to $s$ on the shortest path in $T$ from $t$ to $s$.
We skip a straightforward verification of the following fact, the second part of which follows from the first part in view of Proposition 2.8.
Fact 4.3.
The inverse systems ${\mathcal{S}}_{\Theta}=[\{K_{F}^{*}\},\{f_{FF^{\prime}}\}]$ and [$[\{X_{F}\},\{\pi_{FF^{\prime}}\}]$ are isomorphic. As a consequence, we have
$\lim\Theta\cong\lim_{\longleftarrow}[\{X_{F}\},\{\pi_{FF^{\prime}}\}].$
Expression of a tree of graphs $\lim\Theta$ as the inverse limit of graphs $X_{F}$, as above, turns out to be very useful for establishing various topological properties of trees of graphs. The next corollary records the most obvious such properties (which follow directly from the corresponding well known properties of inverse limits, e.g. the estimate of the topological dimension of the limit by supremum of the topological dimensions of the terms in the inverse system). Further properties, for a subclass of trees of 2-connected graphs, are discussed in the next subsection.
Corollary 4.4.
Any tree of graphs is a compact metrizable topological space of topological dimension $\leq 1$.
4.2. Trees of 2-connected graphs
We now discuss a subclass in the class of trees of graphs, so called trees of 2-connected graphs, which is most interesting from the perspective of studying the topology of Gromov boundaries of hyperbolic groups. As we will see, the spaces from this subclass are connected, locally connected, cutpoint-free, and they have topological dimension equal to 1.
A graph is 2-connected if it is finite, connected, nonempty, not equal to a single vertex, and contains no cutpoint. We say that a tree system of punctured graphs $\Theta=(T,\{\Gamma_{t}^{\circ}\},\{\Sigma_{u}\},\{\varphi_{e}\})$ in which all the graphs $\Gamma_{t}$ are 2-connected is a tree system of punctured 2-connected graphs; likewise, a tree of 2-connected graphs is the limit of a tree system of punctured 2-connected graphs.
In our analysis of trees of 2-connected graphs we will need the following rather well known observation.
Fact 4.5.
Let ${\mathcal{S}}=(\{K_{\alpha}\}_{\alpha},\{f_{\alpha\alpha^{\prime}}\}_{\alpha\leq\alpha^{\prime}})$ be an inverse system of compact metric spaces with surjective bonding maps. Let $C$ be a closed subspace of some $K_{\alpha}$. Then the preimage $f_{\alpha}^{-1}(C)$ of $C$ in the inverse limit of $\mathcal{S}$ is disconnected if and only if $f_{\alpha\hat{\alpha}}^{-1}(C)$ is disconnected for some $\hat{\alpha}\geq\alpha$.
For what follows, we use the notation $\pi_{0}(X)$ to denote the set of connected components of a space $X$.
Lemma 4.6.
Let $X$ be the realization of a 2-connected graph and let $x$ and $y$ be two points of $X$. The inclusion $\iota_{x,y}\colon P_{x}\hookrightarrow X^{\#}(\{x,y\})$ is $\pi_{0}$-surjective. Consequently, the map $\iota_{x,y}$ is $\pi_{0}$-bijective if and only if $\deg(x)$ is equal to the number of components of $X^{\#}(\{x,y\})$.
Proof.
The blow-down map $X^{\#}(\{x,y\})\to X^{\#}(y)$ is the quotient map identifying all points of $P_{x}$. Since $X$ has no cutpoints, the blow-up $X^{\#}(y)$ is connected. It follows that $P_{x}$ intersects every component of $X^{\#}(\{x,y\})$, i.e., the inclusion $\iota_{x,y}$ is $\pi_{0}$-surjective.
The last sentence of the statement follows from the fact that $|P_{x}|=\deg(x)$ and $\deg(x)$ is finite. ∎
Remark 4.7.
Let $X$ be the realization of a finite graph and let $J$ be a finite subset of $X$. The inclusion $X\setminus J\hookrightarrow X^{\#}(J)$ is $\pi_{0}$-bijective. Thus there exists a map $\bigcup_{x\in J}P_{x}\to\pi_{0}(X\setminus J)$ given by composing $\bigcup_{x\in J}P_{x}\to\pi_{0}\bigl{(}X^{\#}(J)\bigr{)}$ with the $\pi_{0}$-inverse $\pi_{0}\bigl{(}X^{\#}(J)\bigr{)}\to\pi_{0}(X\setminus J)$ of the inclusion. By Lemma 4.6, if $X$ is 2-connected and $J=\{x,y\}$ then the restriction $P_{x}\to\pi_{0}(X\setminus\{x,y\})$ is surjective always and is bijective if $\deg(x)=\bigl{|}\pi_{0}(X\setminus\{x,y\})\bigr{|}$.
The description in Section 4.1 of trees of punctured graphs as limits of inverse systems of graphs alludes to the notion of a connected sum of graphs. Here we give an explicit definition. If $X_{1}$ and $X_{2}$ are realizations of finite graphs with $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$ and $\ell\colon P_{x_{1}}\to P_{x_{2}}$ is a bijection between the blow-up divisor $P_{x_{1}}$ of $x_{1}$ in $X_{1}$ and the blow-up divisor $P_{x_{2}}$ of $x_{2}$ in $X_{2}$ then the connected sum $X_{1}\#_{\ell}X_{2}$ is the quotient of $X_{1}^{\#}(x_{1})\sqcup X_{2}^{\#}(x_{2})$ by the equivalence relation identifying each $p\in P_{x_{1}}$ with $\ell(p)\in P_{x_{2}}$. For $i\in\{1,2\}$, the projection $X_{1}\#_{\ell}X_{2}\to X_{i}$ is the map that is identity on $X_{i}\setminus\{x_{i}\}=X_{i}^{\#}(x_{i})\setminus P_{x_{i}}\subset X_{1}\#_{\ell}X_{2}$ and that sends the remaining points of $X_{1}\#_{\ell}X_{2}$ to $x_{i}$.
Lemma 4.8.
Let $X$ be a connected and locally connected $T_{1}$ space, let $x\in X$ and let $C$ be a component of $X\setminus\{x\}$. Then the closure $\bar{C}$ in $X$ of $C$ is $C\cup\{x\}$.
Proof.
Since $X$ is $T_{1}$, the set $X\setminus\{x\}$ is open in $X$ and so, by local connectedness, the components of $X\setminus\{x\}$ are open in $X$. In particular, the component $C$ is open and so, since $X$ is connected and $C$ is a proper subset of $X$, it cannot be that $C$ is also closed in $X$. It suffices then to show that $\bar{C}\subset C\cup\{x\}$. But the complement of $C\cup\{x\}$ is the union of the non-$C$ components of $X\setminus\{x\}$, which is open. Thus $C\cup\{x\}$ is closed and so $\bar{C}$ is indeed contained in $C\cup\{x\}$. ∎
Corollary 4.9.
Let $X$ be a connected, locally connected, cutpoint-free $T_{1}$ space. Let $x,y\in X$ and let $C$ be a component of $X\setminus\{x,y\}$. The closure $\bar{C}$ of $C$ in $X$ is $C\cup\{x,y\}$.
Proof.
If $x=y$ then we can directly apply Lemma 4.8 so we assume that $x\neq y$. Put $X_{y}=X\setminus\{y\}$ and note that $X_{y}$ is connected (because $X$ is cutpoint-free). Then $C$ is a component of $X_{y}\setminus\{x\}$ so, by Lemma 4.8, the closure of $C$ in $X_{y}$ is $C\cup\{x\}$. Then $\bar{C}$ is either $C\cup\{x\}$ or $C\cup\{x,y\}$. Swapping the roles of $x$ and $y$ we see that $\bar{C}=C\cup\{x,y\}$. ∎
Lemma 4.10.
Let $X_{1}$ and $X_{2}$ be realizations of graphs, suppose that $X_{2}$ is 2-connected, and let $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$. Let $\ell\colon P_{x_{1}}\to P_{x_{2}}$ be a bijection between the blow-up divisor $P_{x_{1}}$ of $x_{1}$ in $X_{1}$ and the blow-up divisor $P_{x_{2}}$ of $x_{2}$ in $X_{2}$. Let $A$ be a connected subspace of $X_{1}$. Then the preimage of $A$ under the projection $X_{1}\#_{\ell}X_{2}\to X_{1}$ is connected.
Proof.
If $A$ is disjoint from $x_{1}$ then the preimage is homeomorphic to $A$ so we may assume, without loss of generality, that $x_{1}\in A$. Any connected subspace of a graph is locally connected so, by Lemma 4.8, the closure in $X_{1}$ of every component of $A^{-}=A\setminus\{x_{1}\}$ contains $x_{1}$. Thus we see that the closure in $X_{1}^{\#}(x_{1})$ of every component of $A^{-}$ intersects the blow-up divisor $P_{x_{1}}$. The preimage of $A$ in $X_{1}\#_{\ell}X_{2}$ is the union of $A^{-}$ and $X_{2}^{\#}(x_{2})$, where the latter is connected since $X_{2}$ is 2-connected. But $A^{-}\cup X_{2}^{\#}(x_{2})$ is equal to the union of $X_{2}^{\#}(x_{2})$ and the closures of the components of $A^{-}$ in $A^{-}\cup P_{x_{1}}$, each of which intersects $P_{x_{1}}=P_{x_{2}}\subset X_{2}^{\#}(x_{2})$. ∎
Lemma 4.11.
Let $X_{1}$ and $X_{2}$ be realizations of 2-connected graphs and let $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$. Let $\ell\colon P_{x_{1}}\to P_{x_{2}}$ be a bijection between the blow-up divisor $P_{x_{1}}$ of $x_{1}$ in $X_{1}$ and the blow-up divisor $P_{x_{2}}$ of $x_{2}$ in $X_{2}$. Then the connected sum $X=X_{1}\#_{\ell}X_{2}$ is 2-connected.
Proof.
It is clear that $X$ is not empty and not equal to a single point. So we need only to show that $X$ is connected and has no cutpoint. That $X$ is connected follows from Lemma 4.10 by setting $A=X_{1}$.
To see that $X$ has no cutpoints, let $p\in X$. Up to homeomorphism, either $p\in X_{1}^{\#}(x_{1})\setminus P_{x_{1}}$ or $p\in X_{2}^{\#}(x_{2})\setminus P_{x_{2}}$ so we may assume, without loss of generality, that $p\in X_{1}^{\#}(x_{1})\setminus P_{x_{1}}$. Then $X\setminus\{p\}$ is the preimage of $A=X_{1}\setminus\{p\}$ under the projection map $X_{1}\#_{\ell}X_{2}\to X_{1}$. Since $X_{1}$ has no cutpoints, the subspace $A$ is connected so, by Lemma 4.10, its preimage under the projection map is connected. ∎
Lemma 4.12.
Let $X$ be the realization of a finite graph and let $A\subset X$ be a connected subspace. Then $A$ is an ascending union of compact connected subspaces of $X$.
Proof.
Let $\bar{A}$ be the closure of $A$. Since $X$ is the realization of a finite graph, the set $\bar{A}\setminus A$ must be finite since otherwise there would exist a common accumulation point of $\bar{A}\setminus A$ and $A$ in a closed edge, which would contradict connectedness of $A$.
We metrize $X$ as a geodesic metric space with each edge isometric to the unit interval $[0,1]$. Let $A_{n}=A\setminus N_{\frac{1}{n}}(\bar{A}\setminus A)$ where $N_{\frac{1}{n}}(\bar{A}\setminus A)$ is the set of all points at distance less than $\frac{1}{n}$ to $\bar{A}\setminus A$. Then $A_{n}=\bar{A}\setminus N_{\frac{1}{n}}(\bar{A}\setminus A)$, and so it is closed. Thus the $A_{n}$ are compact. Moreover, since $\bar{A}\setminus A$ is finite, the ascending union $\bigcup_{n}A_{n}$ is equal to $A$. It remains to show that the $A_{n}$ are connected for $n$ large enough.
Let $L$ be large enough that for any $x\in\bar{A}\setminus A$, the metric ball $B_{\frac{1}{L}}(x)$ intersects no vertex of $X$ aside from $x$, in the case where $x$ is itself a vertex. Take $p,q\in A_{n}$ with $n\geq L$. Since $A$ is a connected subspace of a graph, there exists an embedded path $P\subset A$ with endpoints $p$ and $q$. We claim that $P$ is disjoint from $N_{\frac{1}{n}}(\bar{A}\setminus A)$. For the sake of finding a contradiction, assume that $P$ intersects $N_{\frac{1}{n}}(\bar{A}\setminus A)$, then $P$ intersects $B_{\frac{1}{n}}(x)$ for some $x\in\bar{A}\setminus A$. But $x\notin A$ so the intersection $B_{\frac{1}{n}}(x)\cap A$ is contained in $B_{\frac{1}{n}}(x)\setminus\{x\}$, which is a disjoint union of open segments since $n\geq L$. Let $U\cap P$ be the intersection of one of these open segments $U$ with $P$. Since $\bar{U}$ contains $x$ and $P$ is closed and disjoint from $x\notin A$, we have $U\cap P\subsetneq U$. This implies that an endpoint of $P$ is contained in $U\subset B_{\frac{1}{n}}(x)\subset N_{\frac{1}{n}}(\bar{A}\setminus A)$, which contradicts $p,q\in A_{n}$. Thus $P$ is disjoint from $N_{\frac{1}{n}}(\bar{A}\setminus A)$ and so is contained in $A_{n}$. ∎
Lemma 4.13.
Let $\Theta$ be a tree system of punctured 2-connected graphs with limit $Z$. Let ${\mathcal{S}}_{\Theta}=(\{K_{F}^{\ast}:F\in{\mathcal{F}}_{T}^{b}\},\{f_{FF^{\prime}}:F\subset F^{\prime}\})$ be the standard inverse system associated to $\Theta$. Let $F$ be a finite b-subtree of $T$ and let $A$ be a connected subspace of the reduced partial union graph $K_{F}^{\ast}$. Then the preimage of $A$ in $Z$ is connected.
Proof.
By Lemma 4.12, the connected subspace $A$ is a union $\bigcup_{n}C_{n}$ of compact connected subspaces $C_{1}\subset C_{2}\subset\cdots$. Thus the preimage $f_{F}^{-1}(A)$ of $A$ is an ascending union of the preimages $f_{F}^{-1}(C_{n})$ of the $C_{n}$, where $f_{F}\colon Z\to K_{F}^{\ast}$ is the projection. It suffices then to prove that the $f_{F}^{-1}(C_{n})$ are connected or, more generally, that the preimage in $Z$ of any compact connected subspace $C$ of $K_{F}^{\ast}$ is connected.
By Fact 4.5, it suffices to prove that $f_{FF^{\prime}}^{-1}(C)$ is connected for any b-subtree $F^{\prime}\supset F$. We prove this by induction on the size of $F^{\prime}\setminus F$. The base case $F^{\prime}=F$ is immediate. Otherwise, the reduced partial union graph $K_{F^{\prime}}^{\ast}$ is the connected sum of the reduced partial union graph $K_{F^{\prime\prime}}^{\ast}$ and the realization of a 2-connected graph, for some b-subtree $F^{\prime\prime}$ with $F^{\prime}\supsetneq F^{\prime\prime}\supset F$. Then $f_{FF^{\prime}}^{-1}(C)=f_{F^{\prime\prime}F^{\prime}}^{-1}(\hat{C})$ where $\hat{C}=f_{FF^{\prime\prime}}^{-1}(C)$, which is connected by the inductive hypothesis. By Lemma 4.10, the preimage $f_{F^{\prime\prime}F^{\prime}}^{-1}(\hat{C})=f_{FF^{\prime}}^{-1}(C)$ is connected. ∎
Lemma 4.14.
Let $\Theta$ be a tree system of punctured 2-connected graphs. Then the limit $Z$ of $\Theta$ is connected, locally connected and cutpoint-free.
Proof.
Let ${\mathcal{S}}_{\Theta}=(\{K_{F}^{\ast}:F\in{\mathcal{F}}_{T}^{b}\},\{f_{FF^{\prime}}:F\subset F^{\prime}\})$ be the standard inverse system associated to $\Theta$. By Lemma 4.11, any reduced partial union $K_{F}^{\ast}$ is connected. Then, since an inverse limit of connected metric compacta is connected (or by Fact 4.5), $Z$ is connected as well.
Let $p\in Z$. Let $U$ be an open neighborhood of $p$. Then there is an open subneighborhood of $p$ of the form $f_{F}^{-1}(V)$ where $f_{F}\colon Z\to K_{F}^{\ast}$ is the projection to some reduced partial union graph $K_{F}^{\ast}$ and $V$ is an open neighborhood of $f_{F}(p)$ in $K_{F}^{\ast}$. Since graphs are locally connected, there is a connected open subneighborhood $W\subset V$ of $f_{F}(p)$. By Lemma 4.13, the preimage $f_{F}^{-1}(W)$ is connected. Thus $f_{F}^{-1}(W)$ is a connected open neighborhood of $p$ contained in $U$ and this shows that $Z$ is locally connected.
Let $p\in Z$. To establish that $Z$ is cutpoint-free, we will show that $Z\setminus\{p\}$ is connected. Let $F_{1}\subset F_{2}\subset\cdots$ be an ascending sequence of finite b-subtrees with $T=\bigcup_{n}F_{n}$. Since this sequence is cofinal in the poset ${\mathcal{F}}_{T}^{b}$, the set $Z\setminus\{p\}$ is the ascending union of the preimages of the sets $K_{F_{n}}^{\ast}\setminus\{f_{F_{n}}(p)\}$. It is thus sufficient to show that all these preimages are connected. However, by Lemma 4.11 (appropriately iterated), the subspaces $K_{F_{n}}^{\ast}\setminus\{f_{F_{n}}(p)\}$ are all connected and so, by applying Lemma 4.13, their preimages in $Z$ are connected as well. ∎
In the next lemma we show that, for trees of 2-connected graphs, the estimate for topological dimension given in Corollary 4.4 actually yields equality.
Lemma 4.15.
Let $Z$ be a tree of 2-connected graphs. Then $Z$ is of topological dimension 1.
Proof.
Note that by the general estimate for the dimension of inverse limits, and by the fact that $\dim X_{F}=1$ for all $F$, we get that $\dim Z\leq 1$. To prove the converse estimate, we will show that $Z$ contains an embedded copy of the circle $S^{1}$. Consider any sequence $(F_{n})$ of finite b-subtrees of $T$ with the following properties:
  1. (0)
    $F_{1}$ is a subtree consisting of a single black vertex, say $t_{1}$;
  2. (1)
    for each $n\geq 1$ we have $F_{n}\subset F_{n+1}$ and $F_{n+1}$ has exactly one black vertex, say $t_{n+1}$, not contained in $V^{b}_{F_{n}}$;
  3. (2)
    $\bigcup_{n\geq 1}F_{n}=T$.
Obviously, such a sequence $(F_{n})$ always exists, and it is cofinal in the poset of all finite b-subtrees of $T$ (ordered by inclusion). Consequently, if
=[${\mathcal{I}}=[\{K^{*}_{F_{n}}:n\geq 1\},\{f_{F_{n}F_{n+1}}:n\geq 1\}]=[\{X_{F_{n}}:n\geq 1\},\{\pi_{F_{n}F_{n+1}}:n\geq 1\}]$
is the inverse sequence obtained by restriction of ${\mathcal{S}}_{\Theta}$ to the subposet $(F_{n})$, then the corresponding tree of graphs $\lim\Theta$ is homeomorphic to the inverse limit $\lim_{\leftarrow}{\mathcal{I}}$. Note that, since $X_{F_{1}}$ has no cut vertex, it contains a cycle, and we pick one such cycle $C_{1}\subset X_{F_{1}}$. We next describe inductively a sequence of cycles $C_{n}\subset X_{F_{n}}$ such that for each $n\geq 1$ we have $C_{n+1}\subset\pi_{F_{n}F_{n+1}}^{-1}(C_{n})$. Having defined the cycle $C_{n}$, let $x\in X_{F_{n}}$ be the point such that $X_{F_{n+1}}$ is obtained by connected sum of $X_{F_{n}}$ with $\Gamma_{t_{n+1}}$ at $x$. More precisely, view $X_{F_{n+1}}$ as obtained from the blow-ups $X_{F_{n}}^{\#}(x)$ and $\Gamma_{t_{n+1}}^{\#}(y)$ by appropriate gluing of the blow-up divisors. If $x\notin C_{n}$, we may view $C_{n}$ as a subset of $X_{F_{n+1}}$, and we put $C_{n+1}:=C_{n}$ in this case. If $x\in C_{n}$, let $p,q$ be the points of the blow-up divisor $P_{x}\subset X_{F_{n}}^{\#}(x)$ corresponding to these two components into which $x$ locally splits $X_{F_{n}}$ which intersect $C_{n}$. Put also $C_{n}^{\#}=(C_{n}\setminus\{x\})\cup\{p,q\}\subset X_{F_{n}}^{\#}(x)$. Denote by $p^{\prime}$ and $q^{\prime}$, respectively, the images of $p$ and $q$ in $P_{y}\subset\Gamma_{t_{n+1}}^{\#}(y)$ through the gluing map of the corresponding connected sum. Consider also an arc $A\subset\Gamma_{t_{n+1}}^{\#}(y)$ connecting $p^{\prime}$ with $q^{\prime}$ (which exists by the assumption that $\Gamma_{t}$ has no cut vertex). Put $C_{n+1}:=C_{n}^{\#}\cup A$.
Now, we get an inverse sequence
${\mathcal{C}}=[(C_{n})_{n\geq 1},(\pi_{F_{n}F_{n+1}}|_{C_{n+1}})_{n\geq 1}]$
whose limit $\lim_{\leftarrow}{\mathcal{C}}$ is obviously a subspace of $Z$. Observe that, since $\mathcal{C}$ is a sequence of circles, with bonding maps which are near-homeomorphisms (i.e. can be approximated by homeomorphisms), by the result of M. Brown [Bro60] the limit $\lim_{\leftarrow}{\mathcal{C}}$ is also a circle, which completes the proof. ∎
5. Necessary conditions for tree of graphs boundary
This section is devoted to the proofs of the implications (1$\Rightarrow$(3) of Theorem A and (1$\Rightarrow$(3) of Theorem B.
5.1. Proof of the impliciation (1$\Rightarrow$(3) of Theorem A
In this subsection we will introduce subnecklaces, prove some results about sufficiently connected trees of graphs and use these results to prove the impliciation (1$\Rightarrow$(3) of Theorem A.
The following definition will allow us to relate the topology of trees of graphs with the topology of boundaries of $1$-ended hyperbolic groups.
Definition 5.1.
Let $X$ be a connected, locally connected, cutpoint-free, compact Hausdorff space. An infinite subspace $N\subset X$ is a subnecklace if it satisfies the following conditions.
  1. (1)
    For any $p,q\in N$, the space $X\setminus\{p,q\}$ has two components.
  2. (2)
    For any $p\in N$, the space $X\setminus\{p\}$ has two ends (i.e. $p$ is a local cutpoint of degree 2).
Remark 5.2.
Let $G$ be a 1-ended hyperbolic group which is not cocompact Fuchsian. Recall the description in Section 3.1 of the classes of local cutpoints of $\partial G$. The subnecklaces of $\partial G$ are precisely the infinite subsets of the necklace classes of $\partial G$.
We now turn to the analysis of subnecklaces in connected, locally connected and cutpoint-free trees of graphs. We start with terminological preparations, and with some preparatory technical observations. Let $\Theta=(T,\{\Gamma_{t}^{\circ}\},\{\Sigma_{u}\},\{\varphi_{e}\})$ be a tree system of punctured graphs, let ${\mathcal{S}}_{\Theta}=(\{K_{F}^{\ast}:F\in{\mathcal{F}}_{T}^{b}\},\{f_{FF^{\prime}}:F\subset F^{\prime}\})$ be the standard inverse system associated to $\Theta$ and assume that the limit $Z=\lim\Theta=\lim_{\leftarrow}\mathcal{S}_{\Theta}$ is connected and locally connected. As described in Section 4.1, for a finite b-subtree $F$ of $T$, the reduced partial union $K_{F}^{\ast}=K_{F}/\Omega_{F}$ of $\Theta$ is a graph. Let $D_{F}$ be the image of the union of the peripheral subspaces $\Omega_{F}$ in $K_{F}^{\ast}$ and note that $D_{F}$ is dense, countable and contains the vertex set of $K_{F}^{\ast}$. Note that any $p\in K_{F}^{\ast}\setminus D_{F}$ has singleton preimage in the limit $Z$. We thus call any such $p$ the stable point of $K_{F}^{\ast}$. We will casually abuse notation and refer to the preimage in $Z$ or $K_{F^{\prime}}^{\ast}$ of a stable point $p\in K_{F}^{\ast}$ by the same name $p$.
Let $K$ be a space that is homeomorphic to the realization of a graph. A topological edge $e$ of $K$ is a subspace $e\subset K$ satisfying the following conditions:
  1. (1)
    $e$ is homeomorphic to $[0,1]\subset\mathbb{R}$;
  2. (2)
    aside from its endpoints, every point of $e$ has degree two in $K$.
If $K$ is the realization of a graph $\Gamma$ then every edge of $\Gamma$ is a topological edge of $K$.
Lemma 5.3.
Let ${\mathcal{S}}_{\Theta}=(\{K_{F}^{\ast}:F\in{\mathcal{F}}_{T}^{b}\},\{f_{FF^{\prime}}:F\subset F^{\prime}\})$ be the standard inverse system associated to a tree system of punctured graphs $\Theta$ and assume that the limit $Z$ is connected, locally connected and has no cutpoints. Let $F$ be a finite b-subtree of $T$ and let $e$ be a topological edge of $K_{F}^{\ast}$ with stable endpoints. Then $f_{F}^{-1}(e)$ is connected, where $f_{F}\colon Z\to K_{F}^{\ast}$ is the projection.
Remark.
Note that in the case when $Z$ is a tree of 2-connected graphs, the assertion above follows by Lemma 4.13. Here we will get this assertion in a more general setting.
Proof.
Let $p$ and $q$ be the endpoints of $e$. By Fact 4.5, it suffices to prove that $f_{FF^{\prime}}^{-1}(e)$ is connected for every $F^{\prime}\supset F$. For the sake of reaching a contradiction, suppose that $e^{\prime}=f_{FF^{\prime}}^{-1}(e)$ is not connected for some $F^{\prime}\supset F$. Let $C$ be the component of $e^{\prime}$ that contains the endpoint $p$. Since $e^{\prime}$ is closed in $K_{F^{\prime}}^{\ast}$, the set $V=K_{F^{\prime}}^{\ast}\setminus C$ is open and nonempty in $K_{F^{\prime}}^{\ast}$. Since $e^{\prime}$ is obtained from $e$ by performing finitely many connected sum operations, the set $U=C\setminus\{p\}$ must also be nonempty.
We will argue that $q\notin C$ so that $U=C\cap\bigl{(}e^{\prime}\setminus\{p,q\}\bigr{)}=C\cap f_{FF^{\prime}}^{-1}(e\setminus\{p,q\})$. For the sake of reaching a contradiction, suppose $C$ contains $q$ in addition to $p$. Let $C^{\prime}$ be the union of the remaining components of $e^{\prime}$. Then $C^{\prime}$ is open in $f_{FF^{\prime}}^{-1}(e\setminus\{p,q\})$, which is open in $K_{F^{\prime}}^{\ast}$, so $C^{\prime}$ is open in $K_{F^{\prime}}^{\ast}$. On the other hand, since $e^{\prime}$ has finitely many components (being homeomorphic to a finite graph), we see that $C^{\prime}$ is closed in $e^{\prime}$, which is closed in $K_{F^{\prime}}^{\ast}$, so $C^{\prime}$ is closed in $K_{F^{\prime}}^{\ast}$. Then $K_{F^{\prime}}^{\ast}$ is disconnected. Since the bonding maps in ${\mathcal{S}}_{\Theta}$ are obviously surjective, this contradicts connectedness of $Z$ by Fact 4.5. So we have $U=C\cap f_{FF^{\prime}}^{-1}(e\setminus\{p,q\})$.
Since $e^{\prime}$ has finitely many components, the component $C$ must be open in $e^{\prime}$. Then $U=C\cap f_{FF^{\prime}}^{-1}(e\setminus\{p,q\})$ is open in $f_{FF^{\prime}}^{-1}(e\setminus\{p,q\})$, which is open in $K_{F^{\prime}}^{\ast}$. We conclude that $U$ and $V$ are nonempty disjoint open sets of $K_{F^{\prime}}^{\ast}$ with $U\sqcup V=K_{F^{\prime}}^{\ast}\setminus\{p\}$. By taking preimages in $Z$ and recalling that $p$ is a stable point we see that the preimage of $p$ in $Z$ is a cutpoint of $Z$, a contradiction. ∎
Corollary 5.4.
Let ${\mathcal{S}}_{\Theta}=(\{K_{F}^{\ast}:F\in{\mathcal{F}}_{T}^{b}\},\{f_{FF^{\prime}}:F\subset F^{\prime}\})$ be the standard inverse system associated to a tree system of punctured graphs $\Theta$ and assume that the limit $Z$ is connected, locally connected and has no cutpoints. Let $F$ be a finite b-subtree of $T$ and let $e$ be an open interval of a topological edge of $K_{F}^{\ast}$. Then $f_{F}^{-1}(e)$ is connected, where $f_{F}\colon Z\to K_{F}^{\ast}$ is the projection.
Proof.
Every such $e$ is an ascending union $e_{1}\subset e_{2}\subset\cdots$ of topological edges with stable endpoints. Then $f_{F}^{-1}(e)$ is an ascending union of the spaces $f_{F}^{-1}(e_{n})$, which are all connected by Lemma 5.3. ∎
Lemma 5.5.
Let $f\colon X\to Y$ be a continuous surjective map of compact metric spaces and let $p\in Y$ have singleton preimage $\{\hat{p}\}$ under $f$. Let $U$ be an open neighbourhood of $\hat{p}$ in $Z$. Then there exists an open neighborhood $V$ of $p$ such that $f^{-1}(V)\subset U$.
Proof.
Suppose no such $V$ exists. Then, for each $n\in\mathbb{N}$, there exists a point $p_{n}\in X\setminus U$ such that $f(p_{n})\in B_{\frac{1}{n}}(p)$. Since $X$ is compact, there is a convergent subsequence $(p_{n_{k}})_{k}$. The limit $p^{\prime}$ of this subsequence is also contained in the closed set $X\setminus U$. In particular $p^{\prime}\neq\hat{p}$. But, by continuity of $f$, we have $f(p^{\prime})=\lim_{k\to\infty}f(p_{n_{k}})=p$, contradicting $f^{-1}(p)=\{\hat{p}\}$. ∎
Proposition 5.6.
Let ${\mathcal{S}}_{\Theta}=(\{K_{F}^{\ast}:F\in{\mathcal{F}}_{T}^{b}\},\{f_{FF^{\prime}}:F\subset F^{\prime}\})$ be the standard inverse system associated to a tree system of punctured graphs $\Theta$ and assume that the limit $Z$ is connected, locally connected and has no cutpoints. Let $F$ be a finite b-subtree of $T$ and let $e$ be a topological edge of the graph $K_{F}^{\ast}$. Then the preimage $N$ in $Z$ of the set of stable points of $e$ is a subnecklace of $Z$.
Proof.
The stable points of $K_{F}^{\ast}$ are contained in the interiors of edges so, if an endpoint of $e$ is stable we can slightly enlarge $e$ to avoid this. Thus, without loss of generality, the endpoints of $e$ are unstable. We will start by proving that $N$ satisfies condition (1) of Definition 5.1. Let $a$ and $b$ be the endpoints of $e$. Let $p$ and $q$ be distinct stable points of $e$. Then $p,q\in e\setminus\{a,b\}$. For points $x,y\in e$ let $(x,y)$ denote the open interval of $e$ bounded by $x$ and $y$. Without loss of generality, the components of $e\setminus\{a,b,p,q\}$ are $(a,p)$, $(p,q)$ and $(q,b)$. By Corollary 5.4, the preimage $f_{F}^{-1}\bigl{(}(p,q)\bigr{)}$ is connected. So, to ensure that $Z\setminus\{p,q\}$ has two components it suffices to prove that $Y=Z\setminus\Bigl{(}f_{F}^{-1}\bigl{(}(p,q)\bigr{)}\cup\{p,q\}\Bigr{)}$ is connected.
For the sake of reaching a contradiction assume that $Y$ is disconnected. Then there exist open subsets $U$ and $V$ of $Z$ such that $Y\subset U\cup V$, $Y\cap U\neq\emptyset$ and $Y\cap V\neq\emptyset$. Since $Y$ is open in $Z$ we can assume that $U$ and $V$ are subsets of $Y$ so that $U\sqcup V=Y$. By Corollary 5.4, the preimages $f_{F}^{-1}\bigl{(}(p,b)\bigr{)}$ and $f_{F}^{-1}\bigl{(}(q,b)\bigr{)}$ are connected. The latter preimage $f_{F}^{-1}\bigl{(}(q,b)\bigr{)}$ is contained in $Y$ so, without loss of generality, we have $f_{F}^{-1}\bigl{(}(q,b)\bigr{)}\subset V$. The intersection $Y\cap f_{F}^{-1}\bigl{(}(p,b)\bigr{)}$ is equal to $f_{F}^{-1}\bigl{(}(q,b)\bigr{)}\subset V$ so $V^{\prime}=V\cup f_{F}^{-1}\bigl{(}(p,b)\bigr{)}$ is open and disjoint from $U$. But then $U\cup V^{\prime}=Y\cup f_{F}^{-1}\bigl{(}(p,b)\bigr{)}=Z\setminus\{p\}$ which is a contradiction since $Z$ has no cutpoints. Thus $Y$ is connected. This completes the proof of condition (1).
It remains to prove that $N$ satisfies condition (2) of Definition 5.1. Let $p$ be a stable point of $e$. Then $p\in e\setminus\{a,b\}$. Let $(U_{n})_{n\in\mathbb{N}}$ be a descending sequence of open intervals of $e$ that contain $p$ such that $\cap_{n}U_{n}=\{p\}$. By Lemma 5.5 and Corollary 5.4, the sets $\bigl{(}f_{F}^{-1}(U_{n})\bigr{)}_{n}$ form a basis of connected open neighbourhoods of $p\in Z$. Since $U_{n}\setminus\{p\}$ is disconnected and $f_{F}$ is surjective, the preimage $f_{F}^{-1}(U_{n}\setminus\{p\})=f_{F}^{-1}(U_{n})\setminus\{p\}$ in $Z$ must also be disconnected. Since $U_{n}\setminus\{p\}$ has two components, each of which is an open interval of $e$, the preimage $f_{F}^{-1}(U_{n})\setminus\{p\}$ must have exactly two components. It follows that $Z\setminus\{p\}$ has exactly two ends, which completes the proof of condition (2). ∎
The next result presents some property crucial from the perspective of our application of subnecklaces in the main argument of this subsection (provided at its very end).
Proposition 5.7.
Let $Z$ be a tree of graphs. Assume that $Z$ is connected, locally connected and has no cutpoints. Let $A\subset Z$ be closed, connected and of cardinality at least $2$. Then, for some subnecklace $N$ of $Z$, the intersection $A\cap N$ is uncountable.
Proof.
Let ${\mathcal{S}}_{\Theta}=(\{K_{F}^{\ast}:F\in{\mathcal{F}}_{T}^{b}\},\{f_{FF^{\prime}}:F\subset F^{\prime}\})$ be the standard inverse system associated to a tree system of punctured graphs $\Theta$ whose limit is $Z$. Since $A$ has cardinality at least $2$, for some finite b-subtree $F$, the image $f_{F}(A)$ in $K_{F}^{\ast}$ has cardinality at least $2$. Since $A$ is connected, so is $f_{F}(A)$. But $K_{F}^{\ast}$ is a graph so then $f_{F}(A)$ must contain a topological edge $e$ of $K_{F}^{\ast}$. But $e$ contains uncountably many stable points and by Proposition 5.6, the preimages of these stable points in $Z$ form a subnecklace. Since these preimages are also easily seen to be contained in $A$, the assertion follows. ∎
Later in this subsection, in the main argument, we will confront the property of subnecklaces established above, in Proposition 5.7, with the following property of necklaces in Gromov boundaries of hyperbolic groups.
Lemma 5.8.
Let $G$ be a $1$-ended hyperbolic group that is not cocompact Fuchsian. Let $H$ be a rigid cluster factor of $G$ and let $N$ be a necklace class of $\partial G$. Then $\partial H\cap N$ has cardinality at most $2$.
Proof.
The closure of $N$ is the limit set $\partial K$ of a flexible factor $K$ of $G$. Since $H$ and $K$ are stabilizers of vertices of the reduced Bowditch JSJ tree $\mathcal{T}^{r}_{G}$ the intersection $H\cap K$ is contained in an edge stabilizer of $\mathcal{T}^{r}_{G}$. Thus $H\cap K$ is either finite or $2$-ended. Since $H$ and $K$ are both quasiconvex (e.g. by an observation in the paragraph right before Corollary 3.7), we have $\partial H\cap\partial K=\partial(H\cap K)$ as subspaces of $\partial G$ [Swe01, Corollary to Theorem 13]. Thus $\partial H\cap N\subset\partial H\cap\partial K=\partial(H\cap K)$ and so $\partial H\cap N$ has cardinality at most $2$. ∎
Before proving the main result, we recall two more properties of hyperbolic groups which we need for the argument.
Lemma 5.9.
Let $G$ be a hyperbolic group and assume that $G$ is not virtually free (in particular, it is not finite and not virtually cyclic). Then $G$ has a $1$-ended quasi-convex subgroup.
Proof.
By the Dunwoody’s Accessibility Theorem [Dun85], there is a splitting of $G$ as a finite graph of groups whose edge groups are finite and whose vertex groups have at most $1$ end. Since graph of finite groups always yields a virtually free group (see Proposition 11 in [Ser03]), at least one of the vertex groups above must be 1-ended. Finally, this vertex group is quasi-convex e.g. by the discussion in the paragraph right before Corollary 3.7. ∎
Next rather well known result follows e.g. by [KB02, Theorem 2.28].
Lemma 5.10.
Let $G$ be a hyperbolic group. Then the boundary $\partial G$ cannot be a singleton.
We are now ready to prove the main result of this section: the impliciation (1$\Rightarrow$(3) of Theorem A. We restate and prove the implication below.
Theorem 5.11.
Let $G$ be a 1-ended hyperbolic group that is not cocompact Fuchsian. If $\partial G$ is homeomorphic to a tree of graphs then each rigid cluster factor of $G$ is virtually free
Proof.
For the sake of reaching a contradiction, assume that $\partial G$ is a tree of graphs and $G$ has a rigid cluster factor $H$ that is not virtually free. Then, by Lemma 5.9, the rigid cluster factor $H$ has a $1$-ended quasi-convex subgroup $K$. Then, since $H$ is quasi-convex in $G$, the boundary $\partial K$ is a closed subspace of $\partial G$. Since $K$ is $1$-ended, the boundary $\partial K$ is connected. By Lemma 5.10, the boundary $\partial K$ has at least two points. Moreover, the boundary $\partial G$ is connected, locally connected and has no cutpoints [KB02, Proposition 7.1, Theorem 7.2]. Thus we can apply Proposition 5.7 to the subspace $\partial K\subset\partial G$ to conclude that $\partial K$ has uncountable intersection with some subnecklace $N^{\prime}\subset\partial G$. Recalling Remark 5.2, we see that $\partial K$ has uncountable intersection with a necklace class $N$ of $\partial G$. Then, since $\partial K\subset\partial H$, we see that $\partial H$ has uncountable intersection with $N$. This contradicts Lemma 5.8. ∎
5.2. Twin graphs and trees of thick $\theta$-graphs
Recall that a $\theta$-graph is a finite graph with two vertices, at least one edge and no loop edges. Note that a $\theta$-graph is 2-connected if it has at least two edges. We say that a $\theta$-graph is thick if it has at least three edges. In this subsection we will study connectedness properties of finite connected sums of $\theta$-graphs which we call twin graphs and characerize in terms of their Bowditch cut pairs. We will use these results in the following subsection to study trees of thick $\theta$-graphs and to prove the implication (1$\Rightarrow$(3) of Theorem B.
In the remainder of this section we treat graphs as $1$-dimensional cellular complexes and thus do not differentiate between a graph and its geometric realization.
Definition 5.12.
Let $X$ be a 2-connected graph. Two points $x$ and $y$ of $X$ are twins if the degree of $x$, the degree of $y$ and the number of components of $X\setminus\{x,y\}$ all coincide. That is, the points $x$ and $y$ are distinct and they are in the same Bowditch class as described in Section 3.1.
Remark 5.13.
The subspace $X\setminus\{x,y\}$ has the same number of components as the blow-up $X^{\#}(\{x,y\})$. Thus if $\iota_{x,y}\colon P_{x}\to\pi_{0}\bigl{(}X^{\#}(\{x,y\})\bigr{)}$ is the map sending each point of the blow-up divisor to its component in the blow-up and $\iota_{y,x}$ is defined similarly then, by Lemma 4.6, the points $x$ and $y$ are twins if and only if $\iota_{x,y}$ and $\iota_{y,x}$ are injective.
Remark 5.14.
Any degree $2$ point of any 2-connected graph has infinitely many twins. Indeed, any two points in the interior of the same topological edge are twins.
The following proposition, which justifies the terminology twin, is proved by Bowditch in the more general setting of Peano continua without global cutpoints.
Lemma 5.15 (Bowditch [Bow98, Lemma 3.8]).
Let $X$ be a 2-connected graph and let $x\in X$. If $\deg(x)\geq 3$ then $x$ has at most one twin.
Definition 5.16.
A point $x$ of a graph $X$ is an essential vertex if it has degree other thant $2$. Note that in a $2$-connected graph, any essential vertex has degree at least $3$.
Definition 5.17.
A 2-connected graph $X$ is a twin graph if every point in $X$ has a twin. If $x\in X$ is an essential vertex then we denote the unique twin of $x$ by $\bar{x}$ and we call $x$ and $\bar{x}$ essential twins.
Example 5.18.
Any 2-connected $\theta$-graph is a twin graph.
Proposition 5.19.
Let $X_{1}$ and $X_{2}$ be 2-connected graphs with $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$. Let $\ell\colon P_{x_{1}}\to P_{x_{2}}$ be a bijection between the blow-up divisor $P_{x_{1}}$ of $x_{1}$ in $X_{1}$ and the blow-up divisor $P_{x_{2}}$ of $x_{2}$ in $X_{2}$. Let $X$ be the connected sum $X_{1}\#_{\ell}X_{2}$ and let $f_{i}\colon X_{1}\#_{\ell}X_{2}\to X_{i}$ be the projection, for each $i\in\{1,2\}$.
  1. (1)
    Let $i\in\{1,2\}$ and let $x,y\in X_{i}$ with $x\neq x_{i}$ and $y\neq x_{i}$. Then $x$ and $y$ are twins in $X$ if and only if $x$ and $y$ are twins in $X_{i}$.
  2. (2)
    For each $i\in\{1,2\}$, let $y_{i}\in X_{i}$ with $y_{i}\neq x_{i}$, let $\iota_{x_{i},y_{i}}\colon P_{x_{i}}\to\pi_{0}\bigl{(}X_{i}^{\#}(\{x_{i},y_{i}\})\bigr{)}$ be the map sending each point of the blow-up divisor to its component in the blow-up and let $\sigma_{i}$ be the partition of $P_{x_{i}}$ into fibers of $\iota_{x_{i},y_{i}}$. Then $y_{1}$ and $y_{2}$ are twins in $X$ if and only if $\deg(y_{i})=\bigl{|}\pi_{0}\bigl{(}X_{i}\setminus\{x_{i},y_{i}\}\bigr{)}\bigr{|}$, for each $i$, and $\ell(\sigma_{1})=\sigma_{2}$. In this case, for each $i\in\{1,2\}$, the inclusion $X_{i}^{\#}(\{x_{i},y_{i}\})\hookrightarrow X^{\#}(\{y_{1},y_{2}\})$ induces a bijection on connected components.
Proof.
We begin by proving (1). By symmetry, we may assume that $i=1$. Since $X_{2}$ is connected and cutpoint-free, the blow-up $X_{2}^{\#}(x_{2})$ is connected. Then the restriction $X_{1}\#_{\ell}X_{2}\setminus\{x,y\}\to X_{1}\setminus\{x,y\}$ of the projection $X_{1}\#_{\ell}X_{2}\to X_{1}$ (which is the quotient identifying $X_{2}^{\#}(x_{2})$ to a point) preserves the number of connected components. Then, since $x$ and $y$ have the same degrees in $X_{1}\#_{\ell}X_{2}$ as they do in $X_{1}$, they are twins in $X_{1}\#_{\ell}X_{2}$ if and only if they are twins in $X_{1}$.
We now prove (2). Suppose first that $\deg(y_{i})\neq\bigl{|}\pi_{0}\bigl{(}X_{i}\setminus\{x_{i},y_{i}\}\bigr{)}\bigr{|}$. Without loss of generality, we can assume $i=1$. Notice that $\deg(y_{1})$ is equal to the cardinality of the blow-up divisor $P_{y_{1}}$ and that $\bigl{|}\pi_{0}\bigl{(}X_{1}\setminus\{x_{1},y_{1}\}\bigr{)}\bigr{|}$ is equal to the number of components of the blow-up $X_{1}^{\#}(\{x_{1},y_{1}\})$. Then, by Lemma 4.6, we have distinct $p,p^{\prime}\in P_{y_{1}}$ that are contained in the same component $C$ of $X_{1}^{\#}(\{x_{1},y_{1}\})$. But $X_{1}^{\#}(x_{1})$ is a subspace of $X$ with $y_{1}\in X_{1}^{\#}(x_{1})$ and $y_{2}\notin X_{1}^{\#}(x_{1})$ so $X_{1}^{\#}(\{x_{1},y_{1}\})$ is a subspace of $X^{\#}(\{y_{1},y_{2}\})$. Then $C$ is a connected subspace of $X^{\#}(\{y_{1},y_{2}\})$ containing $p,p^{\prime}\in P_{y_{1}}$, which by Remark 5.13, implies that $y_{1}$ and $y_{2}$ are not twins.
Suppose now that $\ell(\sigma_{1})\neq\sigma_{2}$. Without loss of generality (by swapping, if necessary, the indices $i\in\{1,2\}$) there are $p,p^{\prime}\in P_{x_{1}}$ contained in the same component $C_{1}$ of $X_{1}^{\#}(\{x_{1},y_{1}\})$ such that $\ell(p)$ and $\ell(p^{\prime})$ are contained in distinct components $C_{2}$ and $C_{2}^{\prime}$ of $X_{2}^{\#}(\{x_{2},y_{2}\})$. By Lemma 4.6, there are $p_{2},p_{2}^{\prime}\in P_{y_{2}}$ with $p_{2}\in C_{2}$ and $p_{2}^{\prime}\in C_{2}^{\prime}$. We can view $X^{\#}(\{y_{1},y_{2}\})$ as the connected sum of $X_{1}^{\#}(y_{1})$ and $X_{2}^{\#}(y_{2})$ at $x_{1}$ and $x_{2}$, along $\ell\colon P_{x_{1}}\to P_{x_{2}}$. Then $C_{1}\cup C_{2}\cup C_{2}^{\prime}$ is a connected subspace of $X^{\#}(\{y_{1},y_{2}\})$ containing both $p_{2}$ and $p_{2}^{\prime}$ which, by Remark 5.13, implies that $y_{1}$ and $y_{2}$ are not twins. Thus we have proved the forward implication of (2).
To prove the reverse implication of (2) assume, for each $i$, that $\deg(y_{i})=\bigl{|}\pi_{0}\bigl{(}X_{i}\setminus\{x_{i},y_{i}\}\bigr{)}\bigr{|}$ and that $\ell(\sigma_{1})=\sigma_{2}$. By Remark 5.13, it suffices to prove that $X^{\#}(\{y_{1},y_{2}\})$, $X_{1}^{\#}(\{x_{1},y_{1}\})$ and $X_{2}^{\#}(\{x_{2},y_{2}\})$ all have the same number of components. For components $C_{1}$ of $X_{1}^{\#}(\{x_{1},y_{1}\})$ and $C_{2}$ of $X_{2}^{\#}(\{x_{2},y_{2}\})$, let $C_{1}\sim C_{2}$ be the intersection relation: $C_{1}\cap C_{2}\neq\emptyset$ in $X^{\#}(\{y_{1},y_{2}\})$. The equality $\ell(\sigma_{1})=\sigma_{2}$ implies that the intersection relation induces a bijection between $\pi_{0}\bigl{(}X_{1}^{\#}(\{x_{1},y_{1}\}))$ and $\pi_{0}\bigl{(}X_{2}^{\#}(\{x_{2},y_{2}\}))$. Moreover, the $\sim$-pairs are in bijection with the components of $X^{\#}(\{y_{1},y_{2}\})$ so we are done. This also makes clear that $X_{i}^{\#}(\{x_{i},y_{i}\})\hookrightarrow X^{\#}(\{y_{1},y_{2}\})$ is a bijection on components for each $i$. ∎
Corollary 5.20.
Let $X_{1}$ and $X_{2}$ be 2-connected graphs with $x_{1},y_{1}\in X_{1}$ and $x_{2},y_{2}\in X_{2}$. Assume that the $x_{i}$ and $y_{i}$ are all distinct and all have degree $2$. Let $X=X_{1}\#X_{2}$ be a connected sum of $X_{1}$ and $X_{2}$ at $x_{1}$ and $x_{2}$. Then $y_{1}$ and $y_{2}$ are twins in $X$ if and only if $x_{i}$ and $y_{i}$ are twins in $X_{i}$, for each $i$. In this case, for each $i\in\{1,2\}$, the inclusion $X_{i}^{\#}(\{x_{i},y_{i}\})\hookrightarrow X^{\#}(\{y_{1},y_{2}\})$ induces a bijection on connected components.
Proposition 5.21.
Any connected sum of twin graphs is a twin graph.
Proof.
Let $X_{1}$ and $X_{2}$ be twin graphs with points $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$. Let $\ell\colon P_{x_{1}}\to P_{x_{2}}$ be a bijection of blow-up divisors. By Lemma 4.11, the connected sum $X=X_{1}\#_{\ell}X_{2}$ is 2-connected so we need only show that every point has a twin. By Remark 5.14, it suffices to consider an essential vertex, i.e., a point $y_{1}\in X$ of degree at least $3$. Then, without loss of generality, we can assume $y_{1}\in X_{1}\setminus\{x_{1}\}\subset X$. Let $\bar{y}_{1}$ be the twin of $y_{1}$ in $X_{1}$, which is unique by Lemma 5.15. By Proposition 5.19(1), if $\bar{y}_{1}\neq x_{1}$ then $\bar{y}_{1}$ remains the twin of $y_{1}$ in $X$. So we can assume that $\bar{y}_{1}=x_{1}$ so that $\deg(x_{2})=\deg(x_{1})=\deg(y_{1})\geq 3$. Let $y_{2}$ be the twin of $x_{2}$ in $X_{2}$. We will prove that $y_{2}$ is the twin of $y_{1}$ in $X$. Since $x_{1}$ and $y_{1}$ are twins in $X_{1}$ and $x_{2}$ and $y_{2}$ are twins in $X_{2}$, the degree condition of Proposition 5.19(2) is satisfied. Moreover, by Remark 5.13, the fibers of $\iota_{x_{i},y_{i}}$ are singletons so $\ell$ preserves the partitions of the $P_{x_{i}}$ into fibers. Thus the remaining condition of Proposition 5.19(2) is satisfied and $y_{1}$ and $y_{2}$ are twins in $X$. ∎
We will need the following modified version of a result of Bowditch.
Lemma 5.22 (Bowditch [Bow98, Lemma 3.18]).
Let $X$ be a connected, locally connected, cutpoint-free $T_{1}$ space and let $x,x^{\prime}\in X$. Assume that $X\setminus\{x,x^{\prime}\}$ has at least three components. If $y,z\in X$ are contained in distinct components $C_{y}$ and $C_{z}$ of $X\setminus\{x,x^{\prime}\}$ then $X\setminus\{y,z\}$ is connected.
Proof.
No component $C^{\prime}$ of $X\setminus\{y,z\}$ can be disjoint from $\{x,x^{\prime}\}$. Otherwise, by Corollary 4.9, we would have $\bar{C}^{\prime}=C^{\prime}\cup\{y,z\}$ so that $C_{y}\cup\bar{C}^{\prime}\cup C_{z}$ would be a connected subspace of $X\setminus\{x,x^{\prime}\}$ contradicting the fact that $C_{y}$ and $C_{z}$ are distinct components of $X\setminus\{x,x^{\prime}\}$. Thus $X\setminus\{y,z\}$ has at most two components, each of which contains either $x$ or $x^{\prime}$.
Assume for the sake of finding a contradiction that $X\setminus\{y,z\}$ is disconnected. Then, by the previous paragraph, it has exactly two components $C_{x}$ and $C_{x^{\prime}}$ which contain $x$ and $x^{\prime}$, respectively. Since $X\setminus\{x,x^{\prime}\}$ has at least three components, there is a component $C$ of $X\setminus\{x,x^{\prime}\}$ that is distinct from $C_{y}$ and $C_{z}$. By Corollary 4.9, we have $\bar{C}=C\cup\{x,x^{\prime}\}$ so that $C_{x}\cup\bar{C}\cup C_{x^{\prime}}$ is a connected subspace of $X\setminus\{y,z\}$ contradicting the fact that $C_{x}$ and $C_{x^{\prime}}$ are distinct components of $X\setminus\{y,z\}$. ∎
Lemma 5.23.
Let $X$ be a twin graph not homeomorphic to the circle. There exists an essential twin pair $\{x,\bar{x}\}$ of $X$ such that at most a single component of $X\setminus\{x,\bar{x}\}$ contains any essential vertices of $X$.
Proof.
For an essential twin pair $\{x,\bar{x}\}$, let $n_{\{x,\bar{x}\}}$ be the maximal number of essential vertices contained in a component of $X\setminus\{x,\bar{x}\}$. The complexity of $\{x,\bar{x}\}$ is the nonnegative integer $c_{\{x,\bar{x}\}}=n-n_{\{x,\bar{x}\}}-2$, where $n$ is the number of essential vertices of $X$. Note that $c_{\{x,\bar{x}\}}$ is the number of essential vertices in all but a “largest” component of $X\setminus\{x,\bar{x}\}$. So it suffices to find an essential twin pair $\{x,\bar{x}\}$ for which $c_{\{x,\bar{x}\}}=0$. This is the base case of an induction starting with an arbitrary essential twin pair $\{x,\bar{x}\}$. It suffices to show that if $c_{\{x,\bar{x}\}}>0$ then we can find an essential twin pair $\{y,\bar{y}\}$ for which $c_{\{y,\bar{y}\}}<c_{\{x,\bar{x}\}}$. Let $C^{\prime}$ be a component of $X\setminus\{x,\bar{x}\}$ containing $n_{\{x,\bar{x}\}}$ essential vertices. Since $n_{\{x,\bar{x}\}}=n-c_{\{x,\bar{x}\}}-2<n-2$, some other component $C$ of $X\setminus\{x,\bar{x}\}$ contains an essential vertex $y$. We will show that $c_{\{y,\bar{y}\}}<c_{\{x,\bar{x}\}}$.
By Lemma 5.22, the twin $\bar{y}$ of $y$ is also contained in $C$. By Corollary 4.9, the closure $\bar{C}^{\prime}$ of $C^{\prime}$ is $C^{\prime}\cup\{x,\bar{x}\}$. Then $C^{\prime}\cup\{x,\bar{x}\}$ is connected and so must be contained in some component of $X\setminus\{y,\bar{y}\}$. But $C^{\prime}\cup\{x,\bar{x}\}$ contains $n_{\{x,\bar{x}\}}+2$ essential vertices so $n_{\{y,\bar{y}\}}\geq n_{\{x,\bar{x}\}}+2$ and so $c_{\{y,\bar{y}\}}=n-n_{\{y,\bar{y}\}}-2\leq n-n_{\{x,\bar{x}\}}-4<c_{\{x,\bar{x}\}}$, as required. ∎
Theorem 5.24.
Let $X$ be a graph that is not homeomorphic to the circle. Then $X$ is a twin graph if and only if it can be obtained from thick $\theta$-graphs by finitely many connected sum operations.
Proof.
The “if” part follows from Proposition 5.21 and the fact that thick $\theta$-graphs are twin graphs.
To prove the “only if” part, suppose $X$ is a twin graph. By Lemma 5.23, there exists a pair of essential twins $x$ and $\bar{x}$ of $X$ such that at most a single component $C$ of $X\setminus\{x,\bar{x}\}$ contains any essential vertices of $X$. It follows from Corollary 4.9 that any component of $X\setminus\{x,\bar{x}\}$ without essential vertices is a single open topological edge with endpoints $x$ and $\bar{x}$. Thus if $C$ also contains no essential vertices then $X$ is a thick $\theta$-graph. This is the base case of an induction on the number of essential vertices of $X$.
Suppose $C$ contains at least one essential vertex. By Lemma 4.6 there is a topological edge $e_{x}$ with endpoint $x$ and whose interior is contained in $C$. Similarly, there is a topological edge $e_{\bar{x}}$ with endpoint $\bar{x}$ and whose interior is contained in $C$. Let $p_{x}$ be a point in the interior of $e_{x}$ and let $p_{\bar{x}}$ be a point in the interior of $e_{\bar{x}}$. Then $X^{\#}(\{p_{x},p_{\bar{x}}\})$ has two components, one of which contains $x$ and $\bar{x}$ and the other of which is naturally a subspace of $C$. Gluing together the pairs of points of the blow-up divisors $P_{p_{x}}\cup P_{p_{\bar{x}}}\subset X^{\#}(\{p_{x},p_{\bar{x}}\})$ that are contained in the same component results in a disjoint union of two graphs: a thick $\theta$-graph $X_{1}$ with essential vertices $x$ and $\bar{x}$ and a graph $X_{2}$ with two less essential vertices than $X$. Moreover, the original graph $X$ is a connected sum of $X_{1}$ and $X_{2}$. Thus, by induction, it suffices to show that $X_{2}$ is a twin graph. This follows immediately from Proposition 5.19(1). ∎
Lemma 5.25.
Let $X$ be a twin graph and let $x,y\in X$ be distinct and of degree $2$. If $X\setminus\{x,y\}$ is connected then $x$ and $y$ are contained in distinct components of $X\setminus\{v,\bar{v}\}$ for some essential twin pair $v,\bar{v}$.
Proof.
Since $X\setminus\{x,y\}$ is connected, the twin graph $X$ cannot be homeomorphic to the circle. By Theorem 5.24, it is a connected sum of thick $\theta$-graphs. We proceed by induction on the number of $\theta$-graph factors. In the base case $X$ is a $\theta$-graph and so $x$ and $y$ must be contained in distinct components of $X\setminus\{v,\bar{v}\}$ where $\{v,\bar{v}\}$ is the unique essential twin pair of $X$.
In the inductive step, we can write $X$ as a connected sum $X_{1}\#X_{2}$ where $X_{1}$ is a twin graph and $X_{2}$ is a thick $\theta$-graph and the connected sum is performed at points $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$. Since performing a connected sum with a $\theta$-graph at an essential vertex does not change the topology, we can assume that $\deg(x_{1})=\deg(x_{2})=2$. Up to homeomorphism, we can assume that $x$ and $y$ are not contained in the blow-up divisor $P_{x_{1}}=P_{x_{2}}\subset X_{1}\#X_{2}$. In fact, we can assume that $x$ and $y$ are not contained in $e\setminus\{v,\bar{v},x_{2}\}\subset X_{1}\#X_{2}$ where $e$ is the topological edge of $X_{2}$ containing $x_{2}$ and $\{v,\bar{v}\}$ is the unique essential twin pair of $X_{2}$. We consider first the case where one of the points, say $x$, is contained in $X_{2}\setminus e\subset X_{1}\#X_{2}$. Then $y$ cannot be contained in the topological edge of $X_{2}$ that contains $x$ since otherwise $X\setminus\{x,y\}$ would be disconnected. Hence $x$ and $y$ are contained in distinct components of the complement of $\{v,\bar{v}\}$, which is a twin pair of $X$ by Proposition 5.19(1). It remains then to consider the case where $x,y\in X_{1}\setminus\{x_{1}\}\subset X_{1}\#X_{2}$. By Proposition 5.19(1), the degree $2$ points $x$ and $y$ are not twins in $X_{1}$ and so $X_{1}\setminus\{x,y\}$ is connected. Then, by induction, there exists an essential twin pair $\{w,\bar{w}\}$ of $X_{1}$ separating $x$ from $y$. By Proposition 5.19(1), the pair $\{w,\bar{w}\}$ continues to be a twin pair in $X_{1}\#X_{2}$. Since the connected sum operation with the $\theta$-graph $X_{2}$ preserves components of the complement of $\{w,\bar{w}\}$ we see that $\{w,\bar{w}\}$ continues to separate $x$ and $y$ in $X=X_{1}\#X_{2}$. ∎
5.3. Proof of the impliciation (1$\Rightarrow$(3) of Theorem B
We now turn our attention from twin graphs towards trees of thick $\theta$-graphs in order to prove the implication (1$\Rightarrow$(3) of Theorem B, i.e., we will show that if $G$ is a $1$-ended hyperbolic group that is not cocompact Fuchsian and $\partial G$ is a tree of thick $\theta$-graphs then $G$ has no rigid JSJ factors.
Recall from Section 3.1 that the degree of a point $p$ in a compact space $X$ is the number of ends of the locally compact space $X\setminus\{p\}$. Since $X$ is compact, any compact subset of $X\setminus\{p\}$ is the complement of some open neighborhood $U$ of $p$. If $U$ is also connected then the degree of $p$ is at least the number of components of $U\setminus\{p\}$ since, in this case, each component of $U\setminus\{p\}$ contains at least one end of $X\setminus\{p\}$. It follows that, if $p$ has a basis $\{U_{\alpha}\}_{\alpha}$ of connected open neighborhoods then the degree of $p$ is the supremum of the number of components of the $U_{\alpha}\setminus\{p\}$.
Lemma 5.26.
Let $\Theta$ be a tree system of punctured thick $\theta$-graphs, with underlying tree $T$. Let $Z$ be the limit of $\Theta$ and let $x\in Z$. Then $x$ has degree at least $3$ if and only if there exists a finite b-subtree $F_{0}\subset T$ such that, for any finite b-subtree $F\supset F_{0}$, the projection of $x$ to the reduced partial union $K_{F}^{\ast}$ has degree at least $3$. In this case, for any $F\supset F_{0}$, the degree of the projection of $x$ to $K_{F}^{\ast}$ is equal to the degree of $x$.
Proof.
We first prove the contrapositive of the forward implication. Assume that for any finite b-subtree $F_{0}\subset T$ there exists a finite b-subtree $F\supset F_{0}$ such that the projection $x_{F}$ of $x$ to the reduced partial union $K_{F}^{\ast}$ has degree less than $3$. Let $U$ be an open neighborhood of $x$ in $Z$. Then there exists a reduced partial union $K_{F}^{\ast}$ and an open neighborhood $V_{F}$ of the projection $x_{F}$ of $x$ in $K_{F}^{\ast}$ such that the preimage of $V_{F}$ in $Z$ is contained in $U$. By our assumption, we can choose $K_{F}^{\ast}$ so that $x_{F}$ has degree $2$ and so we can choose $V_{F}$ so that it is homeomorphic to an open interval. By Lemma 4.13, the preimage $V$ of $V_{F}$ in $Z$ is connected. By the paragraph preceding the statement of the lemma, it suffices to show that $V\setminus\{x\}$ has at most two components in order to establish that $x$ has degree at most $2$. Now, let $K_{\hat{F}}^{\ast}$ be any reduced partial union with $\hat{F}\supset F$, let $V_{\hat{F}}$ be the preimage of $V_{F}$ in $K_{\hat{F}}^{\ast}$ and let $x_{\hat{F}}$ be the projection of $x$ to $K_{\hat{F}}^{\ast}$. Then $V_{\hat{F}}$ is homeomorphic to the complement of a point of degree $2$ in a finite connected sum of $2$-connected graphs. To see this, view $V_{F}$ as the complement of a point in the circle $S^{1}$ and then perform the connected sum operations required to obtain $K_{\hat{F}}^{\ast}$ from $K_{F}^{\ast}$. It follows then from Lemma 4.6 and Remark 4.7 that $V_{\hat{F}}\setminus\{x_{\hat{F}}\}$ has at most two components. But $V\setminus\{x\}$ is the ascending union of the preimages of the $V_{\hat{F}}\setminus\{x_{\hat{F}}\}\subset K_{\hat{F}}^{\ast}$ as $\hat{F}$ ranges over all finite b-subtrees containing $F$. Hence $V$ itself has at most two components. This establishes the forward implication.
To prove the reverse implication, assume that $F_{0}$ is a finite b-subtree and that in any finite b-subtree $F\supset F_{0}$, the projection $x_{F}$ of $x$ to the reduced partial union $K_{F}^{\ast}$ has degree at least $3$. Since $K_{F}^{\ast}$ is obtained from $K_{F_{0}}^{\ast}$ by a sequence of connected sum operations with thick $\theta$-graphs and the projection of $x$ to each of these has degree at least $3$, we see that the degree of $x_{F}$ is equal to the degree of the projection $x_{F_{0}}$ of $x$ in $K_{F_{0}}^{\ast}$. Indeed, the degree is unchanged when the connected sum is performed away from the projection of $x$ and, if the connected sum is performed at the projection of $x$, then the projection of $x$ in the new reduced partial union can only be at the remaining essential vertex of the $\theta$-graph summand, which has the same degree.
It remains to show that $\deg(x)=\deg(x_{F_{0}})$. Let $U$ be an open neighborhood of $x$ and let $F\supset F_{0}$ be large enough that a normal neighborhood $V_{F}$ of the projection $x_{F}$ of $x$ in $K_{F}^{\ast}$ has the property that the preimage $V$ of $V_{F}$ in $Z$ is contained in $U$. By Lemma 4.13, the preimage $V$ is connected. It suffices to prove that $V\setminus\{x\}$ has the same number of components as $V_{F}\setminus\{x_{F}\}$, which has $\deg(x_{F_{0}})$-many components. As above, we will describe $V\setminus\{x\}$ as an ascending union of preimages of spaces $V_{\hat{F}}\setminus\{x_{\hat{F}}\}$ where $V_{\hat{F}}$ is the preimage of $V_{F}$ in a reduced partial union $K_{\hat{F}}^{\ast}$ with $\hat{F}\supset F$ and $x_{\hat{F}}$ is the projection of $x$ to $K_{\hat{F}}^{\ast}$. But $V_{\hat{F}}$ is obtained from $V_{F}$ by performing a series of connected sum operations with thick $\theta$-graphs. These connected sum operations do not change the topology when performed at the projection of $x$, since we are summing with a thick $\theta$-graph and the new projection of $x$ is at the remaining essential vertex of the thick $\theta$-graph. When the connected sum operation is performed away from the projection of $x$ it does not disconnect (or join) the components of the complement of the projection of $x$, since we are summing with a $\theta$-graph and these are $2$-connected. Thus we see that $V_{\hat{F}}\setminus\{x_{\hat{F}}\}$ has the same number of components as $V_{F}\setminus\{x_{F}\}$ and the inclusion of the preimage of $V_{F}\setminus\{x_{F}\}$ in $V_{\hat{F}}\setminus\{x_{\hat{F}}\}$ is bijective on components. Thus, by Lemma 4.13, the preimages of the $V_{\hat{F}}\setminus\{x_{\hat{F}}\}$ in $Z$ with $\hat{F}$ ranging over a cofinal sequence of finite b-subtrees containing $F$ expresses $V\setminus\{x\}$ as an ascending union of spaces, each having $\deg(x_{F_{0}})$-many components and where the inclusion of one in the next induces a bijection on components. It follows that $V\setminus\{x\}$ has $\deg(x_{F_{0}})$-many components. ∎
Lemma 5.27.
Let $\Theta$ be a tree system of punctured thick $\theta$-graphs, with underlying tree $T$. Let $Z$ be the limit of $\Theta$, let $S\subset Z$ be a finite set of degree $2$ points and let $F_{0}\subset T$ be a finite b-subtree. Then there exists a finite b-subtree $F\supset F_{0}$ such that the projection of each $x\in S$ to the reduced partial union $K_{F}^{\ast}$ has degree $2$.
Proof.
We proceed by induction on $|S|$. For the base case $|S|=0$ we can take $F=F_{0}$. Assume now that $|S|>0$ and $x_{0}\in S$. By induction, we have a finite b-subtree $F_{-}\supset F_{0}$ such that the projection of each $x\in S\setminus\{x_{0}\}$ to the reduced partial union $K_{F_{-}}^{\ast}$ has degree $2$. By Lemma 5.26 there exists a finite b-subtree $F_{+}\supset F_{-}$ such that the projection of $x_{0}$ to the reduced partial union $K_{F_{+}}^{\ast}$ has degree $2$. Recall that $K_{F_{+}}^{\ast}$ is obtained from $K_{F_{-}}^{\ast}$ by a series of connected sum operations, each of which corresponds to growing the b-subtree inside of $F_{+}$. For a point $x\in S$, these operations do not change the degree of the projection of $x$ unless the connected sum is performed at the projection. It follows that if we perform only those operations that occur at the projection of $x_{0}$, we will obtain a reduced partial union $K_{F}^{\ast}$, with $F_{-}\subset F\subset F_{+}$, such that the degrees of projections of the points of $S\setminus\{x_{0}\}$ are as they are in $K_{F_{-}}^{\ast}$ and the degree of the projection of $x_{0}$ is as it is in $K_{F_{+}}^{\ast}$. ∎
Lemma 5.28.
Let $\Theta$ be a tree system of punctured thick $\theta$-graphs, with underlying tree $T$. Let $Z$ be the limit of $\Theta$, let $x,y\in Z$ be degree $2$ points. Then $Z\setminus\{x,y\}$ has at most $2$ components.
Proof.
By Lemma 5.27, there exists a cofinal sequence of b-subtrees $F_{0}\subset F_{1}\subset F_{2}\subset\cdots$ such that $x$ and $y$ project to points $x_{F_{n}}$ and $y_{F_{n}}$ of degree $2$ in the reduced partial unions $K_{F_{n}}^{\ast}$. By Remark 4.7, each $K_{F_{n}}^{\ast}\setminus\{x_{F_{n}},y_{F_{n}}\}$ has at most two components, each of which has connected preimage in $Z$ by Lemma 4.13. But $Z\setminus\{x,y\}$ is the ascending union of the preimages of the $K_{F_{n}}^{\ast}\setminus\{x_{F_{n}},y_{F_{n}}\}$ and so has at most two components. ∎
Lemma 5.29.
Let $\Theta$ be a tree system of punctured thick $\theta$-graphs, with underlying tree $T$. Let $Z$ be the limit of $\Theta$ and let $x,y\in Z$ be degree $2$ points. Then $Z\setminus\{x,y\}$ is connected if and only if there exists a finite b-subtree $F\subset T$ such that the projections $x_{F}$ and $y_{F}$ of $x$ and $y$ to the reduced partial union $K_{F}^{\ast}$ have degree $2$ and $K_{F}^{\ast}\setminus\{x_{F},y_{F}\}$ is connected. In this case, for any finite b-subtree $\hat{F}\supset F$ for which $x,y$ have degree $2$ projections $x_{\hat{F}},y_{\hat{F}}$ in the reduced partial union $K_{\hat{F}}^{\ast}$, the complement $K_{\hat{F}}^{\ast}\setminus\{x_{\hat{F}},y_{\hat{F}}\}$ is connected.
Remark.
In the last sentence of Lemma 5.29, the condition that $x_{\hat{F}}$ and $y_{\hat{F}}$ have degree $2$ can be removed. We avoid proving the more general statement since we do not need it here.
Proof of Lemma 5.29.
By Lemma 5.27 there is a cofinal chain $F_{0}\subset F_{1}\subset F_{2}\subset\cdots$ of finite b-subtrees of $T$ such that the projections $x_{F_{i}}$ and $y_{F_{i}}$ of $x$ and $y$ have degree $2$ in the reduced partial unions $K_{F_{i}}^{\ast}$. Then $Z\setminus\{x,y\}$ is the ascending union of the preimages of the $K_{F_{i}}^{\ast}\setminus\{x_{F_{i}},y_{F_{i}}\}$.
By Remark 4.7, each $K_{F_{i}}^{\ast}\setminus\{x_{F_{i}},y_{F_{i}}\}$ has either $1$ or $2$ components, corresponding to whether or not $x_{F_{i}}$ and $y_{F_{i}}$ are twins in $K_{F_{i}}^{\ast}$. For some $i$, let $V$ be the set of vertices of $F_{i+1}\setminus F_{i}$ that are adjacent to $F_{i}$ and, for each $v\in V$, let $F_{v}$ be the component of $F_{i+1}\setminus\{v\}$ that does not contain $F_{i}$. The reduced partial union $K_{F_{i+1}}^{\ast}$ is obtained from $K_{F_{i}}^{\ast}$ by performing connected sum operations (in arbitrary order) with the reduced partial unions $K_{F_{v}}^{\ast}$ at various points of $K_{F_{i}}^{\ast}$. By Lemma 4.11, the $K_{F_{v}}^{\ast}$ are all $2$-connected so the connected sum operations performed at points of $K_{F_{i}}^{\ast}\setminus\{x_{F_{i}},y_{F_{i}}\}$ preserves the components of the complement of the projection of $\{x,y\}$. By Corollary 5.20, a connected sum operation performed at $x_{F_{i}}$ or $y_{F_{i}}$ can only reduce the number of components of the complement of the projection of $\{x,y\}$ and will otherwise preserve the components. It follows that if $K_{F_{i}}^{\ast}\setminus\{x_{F_{i}},y_{F_{i}}\}$ is connected then so is $K_{F_{i+1}}^{\ast}\setminus\{x_{F_{i+1}},y_{F_{i+1}}\}$ and that if $K_{F_{i+1}}^{\ast}\setminus\{x_{F_{i+1}},y_{F_{i+1}}\}$ has $2$ components then the inclusion of the preimage of $K_{F_{i}}^{\ast}\setminus\{x_{F_{i}},y_{F_{i}}\}$ in $K_{F_{i+1}}^{\ast}\setminus\{x_{F_{i+1}},y_{F_{i+1}}\}$ induces a bijection on components.
We are now ready to prove the statement. To prove the reverse implication, assume that there exists a finite b-subtree $F$ as in the statement, i.e., a finite b-subtree $F\subset T$ such that the projections $x_{F}$ and $y_{F}$ of $x$ and $y$ to the reduced partial union $K_{F}^{\ast}$ have degree $2$ and $K_{F}^{\ast}\setminus\{x_{F},y_{F}\}$ is connected. Then we can obtain a cofinal chain of finite b-subtrees $F_{0}\subset F_{1}\subset F_{2}\subset\cdots$ as above with $F=F_{0}$. By the above argument, the $K_{F_{i}}^{\ast}\setminus\{x_{F_{i}},y_{F_{i}}\}$ are all connected. By Lemma 4.13, the preimages of the $K_{F_{i}}^{\ast}\setminus\{x_{F_{i}},y_{F_{i}}\}$ in $Z$ are then also connected so that their ascending union, which is equal to $Z\setminus\{x,y\}$, must also be connected.
To prove the reverse implication, assume that no such $F$ exists as in the statement. Then the $K_{F_{i}}^{\ast}\setminus\{x_{F_{i}},y_{F_{i}}\}$ all have $2$ components for a cofinal chain of finite b-subtrees $F_{0}\subset F_{1}\subset F_{2}\subset\cdots$ as above. As argued above, the inclusion of the preimage of $K_{F_{i}}^{\ast}\setminus\{x_{F_{i}},y_{F_{i}}\}$ in $K_{F_{i+1}}^{\ast}\setminus\{x_{F_{i+1}},y_{F_{i+1}}\}$ induces a bijection on components. Thus the preimages of the $K_{F_{i}}^{\ast}\setminus\{x_{F_{i}},y_{F_{i}}\}$ in $Z$ express $Z\setminus\{x,y\}$ as a disjoint union of two ascending unions of connected open subspaces. It follows that $Z\setminus\{x,y\}$ has two components.
The last sentence of the statement follows by the same argument given above showing that $K_{F_{i+1}}^{\ast}\setminus\{x_{F_{i+1}},y_{F_{i+1}}\}$ has no more components than $K_{F_{i}}^{\ast}\setminus\{x_{F_{i}},y_{F_{i}}\}$. ∎
Lemma 5.30.
Let $\Theta$ be a tree system of punctured thick $\theta$-graphs, with underlying tree $T$. Let $Z$ be the limit of $\Theta$. Let $F_{0}\subset T$ be a finite b-subtree and let $v$ be an essential vertex of the reduced partial union $K_{F_{0}}^{\ast}$. There exists a unique $\tilde{v}\in Z$ such that
  1. (1)
    $\tilde{v}$ projects to $v$ in $K_{F_{0}}^{\ast}$ and
  2. (2)
    for any finite b-subtree $F\supset F_{0}$, the projection of $\tilde{v}$ to $K_{F}^{\ast}$ has the same degree as $v$.
If $u$ is another essential vertex of $K_{F_{0}}^{\ast}$ then the preimage in $Z$ of each component of $K_{F_{0}}^{\ast}\setminus\{u,v\}$ is contained in a component of $Z\setminus\{\tilde{u},\tilde{v}\}$ and the map $\pi_{0}(K_{F_{0}}^{\ast}\setminus\{u,v\})\to\pi_{0}(Z\setminus\{\tilde{u},\tilde{v}\})$ induced by these inclusions is a bijection.
Proof.
We will define $\tilde{v}$ as the limit of a system of lifts $v_{F}\in K_{F}^{\ast}$ of $v$ where $F\supset F_{0}$. These lifts need to be consistent with respect to the bonding maps so it suffices to show how to lift $v_{F}$ to $v_{\hat{F}}$ for finite b-subtrees $\hat{F}\supset F\supset F_{0}$ for which no b-subtree lies properly between $\hat{F}$ and $F$. Then the reduced partial union $K_{\hat{F}}^{\ast}$ is obtained from $K_{F}^{\ast}$ by performing a connected sum operation with a $\theta$-graph $K_{t}^{\ast}$. In the case that the connected sum operation is performed away from $v_{F}$ then we let $v_{\hat{F}}=v_{F}$ in $K_{\hat{F}}^{\ast}$. In the case that the connected sum operation is performed at $v_{F}$, we let $v_{\hat{F}}$ be the essential vertex of $K_{t}^{\ast}$ that is not involved in the connected sum operation. Notice that, in either case, $v_{\hat{F}}$ is the unique vertex of $K_{\hat{F}}^{\ast}$ of degree $\deg(v_{F})$ that projects onto $v_{F}$ in $K_{F}^{\ast}$. Thus the limit $\tilde{v}\in Z$ of the $v_{F}$ is the unique point in $Z$ projecting to $v$ in $K_{F_{0}}^{\ast}$ and projecting to an essential vertex of degree $\deg(v_{F_{0}})$ in every $K_{F}^{\ast}$ with $F\supset F_{0}$.
We now prove the final part of the lemma. Let $u$ be another essential vertex of $K_{F_{0}}^{\ast}$. As above, consider finite b-subtrees $\hat{F}\supset F\supset F_{0}$ for which no b-subtree lies properly between $\hat{F}$ and $F$ and decompose $K_{\hat{F}}^{\ast}$ as $K_{F}^{\ast}\#K_{t}^{\ast}$. Since $\theta$-graphs are $2$-connected, if the connected sum operation is performed away from $u_{F}$ and $v_{F}$ then each component of $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}$ has connected preimage in $K_{\hat{F}}^{\ast}\setminus\{u_{\hat{F}},v_{\hat{F}}\}=K_{\hat{F}}^{\ast}\setminus\{u_{F},v_{F}\}$. Moreover, the inclusion in $K_{\hat{F}}^{\ast}\setminus\{u_{F},v_{F}\}$ of the preimage of $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}$ is bijective on components. If the connected sum operation is performed at $u_{F}$ or $v_{F}$, say at $v_{F}$, then $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}$ is a subspace of $K_{\hat{F}}^{\ast}\setminus\{u_{\hat{F}},v_{\hat{F}}\}=K_{\hat{F}}^{\ast}\setminus\{u_{F},v_{\hat{F}}\}$. The closure of $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}$ in $K_{\hat{F}}^{\ast}\setminus\{u_{F},v_{\hat{F}}\}$ is $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}\cup P_{v_{F}}$ where $P_{v_{F}}$ is the blow-up divisor of $v_{F}$. We see that $K_{\hat{F}}^{\ast}\setminus\{u_{F},v_{\hat{F}}\}$ is obtained from $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}\cup P_{v_{F}}$ by gluing a copy of the half-open interval $[0,1)$ along 00 to each point in $P_{v_{F}}$. Thus $K_{\hat{F}}^{\ast}\setminus\{u_{F},v_{\hat{F}}\}$ has a deformation retraction onto $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}\cup P_{v_{F}}$ and it follows that the inclusion of $K_{F}^{\ast}\setminus\{u_{F},v_{F}\}$ in $K_{\hat{F}}^{\ast}\setminus\{u_{\hat{F}},v_{\hat{F}}\}$ is bijective on components. Let $F_{0}\subset F_{1}\subset F_{2}\subset\cdots$ be a cofinal sequence of finite b-subtrees such that for each $n\geq 0$ no b-subtree lies properly between $F_{n}$ and $F_{n+1}$. For each $n$, let $X_{n}=K_{F_{n}}^{\ast}\setminus\{u_{F_{n}},v_{F_{n}}\}$ and let $\tilde{X}_{n}$ be the preimage in $Z$ of $X_{n}$. Then $Z\setminus\{\tilde{u},\tilde{v}\}$ is the ascending union of the $\tilde{X}_{n}$. By Lemma 4.13, the preimage in $Z$ of any component of any $X_{n}$ is connected. Consequently, the preimage in $Z$ of any component of $X_{0}=K_{F_{0}}^{\ast}\setminus\{u,v\}$ is contained in a component of $Z\setminus\{\tilde{u},\tilde{v}\}$. By the above argument, the inclusion of $\tilde{X}_{n}$ in $\tilde{X}_{n+1}$ induces a bijection on components. It follows that the map $\pi_{0}(K_{F_{0}}^{\ast}\setminus\{u,v\})\to\pi_{0}(Z\setminus\{\tilde{u},\tilde{v}\})$ sending a component $C$ of $K_{F_{0}}^{\ast}\setminus\{u,v\}$ to the component of $Z\setminus\{\tilde{u},\tilde{v}\}$ that contains the preimage of $C$ is bijective. ∎
Recall the description of Bowditch classes in Section 3.1: a pair of points $\{x,y\}$ is a Bowditch cut pair if $\deg(x)=\deg(y)\geq 2$ and the common degree of $x$ and $y$ is equal to the number of connected components in the complement of $\{x,y\}$. The degree of $\{x,y\}$ is defined as the common degree of $x$ and $y$. In the case of graphs we have used the terminology twin pairs for Bowditch cut pairs. We will also consider Bowditch cut pairs in the case of trees of punctured thick $\theta$-graphs.
Corollary 5.31.
Let $\Theta$ be a tree system of punctured thick $\theta$-graphs, with underlying tree $T$. Let $Z$ be the limit of $\Theta$ and let $x,y\in Z$. Assume that there exists a finite b-subtree $F\subset T$ such that the projections $x_{F}$ and $y_{F}$ of $x$ and $y$ in the reduced partial union $K_{F}^{\ast}$ are contained in distinct components of the complement $K_{F}^{\ast}\setminus\{v,\bar{v}\}$ of an essential twin pair $\{v,\bar{v}\}$. Then there exists a Bowditch cut pair $\{z,z^{\prime}\}$ in $Z$ of degree $\deg(v)$ such that
  1. (1)
    $z$ and $z^{\prime}$ project to $v$ and $\bar{v}$, and
  2. (2)
    $x$ and $y$ are contained in distinct components of $Z\setminus\{z,z^{\prime}\}$.
Proof.
Let $z,z^{\prime}\in Z$ be the lifts of $v$ and $\bar{v}$ given by Lemma 5.30. Lemma 5.26 ensures that $z$ and $z^{\prime}$ have the same degree as $v$. The final sentence of Lemma 5.30 ensures that $Z\setminus\{z,z^{\prime}\}$ has $\deg(z)$-many components and separates $x$ from $y$. ∎
Corollary 5.32.
Let $\Theta$ be a tree system of punctured thick $\theta$-graphs, with underlying tree $T$. Let $Z$ be the limit of $\Theta$. Let $\{x,x^{\prime}\}$ be a Bowditch cut pair of $Z$ of degree $d\geq 3$. Then there exists a finite b-subtree $F_{0}\subset T$ such that, for any finite b-subtree $F\supset F_{0}$, the projections of $x$ and $x^{\prime}$ to the reduced partial union $K_{F}^{\ast}$ are twins of degree $d$.
Proof.
By Lemma 5.26, there exists a finite b-subtree $F_{0}\subset T$ such that, for any finite b-subtree $F\supset F_{0}$, the projections $x_{F}$ and $x^{\prime}_{F}$ of $x$ and $x^{\prime}$ in the reduced partial union $K_{F}^{\ast}$ have degree $d$. Then $x$ and $x^{\prime}$ are the canonical lifts of $x_{F}$ and $x^{\prime}_{F}$ given by Lemma 5.30. It follows from the last sentence of Lemma 5.30 that $K_{F}^{\ast}\setminus\{x_{F},x^{\prime}_{F}\}$ has the same number of components, $d$, as does $Z\setminus\{x,x^{\prime}\}$ and so $x_{F}$ and $x^{\prime}_{F}$ are twins in $K_{F}^{\ast}$. ∎
Lemma 5.33.
Let $\Theta$ be a tree system of punctured thick $\theta$-graphs, with underlying tree $T$. Let $Z$ be the limit of $\Theta$, and let $\{x,x^{\prime}\}$ and $\{y,y^{\prime}\}$ be disjoint Bowditch cut pairs of $Z$. If $\{x,x^{\prime}\}$ has degree at least $3$ and $\{y,y^{\prime}\}$ has degree at least $3$ then there exists a Bowditch cut pair $\{z,z^{\prime}\}$ of degree $2$ such that $\{x,x^{\prime}\}$ and $\{y,y^{\prime}\}$ are contained in distinct components of $Z\setminus\{z,z^{\prime}\}$.
Proof.
It follows from Corollary 5.32 that there exists a finite b-subtree $F_{0}\subset T$ such that, for any finite b-subtree $F\supset F_{0}$, the projection $\{x_{F},x^{\prime}_{F}\}$ of $\{x,x^{\prime}\}$ in the reduced partial union $K_{F}^{\ast}$ is a twin pair of the same degree as $\{x,x^{\prime}\}$ and the projection $\{y_{F},y^{\prime}_{F}\}$ of $\{y,y^{\prime}\}$ in $K_{F}^{\ast}$ is a twin pair of the same degree as $\{y,y^{\prime}\}$. Then $x,x^{\prime},y,y^{\prime}$ are the canonical lifts of $x_{F},x^{\prime}_{F},y_{F},y^{\prime}_{F}$ given by Lemma 5.30 and so $\{x_{F},x^{\prime}_{F}\}$ and $\{y_{F},y^{\prime}_{F}\}$ must be distinct essential twin pairs of $K_{F}^{\ast}$. Recall that $K_{F}^{\ast}$ is obtained from thick $\theta$-graphs by a series of connected sum operations. By first performing the connected sum operations at essential vertices, we can assume that the connected sum operations take place only at degree $2$ points so that, by Proposition 5.19(1), the connected sum operations preserve essential twins. Then $\{x_{F},x^{\prime}_{F}\}$ and $\{y_{F},y^{\prime}_{F}\}$ correspond to essential twin pairs in distinct thick $\theta$-graph summands and, consequently, some blow-up divisor $\{q_{1},q_{2}\}=P_{q}\subset K_{F}^{\ast}$ of a degree $2$ point $q$ at which a connected sum operation was performed separates $\{x_{F},x^{\prime}_{F}\}$ and $\{y_{F},y^{\prime}_{F}\}$ into distinct components. Since connected sums of thick $\theta$-graphs are 2-connected, the complement $K_{F}^{\ast}\setminus\{q_{1},q_{2}\}$ has two components. Since $q_{1}$ and $q_{2}$ are stable points (i.e. no connected sum operation is performed at these points to obtain a larger reduced partial union) they have singleton preimages $\{\tilde{q}_{1}\}$ and $\{\tilde{q}_{2}\}$ in $Z$ and $Z\setminus\{\tilde{q}_{1},\tilde{q}_{2}\}$ is the preimage in $Z$ of $K_{F}^{\ast}\setminus\{q_{1},q_{2}\}$. Then Lemma 4.13 implies that $Z\setminus\{\tilde{q}_{1},\tilde{q}_{2}\}$ has $2$ components and Lemma 5.26 implies that $\tilde{q}_{1}$ and $\tilde{q}_{2}$ have degree $2$ so that we can take $\{z,z^{\prime}\}=\{\tilde{q}_{1},\tilde{q}_{2}\}$. ∎
Recall from Section 3.1 that the reflexive closure of the Bowditch cut pair relation is an equivalence relation. We call the equivalence classes Bowditch classes. Recall that a Bowditch class $\zeta$ of degree at least $3$ has cardinality $|\zeta|=2$, i.e., consists of a single Bowditch cut pair.
We are now ready to prove the main result of this section: the implication (1$\Rightarrow$(3) of Theorem B. We restate and prove the implication below.
Theorem 5.34.
Let $G$ be a $1$-ended hyperbolic group which is not cocompact Fuchsian. If $\partial G$ is homeomorphic to a tree of thick $\theta$-graphs then $G$ has no rigid factor in its (Bowditch) JSJ splitting.
Proof.
Let $T$ be the underlying tree of the tree system of punctured graphs whose limit is homeomorphic to $\partial G$. We identify $\partial G$ with this limit throughout the proof.
Recall from Section 3.1 that a rigid factor of $G$ corresponds to a vertex of type (v3) in the Bowditch JSJ tree $T_{G}$. Such a vertex corresponds to an infinite maximal family $\mathcal{F}$ of pairwise not separated Bowditch classes of $\partial G$. So for any two classes $\zeta,\eta\in\mathcal{F}$ there does not exist a Bowditch class $\xi$ and $x,y\in\xi$ such that $\zeta$ and $\eta$ are contained in distinct components of $\partial G\setminus\{x,y\}$.
We split the rest of the proof into the following two cases: I. $\mathcal{F}$ has two distinct classes $\zeta$ and $\eta$ of degree at least $3$; II. $\mathcal{F}$ has two distinct classes $\zeta$ and $\eta$ of degree $2$.
Case I. $\mathcal{F}$ has two distinct classes $\zeta$ and $\eta$ of degree at least $3$
Then $\zeta$ and $\eta$ each consists of a single Bowditch cut pair of degree at least $3$. Then, by Lemma 5.33, there is a Bowditch cut pair $\{z,z^{\prime}\}$ of degree $2$ such that $\zeta$ and $\eta$ are contained in distinct components of $\partial G\setminus\{z,z^{\prime}\}$, contradicting the fact that $\zeta$ and $\eta$ are not separated.
Case II. $\mathcal{F}$ has two distinct classes $\zeta$ and $\eta$ of degree $2$
Take $x\in\zeta$ an $y\in\eta$. By Lemma 5.28, there are at most two components of $\partial G\setminus\{x,y\}$. But $\zeta$ and $\eta$ are distinct Bowditch classes so the points $x$ and $y$ do not form a Bowditch cut pair. Thus $\partial G\setminus\{x,y\}$ does not have two components and so it must be connected. By Lemma 5.29, there is a finite b-subtree $F\subset T$ for which the projections $x_{F}$ and $y_{F}$ of $x$ and $y$ in the reduced partial union $K_{F}^{\ast}$ have degree $2$ and the complement $K_{F}^{\ast}\setminus\{x_{F},y_{F}\}$ is connected. By Theorem 5.24, the reduced partial sum $K_{F}^{\ast}$ is a twin graph. By Lemma 5.25, there is an essential twin pair $v,\bar{v}$ such that $x_{F}$ and $y_{F}$ are contained in distinct components of $K_{F}^{\ast}\setminus\{v,\bar{v}\}$. Then, by Corollary 5.31, the points $x$ and $y$ are contained in distinct components $C_{x}$ and $C_{y}$ of the complement $\partial G\setminus\{z,z^{\prime}\}$ of a Bowditch cut pair $\{z,z^{\prime}\}$ in $\partial G$ of degree $\deg(v)\geq 3$. Since moreover, by Lemma 5.22, we have $\zeta\subset C_{x}$ and $\eta\subset C_{y}$, this yields a contradiction with the fact that $\zeta$ and $\eta$ are not separated. ∎
6. Regular trees of graphs
In this section we describe a class of tree systems of punctured graphs which depend uniquely on certain finite collections of defining data (which we call graphical connecting systems, see Definition 6.1 below). We call the limits of such tree systems the regular trees of graphs. Although we do not justify this in the present paper (since this is not essential for our results), the regular trees of graphs belong to a larger class of spaces called Markov compacta, which are described by certain recursive procedures based on some finite collections of data (see [Dra07, Paw15, Cza22]). Since, due to the main result from [Paw15], Gromov boundaries of hyperbolic groups are all homeomorphic to Markov compacta, the regular trees of graphs are good candidates for spaces that can be potentially realized as Gromov boundaries of some hyperbolic groups. Actually, the main goal of the present paper is to characterize those 1-ended hyperbolic groups $G$ whose Gromov boundaries $\partial G$ are (homeomorphic to some) trees of graphs, and to show that these trees of graphs are then regular.
Recall that in this paper we allow graphs to have multiple edges and loop edges. Given a graph $\Gamma$ and its vertex $v$, the link of $\Gamma$ at $v$, denoted $\operatorname{Lk}(v)$, is the set of oriented edges of $\Gamma$ issuing from $v$. Equivalently, $\operatorname{Lk}(v)$ can be described as the blow-up divisor $P_{v}$ of the blow-up $|\Gamma|^{\#}(v)$ at $v$ (as described in Section 4.1).
Given an oriented edge $\varepsilon$ of a graph $\Gamma$, we denote by $\bar{\varepsilon}$ the same edge taken with the opposite orientation; we denote by $|\varepsilon|$ the unoriented edge which underlies the oriented edge $\varepsilon$.
Definition 6.1.
A graphical connecting system is a triple $\mathcal{R}=(\Gamma,\mathbf{a},\mathcal{A})$ such that:
  1. (1)
    $\Gamma=\sqcup_{j=1}^{k}\Gamma_{j}$ is the disjoint union of a finite nonempty collection of connected finite graphs $\Gamma_{j}$, each of which is nonempty and not equal to a single vertex (in particular, $k\geq 1$);
  2. (2)
    $\mathbf{a}=(a,\{\alpha_{v}\}_{v\in V_{\Gamma}})$ is a pair consisting of an involution $a:V_{\Gamma}\to V_{\Gamma}$ of the vertex set of $\Gamma$ which preserves degrees of vertices, and of bijections $\alpha_{v}:\operatorname{Lk}(v)\to\operatorname{Lk}(a(v))$ such that for any $v\in V_{\Gamma}$ we have $\alpha_{a(v)}=\alpha_{v}^{-1}$; we will call $\mathbf{a}$ the $V$-involution of $\mathcal{R}$;
  3. (3)
    $\mathcal{A}$ is a set of pairs $(\varepsilon_{1},\varepsilon_{2})$ of (not necessarily distinct) oriented edges of $\Gamma$ closed under the involutions $(\varepsilon_{1},\varepsilon_{2})\to(\varepsilon_{2},\varepsilon_{1})$ and $(\varepsilon_{1},\varepsilon_{2})\to(\bar{\varepsilon}_{1},\bar{\varepsilon}_{2})$, and such that for any oriented edge $\varepsilon$ of $\Gamma$ there is at least one pair in $\mathcal{A}$ containing $\varepsilon$; we will call $\mathcal{A}$ the set of $E$-connections of $\mathcal{R}$;
  4. (4)
    we require also the following transitivity property: for any two distinct components $\Gamma_{i},\Gamma_{j}$ of $\Gamma$ there is a sequence $j(0),j(1),\dots,j(m)$ of indices from $\{1,\dots,k\}$ such that $j(0)=i$, $j(m)=j$, and for each $1\leq n\leq m$ at least one of the following two conditions holds:
    1. (i)
      for some $v\in V_{\Gamma_{j(n-1)}}$ we have $a(v)\in V_{\Gamma_{j(n)}}$;
    2. (ii)
      for some oriented edge $\varepsilon$ of $\Gamma_{j(n-1)}$ there is an oriented edge $\varepsilon^{\prime}$ of $\Gamma_{j(n)}$ such that $(\varepsilon,\varepsilon^{\prime})\in{\mathcal{A}}$.
Before presenting the next crucial definition (Definition 6.2) we need some further terminological and notational preparations. Let ${\mathcal{N}}_{\Gamma}$ be a standard family of normal neighbourhoods of $\Gamma$, as described in Example 2.2, and let $\Gamma^{\circ}:=|\Gamma|\setminus\bigcup\{\operatorname{int}B:B\in{\mathcal{N}}_{\Gamma}\}$ be the space with peripherals, called the punctured graph $\Gamma$, as described in the same example. Denote by $\Omega^{E}_{\Gamma}=\{{\rm bd}(B):B\in{\mathcal{N}}_{\Gamma},B\cap V_{\Gamma}=\emptyset\}$ the set of those peripherals in $\Gamma^{\circ}$ which correspond to those $B\in{\mathcal{N}}_{\Gamma}$ which are normal neighbourhoods of interior points of the edges of $\Gamma$. Similarly, put $\Omega_{\Gamma}^{V}=\{{\rm bd}(B):B\in{\mathcal{N}}_{\Gamma},B\cap V_{\Gamma}\neq\emptyset\}$. We will call the elements of the above sets E-peripherals and V-peripherals, respectively. Observe that each E-peripheral is a set of cardinality 2, i.e., a doubleton. Each such set can be equipped with one of the two possible orders, so that it becomes an ordered pair of points, and we will call any such order an orientation associated to this E-peripheral. We denote by $\widetilde{\Omega}^{E}_{\Gamma}$ the set of all oriented E-peripherals of $\Gamma^{\circ}$, i.e., the set $\{(p,q):\{p,q\}\in\Omega^{E}_{\Gamma}\}$. Observe that each of the oriented E-peripherals may be either compatible or incompatible with the fixed orientation of the edge $e$ of $\Gamma$ in the interior of which it is contained. Actually, in the definition below we use the convention of identifying orientations of E-peripherals of $\Gamma^{\circ}$ with the corresponding compatible orientations of the edges of $\Gamma$ containing those peripherals. For any oriented peripheral $\delta\in\widetilde{\Omega}^{E}_{\Gamma}$, we denote by $|\delta|$ the corresponding underlying (unoriented) peripheral, and by $\bar{\delta}$ the same peripheral $|\delta|$ oriented oppositely to $\delta$.
Definition 6.2.
Given a graphical connecting system $\mathcal{R}=(\Gamma,\mathbf{a},\mathcal{A})$, with $\Gamma$ split into connected components $\Gamma_{j}$, we say that a tree system of punctured graphs $\Theta=(T,\{K_{t}\},\{\Sigma_{u}\},\{\varphi_{e}\})$ is compatible with $\mathcal{R}$, if the following conditions are satisfied:
  1. (1)
    $T$ is the unique tree bipartitioned into black and white vertices such that the degree of any white vertex is 2, and the degree of any black vertex is countable infinite;
  2. (2)
    for any black vertex $t$ of $T$ there is an index $j(t)$ such that $K_{t}\cong\Gamma^{\circ}_{j(t)}$ (i.e. $K_{t}$ is identified with $\Gamma^{\circ}_{j(t)}$, as a space with peripherals);
  3. (3)
    for each white vertex $u$ of $T$, if $t_{1},t_{2}$ are its (black) neighbours, then
    1. (a)
      if the image $\varphi_{[u,t_{1}]}(\Sigma_{u})$ is a V-peripheral of $K_{t_{1}}\cong\Gamma^{\circ}_{j(t_{1})}$, and if it corresponds to a vertex $v\in V_{\Gamma_{j(t_{1})}}$, then $\Gamma_{j(t_{2})}$ is the component of $\Gamma$ which contains the vertex $a(v)$, and the image $\varphi_{[u,t_{2}]}(\Sigma_{u})$ is the V-peripheral of $K_{t_{2}}\cong\Gamma^{\circ}_{j(t_{2})}$ corresponding to $a(v)$; moreover, the composition map $\varphi_{[u,t_{2}]}\circ\varphi_{[u,t_{1}]}^{-1}$ coincides with the bijection $\alpha_{v}$ (after natural identifications of the vertex links with the sets of elements of the V-peripherals, at the vertices $v$ and $a(v)$);
    2. (b)
      if $\varphi_{[u,t_{1}]}(\Sigma_{u})$ is an E-peripheral of $K_{t_{1}}\cong\Gamma^{\circ}_{j(t_{1})}$, then $\varphi_{[u,t_{2}]}(\Sigma_{u})$ is an E-peripheral of $K_{t_{2}}\cong\Gamma^{\circ}_{j(t_{2})}$; moreover, if $e_{i}$ is the edge of $\Gamma$ containing $\varphi_{[u,t_{i}]}(\Sigma_{u})$, for $i=1,2$, and if $\varepsilon_{1}$ is any oriented edge with $|\varepsilon_{1}|=e_{1}$, then denoting by $\varepsilon_{2}$ the oriented edge related to $e_{2}$ whose orientation is induced from that of $\varepsilon_{1}$ via $\varphi_{[u,t_{2}]}\circ\varphi_{[u,t_{1}]}^{-1}$, we have $(\varepsilon_{1},\varepsilon_{2})\in\mathcal{A}$;
  4. (4)
    for any black vertex $t$ of $T$ and for any oriented edge $\varepsilon$ of $\Gamma_{j(t)}$, denote by $\widetilde{\Omega}(\varepsilon)$ the family of those oriented E-peripherals in $K_{t}\cong\Gamma^{\circ}_{j(t)}$ which are contained in $|\varepsilon|$, and which are oriented consistently with the orientation of $\varepsilon$; put also $O_{\varepsilon}:=\{\varepsilon^{\prime}:(\varepsilon,\varepsilon^{\prime})\in\mathcal{A}\}$, and consider the map $f_{\varepsilon}:\widetilde{\Omega}(\varepsilon)\to O_{\varepsilon}$ described as follows: if $\omega\in\widetilde{\Omega}(\varepsilon)$ and $u$ is the white vertex of $T$ for which $|\omega|=\varphi_{[u,t]}(\Sigma_{u})$ and $t^{\prime}$ is the black vertex of $T$ adjacent to $u$ and distinct from $t$ then $f_{\varepsilon}(\omega)$ is the oriented edge $\varepsilon^{\prime}$ of $\Gamma_{j(t^{\prime})}$ containing $\varphi_{[u,t^{\prime}]}(\Sigma_{u})$ and oriented in the way induced by the map $\varphi_{[u,t^{\prime}]}\circ\varphi_{[u,t]}^{-1}$ from the orientation of $\varepsilon$; it is required that for each $\varepsilon^{\prime}\in O_{\varepsilon}$ the union $\bigcup\{|\omega|:\omega\in\widetilde{\Omega}(\varepsilon)\hbox{ and }f_{\varepsilon}(\omega)=\varepsilon^{\prime}\}$ is dense (and in particular nonempty) in $|\varepsilon|\cap\Gamma^{\circ}$.
Though technically involved, the above definitions (i.e. Definitions 6.1 and 6.2) are in fact quite natural, and their significance stems from the following lemma, in view of which to any graphical connecting system $\mathcal{R}$ there is associated some unique up to isomorphism tree system of punctured graphs (which we denote $\Theta[{\mathcal{R}}]$), and some unique up to homeomorphism topological space having the form of a tree of graphs (which we denote $\mathcal{X}(\mathcal{R})$, and which is given as $\mathcal{X}(\mathcal{R}):=\lim\Theta[{\mathcal{R}}]$).
Lemma 6.3.
For any graphical connecting system $\mathcal{R}$ a tree system of punctured graphs compatible with $\mathcal{R}$ exists. Moreover, any two tree systems of punctured graphs compatible with $\mathcal{R}$ are isomorphic.
Proof.
Given any graphical connecting system $\mathcal{R}=(\Gamma,\mathbf{a},\mathcal{A})$, we describe a tree system $\Theta[{\mathcal{R}}]$ of punctured graphs which is by construction compatible with $\mathcal{R}$. After this rather long construction we provide also some less detailed hints for concluding the uniqueness up to isomorphism.
Recall that we denote by $\widetilde{\Omega}_{\Gamma}^{E}$ the set of all oriented E-peripherals in the punctured graph $\Gamma^{\circ}$. We will now describe some auxiliary labelling $\lambda$ of the family $\widetilde{\Omega}_{\Gamma}^{E}$. The labels of this labelling will be in the set $O_{\Gamma}$ of all oriented edges of $\Gamma$. Observe that, by condition (3) of Definition 6.1, for any oriented edge $\varepsilon$ of $\Gamma$ the corresponding set $O_{\varepsilon}$ (as defined in the statement of condition (4) of Definition 6.2) is nonempty. Take as $\lambda$ any labelling of $\widetilde{\Omega}_{\Gamma}^{E}$ such that:
  1. (L1)
    for any $\varepsilon\in O_{\Gamma}$, and for any oriented peripheral $\delta$ contained in the interior of $|\varepsilon|$ whose orientation is compatible with that of $\varepsilon$, the label of $\delta$ is in the subset $O_{\varepsilon}$, i.e., $\lambda(\delta)\in O_{\varepsilon}$;
  2. (L2)
    for any $\delta\in\widetilde{\Omega}_{\Gamma}^{E}$, we have $\lambda(\bar{\delta})=\overline{\lambda(\delta)}$;
  3. (L3)
    for each $\varepsilon\in O_{\Gamma}$ and any $\varepsilon^{\prime}\in O_{\varepsilon}$ the union of the family
    $\{|\delta|\in\Omega^{E}_{\Gamma}:\text{$|\delta|\subset\operatorname{int}|\varepsilon|,\delta$ is oriented compatibly with $\varepsilon$, and $\lambda(\delta)=\varepsilon^{\prime}$}\}$
    is dense in $|\varepsilon|\cap\Gamma^{\circ}$.
It is not hard to realize (and we omit the details) that a labelling $\lambda$ as above exists, and is unique in the sense of the first part of the following claim. We omit the proof, which is a standard back-and-forth argument.
Claim 6.4.
For any two labellings $\lambda,\lambda^{\prime}$ as above, satisfying conditions (L1)(L3), there is a homeomorphism $h:|\Gamma|\to|\Gamma|$ which fixes all vertices of $\Gamma$, which preserves all edges of $\Gamma$, which preserves the family ${\mathcal{N}}_{\Gamma}$, and whose restriction $h^{\circ}=h|_{\Gamma^{\circ}}$ is a homeomorphism of $\Gamma^{\circ}$ such that $\lambda^{\prime}(h^{\circ}(\delta))=\lambda(\delta)$ for any $\delta\in\widetilde{\Omega}_{\Gamma}^{E}$ (where $h^{\circ}(\delta)$ is viewed as an oriented E-peripheral with orientation induced from $\delta$ via $h^{\circ}$). Moreover, if $\delta,\delta^{\prime}\in\widetilde{\Omega}^{E}_{\Gamma}$ are contained in the interior of the same edge of $\Gamma$, if they are consistently oriented, and if $\lambda(\delta)=\lambda^{\prime}(\delta^{\prime})$, then $h$ above can be chosen so that $h^{\circ}(\delta)=\delta^{\prime}$.
We are now ready to describe the tree system $\Theta[{\mathcal{R}}]=(T,\{\Gamma_{t}^{\circ}\},\{\Sigma_{u}\},\{\varphi_{e}\})$. As the underlying tree $T$ take the unique bipartite tree, with the vertex set partitioned into black and white vertices, such that the degree of any white vertex is equal to 2, and the degree of any black vertex is countable infinite. Now, we describe resursively the punctured graphs $\Gamma_{t}^{\circ}$ associated to the black vertices of $T$. Let $t_{0}$ be any black vertex of $T$ (which we will view as the initial vertex). Put $\Gamma_{t_{0}}$ to be any connected component $\Gamma_{i}$ of $\Gamma$. Fix also some auxiliary bijection $\beta_{t_{0}}$ between the set $N_{t_{0}}$ of all (white) vertices of $T$ adjacent to $t_{0}$ and the set of peripherals in the punctured graph $\Gamma_{t_{0}}^{\circ}$. (We will be choosing appropriate bijections $\beta_{t}:t\in V^{b}_{T}$ simultaneously with the description of the constituent spaces $\Gamma_{t}^{\circ}$; these bijections will play an essential role at later stages of the construction of the tree system $\Theta[{\mathcal{R}}]$.) Now, for each $k\geq 0$, denote by $T_{k}$ the b-subtree of $T$ spanned by all vertices of $T$ lying at distance $\leq 2k$ from $t_{0}$. Suppose that for all $t\in V^{b}_{T_{k}}$ we have already chosen the punctured graphs $\Gamma_{t}^{\circ}$ and the corresponding bijections $\beta_{t}$ from the sets $N_{t}$ (of all neighbours of $t$ in the tree $T$) to the sets of peripherals in $\Gamma_{t}^{\circ}$, so that the following conditions hold:
  1. (R1)
    each $\Gamma_{t}$ is a copy of one of the components of $\Gamma$, and we denote this component by $\Gamma_{i(t)}\subset\Gamma$;
  2. (R2)
    for any two black vertices $t,t^{\prime}$ of $T_{k}$ lying at combinatorial distance 2, denoting by $u$ the unique common neighbour in $T$ of those two vertices, identifying the constituent spaces $\Gamma_{t}^{\circ}$, $\Gamma_{t^{\prime}}^{\circ}$ with the corresponding punctured components $\Gamma_{i(t)}^{\circ}$, $\Gamma_{i(t^{\prime})}^{\circ}\subset\Gamma^{\circ}$, viewing them naturally as subspaces in $\Gamma^{\circ}$, and viewing $\lambda$ as the labelling of E-peripherals in both $\Gamma_{t}^{\circ}$ and $\Gamma_{t^{\prime}}^{\circ}$ (via the above mentioned identifications), we have
    1. (a)
      if $\beta_{t}(u)$ is an E-peripheral contained in the interior of an edge $|\varepsilon|$ of $\Gamma_{t}$ (where $\varepsilon\in O_{\Gamma}$) then $\beta_{t^{\prime}}(u)$ is also an E-peripheral; moreover, there is a choice of orientation on the edge containing the peripheral $\beta_{t^{\prime}}(u)$, with which we denote this edge by $\varepsilon^{\prime}$, such that if we orient $\beta_{t}(u)$ and $\beta_{t^{\prime}}(u)$ consistently with $\varepsilon$ and $\varepsilon^{\prime}$, thus getting the oriented peripherals $\delta,\delta^{\prime}$, respectively, we have $\lambda(\delta)=\varepsilon^{\prime}$ and $\lambda(\delta^{\prime})=\varepsilon$;
    2. (b)
      if $\beta_{t}(u)$ is a V-peripheral, i.e., $\beta_{t}(u)={\rm bd}(B)$ for $B$ a normal neighborhood of a vertex $v$ of $\Gamma_{i(t)}$ then $\beta_{t^{\prime}}(u)$ is also a V-peripheral; moreover, if $v^{\prime}$ is the corresponding vertex of $\Gamma_{i(t^{\prime})}$ for $\beta_{t^{\prime}}(u)$, then, viewing $v,v^{\prime}$ as vertices of $\Gamma$, we have $a(v)=v^{\prime}$.
As the main recursive step of the description of constituent spaces, we describe now the spaces $\Gamma_{t}^{\circ}$, and the associated bijections $\beta_{t}$, for all vertices $t\in V^{b}_{T_{k+1}}\setminus V^{b}_{T_{k}}$, so that the above conditions (R1) and (R2) will be satisfied with $k$ replaced by $k+1$. Consider any vertex $t\in V^{b}_{T_{k+1}}\setminus V^{b}_{T_{k}}$, and denote by $s$ the unique vertex in $V^{b}_{T_{k}}$ at combinatorial distance 2 from $t$, and by $u$ the unique common neighbour of both $t$ and $s$ (which is a white vertex). Suppose first that the subspace $\beta_{s}(u)\subset\Gamma_{s}^{\circ}$ is an E-peripheral, i.e., it is contained in the interior of a single edge, say $|\varepsilon|$, of $\Gamma_{s}$, and denote by $\delta$ this same peripheral oriented consistently with $\varepsilon$. Recall that, under our identification conventions, the oriented peripheral $\delta$ has a label, and denote this label by $\varepsilon^{\prime}$. Put $\Gamma_{t}$ to be a copy of this component of $\Gamma$ which contains $|\varepsilon^{\prime}|$, and denote this component by $\Gamma_{i(t)}$. Take as $\beta_{t}$ any bijection from $N_{t}$ to the set of peripherals of $\Gamma_{t}^{\circ}$ for which $\beta_{t}(u)$ is contained in the interior of $|\varepsilon^{\prime}|$ and, after orienting it consistently with $\varepsilon^{\prime}$, it is labeled with $\varepsilon$. Suppose now that $\beta_{s}(u)$ is a V-peripheral, and denote by $v$ this vertex of $\Gamma_{s}$ for which $\beta_{s}(u)=\partial B$, where $B$ is a normal neighbourhood of $v$. Identifying $\Gamma_{s}$ with the component $\Gamma_{i(s)}$ of $\Gamma$, put $v^{\prime}:=a(v)$. Denote by $\Gamma_{i(t)}$ this component of $\Gamma$ which contains $v^{\prime}$, and put $\Gamma_{t}=\Gamma_{i(t)}$. Finally, take as $\beta_{t}$ any bijection from $N_{t}$ to the set of peripherals of $\Gamma_{t}^{\circ}$ for which $\beta_{t}(u)=\partial B^{\prime}$, where $B^{\prime}$ is a normal neighbourhood of $v^{\prime}$. We skip a straightforward verification that this provides data as required.
We now turn to describing the sets $\Sigma_{u}$ and the maps $\varphi_{e}$ of $\Theta[{\mathcal{R}}]$. Let $u$ be any white vertex of $T$, and denote by $t,s$ the two black vertices adjacent to $u$. Suppose first that $\beta_{t}(u)$ is an E-peripheral of $\Gamma_{t}^{\circ}$, i.e., it is contained in the interior of some edge, say $|\varepsilon|$, of $\Gamma_{t}$. Observe that, by the construction, the peripheral $\beta_{s}(u)$ is then also contained in the interior of an edge of the corresponding graph $\Gamma_{s}$, and denote this edge by $|\varepsilon^{\prime}|$. Then $\varepsilon$ and $\varepsilon^{\prime}$ induce orientations on $\beta_{t}(u)$ and $\beta_{s}(u)$ giving us oriented E-peripherals $\delta$ and $\delta^{\prime}$, respectively. Without loss of generality (changing $\varepsilon^{\prime}$ into $\overline{\varepsilon^{\prime}}$ and $\delta^{\prime}$ into $\overline{\delta^{\prime}}$, if necessary) we have $\lambda(\delta)=\varepsilon^{\prime}$ and $\lambda(\delta^{\prime})=\varepsilon$. Put $\Sigma_{u}$ to be a set consisting of two points, with discrete topology, and consider an auxiliary order on this set. Put $\varphi_{[t,u]}:\Sigma_{u}\to|\delta|$ and $\varphi_{[s,u]}:\Sigma_{u}\to|\delta^{\prime}|$ to be the bijections which preserve the distinguished orders.
Finally, suppose that $\beta_{t}(u)$ is a V-peripheral, and denote by $v$ this vertex of $\Gamma_{t}$ for which $\beta_{t}(u)=\partial B$, where $B$ is a normal neighbourhood of $v$. Similarly, denote by $v^{\prime}$ this vertex of $\Gamma_{s}$ for which $\beta_{s}(u)=\partial B^{\prime}$, where $B^{\prime}$ is a normal neighbourhood of $v^{\prime}$. Note that, by (R2)(b), we have $a(v)=v^{\prime}$, and that, identifying $\Gamma_{t}$ and $\Gamma_{s}$ with the corresponding components of $\Gamma$, we have a bijection $\hat{\alpha}_{v}:\beta_{t}(u)\to\beta_{s}(u)$ naturally induced by the bijection $\alpha_{v}$. Take as $\Sigma_{u}$ any set of the same cardinality as $\beta_{t}(u)$ (and as $\beta_{s}(u)$), with discrete topology. Take as the maps $\varphi_{[u,t]}$ and $\varphi_{[u,s]}$ any bijections for which the composition $\varphi_{[u,s]}\circ\varphi_{[u,t]}^{-1}$ coincides with $\hat{\alpha}_{v}$.
Having completed the construction of $\Theta[\mathcal{R}]$, we skip the details of a verification that this system is compatible with $\mathcal{R}$. We also note that an isomorphism between any two tree systems compatible with $\mathcal{R}$, as in the second assertion of the lemma, can be constructed recursively, in a straightforward way, using Claim 6.4 (and more precisely, its final assertion). To start such recursive construction, one needs to refer to the transitivity condition (4) of Definition 6.1. We omit further details. ∎
Example 6.5.
For any finite connected graph $\Gamma$, let ${\mathcal{R}}^{r}_{\Gamma}=(\Gamma,\mathbf{a},\mathcal{A})$ be a graphical connecting system in which $a={\rm id}_{V_{\Gamma}}$, $\alpha_{v}={\rm id}_{\operatorname{Lk}(v)}$ for each $v\in V_{\Gamma}$, and ${\mathcal{A}}=\{(\varepsilon,\varepsilon):\varepsilon\in O_{\Gamma}\}$, where $O_{\Gamma}$ denotes the set of all oriented edges of $\Gamma$. We call ${\mathcal{R}}^{r}_{\Gamma}$ the reflection connecting system for $\Gamma$. The associated tree systems $\Theta[{\mathcal{R}}^{r}_{\Gamma}]$ lead to a class of regular trees of graphs (see Definition 6.6 below) which are called reflection trees of graphs and which have been described and studied earlier in [Ś21].
Definition 6.6.
A regular tree of graphs is any space of the form ${\mathcal{X}}({\mathcal{R}})=\lim\Theta[{\mathcal{R}}]$, where $\mathcal{R}$ is a graphical connecting system, and where $\Theta[{\mathcal{R}}]$ is the (unique up to an isomorphism) tree system of punctured graphs compatible with $\mathcal{R}$. Moreover, we call any space ${\mathcal{X}}({\mathcal{R}})$ as above the regular tree of graphs for $\mathcal{R}$.
Definition 6.7.
A graphical connecting system ${\mathcal{R}}=(\Gamma,\mathbf{a},\mathcal{A})$ is 2-connected if all the connected components $\Gamma_{j}$ of $\Gamma$ are 2-connected.
Note that, if $\mathcal{R}$ is 2-connected, then obviously the corresponding system $\Theta[{\mathcal{R}}]$ is a tree system of punctured 2-connected graphs, and consequently, the regular tree of graphs ${\mathcal{X}}(\mathcal{R})$ is then connected, locally connected, cutpoint-free and has topological dimension 1 (see Lemma 4.14 and Lemma 4.15 above). We call the spaces ${\mathcal{X}}(\mathcal{R})$ as above the regular trees of 2-connected graphs.
7. $V$-trees of graphs and their decompositions into core and arms
In this section we begin preparations for the proof of the implication (3$\Rightarrow$(2) in Theorem A (and the implication (3$\Rightarrow$(2) of Theorem B). The V-tree systems as discussed in this section will appear later as subsystems in some special regular tree systems of graphs that we will be considering, and their closer understanding is thus a first step in understanding the latter, and in relating the latter with Gromov boundaries of hyperbolic groups that are under our interest.
Definition 7.1 (connecting $V$-system).
A connecting $V$-system is a pair ${\mathcal{V}}=(\Gamma,{\mathbf{a}})$ such that:
  1. (1)
    $\Gamma$ is a finite graph which is essential (i.e. is nonempty and has no isolated vertex), and we assume it has no loop edges;
  2. (2)
    ${\mathbf{a}}=(a,\{\alpha_{v}\}_{v\in V_{\Gamma}})$ is a pair consisting of an involution $a:V_{\Gamma}\to V_{\Gamma}$ of the vertex set of $\Gamma$ which preserves degrees of vertices, and of bijections $\alpha_{v}:\hbox{Lk}(v)\to\hbox{Lk}(a(v))$ such that for any $v\in V_{\Gamma}$ we have $\alpha_{a(v)}=\alpha_{v}^{-1}$, where links are taken in $\Gamma$.
A connecting $V$-system is thus a structure consisting of slightly less data than a graphical connecting system, as described in Definition 6.1, and we insist on graphs $\Gamma$ to have a more restrictive form (no loop edges). Intuitively, a $V$-system structure contains only the information about connected sum operations at vertices of the involved copies of graphs, as encoded in the $V$-involution $\mathbf{a}$. Although in our later applications graphs $\Gamma$ of $V$-systems will be typically connected, we do not need to assume this here, in the description of the general concept.
We now pass to describing a certain tree system $\Theta[{\mathcal{V}}]=\Theta[\Gamma,{\mathbf{a}}]$ canonically associated to a connecting $V$-system $\mathcal{V}$. Its limit $\lim\Theta[{\mathcal{V}}]$ will be called the $V$-tree of graphs for $\mathcal{V}$, and it will be denoted ${\mathcal{X}}_{V}({\mathcal{V}})$, or ${\mathcal{X}}_{V}(\Gamma,{\mathbf{a}})$, where the role of the subscript $V$ is just to indicate the $V$-tree nature of the resulting space. Actually, the $V$-tree ${\mathcal{X}}_{V}({\mathcal{V}})$ will be natuarally a space with peripherals, and the description of a family of peripherals will be a part of its definition.
Similarly as in Section 6, we consider a standard family ${\mathcal{N}}_{\Gamma}$ of normal neighbourhoods in $\Gamma$, as described in Example 2.2, and the space with peripherals $\Gamma^{\circ}$, called the punctured graph $\Gamma$, described in the same example. Consider also the set $\Omega^{V}_{\Gamma}$ of $V$-peripherals of $\Gamma^{\circ}$, as described right after Definition 6.1.
As the underlying tree $T$ of the $V$-tree system $\Theta[{\mathcal{V}}]=(T,\{\Gamma_{t}^{\circ}\},\{\Sigma_{u}\},\{\varphi_{e}\})$ we take the unique tree, with the vertex set bipartitioned into black and white vertices, such that the degree of any white vertex is equal to 2, and the degree of any black vertex coincides with the (finite) cardinality of the vertex set $V_{\Gamma}$ of $\Gamma$. The system $\Theta[{\mathcal{V}}]$ is then described uniquely up to an isomorphism of tree systems by the following requirements:
  1. (v1)
    for each black vertex $t$ of $T$ the constituent space with peripherals at $t$ is equal to the pair $(\Gamma^{\circ},\Omega^{V}_{\Gamma})$; this means in particular that the vertices in $N_{t}$ (i.e. the vertices adjacent to $t$) are in fixed bijective correspondence with the family $\Omega_{\Gamma}^{V}$;
  2. (v2)
    for each white vertex $u$ of $T$, denote by $e,e^{\prime}$ the two edges of $T$ adjacent to $u$, and by $P_{e},P_{e^{\prime}}$ the $V$-peripherals of $\Gamma_{b(e)}^{\circ}$ and $\Gamma_{b(e^{\prime})}^{\circ}$ associated to $u$ via the bijective correspondences mentioned above in condition (v1); then, denoting by $v,v^{\prime}$ the vertices of $\Gamma$ corresponding to the $V$-peripherals $P_{e},P_{e^{\prime}}$, respectively, we have $a(v)=v^{\prime}$; moreover, the set $\Sigma_{u}$ and the bijections $\varphi_{e}:\Sigma_{u}\to P_{e},\varphi_{e^{\prime}}:\Sigma_{u}\to P_{e^{\prime}}$ are given so that the composition $\varphi_{e^{\prime}}\circ\varphi_{e}^{-1}:P_{e}\to P_{e^{\prime}}$ coincides with the map $\alpha_{v}$, where we refer to the natural identifications $P_{e}={\rm Lk}(v,\Gamma)$ and $P_{e^{\prime}}={\rm Lk}(v^{\prime},\Gamma)$.
We skip a straightforward justification of the fact that a tree system satisfying the above requirements exists and is unique up to an isomorphism of tree systems (the argument is similar to that presented in Section 6 for the tree system $\Theta[{\mathcal{R}}]$; see the proof of Lemma 6.3).
Now, let ${\mathcal{X}}_{V}({\mathcal{V}}):=\lim\Theta[{\mathcal{V}}]$. We describe the family $\Omega_{V}({\mathcal{V}})$ of natural peripheral subsets for ${\mathcal{X}}_{V}({\mathcal{V}})$. Let $P$ be any $E$-peripheral in any constituent space $\Gamma^{\circ}_{t}$, $t\in V^{b}_{T}$. Clearly, $P$ is then naturally a subset of the limit $\lim\Theta[{\mathcal{V}}]$. Moreover, for any two distinct $E$-peripherals $P,P^{\prime}$ as above, the induced subsets of $\lim\Theta[{\mathcal{V}}]$ are disjoint. It is also not hard to realize (by deducing it e.g. from Fact 2.5 in view of Fact 2.6) that the family $\Omega_{V}({\mathcal{V}})$ of all such subsets $P\subset\lim\Theta[{\mathcal{V}}]$, for all $t\in V^{b}_{T}$, is null, and thus we take it as the family of natural peripherals for ${\mathcal{X}}_{V}({\mathcal{V}})$.
Example 7.2 ($V$-tree of segments).
A graph which coincides with a single edge will be called the segment, and we donote this graph by $I$. The punctured segment $I^{\circ}$ is then a space with peripherals in which two of the peripheral sets are singletons, while the others are doubletons. If we delete from the set of peripherals of $I^{\circ}$ the two singleton peripherals, we get a new space with peripherals (with the same underlying space, but smaller family of peripherals) which we call the internally punctured segment, and which we denote $I^{\circ}_{\rm int}$. Alternatively, $I^{\circ}_{\rm int}$ can be described as obtained by deleting from $I$ the interiors of a dense and null family of pairwise disjoint normal neighbourhoods of points lying in the interior of $I$, and by taking the boundaries of those neighbourhoods as peripherals.
Observe that there are exactly two essentially distinct connecting $V$-systems ${\mathcal{V}}=(\Gamma,{\mathbf{a}})$ with $\Gamma=I$: the first one with the involution $a:V_{I}\to V_{I}$ equal to the identity of the vertex set $V_{I}$, and the second one with $a$ equal to the transposition of the doubleton $V_{I}$. In both cases the resulting $V$-tree ${\mathcal{X}}_{V}({\mathcal{V}})$ will be called a $V$-tree of segments. Moreover, we obviously have the following.
Fact 7.3.
Any $V$-tree of segments is, as a space with peripherals, homeomorphic to the internally punctured segment $I^{\circ}_{\rm int}$.
We now pass to describing some natural decomposition of any $V$-tree of graphs ${\mathcal{X}}_{V}({\mathcal{V}})$ into convenient pieces. These pieces will be called the core and the arms of ${\mathcal{X}}_{V}({\mathcal{V}})$, and they will have a form of spaces with peripherals. We will show (in Lemma 7.12) that decomposition into these pieces has a tree-like nature, i.e., that the space ${\mathcal{X}}_{V}({\mathcal{V}})$ is canonically homeomorphic to a certain tree of these pieces. This new expression of ${\mathcal{X}}_{V}({\mathcal{V}})$ as a tree of spaces does not coincide with the original expression of this space as a $V$-tree of graphs, and this reinterpretation of the spaces of the form ${\mathcal{X}}_{V}({\mathcal{V}})$ will be one of our tools in the proof of the implications (3$\Rightarrow$(2) in Theorem A and (3$\Rightarrow$(2) in Theorem B.
Before describing the pieces and the decomposition, we need to introduce more terminology. A punctured edge of a punctured graph $\Gamma^{\circ}$ is the intersection of $\Gamma^{\circ}$ (viewed as subspace of $\Gamma$) with any edge $e$ of $\Gamma$, and we denote such a punctured edge by $e^{\circ}$. Obviously, since we assume that $\Gamma$ contains no loop edges, each $e^{\circ}$ naturally has the structure of a space with peripherals with which it is homeomorphic to the punctured segment $I^{\circ}$. The endpoints of a punctured edge $e^{\circ}$ are the two points of $e^{\circ}$ which correspond to the two singleton peripherals of this space. The set of all endpoints in all punctured edges of $\Gamma^{\circ}$ is clearly in a natural bijective correspondence with the union of all vertex links of the graph $\Gamma$, and we will identify each such endpoint with the corresponding point in the appropriate vertex link. Now, given a $V$-tree system $\Theta[{\mathcal{V}}]=(T,\{\Gamma_{t}^{\circ}\},\{\Sigma_{u}\},\{\varphi_{e}\})$, we say that punctured edges $e_{1}^{\circ}$ of $\Gamma_{t_{1}}^{\circ}$ and $e_{2}^{\circ}$ of $\Gamma_{t_{2}}^{\circ}$ are $\Theta$-adjacent if some endpoints $x_{1},x_{2}$ in $e_{1}^{\circ},e_{2}^{\circ}$ get identified in the limit space ${\mathcal{X}}_{V}({\mathcal{V}})$, i.e., if $t_{1},t_{2}$ are distinct black vartices of $T$ with common adjacent white vertex, which we denote $u$, and if we have $x_{i}\in\varphi_{[t_{i},u]}(\Sigma_{u})$, for $i=1,2$, and $\varphi_{[t_{1},u]}^{-1}(x_{1})=\varphi_{[t_{2},u]}^{-1}(x_{2})$. We then say that $e_{1}^{\circ}$ and $e_{2}^{\circ}$ are $\Theta$-adjacent at their endpoints $x_{1},x_{2}$. Note that if $t_{1},t_{2},u$ are as above, and if $e_{1}^{\circ}$ is a punctured edge of $\Gamma_{t_{1}}^{\circ}$ having an endpoint $x_{1}\in\varphi_{[t_{1},u]}(\Sigma_{u})$, then there is precisely one punctured edge $e_{2}^{\circ}$ in $\Gamma_{t_{2}}^{\circ}$ which is $\Theta$-adjacent to $e_{1}^{\circ}$.
Definition 7.4 (line in $\Theta[{\mathcal{V}}]$).
A line in a $V$-tree system $\Theta[{\mathcal{V}}]$ is a bi-infinite sequence $(e_{n}^{\circ})_{n\in\mathbb{Z}}$ of punctured edges from the punctured graphs $\Gamma_{t}^{\circ}:t\in V^{b}_{T}$ such that for each $n$ the edge $e_{n}^{\circ}$ is $\Theta$-adjacent to $e_{n-1}^{\circ}$ at one of its endpoints, and it is $\Theta$-adjacent to $e_{n+1}^{\circ}$ at its other endpoint. We do not distinguish sequences as above which differ by a shift or by inversion of the indices (i.e. we view two such sequences as representing the same line).
Obviously, since $\Gamma$ has no loop edges, any punctured edge in any constituent graph $\Gamma_{t}^{\circ}$ belongs to a unique line in $\Theta[{\mathcal{V}}]$. It is also not hard to observe that in each $V$-tree system $\Theta[{\mathcal{V}}]$ for which $\Gamma$ has more than two vertices, there are many (actually countably infinitely many) lines. We denote the set of all lines in $\Theta[{\mathcal{V}}]$ by ${\mathcal{L}}({\mathcal{V}})$. Since each term $e_{n}^{\circ}$ of any line $L$ is a punctured edge in some copy $\Gamma_{t}$ of the graph $\Gamma$, we will often refer to it as an appearance in $L$ of the corresponding (unpunctured) edge of $\Gamma$.
Let $L=(e^{\circ}_{n})_{n\in\mathbb{Z}}$ be a line in $\Theta[{\mathcal{V}}]$, and for each $n$ denote by $t_{n}$ this black vertex of $T$ for which $e_{n}^{\circ}$ is a punctured edge of $\Gamma_{t_{n}}^{\circ}$, and by $u_{n}$ the unique white vertex of $T$ adjacent to both $t_{n-1}$ and $t_{n}$. The bi-infinite sequence
$\dots,u_{n},t_{n},u_{n+1},t_{n+1},\dots$
is then a (bi-infinite) combinatorial geodesic in $T$, which we denote by $\gamma_{L}$ and call the geodesic induced by $L$. (We view such geodesics as subsets in the vertex set $V_{T}$ rather than as sequences, so that a shift of the indices or reversal of the order of terms yields the same geodesic.) We make a record of the following properties of lines and their induced geodesics, whose straightforward proofs we omit. (These proofs are based on the easily seen observation that lines, viewed as sequences of appearances of edges from $\Gamma$, are periodic.)
Lemma 7.5.
  1. (1)
    Given two distinct lines $L_{1}\neq L_{2}$ in $\Theta[{\mathcal{V}}]$, their corresponding geodesics $\gamma_{L_{1}},\gamma_{L_{2}}$ either coincide, or have bounded intersection in $T$ (including the case of empty intersection). In particular, the doubleton subsets $\partial\gamma_{L_{1}},\partial\gamma_{L_{2}}$ in $\partial T$ either coincide or are disjoint.
  2. (2)
    There is a universal finite bound on the cardinality of a set of lines $L$ in $\Theta[{\mathcal{V}}]$ which share the corresponding doubletons $\partial\gamma_{L}$.
Consider the equivalence relation $\sim_{\Theta}$ on the set of edges of the graph $\Gamma$ described as follows: we have $e\sim_{\Theta}e^{\prime}$ iff $e$ and $e^{\prime}$ have appearances in the same line $L\in{\mathcal{L}}({\mathcal{V}})$. Note that this equivalence relation is well defined since if some two lines $L_{1},L_{2}$ contain appearances of the same edge of $\Gamma$, then any edge of $\Gamma$ appears in $L_{1}$ if and only if it appears in $L_{2}$. Thus, we get a partition of the edge set $E_{\Gamma}$ into classes of the relation $\sim_{\Theta}$.
Definition 7.6 (arm of ${\mathcal{X}}_{V}({\mathcal{V}})$).
An arm in a $V$-tree of graphs ${\mathcal{X}}_{V}({\mathcal{V}})$ is its subspace $A_{L}$ of the following form. Given any line $L\in{\mathcal{L}}({\mathcal{V}})$, $L=(e_{n}^{\circ})_{n\in Z}$, take as $A_{L}$ the union of the set $\partial\gamma_{L}\subset\partial T$ (consisting of the two ends of the geodesic $\gamma_{L}$) and the image in #$\#\Theta[{\mathcal{V}}]$ of the union $\bigcup_{n\in Z}e_{n}^{\circ}$. We view $A_{L}$ as a space with peripherals, where the family of peripherals consists of only one peripheral set, namely $\partial\gamma_{L}$. We distinguish also in $A_{L}$ some other subsets, namely the ones corresponding to all doubleton peripherals in all punctured edges $e_{n}^{\circ}$ constituting $A_{L}$; we call these subsets the supplementary peripherals of $A_{L}$, in contrast with the above mentioned subset $\partial\gamma_{L}$, which we call the basic peripheral of $A_{L}$.
For each arm $A_{L}$ we consider the subspace topology induced from ${\mathcal{X}}_{V}({\mathcal{V}})$. It is not hard to observe by referring to Fact 2.5 that, when equipped with this topology, any arm $A_{L}$ naturally gets the form of a $V$-tree of segments. In particular, by Fact 7.3, we obtain the following.
Fact 7.7.
Any arm $A_{L}$, as a space equipped with the family of all supplementary peripherals, is homeomorphic to the internally punctured segment $I^{\circ}_{\rm int}$. Moreover, its basic peripheral subset $\partial\gamma_{L}$ coincides then with the set of endpoints of $I^{\circ}_{\rm int}$.
Definition 7.8 (core of ${\mathcal{X}}_{V}({\mathcal{V}})$).
The core of ${\mathcal{X}}_{V}({\mathcal{V}})=\lim\Theta[{\mathcal{V}}]$ is the subspace $C({\mathcal{V}})$ corresponding to the subset $\partial T\subset\lim\Theta[{\mathcal{V}}]$. It is equipped (as part of its structure) with the family of doubleton subsets $\Omega^{c}_{\mathcal{V}}:=\{\partial\gamma_{L}:L\in{\mathcal{L}}({\mathcal{V}})\}$ (with which it is a space with peripherals, see Remark 7.9.2 below).
Remark 7.9.
  1. 1.
    Using the fact that the tree $T$ is locally finite, it is not hard to see that the subspace topology in the subset $\partial T\subset\lim\Theta[{\mathcal{V}}]$ coincides with the standard topology in $\partial T$, and hence $\partial T$ is a Cantor set (except the elementary cases when $\Gamma$ has two vertices, in which $\partial T$ is a doubleton).
  2. 2.
    The family $\Omega^{c}_{\mathcal{V}}$ of subsets of the core is null. To see this, it is sufficient to note that the distances in $T$ from a fixed base point to the geodesics $\gamma_{L}$ diverge to $\infty$. This, in turn, easily follows from the fact that only finitely many geodesics of the form $\gamma_{L}$ passes through any vertex of $T$, and the fact that $T$ is locally finite. The family $\Omega^{c}_{\mathcal{V}}$ consists also of sets that are pairwise disjoint, see Lemma 7.5(1).
  3. 3.
    The sets in the family $\Omega^{c}_{\mathcal{V}}$, viewed as subsets of the $V$-tree ${\mathcal{X}}_{V}({\mathcal{V}})$, clearly coincide with the basic peripherals in all arms of this $V$-tree.
Example 7.10 (core and arm in a $V$-tree of segments).
Note that for any connecting $V$-system of segments ${\mathcal{V}}=(I,{\mathbf{a}})$ (as described in Example 7.2), there is precisely one line $L$ in this system. The $V$-tree of segments ${\mathcal{X}}_{V}({\mathcal{V}})$ has then precisely one arm, $A_{L}$, and the core $C({\mathcal{V}})$ is a doubleton which coincides with the boundary $\partial\gamma_{L}$ and with the set of endpoints of the arm $A_{L}$. Actually, as a space with peripherals the $V$-tree ${\mathcal{X}}_{V}({\mathcal{V}})$ coincides with the arm $A_{L}$ equipped with its supplementary peripherals, and the basic peripheral of $A_{L}$ coincides with the core $C({\mathcal{V}})$.
We now generalize V-trees of segments to V-trees of $\theta$-graphs. We start with describing the $\theta$-graphs—a family of graphs which appear explicitly in the statement of Theorem B, and which also appear implicitly in Theorem A, since the regular trees of graphs mentioned in conditions (2) and (3) of this result necessarily involve certain $\theta$-graphs (compare Theorem 10.1 and the description of the graphical connecting system ${\mathcal{R}}_{G}$ in Section 10).
Example 7.11 (V-tree of $\theta$-graphs).
A $\theta$-graph is a finite connected graph $\Xi$ which has precisely two vertices and no loop edges. The number of edges is thus a natural number $k\geq 1$, and each edge has both vertices of $\Xi$ as its endpoints. We will denote the $\theta$-graph with $k$ edges by $\Xi_{k}$. Obviously, $\Xi_{1}$ is simply the segment, as described in Example 7.2.
Generalizing the exposition of Example 7.2, we define the internally punctured $\theta$-graph $(\Xi_{k})^{\circ}_{{\rm int}}$ as the space with peripherals obtained from $\Xi_{k}$ by deleting the interiors of any dense family of pairwise disjoint normal neighbourhoods of non-vertex points of $\Xi_{k}$, and by taking the boundaries of those neighbourhoods as peripherals. Such a space is obviously unique up to a homeomorphism respecting peripherals, its peripherals are all doubletons, and vertices of $\Xi_{k}$ belong to this space, and are not contained in any peripheral.
Observe that any connecting $V$-system ${\mathcal{V}}=(\Xi_{k},{\mathbf{a}})$ of $\theta$-graphs has precisely $k$ lines $L_{1},\dots,L_{k}$. Moreover, the arms $A_{1},\dots,A_{k}$ of the corresponding $V$-tree of $\theta$-graphs ${\mathcal{X}}_{V}({\mathcal{V}})$ have a common pair of endpoints, and this pair coincides with the core $C({\mathcal{V}})$ of this $V$-tree. The set of peripherals of ${\mathcal{X}}_{V}({\mathcal{V}})$ coincides with the union of the sets of supplementary peripherals of the arms $A_{i}$. Thus, as a space with peripherals, ${\mathcal{X}}_{V}({\mathcal{V}})$ is homeomorphic to the internally punctured $\theta$-graph $(\Xi_{k})^{\circ}_{\rm int}$. The core $C({\mathcal{V}})$ of this $V$-tree coincides then with the vertex set of $\Xi_{k}$ (viewed as the subset of $(\Xi_{k})^{\circ}_{\rm int}$). The only peripheral of this core coincides with the whole of the core.
For our later purposes, we now describe some special connecting $V$-system ${\mathcal{V}}_{\Xi_{k}}=(\Xi_{k},{\mathbf{a}}_{k})$, which we call the standard connecting $V$-system for $\Xi_{k}$. The $V$-involution ${\mathbf{a}}_{k}=(a_{k},\{\alpha_{u},\alpha_{v}\})$ consists of the transposition $a_{k}$ of the two vertices $u,v$ of $\Xi_{k}$, and both $V$-maps $\alpha_{u}:\hbox{Lk}(u)\to\hbox{Lk}(v)$, $\alpha_{v}:\hbox{Lk}(v)\to\hbox{Lk}(u)$ are “tautological” in the following sense: for any element $p$ in the link $\hbox{Lk}(u)$ or $\hbox{Lk}(v)$ its image through the corresponding $V$-map coincides with this element in the other vertex link which is induced by the same edge of $\Xi_{k}$ as $p$. Observe that all $k$ lines in so described standard system ${\mathcal{V}}_{\Xi_{k}}$ consist of appearances of a single edge of the graph $\Xi_{k}$, and those edges are distinct for distict lines.
Coming back to general connecting $V$-systems $\mathcal{V}$, we now describe the tree system $\Psi[{\mathcal{V}}]$ consisting of the core and the arms of the $V$-tree ${\mathcal{X}}_{V}({\mathcal{V}})$ as the constituent spaces of this system. The underlying tree $T_{\Psi}$ of $\Psi[{\mathcal{V}}]$ is described as follows. There are three kinds of vertices in $T_{\Psi}$:
  1. (1)
    a single “central” black vertex $t_{c}$ corresponding to the core of ${\mathcal{X}}_{V}({\mathcal{V}})$;
  2. (2)
    black vertices $t_{A}$ corresponding to arms $A$ of ${\mathcal{X}}_{V}({\mathcal{V}})$;
  3. (3)
    white vertices $u_{P}$ corresponding to peripherals $P\in\Omega^{c}_{\mathcal{V}}$.
The set of edges is described as follows. For each of the white vertices $u_{P}$ there is an edge $[u_{P},t_{c}]$ in $T_{\Psi}$. For each arm $A=A_{L}$, putting $P=\partial\gamma_{L}$, there is an edge $[t_{A},u_{P}]$ in $T_{\Psi}$. Note that, due to Lemma 7.5(2), the tree $T_{\Psi}$ has finite degree at each white vertex $u_{P}$.
As the constituent spaces for $\Psi[{\mathcal{V}}]$ we take $K_{t_{c}}=C({\mathcal{V}})$ and $K_{t_{A}}=A$. As the peripheral spaces for $\Psi[{\mathcal{V}}]$ we take $\Sigma_{u_{P}}=P$, and as the maps $\varphi_{[u,t]}:\Sigma_{u}\to K_{t}$ we take the natural inclusions of the forms $P\subset C[{\mathcal{V}}]$ and $\partial\gamma_{L}\subset A_{L}$.
We have the following expression of the space ${\mathcal{X}}_{V}({\mathcal{V}})$ as a tree of spaces.
Lemma 7.12.
The inclusions of the form $A_{L}\subset{\mathcal{X}}_{V}({\mathcal{V}})$, for all $L\in{\mathcal{L}}({\mathcal{V}})$, and $C[{\mathcal{V}}]\subset{\mathcal{X}}_{V}({\mathcal{V}})$ induce a well defined map $i_{\mathcal{V}}:\lim\Psi[{\mathcal{V}}]\to{\mathcal{X}}_{V}({\mathcal{V}})$, and this map is a homeomorphism.
Proof.
The induced map $i_{\mathcal{V}}$ is obviously a bijection. Since both spaces $\lim\Psi[{\mathcal{V}}]$ and ${\mathcal{X}}_{V}({\mathcal{V}})$ are compact and metrizable, it is enough to show that the map $i_{\mathcal{V}}$ is continuous. We leave it to the reader as a straightforward exercise concerning the topology on limits of tree systems. ∎
Example 7.13 (wedge of arms).
We present here some alternative expression of the $\mathcal{V}$-tree of $\theta$-graphs ${\mathcal{X}}_{V}[{\mathcal{V}}_{\Xi_{k}}]$ as tree of arms. This expression, a bit simpler than that provided by the tree system $\Psi[{\mathcal{V}}_{\Xi_{k}}]$, will be used in the proof of the implication (3$\Rightarrow$(2) of Theorem A given in Section 13.
Consider the tree system $\Psi_{w}[{\mathcal{V}}_{\Xi_{k}}]$ described as follows. The underlying tree $T_{w}$ of $\Psi_{w}[{\mathcal{V}}_{\Xi_{k}}]$ consists of a single white vertex $u_{c}$ and precisely $k$ black vertices $t_{A}$ corresponding to the arms $A$ of ${\mathcal{X}}_{V}({\mathcal{V}}_{\Xi_{k}})$, and of precisely $k$ edges of the form $[u_{c},t_{A}]$. As the constituent spaces for $\Psi_{w}[{\mathcal{V}}_{\Xi_{k}}]$ we take $K_{t_{A}}=A$, and as the peripheral space we take $\Sigma_{u_{c}}=C[{\mathcal{V}}_{\Xi_{k}}]$. As the maps $\varphi_{[u_{c},t_{A}]}:\Sigma_{u_{c}}\to K_{t_{A}}$ we take the natural inclusions of the form $C[{\mathcal{V}}_{\Xi_{k}}]\subset A$.
Since, by an easy observation, we have a natural identification $\lim\Psi_{w}[{\mathcal{V}}_{\Xi_{k}}]=\lim\Psi[{\mathcal{V}}_{\Xi_{k}}]$, it follows then from Lemma 7.12 that the inclusions $A\subset{\mathcal{X}}_{V}({\mathcal{V}}_{\Xi_{k}})$ induce a natural identification of the limit space $\lim\Psi_{w}[{\mathcal{V}}_{\Xi_{k}}]$ with the $V$-tree ${\mathcal{X}}_{V}({\mathcal{V}}_{\Xi_{k}})$.
As the last subject in this section, we discuss orientability of arms in V-trees of graphs.
Definition 7.14 (orientable line and arm).
A line $L$ in a V-tree system $\Theta[{\mathcal{V}}]$ is orientable if any two appearances $e_{k}^{\circ}$, $e_{m}^{\circ}$ in $L$ of any edge $e$ of $\Gamma$ are consistently oriented. The latter means that in the corresponding arm $A_{L}$, viewed as an internally punctured segment, the orientations of subsegments $e_{k}^{\circ}$ and $e_{m}^{\circ}$ induced from a fixed orientation of $e$ are consistent. A line $L$ is non-orientable, if some edge $e$ of $\Gamma$ has some two inconsistently oriented appearances in $L$. Accordingly, we speak of orientability and non-orientability of the corresponding arms $A_{L}$.
We have the following useful characterization of orientability of lines (and arms).
Lemma 7.15.
Let ${\mathcal{V}}=(\Gamma,{\mathbf{a}})$ be a connecting V-system. Then the V-tree system $\Theta[{\mathcal{V}}]$ contains a nonorientable line if and only if there is $v\in V_{\Gamma}$ such that $a(v)=v$ and there is $p\in\hbox{Lk}(v)$ such that $\alpha_{v}(p)=p$. In particular, if the involution $a:V_{\Gamma}\to V_{\Gamma}$ is fixed point free then all lines in the V-tree system $\Theta[{\mathcal{V}}]$ are orientable.
Proof.
If there are $v$ and $p$ such that $a(v)=v$ and $\alpha_{v}(p)=p$, denote by $e$ this edge of $\Gamma$ issuing from $v$ which corresponds to $p$. The equalities above obviously imply that in any line $L$ containing an appearance of $e$ we have another appearance of $e$, adjacent to this first appearance, and oppositely oriented. Thus any such line $L$ is nonorientable.
To prove the converse implication, observe that all appearances of any oriented edge $\varepsilon$ of $\Gamma$ in lines of $\Theta[{\mathcal{V}}]$ are followed by adjacent appearances of some other oriented edge uniquely determined by $\varepsilon$ (which is not necessarily distinct from $\varepsilon$). Thus, if some line contains oppositely oriented appearances of some edge, it must also contain, somewhere in between, oppositely oriented and adjacent appearances of some other edge, which clearly implies existence of $v$ and $p$ as in the statement. This completes the proof. ∎
Remark/Definition 7.16.
  1. (1)
    If a line $L$ is orientable, and if we fix an order of the doubleton $\partial\gamma_{L}$ (i.e. of the set of endpoints of the arm $A_{L}$), then this order induces orientations at all those edges of $\Gamma$, which have appearances in $L$. We call such orientations of the above mentioned edges the order-induced orientations (with respect to the above fixed order of $\partial\gamma_{L}$).
  2. (2)
    Let ${\mathcal{V}}=(\Gamma,{\mathbf{a}})$ be a connecting $V$-system, and let $P\subset{\mathcal{C}}({\mathcal{V}})$ be a peripheral of the core ${\mathcal{C}}({\mathcal{V}})$ of the $V$-tree ${\mathcal{X}}_{V}({\mathcal{V}})$. Consider the family ${\mathcal{L}}_{P}$ of all lines $L$ in $\Theta[{\mathcal{V}}]$ such that $\partial\gamma_{L}=P$. Note that ${\mathcal{L}}_{P}$ is nonempty and finite. Define the set of edges of $\Gamma$ associated to $P$, denoted $E_{\Gamma}^{P}$, as the set of all edges which have an appearance in any line $L\in{\mathcal{L}}_{P}$.
  3. (3)
    Suppose that the $V$-involution map $a:V_{\Gamma}\to V_{\Gamma}$ of $\mathcal{V}$ is fixed point free. Let $P$ be some doubleton peripheral as in (2) above, and suppose that an edge $e$ of $\Gamma$ has appearances in two distinct lines $L_{1},L_{2}\in\mathcal{L}_{P}$. Fix some order of $P$. The argument similar as in the proof of Lemma 7.15 shows that the order-induced (from the above fixed order of $P$) orientations of $e$ determined by the two above mentioned appearances of $e$ in $L_{1}$ and $L_{2}$ coincide. It follows that any fixed order of the doubleton $P$ uniquely induces orientations on all edges from the set $E_{\Gamma}^{P}$. We again call the so described orientations the order-induced orientations (with respect to the above fixed order of the peripheral $P$).
Observation 7.17.
Note that in any standard connecting $V$-system ${\mathcal{V}}_{\Xi_{k}}$ of $\theta$-graphs (as described in Example 7.11), each line (and consequently each arm in the corresponding $V$-tree ${\mathcal{X}}_{V}({\mathcal{V}}_{\Xi_{k}})$) is orientable. Note that, fixing an order for the common set of endpoints of the arms of ${\mathcal{X}}_{V}({\mathcal{V}}_{\Xi_{k}})$, we get order-induced orientations of the edges of the $\theta$-graph $\Xi_{k}$ which are consistent (in the sense that they induce the same order on the set of two vertices of $\Xi_{k}$).
More complicated examples of connecting $V$-systems and $V$-trees will appear in Section 8.3, in connection with abstract virtually free factors and their Whitehead graphs.
8. Abstract factors
In this section we prepare some further tools and terminology necessary for our proof of the implication (3$\Rightarrow$(2) of Theorem A (and the implication (3$\Rightarrow$(2) of Theorem B). In particular, we discuss the so called Whitehead graphs associated to virtually free factors of the considered groups $G$. Such graphs (or graphs closely related to them) will appear in our expressions of the Gromov boundaries $\partial G$ as regular trees of graphs.
8.1. Abstract factors and their decorations
Definition 8.1 (abstract factors).
An abstract factor is a pair $(K,\{H_{i}\}_{i\in I})$, where $K$ is a nonelementary hyperbolic group and $\{H_{i}\}_{i\in I}$ is a finite collection of maximal 2-ended subgroups of $K$ which are pairwise non-conjugate. Actually, we will view $\{H_{i}\}_{i\in I}$ rather as a family of conjugacy classes of maximal 2-ended subgroups, and we will not distinguish abstract factors which have the same $K$, and for which the conjugacy classes of the subgroups coincide.
Example 8.2.
Given a 1-ended hyperbolic group $G$, any black vertex group $G(v)$ of the Bowditch JSJ splitting of $G$ (as well as of the reduced Bowditch JSJ splitting of $G$), equipped with the family ${\mathcal{H}}_{v}$ of conjugacy classes of its edge subgroups $G([v,u])=G(v)\cap G(u)$ with $u\in N_{v}$, is an abstract factor. In particular, we may apply all of the discussion below to such factors of the corresponding splittings.
It is obviously true that the conjugate subgroup $gH_{i}g^{-1}<K$ depends only on the coset in $K/H_{i}$ to which the conjugating element $g$ belongs. Moreover, since $H_{i}$ is maximal 2-ended, it coincides with the stabilizer of $\partial H_{i}$ in $\partial G$ (the stabilizer is a 2-ended subgroup containing $H_{i}$) which implies that $H_{i}$ is self-normalizing. Thus the cosets and conjugates of $H_{i}$ are in bijective correspondence and we will freely refer to a conjugate $gH_{i}g^{-1}$ by way of its corresponding coset $gH_{i}\in K/H_{i}$. Since it is well known that for distinct maximal 2-ended subgroups $H,H^{\prime}$ of a hyperbolic group $K$ their boundaries $\partial H,\partial H^{\prime}$ viewed as subsets in the Gromov boundary $\partial K$ are disjoint (see e.g. the sentence right before Lemma 1.1 in [Bow98]), it follows that the family of boundaries of the conjugates
(1)1( 1 ) $\partial(gH_{i}g^{-1}):i\in I,g\in K/H_{i}$
consists of pairwise disjoint subsets of $\partial K$. It is also not hard to observe that this family of subsets of $\partial K$ is null. To see it, note that any conjugate $gH_{i}g^{-1}$ lies at finite Hausdorff distance (for any word metric $d_{K}$ in $K$) from the coset $gH_{i}$, and thus we have $\partial(gH_{i}g^{-1})=\partial(gH_{i})$. Moreover, for any fixed $i\in I$ and for the identity element $e\in K$ the distances
$d_{K}(e,gH_{i}):gH_{i}\in K/H_{i}$
clearly diverge to infinity. Since the cosets $gH_{i}$ of $K/H_{i}$ are uniformly quasiconvex in $K$ (in particular, they lie at uniformly finite distance from geodesics in $K$ connecting their boundary points), it follows that the diameters $\hbox{diam}(\partial(gH_{i})):gH_{i}\in K/H_{i}$ converge to 00, which obviously implies that the family (1) is null. This allows us to introduce the following concept.
Definition 8.3 (space with peripherals induced by an abstract factor).
The space with peripherals associated to an abstract factor $(K,\{H_{i}\}_{i\in I})$ is the pair
$(\partial K,\{\partial(gH_{i}g^{-1}):i\in I,gH_{i}\in K/H_{i}\}).$
Peripherals of the form $\partial(gH_{i}g^{-1})$ are called $H_{i}$-peripherals, or peripherals of type $H_{i}$.
We now pass to discussing orientability phenomona for peripheral subgroups of abstract factors.
Definition 8.4.
A 2-ended group $H$ is orientable if each element $h\in H$ fixes $\partial H$ pointwise. Otherwise $H$ is non-orientable. If $H$ is orientable, an orientation for $H$ is a choice of any order on the set $\partial H$.
Note that $H$ as above is non-orientable iff for some $h\in H$ the induced homeomorphism of $\partial K$ transposes the points of the doubleton $\partial H$. Assuming now that $H$ is a maximal 2-ended subgroup of a hyperbolic group $K$, if $H$ is orientable then any conjugate of $H$ in $K$ is also orientable, and a similar property holds for non-orientability. Moreover, if $H$ is orientable, any fixed orientation for $H$ induces orientations for all conjugates $gHg^{-1}$. In particular, an orientation for $H$ induces also orders on the boundary sets $\partial(gHg^{-1})$ (which are doubletons). (Here we use the fact that the stabilizer of the doubleton $\partial H\subset\partial K$ coincides with $H$.) As a matter of facts, $H$ is non-orientable iff for some $h\in H$ the left multiplication by $h$ reverses the ends of $H$, so that non-orientability (and orientability) is in fact an intrinsic property of $H$. However, for the purposes of this paper, the corresponding earlier metioned “extrinsic” manifestation of this property is more important.
Coming back to the space with peripherals associated to an abstract factor $(K,\{H_{i}\})$, we say that an $H_{i}$-peripheral is orientable if the corresponding subgroup $H_{i}$ is orientable; otherwise an $H_{i}$-peripheral is non-orientable. An orientation for an orientable peripheral is a choice of an order, and obviously it can be compatible or incompatible with a fixed orientation of the corresponding 2-ended subgroup $H_{i}$.
Definition 8.5 (decoration of an abstract factor).
Orientation data for an abstract factor $(K,\{H_{i}\}_{i\in I})$ is information about orientability and non-orientability of the subgroups $H_{i}$ (and thus of all subgroups from their conjugacy classes) and, for each orientable subgroup $H_{i}$, a choice of an orientation. Such orientation data induces a decoration of the associated space with peripherals, which consists of information about the types ($H_{i}$-types) of the peripherals, and about their orientability, and if a peripheral is orientable, this decoration indicates also one of the orientations of this peripheral, namely this one which is compatible with the fixed orientation of the corresponding 2-ended subgroup.
We make a record of the following obvious observations, which follow from the well known properties of the action of a hyperbolic group on its boundary.
Lemma 8.6.
Let $(K,\{H_{i}\}_{i\in I})$ be an abstract factor.
  1. (1)
    Each homeomorphism of the action of $K$ on $\partial K$ preserves any decoration of the associated space with peripherals. More precisely, this means that any such homeomorphism preserves the types of the peripherals, it maps non-orientable peripherals to non-orientable ones, and it maps orientable peripherals to orientable ones, and in the latter case, it preserves the orientations indicated by the decoration.
  2. (2)
    For any given type $H_{i}$, $K$ acts transitively on the peripherals of this type.
  3. (3)
    If a peripheral $P$ in the associated space with peripherals is non-orientable, then there is a homeomorphism of this space with peripherals which preserves types $H_{i}$ of all peripherals, preserves any decoration, and maps $P$ to itself via the transposition.
8.2. Flexible factors
Definition 8.7 (abstract flexible factor).
An abstract factor $(K,\{H_{i}\}_{i\in I})$ is called an abstract flexible factor if the associated space with peripherals is homeomorphic (as a space with peripherals) to the punctured circle $(S^{1})^{\circ}$.
Example 8.8.
Let $G(v)$ be a flexible factor of the Bowditch JSJ-splitting of a 1-ended hyperbolic group $G$. Then $(G(v),{\mathcal{H}}_{v})$, i.e., the abstract factor associated to $G(v)$ as in Example 8.2, is an abstract flexible factor, in the sense of Definition 8.7. For justification, see Section 5 in [Bow98], where flexible factors are called maximal hanging Fuchsians (MHF), or the paragraph after Definition 5.9 in [KW22].
Similarly, let $G(v^{\prime})$ be a flexible factor of the reduced Bowditch splitting of a 1-ended hyperbolic group $G$. In view of Fact 3.4, the abstract factor $(G(v^{\prime}),{\mathcal{H}}_{v^{\prime}})$ associated to $G(v^{\prime})$ is also a flexible abstract factor.
Remark 8.9.
  1. (1)
    Since, as a topological space, $(S^{1})^{\circ}$ is a Cantor set, it follows that the underlying group $K$ of an abstract flexible factor is a virtually free group.
  2. (2)
    Recall that the natural embedding of the punctured circle $(S^{1})^{\circ}$ in the circle $S^{1}$ is uniquely described, up to a homeomorphism of the ambient $S^{1}$, by the following condition: for any peripheral doubleton $P$ of $(S^{1})^{\circ}$ there is a segment $I_{P}$ contained in $S^{1}$ whose set of endpoints coincides with $P$ and whose interior is disjoint with $(S^{1})^{\circ}$. We will call this embedding the natural extension of $(S^{1})^{\circ}$ to a circle. Each segment $I_{P}$ as above will be called the segment induced by $P$. Note that any homeomorphism $h$ of $(S^{1})^{\circ}$ preserving the family of peripherals can be extended in an essentially unique way to a homeomorphism $\bar{h}:S^{1}\to S^{1}$ of the natural extension. Thus, it makes sense to speak of orientation-preserving and orientation-reversing homeomorphisms $h$, by referring to the corresponding property of the induced $\bar{h}$.
Remark 8.9(2) leads to the following concept.
Definition 8.10 (orientability of an abstract flexible factor).
An abstract flexible factor $(K,\{H_{i}\}_{i\in I})$ is orientable if for each $g\in K$ the induced homeomorphism of $\partial K=(S^{1})^{\circ}$ is orientation-preserving (in the sense explained in Remark 8.9(2)). An orientation of such a factor is a choice of orientation for the corresponding extended circle $S^{1}$.
Note that an abstract flexible factor $(K,\{H_{i}\}_{i\in I})$ is non-orientable if for some $g\in K$ the induced homeomorphism of $\partial K=(S^{1})^{\circ}$ is orientation-reversing. The next lemma relates orientability properties of an abstract flexible factor $(K,\{H_{i}\}_{i\in I})$ with orientability properties of its peripheral subgroups $H_{i}$. We omit the straightforward proof.
Lemma 8.11.
Let ${\mathcal{F}}=(K,\{H_{i}\}_{i\in I})$ be an abstract flexible factor.
  1. (1)
    If some peripheral subgroup $H_{i}$ is non-orientable (in the sense of Definition 8.4) then $\mathcal{F}$ is also non-orientable.
  2. (2)
    If $\mathcal{F}$ is orientable then all of its peripheral subgroups are also orientable.
The reader should keep in mind that, given an abstract flexible factor $\mathcal{F}$, orientability of all its peripheral subgroups does not necessarily lead to orientability of the factor. For example, if $K$ is the fundamental group of a compact connected hyperbolic non-orientable surface with non-empty boundary, and if we take as its peripherals the subgroups corresponding to boundary components, then the obtained flexible factor is non-orientable, though all its peripheral subgroups are cyclic, and hence orientable.
Definition 8.12 (compatibility of orientations of a flexible factor and of its peripheral subgroup).
Given an orientable abstract flexible factor ${\mathcal{F}}=(K,\{H_{i}\}_{i\in I})$, with a fixed orientation, and an orientable peripheral subgroup $H\in\{H_{i}\}_{i\in I}$, also with a fixed orientation (i.e. an order of the doubleton $P=\partial H$), we say that these orientations are compatible if the orientation induced from $P$ on the induced segment $I_{P}$ in the extended circle $S^{1}$ coincides with the orientation of this segment induced from the orientation of the whole $S^{1}$ which represents the fixed orientation of $\mathcal{F}$.
We make a record of the two observations below, which follow easily from the fact that each orbit of a hyperbolic group in the action on its Gromov boundary is dense.
Fact 8.13.
Let ${\mathcal{F}}=(K,\{H_{i}\}_{i\in I})$ be an abstract flexible factor.
  1. (1)
    Suppose that $\mathcal{F}$ is orientable, with one of the orientations fixed. If we also fix one of the orientations of a peripheral subgroup $H_{i}$, then either the induced orientations of all $H_{i}$-peripherals are compatible with the orientation of $\mathcal{F}$, or they are all incompatible. In any case, the union of all $H_{i}$-peripherals is dense in $\partial K=(S^{1})^{\circ}$.
  2. (2)
    Suppose that $\mathcal{F}$ is non-orientable. Consider the extended circle $S^{1}$ for $\partial K=(S^{1})^{\circ}$, and fix one of its orientations. If a subgroup $H_{i}$ is orientable, and if one of its orientations is fixed, consider the induced orientations for all conjugates $gH_{i}g^{-1}$, and the corresponding orientations of all $H_{i}$-peripherals. Then the family of those $H_{i}$-peripherals whose just described orientations are compatible (respectively, incompatible) with the above fixed orientation of $S^{1}$ is non-empty, and the union of this family is dense in $\partial K=(S^{1})^{\circ}$.
8.3. Abstract virtually free factors and their Whitehead graphs
In this subsection we first recall the concept of a Whitehead graph, as applied to abstract factors that are free. We then extend all of our discussion to abstract virtually free factors. The main result in this subsection, Corollary 8.25, identifies the space with peripherals associated to an abstract virtually free factor with the core of the associated V-tree of Whitehead graphs. This result will play a crucial role in our proof of the implication (3$\Rightarrow$(2) of Theorem A, given in Section 13, and it will be applied to virtually free rigid cluster factors of the reduced Bowditch JSJ splitting of a group $G$.
Definition 8.14.
An abstract free factor is an abstract factor $(F,\{H_{i}\})$ in which $F$ is a finitely generated non-abelian free group.
Remark 8.15.
Note that, since $F$ is torsion free, each of the subgroups $H_{i}$ is a maximal infinite cyclic subgroup of $F$.
To each abstract free factor $(F,{\mathcal{H}})$, equipped with a basis $B$ of free generators for $F$, there is associated some finite graph, called the Whitehead graph of the factor, see Section 2E in [Cas16]. We now recall the description of this graph.
Definition 8.16 (Whitehead graph of an abstract free factor).
Given an abstract free factor ${\mathcal{Q}}=(F,\{H_{i}\})_{1\leq i\leq k}$, let $B=(x_{1},\dots,x_{m})$ be some base set of free generators for $F$. For each $1\leq i\leq k$ fix a generator $h_{i}$ of the cyclic subgroup $H_{i}$, and denote by $w_{i}$ the unique cyclically reduced word over $B\cup B^{-1}$ in the conjugacy class of $h_{i}$. (Here we view $w_{i}$ as a cyclic word, with letters occuring in cyclic order rather than in linear one.) The Whitehead graph for $\mathcal{Q}$ with respect to $B$ is the graph $W=W_{\mathcal{Q}}$ with the vertex set $V_{W_{\mathcal{Q}}}=\{x_{1},\dots,x_{m},x_{1}^{-1},\dots,x_{m}^{-1}\}$ and with one edge connecting distinct vertices $x,y\in V_{W_{\mathcal{Q}}}$ for every occurence of the subword $x^{-1}y$ in the cyclic words $w_{1},\dots,w_{k}$. (We use the convention that if $w_{i}=x$ has length 1 then there is precisely one occurence of $xx$ in $w_{i}$, and no occurence of any other word of length 2; moreover, if $w_{i}=uz$ has length 2, then $w_{i}$ has two subwords of length 2, namely $uz$ and $zu$.)
Remark 8.17.
  1. (1)
    Observe that, since the subgroups $H_{i}<F$ are maximal cyclic, their generators are not the powers in $F$, and hence the cyclic words $w_{i}$ are not periodic. For this reason, any two occurences of the same subword in any fixed $w_{i}$ are essentially distinct. This dispels some potential doubts concerning the above definition.
  2. (2)
    Note that the Whitehead graph $W_{\mathcal{Q}}$ is obviously a graph with no loop edges, but it may contain multiple edges. $W_{\mathcal{Q}}$ depends not only on the factor $\mathcal{Q}$, but also on the choice of a basis $B$ in the free group $F$. The choices for which the number of edges in corresponding Whitehead graphs is minimal lead to so called minimal Whitehead graphs for the factor $\mathcal{Q}$. As we will show later (see Proposition 9.3), in the situations considered in this paper the minimal Whitehead graphs will always be 2-connected (i.e. finite, connected, nonempty, not equal to a single vertex and cutpoint-free), so in particular they will be essential (i.e. nonempty and with no isolated vertices). Since all our further observations apply to any choice of a minimal Whitehead graph for a factor $\mathcal{Q}$, we will not bother about this choice, and we will speak of the Whitehead graph associated to $\mathcal{Q}$, meaning any of the minimal Whitehead graphs as above.
Any Whitehead graph $W_{\mathcal{Q}}$ which is essential is canonically equipped with some $V$-involution ${\mathbf{a}}_{\mathcal{Q}}$, so that it yields an associated connecting $V$-system ${\mathcal{V}}_{\mathcal{Q}}=(W_{\mathcal{Q}},{\mathbf{a}}_{\mathcal{Q}})$. This $V$-involution consists of the involution $a_{\mathcal{Q}}:V_{W_{\mathcal{Q}}}\to V_{W_{\mathcal{Q}}}$ given by $a_{\mathcal{Q}}(x)=x^{-1}$ for each $x\in V_{W_{\mathcal{Q}}}$ (where we use the convention that $(x_{i}^{-1})^{-1}=x_{i}$), and of the maps $\alpha_{x}:\hbox{Lk}(x)\to\hbox{Lk}(x^{-1})$ described as follows. For any $y\in V_{W_{\mathcal{Q}}}$ and any occurence of $y$ in words $w_{1},\dots,w_{k}$ consider the subwords of the form $xy$ and $yz$, one for each of the two forms, which extend this occurence of $y$, and denote by $\sigma,\tau$ the corresponding occurences of these subwords, respectively. Denote by $e_{\sigma}$ and $e_{\tau}$ the edges of $W$ corresponding to the above occurences of $xy$ and $yz$. Denote also by $[e_{\sigma}]_{y}$ and $[e_{\tau}]_{y^{-1}}$ the points in the links $\hbox{Lk}(y)$ and $\hbox{Lk}(y^{-1})$ induced in those links by the edges $e_{\sigma}$ and $e_{\tau}$, respectively. Finally, put $\alpha_{y}([e_{\sigma}]_{y})=[e_{\tau}]_{y^{-1}}$ and $\alpha_{y^{-1}}([e_{\tau}]_{y^{-1}})=[e_{\sigma}]_{y}$; note that the collection of all assignments as above fully describes all the maps $\alpha_{y}:y\in V_{W_{\mathcal{Q}}}$, and that the conditions required in Definition 7.1 are satisfied. We skip straightforward details of the justification.
The next observation follows directly from Lemma 7.15, in view of the fact that the involution $a_{\mathcal{Q}}$, as given above, is fixed point free.
Lemma 8.18.
For any essential Whitehead graph $W_{\mathcal{Q}}$ (associated to an abstract free factor $\mathcal{Q}$), each line in the corresponding V-tree system $\Theta[{\mathcal{V}}_{\mathcal{Q}}]$ is orientable.
We now show that, for any abstract free factor ${\mathcal{Q}}=(F,\{H_{i}\}_{i\in I})$, the core $C({\mathcal{V}}_{\mathcal{Q}})$ of the $V$-tree ${\mathcal{X}}_{V}({\mathcal{V}}_{\mathcal{Q}})$ naturally carries an action of the group $F$. To do this, we will first identify naturally the underlying tree $T$ of the $V$-tree system $\Theta[{\mathcal{V}}_{\mathcal{Q}}]$ with the first barycentric subdivision of the Cayley graph ${\rm Cay}(F,B)$ of the group $F$ with respect to its basis $B=\{x_{1},\dots,x_{m}\}$.
Recall that the tree $T$ is bipartite, with all white vertices of degree 2. Thus $T$ is the first barycentric subdivision of a certain tree, which we denote $T_{*}$. The vertices of $T_{*}$ are naturally identified with the black vertices of $T$. Our goal is to identify $T_{*}$ with ${\rm Cay}(F,B)$. To do this, we associate to each edge $e$ of $T_{*}$ some orientation and some label from the generating set $B$. Let $t,s$ be the endpoints of $e$, and view them as black vertices of $T$. Let $u$ be the barycenter of $e$, which we view as an appropriate white vertex of $T$ (the unique white vertex of $T$ adjacent to both $t$ and $s$). The constituent spaces at $t$ and $s$ of the tree system $\Theta[{\mathcal{V}}_{\mathcal{Q}}]$, denoted $\Gamma_{t}^{\circ}$ and $\Gamma_{s}^{\circ}$, are the copies of the punctured Whitehead graph $W_{\mathcal{Q}}$. The images in $\Gamma_{t}^{\circ}$ and in $\Gamma_{s}^{\circ}$ of the peripheral space $\Sigma_{u}$ are some $V$-peripherals related to vertices of $\Gamma_{t}$ and $\Gamma_{s}$, and we denote these vertices by $z$ and $y$, respectively. Identifying both $z$ and $y$ as vertices of the Whitehead graph $W_{\mathcal{Q}}$, we have $z,y\in B\cup B^{-1}$, and by the form of the associated $V$-involution $a_{\mathcal{Q}}$, we have $y=z^{-1}$. Without loss of generality, assume that $z\in B$. Orient then the edge $e$ consistently with the order $(z,y)$, and label it with $z$. It is then easy to observe that for any vertex $v$ of $T_{*}$ we have exactly $m$ oriented edges issuing from $v$, labelled bijectively by the elements from $B$, and exactly $m$ oriented edges incoming to $v$, also labelled bijectively with the elements of $B$. Thus the tree $T_{*}$ is locally isomorphic to the Cayley graph ${\rm Cay}(F,B)$, which is also a tree. To get an explicit identification of $T_{*}$ with ${\rm Cay}(F,B)$, fix a vertex of $T_{*}$ and identify it with the vertex of ${\rm Cay}(F,B)$ corresponding to the unit of $F$. This identification obviously extends to a unique label preserving isomorphism of oriented graphs, and we take this isomorphism as the desired identification.
Given the identification of $T_{*}$ with ${\rm Cay}(F,B)$ as above, the natural action of $F$ on ${\rm Cay}(F,B)$ induces an action of $F$ on $T_{*}$. This clearly yields also an action of $F$, by automorphisms, on $T$. Since the core $C({\mathcal{V}}_{\mathcal{Q}})$ coincides with the boundary $\partial T$, we equip it with the induced action of $F$ on $\partial T$. Observe that, being orientation and label preserving, the above described action of $F$ on $T$ preserves also the lines of the $V$-tree system $\Theta[{\mathcal{V}}_{\mathcal{Q}}]$, as explained in some detail in the proof of Lemma 8.19. As a consequence, the induced action of $F$ on the core ${\mathcal{C}}({\mathcal{V}}_{\mathcal{Q}})=\partial T$ preserves the peripherals in this core.
Note that, since $F$ is a free group, its Gromov boundary $\partial F$ is the Cantor set. Thus, the space with peripherals associated to an abstract free factor is a Cantor set equipped with some family of doubleton subsets. The next lemma exhibits an important relationship between this space with peripherals and the $V$-tree of Whitehead graphs induced by the corresponding factor.
Lemma 8.19.
Let ${\mathcal{Q}}$ be an abstract free factor equipped with an essential Whitehead graph $W_{\mathcal{Q}}$, and let ${\mathcal{V}}_{\mathcal{Q}}=(W_{\mathcal{Q}},{\mathbf{a}}_{\mathcal{Q}})$ be the connecting $V$-system associated to it (as described above). Then the space with peripherals associated to $\mathcal{Q}$ is homeomorphic (as a space with peripherals) to the core of the $V$-tree of graphs ${\mathcal{X}}_{V}({\mathcal{V}}_{\mathcal{Q}})$, through an $F$-equivariant homeomorphism (i.e. a homeomorphism which respects the $F$-actions on the two spaces).
Proof.
Recall that the space with peripherals associated to ${\mathcal{Q}}=(F,\{H_{i}\}_{i\in I})$ is the pair
$(\partial F,\{\partial(gH_{i}g^{-1}):i\in I,g\in F/H_{i}\}),$
and that the peripherals of the core $C({\mathcal{V}}_{\mathcal{Q}})$ have the form $\partial\gamma_{L}$, where $L$ runs through the set of all lines of the system $\Theta[{\mathcal{V}}_{\mathcal{Q}}]$. As a required homeomorphism between the core $C({\mathcal{V}}_{\mathcal{Q}})=\partial T$ and the associated space with peripherals $\partial F=\partial[{\rm Cay}(F,B)]$ take the homeomorphism induced on the boundaries by the indentification of $T$ with the first barycentric subdivision of ${\rm Cay}(F,B)$, as described right above the statement of the lemma. Denote this homeomorphism by $h$. It is obvious that $h$ is $F$-equivariant, by the definition of the action of $F$ on $\partial T$. We need to show that $h$ preserves the families of the peripherals.
Let $L$ be a line of the $V$-tree system $\Theta[{\mathcal{V}}_{\mathcal{Q}}]$. Observe that any two consecutive edges of $L$ correspond then to some consecutive occurences of subwords of form $xy$ and $yz$ in precisely one of the cyclically reduced words $w_{i}$. We will emphasize this fact by saying that the line $L$ corresponds to this generator $w_{i}$ (of the corresponding peripheral subgroups $H_{i}$). Let $P=\partial\gamma_{L}$ be the corresponding peripheral of the core $C({\mathcal{V}}_{\mathcal{Q}})$. Observe that the geodesic $\gamma_{L}$ coincides then, via the identification of $T_{*}$ with ${\rm Cay}(F,B)$, with the axis of the automorphism of ${\rm Cay}(F,B)$ given by left multiplication through some conjugate $gw_{i}g^{-1}$. Since this axis stays at finite Hausdorff distance in ${\rm Cay}(F,B)$ from the subgroup $gH_{i}g^{-1}<F$, it has the same boundary at infinity as this subgroup, and thus we get
$h(\partial\gamma_{L})=\partial(gH_{i}g^{-1}).$
So $h$ maps peripherals to peripherals. A similar argument shows that $h^{-1}$ also maps peripherals to peripherals, and thus $h$ is a homeomorphism of spaces with peripherals, which completes the proof. ∎
Remark 8.20.
Observe that the choices of the generators $h_{i}$ in the subgroups $H_{i}<F$ determine (uniquely) the choices of the generators $gh_{i}g^{-1}$ in the conjugate subgroups $gH_{i}g^{-1}<F$, and hence also the orders on all peripherals in the space of peripherals $\partial F$ associated to $\mathcal{Q}$. On the other hand, each $w_{i}$ induces orientations on all edges of any line $L$ in $\Theta[{\mathcal{V}}_{\mathcal{Q}}]$ corresponding to $w_{i}$. It follows that the choices of $w_{i}$ induce orientations on all geodesics $\gamma_{L}$, and thus they induce orders on all peripherals in the core $C({\mathcal{V}}_{\mathcal{Q}})$. It is not hard to see that the homeomorphism $h$ from the above proof of Lemma 8.19 preserves the above mentioned orders of the peripherals.
We now pass to virtually free abstract factors.
Definition 8.21.
An abstract virtually free factor is an abstract factor $\hat{\mathcal{Q}}=(\hat{F},\{\hat{H}_{i}\})$ in which $\hat{F}$ is a finitely generated infinitely ended virtually free group.
Consider the associated space with peripherals for $\hat{\mathcal{Q}}$
$(\partial\hat{F},\{\partial\hat{H}_{i}^{g}:1\leq i\leq k,\,g\in\hat{F}/\hat{H}_{i}\})$
(here we again use the convention of conjugating through a coset $g\in\hat{F}/\hat{H}_{i}$, since the result of the conjugation does not depend on a choice of a coset representative). Our goal is to give an appropriate extension of the concept of Whitehead graph associated to a factor, which will apply to virtually free factors as well. To do this, we simply associate to $\hat{\mathcal{Q}}$ some appropriate free factor structure for a finite index free subgroup of $\hat{F}$. The details are provided below, with culmination in Corollary 8.25.
Let $F$ be a free finite index subgroup of $\hat{F}$. For each $i\in\{1,\dots,k\}$ and any $\hat{g}\in\hat{F}$ consider the subgroup $\hat{H}_{i}^{\hat{g}}\cap F<F$, which is torsion-free and of finite index in $\hat{H}_{i}^{\hat{g}}$, and which is hence infinite cyclic. Since we clearly have $\hbox{Stab}_{F}\partial\hat{H}_{i}^{\hat{g}}=\hat{H}_{i}^{\hat{g}}\cap F$, this subgroup is a maximal 2-ended subgroup of $F$. Note that $F$ acts on the set $\{\hat{H}_{i}^{\hat{g}}\cap F:\hat{g}\in\hat{F}/\hat{H}_{i}\}$ of such subgroups by conjugations, as in
$f(\hat{H}_{i}^{\hat{g}}\cap F)f^{-1}=\hat{H}_{i}^{f\cdot\hat{g}}\cap F.$
Since $F$ has finite index in $\hat{F}$, the above action by conjugations has finitely many orbits. Denote by $g_{i,j}:1\leq j\leq k_{i}$ some elements of $\hat{F}$ such that
$\hat{H}_{i}^{g_{i,j}}\cap F:1\leq j\leq k_{i}$
is a family of representatives of all orbits of the above action of $F$ by conjugation. Since the stabilizer of each such representative satisfies
$\hbox{Stab}_{F}(\hat{H}_{i}^{g_{i,j}}\cap F)=\hat{H}_{i}^{g_{i,j}}\cap F,$
its orbit can be described as the set
$\{\hat{H}_{i}^{f\cdot g_{i,j}}\cap F:f\in F/(\hat{H}_{i}^{g_{i,j}}\cap F)\}.$
Put
[…]
Then $(F,{\mathcal{P}})$ is obviously an abstract free factor. Since we clearly have $\partial F=\partial\hat{F}$ and $\partial(\hat{H}_{i}^{g_{i,j}}\cap F)=\partial\hat{H}_{i}^{g_{i,j}}$, we get the following.
Lemma 8.22.
The spaces with peripherals
$(\partial\hat{F},\{\partial\hat{H}_{i}^{g}:1\leq i\leq k,\,g\in\hat{F}/\hat{H}_{i}\})\hbox{\quad and \quad}(\partial F,\{\partial H^{f}:H\in{\mathcal{P}},\,f\in F/H\})$
are canonically homeomorphic.
We now extend the notion of a Whitehead graph to virtually free abstract factors.
Definition 8.23.
Let $\hat{\mathcal{Q}}=(\hat{F},\{\hat{H}_{i}\})$ be a virtually free abstract factor. The Whitehead graph $W=W_{\hat{\mathcal{Q}}}$ of this factor is any graph of the following form. Choose any free subgroup $F<\hat{F}$ of finite index, and a basis $B$ for $F$, and consider the induced peripheral system $\mathcal{P}$ of subgroups of $F$, as described above. Take as $W_{\hat{\mathcal{Q}}}$ the Whitehead graph for the free factor ${\mathcal{Q}}=(F,{\mathcal{P}})$ with respect to $B$. Take also as the accompanying connecting $V$-system ${\mathcal{V}}_{\hat{\mathcal{Q}}}$ the $V$-system ${\mathcal{V}}_{\mathcal{Q}}$ for $W_{\mathcal{Q}}$.
Remark 8.24.
Similarly as in the case of a free abstract factor, the Whitehead graph $W_{\hat{\mathcal{Q}}}$ for $\hat{\mathcal{Q}}$ is not uniquely determined, and depends on the choices of $F$ and $B$. In our later applications of this concept, we will deal with any choice of a subgroup $F$, and such a choice of a basis $B$ for $F$ that the corresponding Whitehead graph for $(F,{\mathcal{P}})$ is minimal. As we will show (see Propositions 9.2 and 9.3), under such choices the Whitehead graphs appearing in our context will always be 2-connected, and in particular essential. Since all further statements concerning the graph $W_{\hat{\mathcal{Q}}}$ apply equally well to all choices as above, we will speak of the Whitehead graph associated to the factor $\hat{\mathcal{Q}}$, meaning any of the graphs as described above.
As a consequence of the above definition, and in view of Lemma 8.22, we have the following corollary to Lemma 8.19.
Corollary 8.25.
Let $\hat{\mathcal{Q}}$ be a virtually free abstract factor, and suppose that the associated Whitehead graph $W_{\hat{\mathcal{Q}}}$ is essential. Let ${\mathcal{V}}_{\hat{\mathcal{Q}}}=(W_{\hat{\mathcal{Q}}},{\mathbf{a}})$ be the associated connecting $V$-system. Then the associated space with peripherals for $\hat{\mathcal{Q}}$ is homeomorphic (as a space with peripherals) to the core ${\mathcal{C}}({\mathcal{V}}_{\hat{\mathcal{Q}}})$ of the corresponding $V$-tree of graphs ${\mathcal{X}}_{V}({\mathcal{V}}_{\hat{\mathcal{Q}}})$.
We now discuss some further features of Whitehead graphs associated to virtually free abstract factors. We refer to the notation as above in this section, additionally denoting by $\mathcal{H}$ the set of conjugacy classes of the peripheral subgroups in an abstract virtually free factor $\hat{\mathcal{Q}}=(\hat{F},\{\hat{H}_{i}\})$. We first describe some natural labelling $\lambda:E_{W_{\hat{\mathcal{Q}}}}\to{\mathcal{H}}$ of the edges of the Whitehead graph $W_{\hat{\mathcal{Q}}}$, with values in the set $\mathcal{H}$. Any edge $e$ in $W_{\hat{\mathcal{Q}}}$ corresponds to an occurence of a subword of length 2 in a cyclically reduced word $w$ from the conjugacy class of a generator of precisely one cyclic peripheral subgroup $\hat{H}_{i}^{g_{i,j}}\cap F\in{\mathcal{P}}$. We then put $\lambda(e)=[\hat{H}_{i}]$, where $[\hat{H}_{i}]\in{\mathcal{H}}$ denotes the conjugacy class of the subgroup $\hat{H}_{i}$ in $\hat{F}$. So described $\lambda$ is clearly surjective, and it splits the set $E_{W_{\hat{\mathcal{Q}}}}$ into subsets corresponding to the conjugacy classes in $\mathcal{H}$. Typically, the subset corresponding to a fixed label $[\hat{H}_{i}]$ consists of more than one edge, and actually its cardinality is equal to the sum of the lengths $|w|$ of those cyclically reduced words $w$ which correspond to the generators of the subgroups $\hat{H}_{i}^{g_{i,j}}\cap F:1\leq j\leq k_{i}$. We make an obvious observation that each of the maps $\alpha_{x}$ from the involution system ${\mathbf{a}}=(a,\{\alpha_{x}\})$ associated to $W_{\hat{\mathcal{Q}}}$ preserves the labelling $\lambda$, in the following sense. If $e$ is an edge of $W_{\hat{\mathcal{Q}}}$ adjacent to $x$, and $e^{\prime}$ is an edge adjacent to $a(x)$, then denoting by $e_{x}$ and $e^{\prime}_{a(x)}$ the points in the corresponding vertex links induced by the corresponding edges, the equality $\alpha_{x}(e_{x})=e^{\prime}_{a(x)}$ implies that $\lambda(e)=\lambda(e^{\prime})$.
The following observations concern some further features of the $V$-tree system $\Theta[{\mathcal{V}}_{\hat{\mathcal{Q}}}]$ associated to a virtually free abstract factor $\hat{\mathcal{Q}}$.
Observation 8.26.
  1. (1)
    Since the subgroups $\hat{H}_{i}^{g_{i,j}}\cap F$ are maximal cyclic in $F$, no two lines in $\Theta[{\mathcal{V}}_{\hat{\mathcal{Q}}}]$ have the same ends, i.e., for any peripheral $P$ of the core $C({\mathcal{V}}_{\hat{\mathcal{Q}}})$ the set ${\mathcal{L}}_{P}=\{L:\partial\gamma_{L}=P\}$ consists of a single line.
  2. (2)
    It follows from the above observation (1) that for any peripheral $P$ of the core $C({\mathcal{V}}_{\hat{\mathcal{Q}}})$ there is precisely one arm in the V-tree ${\mathcal{X}}_{V}({\mathcal{V}}_{\hat{\mathcal{Q}}})$ which is attached to $C({\mathcal{V}}_{\hat{\mathcal{Q}}})$ along $P$.
  3. (3)
    Since, due to Lemma 8.18, the line $L\in\mathcal{L}_{P}$, as above in (1), is orientable in $\Theta[{\mathcal{V}}_{\hat{\mathcal{Q}}}]$, any fixed order of the peripheral $P$ induces orientations on all the edges of $W_{\hat{\mathcal{Q}}}$ associated to $P$ (i.e. all edges having appearances in $L$). In accordance with Remark/Definition 7.16(1), we call such orientations order-induced (with respect to the fixed order of $P$).
9. The (extended) Whitehead graph of a virtually free rigid cluster factor of $G$
Let $v$ be a rigid cluster vertex in the tree $T^{r}_{G}$ of the reduced Bowditch JSJ splitting of a hyperbolic group $G$ (see Section 3). We then call the group $G(v)$ a rigid cluster factor of $G$. Denote by $N_{v}$ the set of vertices in $T^{r}_{G}$ adjacent to $v$, and denote by ${\mathcal{H}}_{v}$ the set of the $G(u)$-conjugacy classes of the subgroups $G(u)\cap G(v):u\in N_{v}$. Since the family ${\mathcal{H}}_{v}$ is finite, the pair $(G(v),{\mathcal{H}}_{v})$ is an abstract factor, which we denote briefly by ${\mathcal{Q}}_{v}$. Denote also by $(X_{v},{\mathcal{D}}_{v})$ the space with peripherals associated to ${\mathcal{Q}}_{v}$, where ${\mathcal{D}}_{v}$ is the family of peripheral (doubleton) subsets of $X_{v}=\partial G(v)$.
The next lemma exhibits some consequence of Fact 3.5 (which says that the rigid cluster vertices of ${T}^{r}_{G}$ are isolated) for the structure of the reduced Bowditch JSJ splitting of $G$.
Lemma 9.1.
Let $v$ be a rigid cluster vertex of the tree ${T}^{r}_{G}$. Then for any $u\in N_{v}$ we have $G(u)\cap G(v)=G(u)$. As a consequence, each $G(u)$ as above is a subgroup in $G(v)$, and the family ${\mathcal{H}}_{v}$ actually consists of the conjugacy classes of the subgroups $G(u):u\in N_{v}$.
Proof.
Suppose on the contrary that $G(u)\cap G(v)$ is a proper subgroup of $G(u)$. This implies that the orbit of the vertex $v$ under the action of $G(u)$ on $T^{r}_{G}$ contains at least one vertex $v^{\prime}\neq v$. By $G$-invariance, this vertex $v^{\prime}$ is also a rigid cluster vertex, and since it is adjacent to $u$, we get a contradiction with the fact that $v$ is isolated (i.e. with Fact 3.5). ∎
9.1. The Whitehead graph $W_{v}$ of a virtually free rigid cluster factor $G(v)$
Suppose now that the group $G(v)$ (at a rigid cluster vertex $v$ of $T^{r}_{G}$) is a virtually free group, and recall that it is then necessarily non-elementary (see Corollary 3.7). Obviously, ${\mathcal{Q}}_{v}=(G(v),{\mathcal{H}}_{v})$ is then an abstract virtually free factor. Denote by $W_{v}$ any minimal Whitehead graph for ${\mathcal{Q}}_{v}$, as described in Section 8.3 (see Definition 8.23 and Remark 8.24), for any choice of a finite index free subgroup of $G(v)$.
In the next two propositions we derive basic properties of the graphs $W_{v}$ as above.
Proposition 9.2.
Let ${\mathcal{Q}}=(F,{\mathcal{H}})$ be a free abstract factor, and let $W_{\rm min}$ be a minimal Whitehead graph for $\mathcal{Q}$. Then the following conditions are equivalent:
  1. (1)
    $W_{\rm min}$ is a 2-connected graph (i.e. finite, connected, nonempty, not equal to a single vertex, and cutpoint-free);
  2. (2)
    $F$ has no splitting as a free product relative to $\mathcal{H}$ (i.e. a splitting as a free product with nontrivial factors where the classes in $\mathcal{H}$ are all elliptic).
Proof.
We say that a vertex $x$ of a graph $X$ is a cut vertex if the space $|X|\setminus\{x\}$ has more connected components than $|X|$. It is an easy observation that any minimal Whitehead graph of a free abstract factor has no cut vertex (see e.g. Lemma 2.3 in [Cas17]) so any component is either cutpoint-free or an isolated edge. On the other hand, by Proposition 2.6 from [Cas16], the following three conditions are equivalent:
  1. (1)
    $F$ splits as a free product relative to $\mathcal{H}$,
  2. (2)
    every minimal Whitehead graph for $\mathcal{Q}$ is not connected,
  3. (3)
    some minimal Whitehead graph for $\mathcal{Q}$ is not connected.
The proposition follows directly from these two observations and the fact that $F$ is non-elementary. ∎
Proposition 9.3.
Let $G$ be a 1-ended hyperbolic group which is not cocompact Fuchsian, and let ${\mathcal{Q}}_{v}=(G(v),{\mathcal{H}}_{v})$ be any rigid cluster factor of the Bowditch JSJ splitting for $G$. If the group $G(v)$ is virtually free, then any minimal Whitehead graph for ${\mathcal{Q}}_{v}$ (for any choice of a finite index free subgroup $F<G(v)$) is 2-connected.
Proof.
Suppose on the contrary that some minimal Whitehead graph $W$ for ${\mathcal{Q}}_{v}$ (related to a finite index free subgroup $F$ of $G(v)$, and to the family $\mathcal{P}$ of peripheral subgroups of $F$ induced from ${\mathcal{H}}_{v}$, as described in Section 8.3, right before Lemma 8.22) is not 2-connected. It follows from Proposition 9.2 that $F$ splits as a free product relative to $\mathcal{P}$. Our next goal is to deduce from this fact the following.
Claim 9.4.
$G(v)$ splits relative to ${\mathcal{H}}_{v}$ over a finite subgroup (i.e. acts on a tree without global fixed points such that edge stabilizers are finite and subgroups from ${\mathcal{H}}_{v}$ are conjugate into vertex stabilizers).
To prove the claim, note that since $G(v)$ is hyperbolic and the subgroups from ${\mathcal{H}}_{v}$ are maximal 2-ended and pairwise nonconjugate, $G(v)$ is relatively hyperbolic with respect to ${\mathcal{H}}_{v}$, and the same is true for the pair $(F,{\mathcal{P}})$. We will argue by referring to some results from the theory of relatively hyperbolic groups. In particular, we will refer to the properties of the so called Bowditch boundary $\partial_{B}(G^{\star},{\mathcal{Q}}^{\star})$ of a relatively hyperbolic group $G^{\star}$ (with respect to a family ${\mathcal{Q}}^{\star}=\{Q_{i}\}$ of subgroups). By a result of Tran [Tra13], if $G^{\star}$ is hyperbolic then $\partial_{B}(G^{\star},{\mathcal{Q}}^{\star})$ is homeomorphic to the quotient of the Gromov boundary $\partial G^{\star}$ through a decomposition provided by the family of subspaces
$\partial(gQ_{i}g^{-1}):Q_{i}\in{\mathcal{Q}^{\star}},\,\,g\in G^{\star}.$
Note that, in view of Lemma 8.22, this means that the Bowditch boundaries $\partial_{B}(G(v),{\mathcal{H}}_{v})$ and $\partial_{B}(F,{\mathcal{P}})$ are canonically homeomorphic.
By Proposition 10.1 in [Bow12], the Bowditch boundary $\partial_{B}(G^{\star},{\mathcal{Q}}^{\star})$ is connected iff $G^{\star}$ does not split over a finite subgroup relative to ${\mathcal{Q}}^{\star}$. Since $F$ splits over the trivial group relative to $\mathcal{P}$, it follows that the Bowditch boundary $\partial_{B}(F,{\mathcal{P}})$ is not connected. As a consequence, the Bowditch boundary $\partial_{B}(G(v),{\mathcal{H}}_{v})$ (which is homeomorphic to $\partial_{B}(F,{\mathcal{P}})$) is also not connected. By the same Proposition 10.1 in [Bow12], $G(v)$ splits over a finite subgroup relative to ${\mathcal{H}}_{v}$, which completes the proof of the claim.
Coming back to the proof of Proposition 9.3, it is not hard to observe that if $G$ contains a rigid cluster factor which splits over a finite subgroup relative to its peripheral subgroups, the group $G$ itself also splits nontrivially over a finite subgroup (as an amalgameted free product or an HNN-extension). Due to a well known result of Stallings, the latter implies that $G$ is not 1-ended, contradicting one of our assumptions, and thus the proposition follows. ∎
As we have already mentioned in Remark 8.24, in our further considerations we will use the convention of choosing as “the Whitehead graph $W_{v}$” any minimal Whitehead graph for $\mathcal{Q}_{v}$. In particular, $W_{v}$ will be always 2-connected.
9.2. The extended Whitehead graph $\widetilde{W}_{v}$ of a virtually free rigid cluster factor $G(v)$
Under the notation of Section 9.1, each conjugacy class $\xi\in{\mathcal{H}}_{v}$ is represented by a group $G(u)$ for some $u\in N_{v}$ (see Lemma 9.1). Denote by $n_{\xi}$ the degree of the vertex $u$ in the tree $T^{r}_{G}$. This number $n_{\xi}$ is independent of the choice of $G(u)$ as above. Recall also (e.g. from the description in Sections 3.1 and 3.2) that for each $\xi\in{\mathcal{H}}_{v}$ we have $n_{\xi}\geq 2$.
Let $W_{v}$ be the Whitehead graph for ${\mathcal{Q}}_{v}$, as described in Section 9.1. The extended Whitehead graph $\widetilde{W}_{v}$ for ${\mathcal{Q}}_{v}$ is the graph obtained from the $W_{v}$ as follows. As the vertex set $V_{\widetilde{W}_{v}}$ we take the set $V_{W_{v}}$ of vertices of $W_{v}$. To describe the edge set, consider the labelling $\lambda$ of edges of $W_{v}$, as described in Section 8.3, right after Corollary 8.25, and denote it by $\lambda_{v}$ (in order to distinguish it from the analogous labellings for Whitehead graphs $W_{v^{\prime}}$ related to other rigid cluster vertices $v^{\prime}$ of $T^{r}_{G}$). For any $\xi\in{\mathcal{H}}_{v}$, replace each edge $e$ of $W_{v}$ labelled with $\xi$ (i.e. such that $\lambda_{v}(e)=\xi$) by $n_{\xi}-1$ edges $e_{1},\dots,e_{n_{\xi}-1}$ having the same endpoints as $e$. We will call $e_{1},\dots,e_{n_{\xi}-1}$ as above the edges of $\widetilde{W}_{v}$ corresponding to an edge $e$ of $W_{v}$. (As we will see later, the reason for choosing the number $n_{\xi}-1$ comes from the fact that the vertex $u$ of $T^{r}_{G}$, as above, has precisely $n_{\xi}-1$ flexible neighbours, in addition to the single rigid neighbour $v$.) As part of its structure, equip the so described graph $\widetilde{W}_{v}$ with its own involution system $\tilde{\mathbf{a}}_{v}=(\tilde{a},\{\tilde{\alpha}_{x}\})$ obtained from the corresponding involution system ${\mathbf{a}}_{v}=(a,\{\alpha_{x}\})$ for $W_{v}$ as follows. Identifying naturally the vertices of $\widetilde{W}_{v}$ with the vertices of $W_{v}$, put $\tilde{a}:=a$. Moreover, referring to the notation concerning vertex links as introduced right before Observation 8.26, for any pair of edges $e,e^{\prime}$ of $W_{v}$ such that $\alpha_{x}(e_{x})=e^{\prime}_{a(x)}$, put $\tilde{\alpha}_{x}((e_{i})_{x}):=(e^{\prime}_{i})_{a(x)}$ for $i=1,\dots,n_{\xi}-1$ (where $\xi=\lambda_{v}(e)$, and at the same time $\xi=\lambda_{v}(e^{\prime})$, due to invariance of the labelling under the maps $\alpha_{x}$). Finally, as the last part of the structure of the extended Whitehead graph, we equip it with the labelling $\tilde{\lambda}_{v}$ of the edge set $E_{\widetilde{W}_{v}}$, induced from the labelling $\lambda_{v}:E_{W_{v}}\to{\mathcal{H}}_{v}$ as follows. As the set of values for this new labelling we take the set ~$\widetilde{\mathcal{H}}_{v}:=\{(\xi,j):\xi\in{\mathcal{H}}_{v},1\leq j\leq n_{\xi}-1\}$, and we put $\tilde{\lambda}_{v}(e_{i}):=(\lambda_{v}(e),i)$ for any edge $e$ of $W_{v}$ and any $i=1,\dots,n_{\xi}-1$. Obviously, this induced labelling $\tilde{\lambda}_{v}$ is invariant under the maps $\tilde{\alpha}_{x}$ of the involution system $\tilde{\mathbf{a}}_{v}$, in the same sense as $\lambda_{v}$ is invariant under the maps $\alpha_{x}$. Moreover, again similarly as with $\lambda_{v}$, this induced labelling $\tilde{\lambda}_{v}$ associates the same label to some two edges of $\widetilde{W}_{v}$ if and only if copies of these two edges are involved in a common arm of the $V$-tree ${\mathcal{X}}_{V}((\widetilde{W}_{v},\tilde{\mathbf{a}}_{v}))$. In particular, we may say that $\tilde{\lambda}_{v}$ uniquely associates the labels from ~$\widetilde{\mathcal{H}}_{v}$ to arms in ${\mathcal{X}}_{V}((\widetilde{W}_{v},\tilde{\mathbf{a}}_{v}))$.
The next observation is a fairly direct consequence of Proposition 9.3.
Fact 9.5.
The extended Whitehead graph $\widetilde{W}_{v}$ of any virtually free rigid cluster factor of the group $G$ is 2-connected (in particular, it is essential).
We now pass to basic observations concerning $V$-trees of extended Whitehead graphs.
Lemma 9.6.
Let $v$ be a rigid cluster vertex in the tree $T^{r}_{G}$ of the reduced Bowditch JSJ splitting for $G$, and let ${\mathcal{V}}_{v}=(W_{v},{\mathbf{a}}_{v})$ and $\widetilde{\mathcal{V}}_{v}=(\widetilde{W}_{v},\tilde{\mathbf{a}}_{v})$ be the associated connecting $V$-systems.
  1. (1)
    The core of the $V$-tree of graphs ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{v})$ has a canonical identification (as a space with peripherals) with the core of ${\mathcal{X}}_{V}({\mathcal{V}}_{v})$. As a consequence, the associated space with peripherals for the factor ${\mathcal{Q}}_{v}=(G(v),{\mathcal{H}}_{v})$ is homeomorphic (as a space with peripherals) to the core of the $V$-tree of extended Whitehead graphs ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{v})$.
  2. (2)
    The $V$-tree ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{v})$ can be obtained from the $V$-tree ${\mathcal{X}}_{V}({\mathcal{V}}_{v})$ by leaving the core unchanged and by replacing each arm $A$ in the latter with $n_{\xi}-1$ copies of this arm (sharing only their endpoints, which coincide with the endpoints of $A$), where $\xi=\lambda_{v}(e)$ for any edge $e$ of $W_{v}$ involved in the arm $A$. In particular, the $V$-tree ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{v})$ has a description as a tree of arms and the core, $\Psi[\widetilde{\mathcal{V}}_{v}]$, as presented in Lemma 7.12 and in the paragraphs right before it.
Proof.
Since the vertex sets of the graphs $W_{v}$ and $\widetilde{W}_{v}$ canonically coincide, and since the involutions $a:V_{W_{v}}\to V_{W_{v}}$ and $\tilde{a}_{v}:V_{\widetilde{W}_{v}}\to V_{\widetilde{W}_{v}}$ also coincide, there is a canonical identification of the underlying trees $T$ and $\widetilde{T}$ of the tree systems $\Theta[{\mathcal{V}}_{v}]$ and $\Theta[\widetilde{\mathcal{V}}_{v}]$ associated to the connecting $V$-systems ${\mathcal{V}}_{v}$ and $\widetilde{\mathcal{V}}_{v}$. This identification results from the fact that the vertices in both trees correspond to (punctured) copies of the graphs, and edges correspond to gluings of the $V$-peripherals associated to vertices, according to the schemes provided by the $V$-involutions. Now, let $L=(e_{n}^{\circ})_{n}$ be a line in $\Theta[\widetilde{\mathcal{V}}_{v}]$, and recall that $\xi:=\lambda_{v}(e_{n})$ is then independent of $n$. Under the above identification of $T$ with $\widetilde{T}$, it is not hard to observe that to each line $(e_{n}^{\circ})_{n}$ in $\Theta[{\mathcal{V}}_{v}]$ there correspond $n_{\xi}-1$ lines $((e_{n})_{i}^{\circ})_{n}:i=1,\dots,n_{\xi}-1$ in the tree system $\Theta[\widetilde{\mathcal{V}}_{v}]$, and that there are no other lines in the latter tree system. Moreover, the geodesics in the tree $\widetilde{T}$ corresponding to the lines $((e_{n})_{i}^{\circ})_{n}:i=1,\dots,n_{\xi}-1$ all coincide, and they are equal to the geodesic corresponding to the line $(e_{n}^{\circ})_{n}$ under our identification of $T$ with $\widetilde{T}$. Since the cores under our interest coincide with the Cantor sets appearing as boundaries of the corresponding trees, and since the peripherals in the cores correspond to doubletons given by ends of geodesics induced by lines, part (1) of the lemma clearly follows. Part (2) follows similarly due to the fact that arms in the $V$-trees are induced by lines. ∎
We now pass to discussing some properties of the decomposition of the V-tree ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{v})$ into core and arms. We begin with a natural concept of types of the peripherals in the core $C(\widetilde{\mathcal{V}}_{v})$.
Definition 9.7.
Under notation as in Definition 7.8, let $P=\partial\gamma_{L}$ be a peripheral in the core $C(\widetilde{\mathcal{V}}_{v})$ (which is identified with the core $C(\mathcal{V}_{v})$ as in Lemma 9.6(1)), and let ${\mathcal{H}}_{v}$ be the set of conjugacy classes of the peripheral subgroups of the abstract factor $G(v)$. Recall that ${\mathcal{L}}_{P}$ is the set of lines of $\mathcal{V}_{v}$ whose corresponding arms contain $P$ and that there is necessarily only one such line so that ${\mathcal{L}}_{P}$ is a singleton (see Observation 8.26). The type of $P$ is defined as the element $\xi=\lambda_{v}(e)\in{\mathcal{H}}_{v}$, where $e$ is any edge of $W_{v}$ whose copy appears in the unique line $L\in{\mathcal{L}}_{P}$ (all such edges have the same label, and the line is unique by Observation 8.26(1)).
This assignment of types to the peripherals of the core $C(\widetilde{\mathcal{V}}_{v})$ is natural in the sense provided by the following observation, which can be deduced directly from the definitions and descriptions of the involved objects.
Fact 9.8.
Let $(X_{v},{\mathcal{D}}_{v})$ be the space with peripherals associated to a virtually free rigid cluster factor $(G(v),{\mathcal{H}}_{v})$. Under the identification of $(X_{v},{\mathcal{D}}_{v})$ with the core $C(\widetilde{\mathcal{V}}_{v})$, the types (from ${\mathcal{H}}_{v}$) of the peripherals in ${\mathcal{D}}_{v}$, as described in Definition 8.3, coincide with the types of the corresponding peripherals of $C(\widetilde{\mathcal{V}}_{v})$, as described in Definition 9.7 above.
As a consequence of Observation 8.26(1)(2) and Lemma 9.6(2), we get the following.
Corollary 9.9.
Let $P$ be any peripheral in the core $C(\widetilde{\mathcal{V}}_{v})$, and let $\xi\in{\mathcal{H}}_{v}$ denote its type. Then there is precisely $n_{\xi}-1$ arms in ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{v})$ which are attached to $C(\widetilde{\mathcal{V}}_{v})$ along $P$. Moreover, among these arms we have precisely one arm for each label $(\xi,j):1\leq j\leq n_{\xi}-1$ (for the labelling $\tilde{\lambda}_{v}$).
As a final element in this section, we make a record of a few further rather obvious observations concerning orientation phenomena in the extended Whitehead graphs and in their associated V-tree systems.
Remark 9.10.
  1. (1)
    Note that the edges $e_{1},\dots,e_{n_{\xi}-1}$ which correspond to an edge $e$ of $W_{v}$ in the above description of the extended Whitehead graph $\widetilde{W}_{v}$ have the same endpoints as $e$ (under the canonical identification of the vertex sets of $W_{v}$ and $\widetilde{W}_{v}$). It follows that any orientation for $e$ induces naturally orientations for all $e_{1},\dots,e_{n_{\xi}-1}$.
  2. (2)
    Let $P$ be a peripheral of the core $C(\widetilde{\mathcal{V}}_{v})$. Then it is also a peripheral of the core $C({\mathcal{V}}_{v})$. Recall that $E_{W_{v}}^{P}$ and $E_{\widetilde{W}_{v}}^{P}$ denote then the sets of edges associated to $P$ in the corresponding graphs $W_{v}$ and $\widetilde{W}_{v}$, as described in Remark/Definition 7.16(2). Using the same notation as in the previous remark, we have that $e\in E_{W_{v}}^{P}$ if and only if all edges $e_{1},\dots,e_{n_{\xi}-1}$ belong to $E_{\widetilde{W}_{v}}^{P}$.
  3. (3)
    Under notation as in the above remark (2), if we fix an order of the peripheral $P$, we get order-induced orientations at all edges of $\widetilde{W}_{v}$ associated to $P$ (see Observation 8.26(3)). Moreover, these orientations are compatible (in the sense of remark (1) above) with the order-induced orientations of the edges of $W_{v}$ associated to $P$ (with respect to the same fixed order of $P$).
10. Structure of a regular tree of graphs for $\partial G$ in terms of the reduced Bowditch JSJ splitting of $G$
Let $G$ be a 1-ended hyperbolic group which is not cocompact Fuchsian, and suppose it satisfies condition (3) of Theorem A (i.e. suppose that each rigid cluster factor of the reduced Bowditch JSJ splitting of $G$ is virtually free). In this section we associate to $G$ some graphical connecting system ${\mathcal{R}}_{G}=(\Gamma,{\mathbf{a}},{\mathcal{A}})$ (see Definition 6.1) for which, as we will show in Section 13, the regular tree of graphs ${\mathcal{X}}({\mathcal{R}}_{G})$ is homeomorphic to the Gromov boundary $\partial G$. This connecting system ${\mathcal{R}}_{G}$ will be explicitly determined by various features of the reduced Bowditch JSJ splitting of $G$, and its description constitutes a natural complement to condition (2) of Theorem A, and to condition (2) of Theorem B. We thus rephrase the implication (3$\Rightarrow$(2) in Theorem A as a more precise statement, presented below as Theorem 10.1. (We postpone the proof of Theorem 10.1, and hence of the corresponding implication (3$\Rightarrow$(2) in Theorem A, until Section 13.) We start by describing the graph $\Gamma$ appearing as part of the data in ${\mathcal{R}}_{G}$.
Let $Z$ be a fixed set of representatives of orbits of the action of $G$ on the subset of the vertex set of $T^{r}_{G}$ consisting of the vertices of the following two kinds:
  1. (1)
    rigid cluster vertices;
  2. (2)
    white vertices of type (v2) (as described in Sections 3.1 and 3.2) which are not adjacent to any rigid cluster vertex.
If $z\in Z$ is of type (v2), put $\Gamma_{z}:=\Xi_{k(z)}$ (the $\theta$-graph), where $k(z)$ is the degree of $z$ in $T^{r}_{G}$ (which is easily seen to coincide with the degree of the corresponding vertex in $T_{G}$); note that, by the description of $T_{G}$ and $T^{r}_{G}$ given in Section 3, for each such $z$ we necessarily have $k(z)\geq 3$. If $z\in Z$ is a rigid cluster vertex, put $\Gamma_{z}:=\widetilde{W}_{z}$ (the extended Whitehead graph for $G(z)$, as described in Section 9). Finally, put $\Gamma:=\bigsqcup_{z\in Z}\Gamma_{z}$ (the disjoint union of the graphs $\Gamma_{z}$).
We now describe the $V$-involution $\mathbf{a}$ appearing as part of data in ${\mathcal{R}}_{G}$. Recall that the extended Whitehead graphs $\widetilde{W}_{z}$ come equipped with $V$-involutions ${\mathbf{\tilde{a}}}_{z}=(\tilde{a}_{z},\{(\tilde{\alpha}_{z})_{v}:v\in V_{\Gamma_{z}}\})$. Equip also all the $\theta$-graphs $\Gamma_{z}=\Xi_{k(z)}\subset\Gamma$ with their standard $V$-involutions (with which they form standard connecting $V$-systems of $\theta$-graphs), as described in the last paragraph of Example 7.11, and denote these involutions also by ${\mathbf{\tilde{a}}}_{z}=(\tilde{a}_{z},\{(\tilde{\alpha}_{z})_{v}:v\in V_{\Gamma_{z}}\})$. Define the $V$-involution ${\mathbf{a}}=(a,\{\alpha_{v}\})$ for $\Gamma$ as follows. The involution $a:V_{\Gamma}\to V_{\Gamma}$ is given by $a(v):=\tilde{a}_{z}(v)$ for each $v\in V_{\Gamma_{z}}\subset V_{\Gamma}$. Moreover, for each $v\in V_{\Gamma}$, if $v$ is a vertex of a component $\Gamma_{z}$, put $\alpha_{v}=(\tilde{\alpha}_{z})_{v}$.
It remains to describe the set $\mathcal{A}$ of $E$-connections for $\Gamma$ appearing in ${\mathcal{R}}_{G}$. For this we need some preparations which concern the action of $G$ on the tree $T^{r}_{G}$. For each vertex $z\in Z$ of type (v2), fix any bijective correspondence between the set of edges of the $\theta$-graph $\Gamma_{z}=\Xi_{k(z)}$ and the set $N_{z}$ of vertices of $T^{r}_{G}$ adjacent to $z$ (which are all flexible). Denote this bijection by $\lambda_{z}^{+}$. Fix also an identification of the vertex set of $\Gamma_{z}$ (which consists of 2 vertices) with the doubleton $\partial G(z)\subset\partial G$. Next, recall that for each rigid cluster vertex $z\in Z$, we have a natural bijective correspondence between the set of orbits in $N_{z}$ under the action of $G(z)$ (i.e. the quotient $N_{z}/G(z)$) and the set ${\mathcal{H}}_{z}$ (of conjugacy classes of peripheral subgroups of $G(z)$, under the view of $G(z)$ as an abstract factor, see Example 8.2). This correspondence is given by assigning to any orbit $[u]\in N_{z}/G(z)$ (with $u\in N_{z}$ as any representative of this orbit) the conjugacy class of the subgroup $G(u)<G(z)$ (note that, due to Lemma 9.1, the groups $G(u)$ are indeed contained in $G(z)$). For any $\xi\in{\mathcal{H}}_{z}$, we fix some representative $u_{\xi}\in N_{z}$ of the orbit in $N_{z}/G(z)$ corresponding to $\xi$, and we also fix a bijective correspondence between the vertices in the set $N_{u_{\xi}}\setminus\{z\}$ (which are all flexible, by the fact that $z$ is isolated, see Fact 3.5) and the set of numbers $11,2,\dots,n_{\xi}-$ 1, where $n_{\xi}$ is the degree of $u_{\xi}$ in $T^{r}_{G}$. This induces a bijective correspondence between the set ~$\widetilde{\mathcal{H}}_{z}=\{(\xi,j):\xi\in{\mathcal{H}}_{z},1\leq j\leq n_{\xi}-1\}$ and the set $N^{2}_{z}$ of all those vertices $y$ in $T^{r}_{G}$ lying at distance 2 from $z$, for which the shortest path between $y$ and $z$ in $T^{r}_{G}$ passes through one of the vertices $u_{\xi}$. We denote by $\lambda^{+}_{z}:E_{\widetilde{W}_{z}}\to N^{2}_{z}$ the labelling obtained from $\tilde{\lambda}_{z}$ (where $\tilde{\lambda}_{z}$ is the labelling described in Section 9, right before Fact 9.5) by replacing the labels from ~$\widetilde{\mathcal{H}}_{z}$ with the corresponding elements from $N^{2}_{z}$.
To describe the set $\mathcal{A}$, consider the set $Y$ of orbits of the action of $G$ on the set of all flexible vertices of $T^{r}_{G}$. The set $\mathcal{A}$ will be given as the (disjoint) union of the sets ${\mathcal{A}}_{y}:y\in Y$, which we call blocks. We now pass to describing these blocks ${\mathcal{A}}_{y}$. Fix any orbit $y\in Y$. For any $z\in Z$ of type $(v2)$, let $y^{z}_{i}:i\in I_{y,z}$ be the set of all vertices adjacent to $z$ in the orbit $y$. Similarly, for any rigid cluster vertex $z\in Z$, let $y^{z}_{i}:i\in I_{y,z}$ be the set of all vertices from the set $N^{2}_{z}$ in the orbit $y$. (Note that for some $z$, the index sets $I_{y,z}$ may be empty.) Now, if $z\in Z$ is of type (v2), each $y_{i}^{z}$ as above belongs to $N_{z}$, and hence the preimage $(\lambda^{+}_{z})^{-1}(y_{i}^{z})$ is a subset of the edge set of $\Gamma_{z}=\Xi_{k(z)}$ consistsing of a single edge, namely the edge labelled with $y_{i}^{z}$. If $z\in Z$ is a rigid cluster vertex, each $y_{i}^{z}$ as above belongs to the set $N^{2}_{z}$, and hence the preimage $(\lambda^{+}_{z})^{-1}(y_{i}^{z})$ is a subset of the edge set of the graph $\Gamma_{z}=\widetilde{W}_{z}$, namely it also consists of these edges of $\widetilde{W}_{z}$ which are labelled with $y_{i}^{z}$ (this time there may be more than one such edge). Put
$E_{y}=\bigcup_{z\in Z}\bigcup_{i\in I_{y,z}}(\lambda^{+}_{z})^{-1}(y_{i}^{z}).$
It is not hard to observe that the family $E_{y}:y\in Y$ is then a partition of the edge set of $\Gamma$, i.e., each edge of $\Gamma$ belongs to precisely one of the sets $E_{y}$.
For any orbit $y\in Y$, fix any vertex $v_{y}\in y$, and note that the group $G(v_{y})$ is, up to an isomorphism (a conjugation in $G$), independent of this choice. If $G(v_{y})$ is non-orientable (when viewed as a flexible factor), denote by $E^{\pm}_{y}$ the set of all oriented edges of $\Gamma$ (with both orientations) whose underlying non-oriented edges are in $E_{y}$. Put ${\mathcal{A}}_{y}:=E^{\pm}_{y}\times E^{\pm}_{y}$. On the other hand, if $G(v_{y})$ is orientable, in order to describe the set ${\mathcal{A}}_{y}$, we need to take more subtle aspects of orientations into account. In particular, we will now describe how to associate to any edge $\varepsilon\in E_{y}$ some distinguished orientation, with which it will be denoted $\varepsilon^{\#}$. To this aim, fix in advance, in a $G$-invariant way, orientations for all flexible factors $G(v)$ with $v\in y$. For any edge $\varepsilon\in E_{y}$, let $\Gamma_{z}$ be the component of $\Gamma$ to which $\varepsilon$ belongs. Put $y_{\varepsilon}=\lambda_{z}^{+}(\varepsilon)$, note that it belongs to the orbit $y$, and consider the orientation of the flexible factor $G(y_{\varepsilon})$ as fixed above. This orientation induces orientations of all peripheral subgroups of this factor (in the sense of compatibility of orientations as in Definition 8.12; see also Lemma 8.11(2)). To describe the oriented edge $\varepsilon^{\#}$, we proceed separately in the following two cases.
Case 1. If $z$ is a rigid cluster vertex (so that $\Gamma_{z}$ is the extended Whitehead graph for the rigid cluster factor ${\mathcal{Q}}_{z}=(G(z),\mathcal{H}_{z})$), put $\xi=\lambda_{z}(\varepsilon)\in{\mathcal{H}}_{z}$. We then have $y_{\varepsilon}\in N_{u_{\xi}}\setminus\{z\}$, so the fixed orientation of $G(y_{\varepsilon})$ induces an orientation of its peripheral subgroup $G(y_{\varepsilon})\cap G(u_{\xi})$. This obviously determines an order of the doubleton $\partial G(u_{\xi})$. Since $G(u_{\xi})$ is a peripheral subgroup of the factor $G(z)$ (compare Lemma 9.1), $\partial G(u_{\xi})$ is a peripheral of the space with peripherals associated to $\mathcal{Q}_{z}$. Identifying the latter naturally with the core $C(\widetilde{\mathcal{V}}_{z})$ of the associated $V$-tree ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{v})$ (as in Lemma 9.6(1)), $\partial G(u_{\xi})$ is a peripheral of this core. Since $\varepsilon$ is an edge of the extended Whitehead graph $\widetilde{W}_{z}=\Gamma_{z}$, and since it belongs to edges associated to the peripheral $\partial G(u_{\xi})$, it carries the order-induced orientation (with respect to the above described order of $\partial G(u_{\xi})$). (See Remark 9.10(3).) We denote by $\varepsilon^{\#}$ the edge $\varepsilon$ with this order-induced orientation.
Case 2. If $z$ is of type (v2) (and non-adjacent to any rigid cluster vertex), $\Gamma_{z}=\Xi_{k(z)}$ is a $\theta$-graph. The edge $\varepsilon$ corresponds then to the unique vertex $y_{i}^{z}\in N_{z}$. Again, since $G(y_{i}^{z})$ is oriented, it induces the order on the boundary set $\partial G(z)$, and hence (by our earlier fixed identification) on the vertex set of $\Gamma_{z}$. The latter induces an orientation on $\varepsilon$, and we again denote by $\varepsilon^{\#}$ the edge $\varepsilon$ with this induced orientation.
Put $E_{y}^{\#}:=\{\varepsilon^{\#}:\varepsilon\in E_{y}\}$, and put also
${\mathcal{A}}_{y}:=\{(o_{1},\bar{o}_{2}):o_{1},o_{2}\in E_{y}^{\#}\}\cup\{(\bar{o}_{1},o_{2}):o_{1},o_{2}\in E_{y}^{\#}\}$
where $\bar{o}_{i}$ denotes the oppositely oriented edge $o_{i}$, for $i=1,2$.
Having defined all the blocks ${\mathcal{A}}_{y}$, we put ${\mathcal{A}}=\bigcup_{y\in Y}{\mathcal{A}}_{y}$. We skip the straightforward verification of the fact that the so described set ${\mathcal{A}}$ does not depend on the choices of the set of representatives $Z$, and the representatives $v_{y}$ of the orbits $y\in Y$. This completes the description of the graphical connecting system ${\mathcal{R}}_{G}$. The fact that ${\mathcal{R}}_{G}$ satisfies the transitivity property (4) from Definition 6.1 follows, roughly speaking, from the fact that the tree $T^{r}_{G}$ is connected, and we omit the details.
Given the graphical connecting system ${\mathcal{R}}_{G}$, as introduced above, we may state the following result, which describes explicitly the form of a regular tree of graphs which appears only implicitly in our statement of Therorem A (in condition (2)). This result, Theorem 10.1 below, may be viewed as the third main result of the paper. Its proof, after necessary preparations provided in the next two sections, will be presented in Section 13.
Theorem 10.1.
Let $G$ be a 1-ended hyperbolic group which is not cocompact Fuchsian, and suppose that each rigid cluster factor of its reduced Bowditch JSJ splitting is virtually free. Let ${\mathcal{R}}_{G}$ be the graphical connecting system associated to $G$ (as described above in this section). Then the Gromov boundary $\partial G$ is homeomorphic to the regular tree of graphs ${\mathcal{X}}({\mathcal{R}}_{G})$.
We finish this section with some further observations concerning the edges in the graph $\Gamma$ of the connecting system ${\mathcal{R}}_{G}$. These observations will be used in the proof of Theorem 10.1 given in Section 13.
Let $z\in Z$ be one of the orbit representatives fixed at the beginning of the above description of the graphical connecting system ${\mathcal{R}}_{G}$. Suppose first that $z$ is a rigid cluster vertex of $T^{r}_{G}$ (the remaining case when $z$ is of type (v2) will be discussed later). A bunch of edges in the extended Whitehead graph $\widetilde{W}_{z}$ is any collection $\{e_{1},\dots,e_{n_{\xi}-1}\}$ of edges that replace an edge $e$ of the Whitehead graph $W_{z}$, as in the description of $\widetilde{W}_{z}$ given in Section 9, where $\xi=\lambda_{z}(e)\in{\mathcal{H}}_{z}$ is the label of $e$. Clearly, all edges in any bunch share the set of endpoints. We assign to a bunch as above the same label $\xi$.
Now, as mentioned in the above description of ${\mathcal{R}}_{G}$, any bunch of edges $\mathcal{M}$ labelled with $\xi$ is in a fixed bijective correspondence (provided by the labelling $\lambda_{z}^{+}$) with the vertex set $N_{u_{\xi}}\setminus\{z\}$. This correspondence allows us to split $\mathcal{M}$ into subsets corresponding to the orbits of the action of $G(u_{\xi})$ on $N_{u_{\xi}}\setminus\{z\}$. More precisely, observe that due to Lemma 9.1, $z$ is a global fixed point of the action of $G(u_{\xi})$ on $N_{u_{\xi}}$ (and an alternative justification follows by the fact that $z$ is the unique rigid cluster vertex in $N_{u_{\xi}}$, due to its isolatedness). Consequently, the $G(u_{\xi})$-orbit $[x]$ of any vertex $x\in N_{u_{\xi}}\setminus\{z\}$ is contained in $N_{u_{\xi}}\setminus\{z\}$. The subset of $\mathcal{M}$ corresponding to any such orbit $[x]$ will be called a packet of edges in $\widetilde{W}_{z}$.
We will later need the following easy but a bit technical observation, the proof of which we omit.
Fact 10.2.
Let $z\in Z$ be a rigid cluster vertex, let $\xi\in{\mathcal{H}}_{z}$, let $\mathcal{M}$ be a bunch of edges in $\Gamma_{z}=\widetilde{W}_{z}$ labelled with $\xi$, and let $M\subset{\mathcal{M}}$ be a packet corresponding to the $G(u_{\xi})$-orbit $[x]$ of some vertex $x\in N_{u_{\xi}}\setminus\{z\}$.
  1. (1)
    The cardinality of the packet $M$ coincides with the index $[G(u_{\xi}):G(u_{\xi})\cap G(x)]=[G(u_{\xi}):G([u_{\xi},x])]$.
  2. (2)
    Having fixed $\mathcal{M}$, for all $M$ and $x$ as above for which the corresponding flexible factor $G(x)$ is orientable, we have
    1. (a)
      if the group $G(u_{\xi})$ is orientable then the orientations of all (oriented) edges $\varepsilon^{\#}$ with $\varepsilon$ in $M$ (as defined in the above description of the set of E-connections for ${\mathcal{R}}_{G}$) are compatible, i.e., they induce the same order on the common set of endpoints of these edges;
    2. (b)
      if the group $G(u_{\xi})$ is non-orientable then the cardinality of the packet $M$ is even, and half of the orientations of the edges $\{\varepsilon^{\#}:\varepsilon\in M\}$ are compatible with one of the orders in the common set of endpoints, while the other half of these orientations are compatible with the opposite order.
Let now $z\in Z$ be a vertex of type (v2). Recall that, by the description of $Z$, the set $N_{z}$ of all vertices in $T^{r}_{G}$ adjacent to $z$ contains no rigid cluster vertex. Recall also that the component $\Gamma_{z}$ of the graph $\Gamma$ is isomorphic to the $\theta$-graph $\Xi_{k(z)}$, where $k(z)$ is the degree of $z$ in $T^{r}_{G}$, and that we have fixed a bijective correspondence between the edge set $E_{\Gamma_{z}}$ and the set $N_{z}$. For any $G(z)$-orbit $[x]$ (where $x$ is any vertex of $N_{z}$), the corresponding set of edges of the $\theta$-graph $\Gamma_{z}$ will be again called a packet of edges in $\Gamma_{z}$.
We have the following observation similar to Fact 10.2, the proof of which we also skip.
Fact 10.3.
Let $z\in Z$ be a vertex of type (v2), and let $M$ be the packet of edges in the $\theta$-graph $\Gamma_{z}$ corresponding to the $G(z)$-orbit $[x]$ of some vertex $x\in N_{z}$.
  1. (1)
    The cardinality of the packet $M$ coincides with the index
    $[G(z):G(z)\cap G(x)]=[G(z):G([z,x])].$
  2. (2)
    For all $M$ and $x$ as above for which the corresponding flexible factor $G(x)$ is orientable, we have
    1. (a)
      if $G(z)$ is orientable then the orientations of all (oriented) edges $\varepsilon^{\#}$ with $\varepsilon$ in any such $M$ (as defined in the above description of the set of E-connections for ${\mathcal{R}}_{G}$) are compatible, i.e., they induce the same order on the common set of endpoints of these edges;
    2. (b)
      if $G(z)$ is non-orientable then the cardinality of the packet $M$ is even, and half of the orientations of the edges $\{\varepsilon^{\#}:\varepsilon\in M\}$ are compatible with one of the orders in the common set of endpoints, while the other half of these orientations are compatible with the opposite order.
11. Regular tree systems of spaces with doubleton peripherals
In this section we prepare some tools and terminology needed in our proof of Theorem 10.1 (from which we then deduce the implications (3$\Rightarrow$(2) in Theorem A and (3$\Rightarrow$(2) in Theorem B). By a space with doubleton peripherals we mean a space with peripherals in which each peripheral subspace consists of precisely two points (i.e. is a doubleton). We start with a rather long definition.
Definition 11.1.
  1. (1)
    A decoration $\mathcal{D}$ of a space with doubleton peripherals $X$ is a collection of the following data:
    • a labelling $\tau$ which to any peripheral subspace $P$ of $X$ associates an element $\tau(P)$ from a finite set $\mathcal{T}$; $\tau(P)$ is then called the type of $P$;
    • for each peripheral $P$, either a single fixed order of the set of two elements of $P$, called the orientation of $P$ (in which case $P$ is called orientable), or an artificial assignment to $P$ of both orders of its two elements (in which case we say that $P$ is non-orientable); we also assume that the peripherals of a fixed type are either all orientable or all non-orienbtable, so that we can speak of orientable and non-orientable types.
    A pair $(X,{\mathcal{D}})$ as above will be called a decorated space with doubleton peripherals.
  2. (2)
    Given a space with doubleton peripherals $X$ and its decoration $\mathcal{D}$, the set of directed peripherals of $(X,{\mathcal{D}})$ consists of all peripherals with all assigned orientations (one directed peripheral for each orientable $P$, and two directed peripherals for each non-orientable $P$). A type of a directed peripheral is the type of its underlying counterpart (i.e. of the same peripheral with its orientation ignored).
  3. (3)
    A decoration-preserving homeomorphism of $(X,{\mathcal{D}})$ is a homeomorphism $h$ of $X$ which maps peripherals to peripherals, which preserves the types (i.e. $\tau(h(P))=\tau(P)$ for any peripheral $P$), and which preserves the orientations of the peripherals. The latter means that $h$ maps non-orientable peripherals to non-orientable ones, and it maps orientable peripherals to orientable ones, preserving their assigned orientations (in other words, the set of all directed peripherals of $(X,{\mathcal{D}})$ is preserved by $h$). Note that the inverse $h^{-1}$ of any decoration-preserving homeomorphism $h$ of $(X,{\mathcal{D}})$ is also a decoration-preserving homeomorphism of $(X,{\mathcal{D}})$.
  4. (4)
    A pair $(X,{\mathcal{D}})$ as above is called a regular space with doubleton peripherals if it satisfies the following regularity condition: for any two directed peripherals of $(X,{\mathcal{D}})$ having the same type there is a decoration-preserving homeomorphism of $(X,{\mathcal{D}})$ which maps the first of these two directed peripherals to the second one. In particular, if $P$ is a non-orientable peripheral, there is a decoration-preserving homeomorphism which maps $P$ to $P$ and for which the restriction $h|_{P}$ is a transposition.
We now present a few classes of examples of regular spaces with doubleton peripherals. These examples will be used in our later considerations and arguments concerning the proof of Theorem 10.1.
Example 11.2.
Recall that to any abstract factor ${\mathcal{F}}=(K,\{H_{i}\}_{i\in I})$ (as in Section 8.1) there is associated the induced space with doubleton peripherals, as described in Definition 8.3; we denote this space with peripherals by $X_{\mathcal{F}}$. We describe a decoration ${\mathcal{D}}={\mathcal{D}}_{\mathcal{F}}$ naturally associated to $X_{\mathcal{F}}$, which turns out to satisfy the regularity condition of Definition 11.1(4), and which yields what we call the regular space with doubleton peripherals associated to an abstract factor.
The labelling $\tau$ of $\mathcal{D}$ has values in the set ${\mathcal{T}}:=I$ and is given by the formula
$\tau(\partial(gH_{i}g^{-1})):=i.$
The orientability (and non-orientability) data for the peripherals, as well as the assigned orientations, are described by inducing in the obvious way the similar data for the corresponding peripheral subgroups (provided by fixed orientation data for the factor, as given in Definition 8.5). By Lemma 8.6, the so described decoration ${\mathcal{D}}_{\mathcal{F}}$ satisfies the regularity condition.
Example 11.3 (densely decorated punctured circle).
Let $(K,\Omega)$ be a punctured circle, i.e., a space with peripherals (as described right before Definition 2.3) in which $K=(S^{1})^{\circ}$, with its standard family $\Omega$ of peripheral subspaces, which are doubletons. Fix an arbitrary auxiliary orientation of $(K,\Omega)$ (in the sense of Remark 8.9). Let ${\mathcal{D}}$ be any decoration for $(K,\Omega)$ which is dense, i.e., satisfies one of the following two conditions:
  1. (a)
    for each type of peripheral, the union of all directed peripherals of this type in $((K,\Omega),{\mathcal{D}})$, with orientations compatible (respectively, incompatible) with the fixed orientation of $(K,\Omega)$, is dense in $K$; that is, if a type of peripherals is non-orientable, we require simply that the union of peripherals of this type is dense in $K$, and if a type is orientable, we require both that the union of all those peripherals of this type whose orientations are compatible with the fixed orientation of $(K,\Omega)$ is dense in $K$ and that the union of all those peripherals of this type whose orientations are incompatible with the fixed orientation of $(K,\Omega)$ is dense in $K$;
  2. (b)
    all peripherals are orientable, and for each type either all peripherals of this type are oriented consistently with the fixed orientation of $(K,\Omega)$, or they are all oriented inconsistently with this orientation; in any case, the union of the peripherals of any fixed type is dense in $K$.
We then say that the pair $((K,\Omega),{\mathcal{D}})$ is a densely decorated punctured circle. We will speak of two kinds of densely decorated punctured circles: those satisfying condition (a) above will be called non-orientable, and those satisying condition (b) will be called orientable. The next lemma collects some useful properties of densely decorated punctured circles, and we omit its straightforward proof which can be performed by a standard back and forth argument.
Lemma 11.4.
  1. (1)
    Any densely decorated punctured circle is a regular space with doubleton peripherals.
  2. (2)
    Any two orientable densely decorated punctured circles with the same number of types of peripherals are isomorphic (via a decoration-preserving homeomorhism $h$), after appropriate identification of the types, and after possibly reversing orientations of all peripherals of some of the types. Moreover, for any bijection $\beta$ between the sets of types of the punctured circles above (which we view as the identification of the types) there is a homeomorphism $h$ as above which is compatible with $\beta$ (i.e. such that for any peripheral $P$ we have $\tau(h(P))=\beta(\tau(P))$).
  3. (3)
    Any two non-orientable densely decorated punctured circles with the same number of types of orientable peripherals, and the same number of types of non-orientable peripherals, are isomorphic, after appropriate identification of the types. Moreover, for any bijection $\beta$ between the sets of types of the punctured circles above, if $\beta$ respects orientability and non-orientability of the types, then there is an isomorphism $h$ between those decorated punctured circles which is compatible with $\beta$.
Remark 11.5.
We should note here that if ${\mathcal{F}}=(K,\{H_{i}\}_{i\in I})$ is an abstract flexible factor, then the space with peripherals $(X_{\mathcal{F}},{\mathcal{D}}_{{\mathcal{F}}})$ associated to $\mathcal{F}$ (as described in Example 11.2), is a densely decorated punctured circle (see Fact 8.13 for the justification). This densely decorated punctured circle is easily seen to be orientable if and only if the flexible factor $\mathcal{F}$ itself is orientable. The numbers of types of orientable and non-orientable peripherals in $(X_{\mathcal{F}},{\mathcal{D}}_{{\mathcal{F}}})$ coincide then with the numbers of (the conjugacy classes of) the peripheral subgroups $H_{i}$ which are orientable and non-orientable, respectively.
We now proceed one level further, and discuss some tree systems made out of regular spaces with doubleton peripherals. As we will see, such tree systems are induced, in an essentially unique way, by rather simple data described in the next definitions.
Definition 11.6.
Given an arbitrary regular (decorated) space with doubleton peripherals $((K,\Omega),{\mathcal{D}})$, a degree function for this space is a positive integer valued function $d:\Omega\to\mathbb{Z}$ which is constant on types, i.e., such that $d(P_{1})=d(P_{2})$ for any two peripherals $P_{1},P_{2}$ of the same type. We call a collection of data ${\mathcal{D}}^{+}=({\mathcal{D}},d)$ consisting of a decoration $\mathcal{D}$ and a degree function $d$ an extra-decoration for $(K,\Omega)$. The pair $((K,\Omega),{\mathcal{D}}^{+})$ is then called an extra-decorated regular space with doubleton peripherals.
Definition 11.7.
Given a finite family ${\mathcal{K}}=((K_{j},\Omega_{j}),{\mathcal{D}}_{j}^{+})_{j\in J}$ of regular extra-decorated spaces with doubleton peripherals, denote by ${\mathcal{T}}={\mathcal{T}}_{\mathcal{K}}=\bigsqcup_{j\in J}{\mathcal{T}}_{j}$ the disjoint union of the sets ${\mathcal{T}}_{j}$ of types of the peripherals in the families $\Omega_{j}$. A connecting system for regular spaces with doubleton peripherals is a pair ${\mathcal{P}}=({\mathcal{K}},{\mathcal{B}})$, where $\mathcal{K}$ is a family as above, and where $\mathcal{B}$, called a block structure, consists of the following data:
  1. (1)
    a partition ${\mathcal{T}}_{\mathcal{K}}=\bigsqcup_{a\in A}B_{a}$ of the set of types ${\mathcal{T}}_{\mathcal{K}}$ into subsets $B_{a}$ called blocks;
  2. (2)
    for each block $B_{a}$, additional on information whether $B_{a}$ is orientable or non-orientable; we require that if $B_{a}$ is orientable then it necessarily consists of orientable types only, but if $B_{a}$ is non-orientable, there is no restriction on orientability and non-orientability of the types that it contains.
We also require in the definition that the block structure $\mathcal{B}$ fulfills the following transitivity condition: for any two indices $j,j^{\prime}\in J$ there is a sequence of indices $j=j_{0},j_{1},\dots,j_{k}=j^{\prime}$ such that for each $i=1,\dots,k$ there exist types $\tau\in{\mathcal{T}}_{j_{i-1}}$ and $\tau^{\prime}\in{\mathcal{T}}_{j_{i}}$ which belong to the same block.
Definition 11.8.
Let $\Theta=(T,\{L_{t}\}),\{\Sigma_{u}\},\{\varphi_{e}\})$ be a tree system, and let
${\mathcal{P}}=(((K_{j},\Omega_{j}),{\mathcal{D}}_{j}^{+})_{j\in J},{\mathcal{B}})$
be a connecting system for regular spaces with doubleton peripherals. For each $j\in J$, denote by $d_{j}:\Omega_{j}\to\mathbb{Z}$ the degree function of the extra-decoration ${\mathcal{D}}_{j}^{+}$. We say that $\Theta$ is modelled on $\mathcal{P}$ if each constituent space $L_{t}$ of $\Theta$ can be equipped with an identification (as a space with peripherals) with one of the spaces $K_{j}$ of $\mathcal{P}$ (and hence also equipped with the corresponding extra-decoration ${\mathcal{D}}_{j}^{+}$) so that the following conditions are satisfied:
  1. (1)
    for each white vertex $u$ of $T$ there is a block $B$ of ${\mathcal{B}}$ such that the set of types of the peripherals of the form $\varphi_{[u,t]}(\Sigma_{u})\subset L_{t}$ (under the identification of $L_{t}$ with one of the $K_{j}$, as above), where $t$ runs through the set of all vertices of $T$ adjacent to $u$, coincides with $B$;
  2. (2)
    for each $u$ and $B$ as in (1), if the block $B$ is orientable, then for any two edges $e_{1}=[u,t_{1}]$ and $e_{2}=[u,t_{2}]$ adjacent to $u$ the map $\varphi_{e_{2}}\varphi_{e_{1}}^{-1}$ between the appropriate peripherals in $L_{t_{1}}$ and $L_{t_{2}}$ preserves the orientations of these peripherals (provided by the identifications of both $L_{t_{i}}$’s with the corresponding $K_{j}$’s, as above); moreover, we assume in this case that for any $\lambda\in B$, denoting by $K_{j}$ this unique space of $\mathcal{P}$ which contains peripherals of type $\lambda$, the number of (black) vertices $t$ adjacent to $u$ for which the spaces $L_{t}$ are identified with $K_{j}$ and the peripherals $\varphi_{[u,t]}(\Sigma_{u})$ are of type $\lambda$ is equal to the degree $d_{j}(\varphi_{[u,t]}(\Sigma_{u}))$ (for any $t$ as above);
  3. (3)
    for any $u$ and $B$ as in (1), if the block $B$ is non-orientable, and if $\lambda\in B$ is a non-orientable type then, denoting by $K_{j}$ this unique space of $\mathcal{P}$ which contains peripherals of type $\lambda$, the number of (black) vertices $t$ adjacent to $u$ for which the spaces $L_{t}$ are identified with $K_{j}$ and the peripherals $\varphi_{[u,t]}(\Sigma_{u})$ are of type $\lambda$ is equal to the degree $d_{j}(\varphi_{[u,t]}(\Sigma_{u}))$ (for any $t$ as above);
  4. (4)
    for any $u$ and $B$ as in (3) (i.e. with $B$ non-orientable), if $\lambda\in B$ is an orientable type then, again denoting by $K_{j}$ this unique space of $\mathcal{P}$ which contains peripherals of type $\lambda$, the number of (black) vertices $t$ adjacent to $u$ for which the spaces $L_{t}$ are identified with $K_{j}$ and the peripherals $\varphi_{[u,t]}(\Sigma_{u})$ are of type $\lambda$ is equal to twice the degree, i.e., to $2\cdot d_{j}(\varphi_{[u,t]}(\Sigma_{u}))$ (for any $t$ as above); moreover, if we choose an auxiliary orientation of the doubleton $\Sigma_{u}$ (i.e. an order for its two points) then for exactly half of the vertices $t$ as above (i.e. for $d_{j}(\varphi_{[u,t]}(\Sigma_{u}))$ of those vertices) the map $\varphi_{[u,t]}$ preserves the orientations, and for another half of such vertices it reverses the orientations.
Informally speaking, the above conditions mean that the constituent spaces of $\Theta$, after their identifications with the spaces $K_{j}$, are glued to each other in $\Theta$ along the peripherals belonging to the same blocks $B$ of $\mathcal{B}$, in numbers appropriately determined by the degrees. Moreover, if a block $B$ is orientable, all the gluings preserve the orientations, and if $B$ is non-orientable, the condition is a bit more complicated, namely: gluings along non-orientable peripherals are arbitrary, and those along orientable peripherals are performed in equal numbers with both possible orientations, for each type.
Any tree system $\Theta$ which is modelled on some connecting system for regular spaces with doubleton peripherals will be called a regular tree system of spaces with doubleton peripherals. Any family of identifications of the constituent spaces $(L_{t},\Omega_{t})$ of $\Theta$ with the spaces $(K_{j},\Omega_{j})$ of $\mathcal{P}$ for which conditions (1)-(4) above are satisfied will be called a $\mathcal{P}$-atlas for $\Theta$.
The significance of the concept of being modelled on a connecting system for regular spaces with doubleton peripherals for our arguments in the proof of Theorem 10.1, as given in Section 13 below, comes from the following observation.
Lemma 11.9.
Any two tree systems modelled on the same connecting system $\mathcal{P}$ for regular spaces with doubleton peripherals are isomorphic (as tree systems).
We omit a rather straightforward (though technically inconvenient to expose) proof of this lemma, indicating only that an isomorphism from the assertion can be constructed recursively, in a way respecting the identifications from the $\mathcal{P}$-atlases, by refering to the regularity of the involved decorated spaces with doubleton peripherals (see part (4) of Definition 11.1). To start such a recursive construction, one needs to refer to the transitivity condition in Definition 11.7 to find some initial constituent spaces in both systems which are identified with the same space $K_{j}$ of the corresponding connecting system.
The next example will play a central role in the final step of the proof of Theorem 10.1, as given in Section 13.
Example 11.10 ($\Theta^{r}(G)$ as a regular tree system of spaces with doubleton peripherals).
We will show that any tree system $\Theta^{r}(G)$, as described in Section 3.2, has a natural structure of a regular tree system of spaces with doubleton peripherals. We will do this by explicitly describing a connecting system ${\mathcal{P}}={\mathcal{P}}^{r}_{G}$ on which the corresponding tree system $\Theta^{r}(G)$ is modelled. We refer to the terminology and notation as in Section 3.2.
To describe the appropriate family $(K_{j},\Omega_{j})_{j\in J}$ of regular spaces with doubleton peripherals in the connecting system ${\mathcal{P}}^{r}_{G}$, consider a set $v_{j}:j\in J$ of representatives of orbits of the action of $G$ on the set of black vertices of $T^{r}_{G}$ (i.e. on the set of flexible vertices and rigid cluster vertices). For each $j\in J$ put
$K_{j}:=\partial G(v_{j})\,\,\,\,\hbox{and}\,\,\,\,\Omega_{j}:=\{\partial G(u):u\in N^{T^{r}_{G}}_{v_{j}}\},$
where $N^{T^{r}_{G}}_{v_{j}}$ is the set of (white) vertices in $T^{r}_{G}$ adjacent to $v_{j}$. (In other words, put $(K_{j},\Omega_{j})$ to be the space with peripherals associated to the group $G(v_{j})$ viewed as an abstract factor in a way described in Example 8.2.)
In order to describe appropriate decorations ${\mathcal{D}}_{j}$ for the spaces $(K_{j},\Omega_{j})$, we need some preparations. Fix in advance, in a conjugacy invariant way, orientations for all those 2-ended groups $G(u)$ (at the white vertices $u$ of $T^{r}_{G}$) which are orientable. Note that this induces the orientations of all those edge groups $G(e)$ for which the edge $e$ has as its endpoint a white vertex $u$ with orientable $G(u)$. Fix also in a conjugacy invariant way orientations for all orientable groups $G(e)$ at those edges $e$ whose white endpoint groups $G(u)$ are non-orientable. Altogether, this results with some conjugacy invariant and consistent choices of orientations for all edge groups $G(e)$ which are orientable. As a consequence, this induces also the choices of orientations (in a conjugacy invariant way) for all orientable peripheral subgroups of all black vertex groups $G(v_{j})$ viewed as abstract factors.
For any $(K_{j},\Omega_{j})$ as above, take as a decoration ${\mathcal{D}}_{j}$ all the data as in Example 11.2, under viewing $G(v_{j})$ as an abstract factor, with the orientations of orientable peripheral subgroups chosen in a way described in the previous paragraph. To describe the appropriate degree function $d_{j}$, consider any peripheral $\partial G(u)$ in $K_{j}=\partial G(v_{j})$, where $u$ is some (white) vertex of $T^{r}_{G}$ adjacent to $v_{j}$. Recall that the peripheral subgroup corresponding to this peripheral of $K_{j}$ is $G([u,v_{j}])=G(u)\cap G(v_{j})$, which we denote briefly by $H_{u}$. If $G(u)$ (which is a 2-ended group) is orientable, we take as $d_{j}(\partial G(u))$ the index $[G(u):H_{u}]$. If both $G(u)$ and $H_{u}$ are non-orientable, we similarly put $d_{j}(\partial G(u)):=[G(u):H_{u}]$. However, if $G(u)$ is non-orientable, and $H_{u}$ is orientable, we notice that the index $[G(u):H_{u}]$ is an even number, and we take as $d_{j}(\partial G(u))$ half of this number. Note that, due to the final comment in Example 11.2, with the so described extra-decoration ${\mathcal{D}}_{j}^{+}=({\mathcal{D}}_{j},d_{j})$, each of the spaces $(K_{j},\Omega_{j})$ becomes an extra-decorated regular space with doubleton peripherals.
We now pass to the description of appropriate data =${\mathcal{B}}=\{B_{a}:a\in A\}$. Take as $A$ the set of $G$-orbits of the white vertices of $T^{r}_{G}$. For any orbit $a\in A$, let $u$ be any vertex in this orbit. Note that for any (black) vertex $t$ adjacent to $u$, the corresponding space with peripherals $\partial G(t)$ is identified, uniquely up to a decoration preserving homeomorphism, with the appropriate space $K_{j}=\partial G(v_{j})$, where $v_{j}$ is in the same orbit as $t$, via any of the maps from the $G$-action of $G$ on $\partial G$ which sends $\partial G(v_{j})$ to $\partial G(t)$. In particular, under the latter identifications, all peripherals in all such spaces $\partial G(t)$ have well defined types. Define $B_{a}$ to be the set of all types appearing among the peripherals of the form $\varphi_{[u,t]}(\Sigma_{u})$ (in the corresponding spaces $\partial G(t)$), where $t$ runs through all (black) vertices of $T^{r}_{G}$ adjacent to $u$. Furthermore, we declare that the block $B_{a}$ is orientable if and only if the corresponding 2-ended group $G(u)$ is orientable.
We skip the straightforward verification of the fact that all the data above are well defined. It is also not hard to verify that the above described data satisfy all requirements for being a connecting system for regular spaces with doubleton peripherals, and we denote this system by ${\mathcal{P}}^{r}_{G}$.
It remains to show that the tree system $\Theta^{r}(G)$ is indeed modelled on the connecting system ${\mathcal{P}}^{r}_{G}$. To do this, we will now describe some ${\mathcal{P}}^{r}_{G}$-atlas ${\mathbf{A}}^{r}_{G}$ for $\Theta^{r}(G)$. Given any constituent space $\partial G(t)$ of $\Theta^{r}(G)$ (where $t$ is a black vertex of $T^{r}_{G}$), consider the unique vertex $v_{j}$ which lies in the same $G$-orbit as $t$. Consider then any $g\in G$ which sends $t$ to $v_{j}$, and take as an identification in ${\mathbf{A}}^{r}_{G}$ the restriction to $\partial G(t)$ of the homeomorphism of $\partial G$ induced by $g$ (which sends $\partial G(t)$ to $\partial G(v_{j})=K_{j}$). The verification that ${\mathbf{A}}^{r}_{G}$ is indeed a ${\mathcal{P}}^{r}_{G}$-atlas (i.e. that it satisfies conditions (1)-(4) of Definition 11.8) becomes (nearly tautologically) clear in view of the following observation, the easy proof of which we also skip.
Observation 11.11.
Under the identifications from the atlas ${\mathbf{A}}^{r}_{G}$, the data from the connecting system ${\mathcal{P}}^{r}_{G}$ can be a posteriori interpreted in the following terms corresponding to the features of the action of $G$ on $T^{r}_{G}$:
  • the set $\mathcal{T}_{G}$ of all types of the peripherals in all constituent spaces in the system $\Theta^{r}(G)$ can be identified with the set of all $G$-orbits $[e]$ of the edges $e$ of $T^{r}_{G}$; under this identification, each peripheral $\varphi_{e}(\Sigma_{u})$ is assigned with the type $[e]\in\mathcal{T}_{G}$;
  • a type $[e]\in\mathcal{T}_{G}$ is orientable iff the corresponding 2-ended group $G(e)$ is orientable;
  • the set $\mathcal{B}$ of blocks is indexed by the set of $G$-orbits $[u]$ in the set of all white vertices $u$ in $T^{r}_{G}$, and each block $B_{[u]}$ consists of those types $[e]$ which correspond to the edges $e$ adjacent to $u$;
  • the set of types $\mathcal{T}_{G}$ is partitioned into subsets
    ${\mathcal{T}}_{j}=\{[e]:\text{ $e$ is an edge of $T^{r}_{G}$ adjacent to $v_{j}$}\},$
    and each ${\mathcal{T}}_{j}$ is the set of types of peripherals in $(K_{j},\Omega_{j})$;
  • if $[e]\in{\mathcal{T}}_{j}$, and if $u$ denotes the white endpoint of $e$, then
    1. (a)
      if both groups $G(e),G(u)$ are orientable, or both are non-orientable, the degree $d_{j}([e])$ coincides with the number of edges adjacent to $u$ in the $G$-orbit of $e$;
    2. (b)
      if $G(e)$ is orientable and $G(u)$ is non-orientable, the degree $d_{j}([e])$ is equal to half of the number of edges adjacent to $u$ in the $G$-orbit of $e$ (where the latter number is easily seen to be even).
Remark 11.12.
Note that, under the notation of the above Example 11.10, if $v_{j}$ is a flexible vertex, the decorated space with doubleton peripherals $((K_{j},\Omega_{j}),{\mathcal{D}}_{j})$ is a densely decorated punctured circle, as described in Example 11.3 (for justification, see Example 8.8 and the last paragraph in Example 11.3). This densely decorated punctured circle is easily seen to be orientable if and only if the flexible factor corresponding to the group $G(v_{j})$ is orientable. The numbers of types of orientable and non-orientable peripherals in $((K_{j},\Omega_{j}),{\mathcal{D}}_{j})$ coincide with the numbers of conjugacy classes in $G(v_{j})$ among the orientable (and non-orientable, respectively) subgroups of the form $G(u)\cap G(v_{j}):u\in N^{T^{r}_{G}}_{v_{j}}$. As we have already noticed before, the information from the last two sentences determines the decorated space $((K_{j},\Omega_{j}),{\mathcal{D}}_{j})$ uniquely up to an isomorphism (see Lemma 11.4(2)(3)).
12. Trees of internally punctured segments
In this section we describe one more ingredient necessary in our proof of Theorem 10.1, as given in Section 13. This ingredient consists of expressing various decorated punctured circles as trees of appropriate internally punctured segments. The section is split into two subsections corresponding to two kinds of the described objects.
12.1. Orientable trees of segments
Fix any positive integer $m\geq 1$, and let ${\mathcal{T}}$ be a set of cardinality $m$, which we will view as the set of types. For each type $\tau\in{\mathcal{T}}$, denote by $I^{\circ}_{\tau}$ the internally punctured segment $I^{\circ}_{\rm int}$, as described in Example 7.2, equipped with the structure consisting of the following data:
  1. (1)
    except for the standard peripheral subspaces in $I^{\circ}_{\rm int}$, which are all doubletons, and which we call the ordinary peripherals of $I^{\circ}_{\tau}$, we consider also an additional subspace, denoted $\partial I^{\circ}_{\tau}$, consisting of the two endpoints of $I^{\circ}_{\rm int}$, and we call it the special peripheral of $I^{\circ}_{\tau}$;
  2. (2)
    identifying naturally $I^{\circ}_{\rm int}$ with the punctured circle $(S^{1})^{\circ}$, and fixing one of its orientations (in the sense of the last sentence of Definition 8.10), we equip all its peripherals (both ordinary and special) with the orientations compatible with the just fixed orientation of the whole space;
  3. (3)
    we associate the type $\tau$ to the special peripheral of $I^{\circ}_{\tau}$, and we do not associate any types to its ordinary peripherals.
Define the tree system $\Psi^{\text{o}}_{m}(I^{\circ}_{\rm int})$ by requiring the following (the superscript “o” on $\Psi^{\text{o}}_{m}$ indicates “orientable”):
  1. (a)
    its constituent spaces are the copies of the internally punctured segments $I_{\tau}^{\circ}:\tau\in{\mathcal{T}}$, with ordinary peripherals as the peripherals of these spaces (this means that the special peripherals are not among the peripherals of those constituent spaces);
  2. (b)
    each gluing in the system involves precisely two constituent spaces (equivalently, each white vertex of the underlying tree of the system has degree 2);
  3. (c)
    gluing maps respect the orientations of the glued peripherals, which means that those maps reverse the orientations of the glued peripherals;
  4. (d)
    for each $\tau,\tau^{\prime}\in{\mathcal{T}}$ (not necessarily distinct), and for any constituent space $K_{t}$ being a copy of $I^{\circ}_{\tau}$, the union of those peripherals of $K_{t}$ along which $K_{t}$ is glued to a copy of $I^{\circ}_{\tau^{\prime}}$, is dense in $K_{t}$.
The fact that such a tree system is unique up to an isomorphism follows by discussion and arguments similar to those in Example 11.3, Lemma 11.4 and Lemma 11.9, but this fact is not essential for our further considerations.
Lemma 12.1.
Consider the limit $X$ of the tree system $\Psi^{\text{o}}_{m}(I^{\circ}_{\rm int})$ of punctured segments, and view it as a space with peripherals, where peripherals are the images of all special peripherals in all constituent punctured segments. Equip also all those peripherals in a natural way with the types from the set $\mathcal{T}$, and with the naturally induced orientations. Then $X$ is homeomorphic, as a space with peripherals, to an orientable densely decorated punctured circle, as described in Example 11.3, and it has precisely $m$ types of peripherals.
Proof.
At first, we will ignore the types and the orientations of the peripherals, and show that, as a pure space with peripherals, the limit of the tree system $\Psi^{\text{o}}_{m}(I^{\circ}_{\rm int})$ is homeomorphic to the punctured circle $(S^{1})^{\circ}$. To do this, consider a related tree system $\overline{\Psi}^{\text{o}}_{m}(I^{\circ}_{\rm int})$ with the same underlying tree $T$, in which the constituent spaces $\overline{K}_{t}$ are obtained from the corresponding constituent spaces $K_{t}$ of $\Psi^{\text{o}}_{m}(I^{\circ}_{\rm int})$ (which are punctured segments) as follows. Glue to each $K_{t}$ a segment $I_{t}$ via a gluing map which sends the special peripheral of $K_{t}$ bijectively to the boundary $\partial I_{t}$. Identify the (ordinary) peripherals in $K_{t}$ as subspaces of $\overline{K}_{t}$, and view them as the peripherals of $\overline{K}_{t}$. For the gluing maps (or more formally, connecting maps) in the new system $\overline{\Psi}^{\text{o}}_{m}(I^{\circ}_{\rm int})$ take then the same maps which appeared as connecting maps in the system $\Psi^{\text{o}}_{m}(I^{\circ}_{\rm int})$.
As in Section 2.6, we associate to the tree system $\overline{\Psi}:=\overline{\Psi}^{\text{o}}_{m}(I^{\circ}_{\rm int})$ the inverse system ${\mathcal{S}}_{\overline{\Psi}}$ such that, by Proposition 2.8, $\lim_{\leftarrow}{\mathcal{S}}_{\overline{\Psi}}\cong\lim\overline{\Psi}$. Now, since all the spaces in the inverse system ${\mathcal{S}}_{\overline{\Psi}}$ are clearly homeomorphic to the circle $S^{1}$, and since all bonding maps in this system consist of shrinking finitely many pairwise disjoint closed subsegments of $S^{1}$ to points, these maps are obviously near homeomorphisms, i.e., they can be approximated by homeomorphisms. Since, by a result of M. Brown ([Bro60, Theorem 4]), the inverse limit maps in an inverse system consisting of near homeomorphisms are near homeomorphisms, it follows that the inverse limit $\lim_{\leftarrow}{\mathcal{S}}_{\overline{\Psi}}$ is homeomorphic to $S^{1}$, and thus we also have $\lim\overline{\Psi}\cong S^{1}$.
The segments $I_{t}$ from the constituent spaces $\overline{K}_{t}$ obviously become subsegments in the limit circle $\lim\overline{\Psi}$, and moreover they are easily seen to be pairwise disjoint. A slightly less clear fact, the straightforward proof of which we omit, is that the union of the segments $I_{t}$ (viewed as subsegments in the circle $\lim\overline{\Psi}$) is dense in $\lim\overline{\Psi}$. On the other hand, the limit $\lim\Psi$ of the original tree system $\Psi=\Psi^{\text{o}}_{m}(I^{\circ}_{\rm int})$ can be identified with a subspace in the limit $\lim\overline{\Psi}$, namely the subspace obtained by deleting from $\lim\overline{\Psi}$ the interiors of all subsegments $I_{t}$. Moreover, the peripherals in $\lim\Psi$ correspond to the pairs of endpoints of the deleted subsegments $I_{t}$. Consequently, as a space with peripherals, $\lim\Psi$ is homeomorphic to the punctured circle, as required.
It remains to check the conditions related to the types and orientations of the peripherals. The condition which requires that the orientations of all peripherals in the limit space $\lim\Psi$ are consistent (as in condition (b) in Example 11.3) follows directly from condition (c) above in this subsection. The condition which requires that the union of the peripherals of any fixed type is dense follows easily from condition (d) above. This completes the proof. ∎
12.2. Non-orientable trees of segments
Fix any non-negative integers $m,n$ such that $m+n\geq 1$, and let ${\mathcal{T}}={\mathcal{T}}_{1}\sqcup{\mathcal{T}}_{2}$ be a set of two kinds of types, with cardinalities $|{\mathcal{T}}_{1}|=m$ and $|{\mathcal{T}}_{2}|=n$. For each type $\tau\in{\mathcal{T}}$, denote by $I^{\circ}_{\tau}$ the internally punctured segment $I^{\circ}_{\rm int}$, as described in Example 7.2, equipped with the structure consisting of the following data:
  1. (1)
    as in the orientable case, we consider the ordinary peripherals and the special peripheral in $I^{\circ}_{\rm int}$, and we equip the special peripheral with the type $\tau$ (and we do not associate any types to the ordinary peripherals);
  2. (2)
    for each type $\tau\in{\mathcal{T}}_{1}$, we equip the special peripheral of $I^{\circ}_{\tau}$ with one of the orientations, and for each type $\tau\in{\mathcal{T}}_{2}$ we associate to the special peripheral of $I^{\circ}_{\tau}$ both of its orientations.
Define the tree system $\Psi^{\text{n-o}}_{m,n}(I^{\circ}_{\rm int})$ by requiring the following (the superscript “n-o” on $\Psi^{\text{n-o}}_{m,n}$ indicates “non-orientable”):
  1. (a)
    its constituent spaces are the copies of the internally punctured segments $I_{\tau}^{\circ}:\tau\in{\mathcal{T}}$, with ordinary peripherals as the peripherals of these spaces;
  2. (b)
    each gluing in the system involves precisely two constituent spaces (equivalently, each white vertex of the underlying tree of the system has degree 2);
  3. (c)
    for each type $\tau\in{\mathcal{T}}$, and each type $\tau^{\prime}\in{\mathcal{T}}_{2}$, for any constituent space $K_{t}$ being a copy of $I^{\circ}_{\tau}$, the union of those peripherals of $K_{t}$ along which $K_{t}$ is glued to a copy of $I^{\circ}_{\tau^{\prime}}$ is dense in $K_{t}$;
  4. (d)
    for each type $\tau\in{\mathcal{T}}$, and each type $\tau^{\prime}\in{\mathcal{T}}_{1}$, any constituent space $K_{t}$ being a copy of $I^{\circ}_{\tau}$, after fixing an auxiliary orientation, satisfies the following condition:
      consider any peripheral $P$ of $K_{t}$ along which $K_{t}$ is glued to a copy of $I^{\circ}_{\tau^{\prime}}$; then this copy of $I^{\circ}_{\tau^{\prime}}$ inherits an orientation from its special peripheral, and thus induces one of the orientations on the peripheral $P$; we will call the latter orientation of $P$ the induced orientation; we require that the union of those peripherals of $K_{t}$ along which $K_{t}$ is glued to a copy of $I^{\circ}_{\tau^{\prime}}$, and whose induced orientations are compatible (respectively, incompatible) with the earlier fixed orientation of $K_{t}$, is dense in $K_{t}$.
Similarly as we have mentioned in the previous subsection, the fact that a tree system as above is unique up to an isomorphism follows by discussion and arguments similar to those in Example 11.3, Lemma 11.4 and Lemma 11.9. However, this fact is not essential for our further considerations.
Lemma 12.2.
Consider the limit $X$ of the tree system $\Psi^{\text{n-o}}_{m,n}(I^{\circ}_{\rm int})$ of punctured segments, and view it as a space with peripherals, where peripherals are the images of all special peripherals in all constituent punctured segments. Equip also all those peripherals in a natural way with the types from the set $\mathcal{T}$, and with the naturally induced orientations (one orientation if $\tau\in{\mathcal{T}}_{1}$, and both orientations if $\tau\in{\mathcal{T}}_{2}$). Then $X$ is homeomorphic, as a space with peripherals, to a non-orientable densely decorated punctured circle, as described in Example 11.3, and it has precisely $m$ types of orientable peripherals, and $n$ types of non-orientable peripherals.
The proof of this lemma is similar to the proof of Lemma 12.1, and we omit it.
13. Proof of Theorem 10.1 and of the missing parts of Theorems A and B
The major part of this section is devoted to the proof of Theorem 10.1. After doing this, we also provide arguments for getting (as rather immediate corollaries to Theorem 10.1) the proofs of the implications (3$\Rightarrow$(2) in Theorem A and (3$\Rightarrow$(2) in Theorem B.
The outline of the proof of Theorem 10.1 is as follows. Given a group $G$ as in the assumption of Theorem 10.1, we have described two tree systems canonically associated to $G$. The first one is the tree system $\Theta^{r}(G)$ associated to the reduced Bowditch JSJ splitting of $G$, as described in Section 3.2, and it has the property that $\lim\Theta^{r}(G)\cong\partial G$ (see Theorem 3.8). The second tree system associated to $G$ is the (regular) tree system of punctured graphs $\Theta[{\mathcal{R}}_{G}]$ associated to the graphical connecting system ${\mathcal{R}}_{G}$, as described in Section 4 (together with the description of ${\mathcal{R}}_{G}$ given in Section 10), and as appearing in the statement of Theorem 10.1. In view of the equality $\partial G\cong\lim\Theta^{r}(G)$, our goal is to show that $\lim\Theta^{r}(G)\cong\lim\Theta[{\mathcal{R}}_{G}]$.
In order to show the latter equality, we shall describe three more tree systems, denoted $\Theta^{V}[{\mathcal{R}}_{G}]$, $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ and $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$, with the following properties. The tree system $\Theta^{V}[{\mathcal{R}}_{G}]$, called the tree system of V-trees associated to $G$, will be described as obtained from $\Theta[{\mathcal{R}}_{G}]$ via a certain consolidation (an operation described in Section 2.7, which does not affect the limit space of the system). The constituent spaces of $\Theta^{V}[{\mathcal{R}}_{G}]$ will be the V-trees of graphs given as limits of certain V-tree systems of graphs naturally appearing inside the system $\Theta[{\mathcal{R}}_{G}]$. The tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$, called the tree system of cores and arms associated to $G$, will be described as obtained from $\Theta^{V}[{\mathcal{R}}_{G}]$ by an operation of splitting (i.e. operation opposite to some operation of consolidation) consisting of appropriate splittings of all V-tree constituent spaces of $\Theta^{V}[{\mathcal{R}}_{G}]$ into cores and arms (as described in Section 7). Finally, the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$, called the tree system of cores and consolidated arms associated to $G$, will be described as obtained from $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ via one more operation of consolidation, which will join together certain configurations of arms in the latter system into bigger constituent spaces which have the form of punctured circles (here we will refer to the results of Section 12).
By Lemma 7.12 and Theorem 2.9, we will then have the natural identifications
$\lim\Theta[{\mathcal{R}}_{G}]\cong\lim\Theta^{V}[{\mathcal{R}}_{G}]\cong\lim\Theta^{V}_{*}[{\mathcal{R}}_{G}]\cong\lim\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}.$
As the last step in the proof of Theorem 10.1, we will show that the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ is isomorphic to the tree system $\Theta^{r}(G)$, thus yielding the desired assertion by referring to the fact that isomorphic tree systems have homeomorphic limits. The existence of an isomorphism between the two just mentioned tree systems will be deduced from Lemma 11.9, by showing that the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ is modelled on the same connecting system (for regular spaces with doubleton peripherals) as the tree system $\Theta^{r}(G)$, namely on the system ${\mathcal{P}}^{r}_{G}$ described in Example 11.10.
We now pass to the details of the above outlined proof. We split the exposition into subsections which correspond to the steps of this proof.
13.1. Description of $\Theta^{V}[{\mathcal{R}}_{G}]$ the tree system of V-trees associated to $G$
Recall that the tree system $\Theta[{\mathcal{R}}_{G}]$ is a tree system of punctured graphs, with
$\Theta[{\mathcal{R}}_{G}]=(T,\{\Gamma^{\circ}_{t}\},\{\Sigma_{u}\},\{\varphi_{e}\}),$
where each graph $\Gamma_{t}$ is either the extended Whitehead graph associated to some rigid cluster factor of $G$ or some $\theta$-graph. Recall also that the set of white vertices $u$ of the underlying tree $T$ is partitioned into two subsets according to whether the corresponding subspaces $\varphi_{e}(\Sigma_{u})$ are V-peripherals or E-peripherals (compare the second paragraph after Definition 6.1). View the tree $T$ as the first barycentric subdivision of another tree $T^{\bullet}$, so that black vertices of $T$ are the vertices of $T^{\bullet}$ and the white vertices of $T$ are the barycenters of the edges of $T^{\bullet}$. Then, by what was said above about white vertices of $T$, we can speak of V-edges and E-edges of the tree $T^{\bullet}$. Note also that, under this interpretation of $T$, the b-subtrees of $T$ correspond to arbitrary subtrees of $T^{\bullet}$.
Call any subtree of the tree $T^{\bullet}$ whose edge set consists only of V-edges a V-subtree of $T^{\bullet}$. Let $\Pi^{V}$ be the partition of the tree $T$ (in the sense explained in Section 2.7) into b-subtrees which correspond to all maximal V-subtrees of $T^{\bullet}$. Take as $\Theta^{V}[{\mathcal{R}}_{G}]$ the tree system obtained from $\Theta[{\mathcal{R}}_{G}]$ by consolidation with respect to the partition $\Pi^{V}$, i.e.,
$\Theta^{V}[{\mathcal{R}}_{G}]:=(\Theta[{\mathcal{R}}_{G}])^{\Pi^{V}}.$
We will now describe in more detail the constituent spaces, the peripherals, and the connecting maps of the tree system $\Theta^{V}[{\mathcal{R}}_{G}]$. Suppose first that a b-subtree $S\in\Pi^{V}$ contains a (black) vertex $t$ for which $\Gamma_{t}$ is the extended Whitehead graph, say $\widetilde{W}_{z}$, associated to some virtually free rigid cluster factor $G(z)$ of $G$. It follows from the form of the V-involution in the graphical connecting system ${\mathcal{R}}_{G}$ that the restriction of the system $\Theta[{\mathcal{R}}_{G}]$ to $S$ is then isomorphic (as a tree system) to the V-tree system $\Theta[\widetilde{\mathcal{V}}_{z}]$ of the extended Whitehead graph $\widetilde{W}_{z}$. In particular, for any black vertex $s$ of $S$ the corresponding graph $\Gamma_{s}$ is isomorphic to $\widetilde{W}_{z}$, and the constituent space $K_{S}$ of the consolidated system $\Theta^{V}[{\mathcal{R}}_{G}]$ is homeomorphic (as a space with peripherals) to the V-tree ${\mathcal{X}}_{V}((\widetilde{W}_{z},\tilde{\mathbf{a}}_{z}))$ (where $\tilde{\mathbf{a}}_{z}$ is the V-involution canonically associated to $\widetilde{W}_{z}$).
Consider now the remaining case when a b-subtree $S\in\Pi^{V}$ contains a (black) vertex $t$ for which $\Gamma_{t}$ is some $\theta$-graph $\Xi_{k}$. It follows from the form of the V-involution in the graphical connecting system ${\mathcal{R}}_{G}$ that the restriction of the system $\Theta[{\mathcal{R}}_{G}]$ to $S$ is then isomorphic (as a tree system) to the V-tree system $\Theta[{\mathcal{V}}_{\Xi_{k}}]$ induced by the connecting V-system ${\mathcal{V}}_{\Xi_{k}}$ (i.e. the standard connecting V-system for $\Xi_{k}$, as described in Example 7.11). In particular, for any black vertex $s$ of $S$ the corresponding graph $\Gamma_{s}$ is isomorhic to $\Xi_{k}$, and the constituent space $K_{S}$ of the consolidated system $\Theta^{V}[{\mathcal{R}}_{G}]$ is homeomorphic (as a space with peripherals) to the internally punctured graph $\Xi_{k}$, i.e., to the space with peripherals $(\Xi_{k})^{\circ}_{int}$ (as in Example 7.11).
Recall again that we have two kinds of peripherals in the initial tree system $\Theta[{\mathcal{R}}_{G}]$, namely V-peripherals and E-peripherals. All V-peripherals disappear in the process of passing, via consolidation, to the consolidated tree system $\Theta^{V}[{\mathcal{R}}_{G}]$, while all E-peripherals survive this process. Thus, we can naturally identify the peripherals of the system $\Theta^{V}[{\mathcal{R}}_{G}]$ with the E-peripherals of the system $\Theta[{\mathcal{R}}_{G}]$. Under this identification, the connecting maps in the system $\Theta^{V}[{\mathcal{R}}_{G}]$ are induced in the obvious way from those maps in the system $\Theta[{\mathcal{R}}_{G}]$ which concern E-peripherals.
13.2. Description of $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ the tree system of cores and arms associated to $G$
Recall that if $S\in\Pi^{V}$ is a subtree such that for its black vertices $t$ the graphs $\Gamma_{t}$ are the extended Whitehead graphs, then the constituent space $K_{S}$ of the system $\Theta^{V}[{\mathcal{R}}_{G}]$ has the form of a V-tree of extended Whitehead graphs ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{z})={\mathcal{X}}_{V}((\widetilde{W}_{z},\tilde{\mathbf{a}}_{z}))$, and that this V-tree can be expressed as the limit of the tree system $\Psi_{S}:=\Psi[\widetilde{\mathcal{V}}_{z}]$ in which the constituent spaces are the core and the arms of ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{z})$, see Lemma 9.6(2) and Lemma 7.12. Fact 7.7 and Lemma 9.6(1) explain then the nature of the core and the arms appearing in this expression. In the remaining case when $S\in\Pi^{V}$ is a subtree with black vertex graphs $\Gamma_{t}$ isomorphic to a $\theta$-graph $\Xi_{k}$ (for some $k$ depending only on $S$), the constituent space $K_{S}$ coincides with the V-tree of $\theta$-graphs ${\mathcal{X}}_{V}({\mathcal{V}}_{\Xi_{k}})$, as described in Example 7.11. In particular, in this case the constituent space $K_{S}$ is isomorphic, as a space with peripherals, to the internally punctured $\theta$-graph $(\Xi_{k})^{\circ}_{int}$. Consequently, this space $K_{S}$ can be expressed as the wedge of ($k$ copies of) punctured segments, $K_{S}=\lim\Psi_{w}[{\mathcal{V}}_{\Xi_{k}}]$, as described in Example 7.13. For $S$ as in this case, we put $\Psi_{S}=\Psi_{w}[{\mathcal{V}}_{\Xi_{k}}]$, and we recall that this system $\Psi_{S}$ consists of precisely one white vertex, $u_{S}$, having degree $k$ in the corresponding underlying tree, and of precisely $k$ black vertices, all adjacent to $u_{S}$, the constituent spaces at which are punctured segments. We call these constituent spaces the wedge arms, and we call their common peripheral, which is a doubleton, the wedge core.
Since for any $S\in\Pi^{V}$, each peripheral of $K_{S}$ is contained in precisely one arm (or wedge arm), and disjoint with the core (respectively, wedge core), we can naturally view the tree system $\Theta^{V}[{\mathcal{R}}_{G}]$ as obtained by consolidation from another tree system, denoted $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$, for which the constituent spaces are the cores, the arms and the wedge arms in all spaces $K_{S}:S\in\Pi^{V}$. The above described systems $\Psi_{S}$ correspond then to the restricted subsystems of such a new system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ along which one consolidates the latter to obtain back the system $\Theta^{V}[{\mathcal{R}}_{G}]$. We skip the straightforward further details of this construction.
Note that all peripherals of the system $\Theta^{V}[{\mathcal{R}}_{G}]$ (which we naturally identify with the E-peripherals of the initial system $\Theta[{\mathcal{R}}_{G}]$) can be naturally identified as some of the peripherals in the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$. Thus, we will likewise refer to the peripherals in $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ of the latter form as E-peripherals. Obviously, there are also some other peripherals in $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$, and they correspond precisely to the peripherals in the cores and to the special peripherals in the arms and wedge arms. We will call the peripherals in $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ of the latter forms C-peripherals.
13.3. U-types, Y-labels and ##\##-orientations of the arms and wedge arms of $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$
Before getting to the further modifications of the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$, we need to describe some natural decorations for arms and wedge arms of this system, and to record some of their properties. We start with describing the so called U-types.
Let $U$ be the set of $G$-orbits in the set of all those edges of the reduced Bowditch JSJ-tree $T^{r}_{G}$ which have a flexible vertex as one of their endpoints. We will use $U$ as a set of types for arms and wedge arms in $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$, and we will say that a type $\varepsilon\in U$ is orientable (respectively, non-orientable) if the edge stabilizing group $G(e)$ for any $e\in\varepsilon$ is orientable (respectively, non-orientable). Now, let $A$ be any arm in the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$. Recall that $A$ corresponds to a class (or a sequence of copies) of edges in one of the extended Whitehead graphs $\Gamma_{z}=\widetilde{W}_{z}$, and the set of edges in $\widetilde{W}_{z}$ whose copies are involved in $A$ (which we denote by $E_{A}$) coincides with the preimage $(\lambda_{z}^{+})^{-1}(x)$ for some (precisely one) vertex $x\in N_{z}^{2}$ (see the notation introduced in the fourth paragraph of Section 10). The vertex $x$ as above is adjacent to precisely one vertex $u_{\xi}$ (as described in the same Section 10), and we put $e=[x,u_{\xi}]$. We associate to $A$ (and call the U-type of $A$) the type $\varepsilon=[e]\in U$, i.e., the $G$-orbit of $e$. This orbit clearly belongs to $U$, since $x$ is a flexible vertex of $T^{r}_{G}$ (which follows from the fact that rigid cluster vertices in the tree $T^{r}_{G}$ are isolated, see Fact 3.5). Consider now the remaining case when $A$ is a wedge arm in $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$, and recall that $A$ is composed of punctured edges obtained by puncturing precisely one edge $\sigma$ in precisely one $\theta$-graph $\Gamma_{z}=\Xi_{k(z)}$. (We then put $E_{A}:=\{\sigma\}$.) Since the edges of $\Gamma_{z}$ are in a fixed bijective correspondence with the vertices of $T^{r}_{G}$ adjacent to $z$ (via the bijection $\lambda_{z}^{+}$ described in the fourth paragraph of Section 10), we denote by $e_{\sigma}=[z,\lambda_{z}^{+}(\sigma)]$ this edge joining $z$ to the vertex which corresponds to $\sigma$. We associate to $A$ (and call the U-type of A) the type $\varepsilon=[e_{\sigma}]\in U$, i.e., the $G$-orbit of $e_{\sigma}$.
Accordingly with the above described assignment of U-types, we call an arm or wedge arm U-orientable if its U-type is orientable, and in the remaining case we call it U-non-orientable.
If some U-type $\varepsilon\in U$ is orientable, its orientation is a choice of an orientation for any 2-ended group $G(e)$ with $e\in\varepsilon$ (and such a choice extends, via conjugations in $G$, to the choices of orientations for all groups $G(e)$ with $e\in\varepsilon$). We will now explain how the choice of an orientation of a U-type induces some orientation on any arm or wedge arm equipped with this U-type. If $\varepsilon$ is the U-type of some arm $A$, there is precisely one rigid cluster vertex $z\in Z$ such that $A$ is an arm in a V-tree of extended Whitehead graphs $\Gamma_{z}$. Moreover, there is precisely one $x\in N^{2}_{z}$ such that $A$ is composed of copies of edges from $(\lambda_{z}^{+})^{-1}(x)$, and if $u_{\xi}$ denotes the white vertex on the path from $x$ to $z$ then $[u_{\xi},x]\in\varepsilon$. If we fix an orientation of $\varepsilon$, it induces an orientation of the group $G([u_{\xi},x])$, and this induces an order of the doubleton $P=\partial G([u_{\xi},x])=\partial G(u_{\xi})=\partial G([u_{\xi},z])\subset\partial G(z)=C_{z}$, where $C_{z}$ is the core of the V-tree ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{z})$ (for the last equality, compare Lemma 9.6). Note that $P$ coincides with the set of endpoints of precisely one arm $A^{\prime}$ in the V-tree ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{z})$ that is made of punctured copies of the same edges of $\widetilde{W}_{z}$ as the arm $A$. Moreover, the above order of $P$ uniquely determines the order-induced orientations of all edges of $\widetilde{W}_{z}$ whose copies are involved in $A^{\prime}$ (compare Remark/Definition 7.16, Lemma 8.18 and Remark 9.10(3)). Consequently, the above order-induced orientations of the edges uniquely induce an orientation of the arm $A$, and we will call it the orientation of $A$ induced from the orientation of $\varepsilon$. Obviously, any orientation of $\varepsilon$ uniquely induces orientations of all peripherals in any arm $A$ of the U-type $\varepsilon$.
Now, suppose that $\varepsilon$ is the U-type of some wedge arm $A$ of the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$, and that this type is orientable, with one of the orientations fixed. Let $z\in Z$ be the (unique) vertex of type (v2) for which $A$ is an arm in a V-tree of $\theta$-graphs $\Gamma_{z}=\Xi_{k(z)}$. Then there is precisely one $x\in N_{z}$ such that $A$ is composed of copies of the edge $(\lambda_{z}^{+})^{-1}(x)$. The fixed orientation of the U-type $\varepsilon$ induces then an orientation of the group $G([z,x])$, and this induces an order on the doubleton $\partial G([z,x])=\partial G(z)$. By our convention from Section 10, we have fixed an identification of the set $\partial G(z)$ with the vertex set of the $\theta$-graph $\Gamma_{z}=\Xi_{k(z)}$, so the above orientation of $\varepsilon$ induces an order on the vertex set of $\Gamma_{z}$. This in turn induces, in the obvious way, an orientation of the edge $(\lambda_{z}^{+})^{-1}(x)$ in $\Gamma_{z}$, and since all punctured copies of this edge appearing inside the arm $A$ are oriented consistently, this induces also an orientation of $A$.
We now pass to describing the Y-labels associated to arms and wedge arms. To describe them, we need to look closer at the set $\mathcal{A}$ of E-connections in the graphical connecting system ${\mathcal{R}}_{G}$. Recall that $\mathcal{A}$ is given as the disjoint union of the blocks ${\mathcal{A}}_{y}:y\in Y$, where $Y$ is the set of $G$-orbits in the set of flexible vertices of $T^{r}_{G}$. Each block ${\mathcal{A}}_{y}$ is described in terms of some set $E_{y}$, where the family $E_{y}:y\in Y$ is a certain partition of the set of edges of $\Gamma$ (see Section 10). It follows from the descriptions of the blocks ${\mathcal{A}}_{y}$ and the sets $E_{y}$ (in Section 10) that for any arm or wedge arm $A$ the corresponding set $E_{A}$ is contained in precisely one of the sets $E_{y}$; we then say that $y$ is the Y-label associated to the arm $A$. Observe that any two arms (or wedge arms) with the same U-type also have the same Y-label, so that the partition of the set of arms according to the U-types is a refinement of the partition according to the Y-labels. Moreover, it is not hard to see that for any $y\in Y$ the set
$U_{y}:=\{\varepsilon\in U:\hbox{ $y$ is the Y-label of any arm or wedge arm having U-type $\varepsilon$}\}$
coincides with the set of $G$-orbits of all those edges of $T^{r}_{G}$ which are adjacent to any fixed vertex $v\in y$.
Now, consider any label $y\in Y$ such that the flexible factors $G(v):v\in y$ (in the reduced Bowditch JSJ splitting of $G$) are orientable. Then to any arm or wedge arm labelled with $y$ we associate the so called ##\##-orientation in the following way. Recall that in our description of the connecting system ${\mathcal{R}}_{G}$ as given in Section 10, we fixed in advance, in a $G$-invariant way, orientations for all flexible factors $G(v)$ with $v\in y$, and that this choice of orientations led to the assignment of an orientation to any edge $\varepsilon\in E_{y}$, yielding a collection of oriented edges $E_{y}^{\#}$. Note that for any arm (or wedge arm) $A$ labelled with $y$, the copies of edges from $E_{y}$ appearing in this arm, viewed with the orientations as in $E_{y}^{\#}$, are oriented consistently. Consequently, these orientations induce an orientation of the entire arm $A$. Accordingly, we will denote by $A^{\#}$ the arm $A$ equipped with the just described orientation, and we will call this orientation the ##\##-orientation of $A$. We will then call any arm or wedge arm $A$ as above ##\##-oriented. The arms or wedge arms not equipped with ##\##-orientations (i.e. those having the Y-label $y$ such that the groups $G(v):v\in y$ are non-orientable, as flexible factors) will be called ##\##-non-orientable.
The next observation follows directly from the description of the connecting system ${\mathcal{R}}_{G}$, and from the above descriptions of U-types, Y-labels and ##\##-orientations. To justify part (3), one also needs to refer to the density condition (4) in Definition 6.2, as applied to the tree system $\Theta[{\mathcal{R}}_{G}]$ (see also condition (L3) in the construction from the proof of Lemma 6.3). In part (3) we use the following terminological convention: under notation as in the whole of Fact 13.1, given a constituent space $K^{*}_{t}$ of $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ which is an arm or a wedge arm, and one of its E-peripherals $P=\varphi^{*}_{[u,t]}(\Sigma^{*}_{u})$, the peripheral adjacent to $P$ is the peripheral $P^{\prime}=\varphi^{*}_{[u,t^{\prime}]}(\Sigma^{*}_{u})$ in the constituent space $K^{*}_{t^{\prime}}$, where $t^{\prime}$ is the unique black vertex of $T_{*}$ adjacent to $u$ and distinct from $t$.
Fact 13.1.
Denote the data of the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ by $(T_{*},\{K_{t}^{*}\},\{\Sigma_{u}^{*}\},\{\varphi_{e}^{*}\})$. Let $A=K_{t}^{*}$ and $\hat{A}=K_{\hat{t}}^{*}$ be some two constituent spaces in $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$, each of which is an arm or a wedge arm, and suppose that there is a white vertex $u$ in $T_{*}$ which is adjacent to both $t$ and $\hat{t}$.
  1. (1)
    If the Y-label of $A$ is $y$, then the Y-label of $\hat{A}$ is also $y$.
  2. (2)
    If $A$ is ##\##-oriented, then $\hat{A}$ is also ##\##-oriented. Moreover, if we induce the ##\##-orientations of $A$ and $\hat{A}$ to the peripherals $\varphi^{*}_{[u,t]}(\Sigma^{*}_{u})\subset A$ and $\varphi^{*}_{[u,\hat{t}]}(\Sigma^{*}_{u})\subset\hat{A}$, then the gluing map $\varphi^{*}_{[u,\hat{t}]}(\varphi^{*}_{[u,t]})^{-1}$ between those peripherals respects the ##\##-orientations, i.e., it reverses the induced orientations of the peripherals.
  3. (3)
    If the Y-label of $A$ is $y$, and if $\varepsilon$ is any U-type from $U_{y}$, then:
    1. (A)
      the union of all those E-peripherals $P$ of $A$ for which the peripheral $P^{\prime}$ adjacent to $P$ is contained in an arm or a wedge arm of U-type $\varepsilon$, is dense in $A$;
    2. (B)
      moreover, if $A$ is ##\##-non-orientable and $\varepsilon$ is orientable then, after fixing an auxiliary orientation of $A$ and some orientation for $\varepsilon$, the union of all those E-peripherals $P$ of $A$ for which the peripheral $P^{\prime}$ adjacent to $P$ is contained in an arm or wedge arm $A^{\prime}$ of U-type $\varepsilon$, and for which the orientation induced from the fixed orientation of $\varepsilon$ (via the induced orientations of $A^{\prime}$ and $P^{\prime}$) is compatible (respectively, incompatible) with the fixed orientation of $A$, is dense in $A$.
We finish this subsection with another two observations concerning the concepts just introduced, which describe some features of the C-peripherals of the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$. These observations follow rather easily from Facts 10.2 and 10.3, respectively, and we omit their proofs.
Fact 13.2.
Under the notation $\Theta^{V}_{*}[{\mathcal{R}}_{G}]=(T_{*},\{K_{t}^{*}\},\{\Sigma_{u}^{*}\},\{\varphi_{e}^{*}\})$, let $t$ be a vertex of the tree $T_{*}$ such that $K^{*}_{t}$ is a constituent space which is a copy of the core $C_{z}$ of the V-tree of extended Whitehead graphs $\widetilde{W}_{z}$, for some rigid cluster vertex $z\in Z$. Let $u$ be any white vertex adjacent to $t$, and denote by $\xi\in{\mathcal{H}}_{z}$ the type of the peripheral $\varphi^{*}_{[u,t]}(\Sigma^{*}_{u})$ (as introduced in Definition 9.7).
  1. (0)
    The degree of the vertex $u$ in $T_{*}$ is equal to $n_{\xi}$, and there is a natural bijective correspondence $b:N_{u}^{T_{*}}\to N_{u_{\xi}}^{T^{r}_{G}}$ between the sets of vertices adjacent to $u$ and $u_{\xi}$ in the corresponding trees, such that $b(t)=z$ and for each arm $K^{*}_{s}$ with $s\in N_{u}^{T_{*}}\setminus\{t\}$ the U-type of $K_{s}^{*}$ is equal to the $G$-orbit of the edge $[u_{\xi},b(s)]$.
  2. (1)
    For any $s\in N_{u}^{T_{*}}\setminus\{t\}$, denoting by $\varepsilon$ the U-type of the arm $K_{s}^{*}$, the cardinality of the set
    $\{s^{\prime}\in N_{u}^{T_{*}}\setminus\{t\}:\varepsilon\hbox{ is the U-type ofthe arm }K^{*}_{s^{\prime}}\}$
    coincides with the index $[G(u_{\xi}):G([u_{\xi},b(s)])]=[G(u_{\xi}):G(u_{\xi})\cap G(b(s))]$.
  3. (2)
    Given $s\in N_{u}^{T_{*}}\setminus\{t\}$, consider the subset $[s]\subset N_{u}^{T_{*}}\setminus\{t\}$ corresponding via the bijection $b$ to the $G(u_{\xi})$-orbit $[b(s)]$ of the vertex $b(s)\in N^{T^{r}_{G}}_{u_{\xi}}\setminus\{z\}$. (Equivalently, in accordance with the above part (0), $[s]$ is precisely the set of those $s^{\prime}\in N_{u}^{T_{*}}\setminus\{t\}$ for which the arm $K_{s^{\prime}}^{*}$ has the same U-label as $K^{*}_{s}$.) Suppose also that the group $G(b(s))$ (as well as $G(b(s^{\prime}))$ for any s[$s^{\prime}\in[s]$) is orientable, so that all arms $K^{*}_{s^{\prime}}:s^{\prime}\in[s]$ are ##\##-oriented. Then:
    1. (A)
      if $G(u_{\xi})$ is orientable, then the ##\##-orientations of all arms $K^{*}_{s^{\prime}}:s^{\prime}\in[s]$ are compatible, i.e., they all induce (via the maps $(\varphi_{[s,u]}^{*})^{-1}$) the same order on the doubleton $\Sigma^{*}_{u}$;
    2. (B)
      if $G(u_{\xi})$ is non-orientable, then the cardinality of $[s]$ is even, and the ##\##-orientations of half of the arms $K^{*}_{s^{\prime}}:s^{\prime}\in[s]$ are compatible with one of the orders of the doubleton $\Sigma_{u}^{*}$, while the ##\##-orientations of the other half of these arms are compatible with the opposite order of $\Sigma^{*}_{u}$.
Fact 13.3.
Under the notation $\Theta^{V}_{*}[{\mathcal{R}}_{G}]=(T_{*},\{K_{t}^{*}\},\{\Sigma_{u}^{*}\},\{\varphi_{e}^{*}\})$, let $u$ be a vertex of the tree $T_{*}$ such that $\Sigma^{*}_{u}$ is a peripheral space of the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ which is a copy of the wedge core of the V-tree of the $\theta$-graphs $\Gamma_{z}=\Xi_{k(z)}$, for some $z\in Z$ of type (v2).
  1. (0)
    The degree of $u$ in $T_{*}$ coincides with the degree $k(z)$ of $z$ in $T^{r}_{G}$, and there is a natural bijective correspondence $b:N_{u}^{T_{*}}\to N_{z}^{T^{r}_{G}}$ such that for each wedge arm $K_{s}^{*}$ with $s\in N_{u}^{T_{*}}$ the U-type of $K_{s}^{*}$ is equal to the $G$-orbit of the edge $[z,b(s)]$.
  2. (1)
    For any $s\in N_{u}^{T_{*}}$, denoting by $\varepsilon$ the U-type of the arm $K_{s}^{*}$, the cardinality of the set
    $\{s^{\prime}\in N_{u}^{T_{*}}:\varepsilon\hbox{ is the U-type of the wedge arm}K^{*}_{s^{\prime}}\}$
    coincides with the index $[G(z):G([z,b(s)])]=[G(z):G(z)\cap G(b(s))]$.
  3. (2)
    Given $s\in N_{u}^{T_{*}}\setminus\{t\}$, consider the subset $[s]\subset N_{u}^{T_{*}}$ corresponding via the bijection $b$ to the $G(z)$-orbit $[b(s)]$ of the vertex $b(s)\in N^{T^{r}_{G}}_{z}$. (Equivalently, in accordance with the above part (0), $[s]$ is precisely the set of those $s^{\prime}\in N_{u}^{T_{*}}$ for which the wedge arm $K_{s^{\prime}}^{*}$ has the same U-label as $K^{*}_{s}$.) Suppose also that the group $G(b(s))$ (as well as $G(b(s^{\prime}))$ for any s[$s^{\prime}\in[s]$) is orientable, so that all wedge arms $K^{*}_{s^{\prime}}:s^{\prime}\in[s]$ are ##\##-oriented. Then:
    1. (A)
      if $G(z)$ is orientable, then the ##\##-orientations of all arms $K^{*}_{s^{\prime}}:s^{\prime}\in[s]$ are compatible, i.e., they all induce (via the maps $(\varphi_{[s,u]}^{*})^{-1}$) the same order on the doubleton $\Sigma^{*}_{u}$;
    2. (B)
      if $G(z)$ is non-orientable, then the cardinality of $[s]$ is even, and the ##\##-orientations of half of the arms $K^{*}_{s^{\prime}}:s^{\prime}\in[s]$ are compatible with one of the orders of the doubleton $\Sigma_{u}^{*}$, while the ##\##-orientations of the other half of these arms are compatible with the opposite order of $\Sigma^{*}_{u}$.
13.4. Consolidation $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ of the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ along arm-subtrees
In this step of the proof, we describe some consolidation of the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$. The proof of Theorem 10.1 will be then concluded by showing, in the next step, that the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ obtained by this consolidation is isomorphic to the tree system $\Theta^{r}(G)$.
Recall that we denote the underlying tree of the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ by $T_{*}$. Let $\Pi^{a}$ be the partition of $T_{*}$ into b-subtrees described as follows. The first kind of b-subtrees in $\Pi^{a}$ are the trivial subtrees $R$ which consist of a single “core vertex”, i.e., a vertex $r$ with one of the cores (of a V-tree of extended Whitehead graphs) as the constituent space $K_{r}^{*}$. The b-subtrees in $\Pi^{a}$ of the second kind are described as follows. Cut $T_{*}$ along all core vertices (as mentioned above), and along all wedge core vertices (which are white), and consider the b-subtrees $R$ spanned on all black vertices in any of the connected components obtained by cutting $T_{*}$ in this way. Note that for such subtrees $R$ all constituent spaces $K_{r}^{*}$ at the black vertices $r$ in $R$ are arms (in some V-trees of exteneded Whitehead graphs) or wedge arms. For this reason the b-subtrees $R\in\Pi^{a}$ of the latter form will be called arm-subtrees of $T_{*}$. Define the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ as the consolidation of $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ with respect to the partition $\Pi^{a}$, i.e., as $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}:=(\Theta^{V}_{*}[{\mathcal{R}}_{G}])^{\Pi^{a}}$.
An alternative way of describing the arm-subtrees $R$ of $T_{*}$ is by saying that they are the maximal b-subtrees of $T_{*}$ in which all white vertices correspond to E-peripherals of the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$. It becomes clear from this perspective, that the peripherals of the final tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ correspond precisely to (and are naturally identified with) the C-peripherals of the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$. These C-peripherals can be further split into peripherals of the following two kinds. Peripherals of the first kind are those which appear in the cores (and which are identified via the connecting maps with the special peripherals in the arms, which we view as being of the same kind); we will call them core-peripherals. Peripherals of the second kind correspond to the wedge cores, and appear as the special peripherals in the wedge arms; we will call them wedge-peripherals.
We now pass to describing closer those constituent spaces $K_{R}$ of the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ which correspond to the arm-subtrees $R$ of $T_{*}$. We will show that, after a suitable interpretation of types, the restriction of the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ to any arm-subtree $R$ is isomorphic to a tree system of internally punctured segments of one of the kinds described in Section 12. As a concequence, by referring to Lemma 12.1 or Lemma 12.2, we will identify each of the constituent spaces $K_{R}$ corresponding to arm-subtrees with an appropriate densely decorated punctured circle. More precisely, note first that each arm or wedge arm (viewed as a constituent space in the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$) contains precisely one C-peripheral. Obviously, we will view this arm (or wedge arm) as an internally punctured segment, with its C-peripheral viewed as a special peripheral, and with its E-peripherals viewed as ordinary peripherals. We associate to the special peripheral of any arm (or wedge arm) $A$ the type $\tau$ coinciding with the U-type of $A$ (as described in the previous subsection). Moreover, if the Y-label $y$ associated to $A$ (as in the previous subsection) is orientable (i.e. if the flexible factor $G(v)$, for any $v\in y$, is orientable), we associate to $A$ the orientation which coincides with the ##\##-orientation (as described also in the previous subsection). Observe that, due to Fact 13.1(1), for any arm-subtree $R$, all arms and wedge arms in the restricted tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]|_{R}$ have the same Y-label $y$, which we call therefore the Y-label associated to $R$. It is not hard to observe that the set of U-types of arms and wedge arms appearing in the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]|_{R}$ coincides then with the set $U_{y}$ described in the previous subsection. For any such $R$, we take $U_{y}$ as the corresponding set $\mathcal{T}$ of types. It follows then from Fact 13.1(2)(3) that for any arm-subtree $R$ the restricted tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]|_{R}$ is isomorphic, as equipped with the above mentioned types and orientations, to one of the tree systems of internally punctured segments described in Section 12. More precisely, if $y$ is the Y-label associated to $R$, and if the corresponding flexible factor $G(v)$ for any $v\in y$ is orientable, the restricted tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]|_{R}$ with the above described types and orientations is isomorphic to the tree system $\Psi^{\text{o}}_{m}(I^{\circ}_{\rm int})$, as described in Section 12.1, where $m$ is the cardinality of the subset $U_{y}\subset U$. Similarly, if the corresponding flexible factor $G(v)$ for any $v\in y$ is non-orientable, then the restricted tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]|_{R}$ is isomorphic (via type and orientation respecting isomorphism) with the tree system $\Psi^{\text{n-o}}_{m,n}(I^{\circ}_{\rm int})$, as described in Section 12.2, where $m$ and $n$ are the cardinalities of the subsets $U_{y,1},U_{y,2}$ consisting of those types from $U_{y}$ which are orientable and nonorientable, respectively. This completes our analysis of the constituent spaces $K_{R}$ corresponding to the arm-subtrees $R$ in the consolidated tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$.
13.5. The final step of the proof of Theorem 10.1
We need to show that the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ is isomorphic to the tree system $\Theta^{r}(G)$. We will argue by showing that $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ can be naturally viewed as a regular tree system with doubleton peripherals, as defined in Section 11, and that it is modelled on the same connecting system (for regular spaces with doubleton peripherals) as the system $\Theta^{r}(G)$, namely on the connecting system ${\mathcal{P}}^{r}_{G}$ described in Example 11.10. The proof will be then concluded by referring to Lemmas 11.9 and 2.7.
Let ${\mathcal{P}}^{r}_{G}$ be the connecting system described in Example 11.10, and denote the data which form this system as ${\mathcal{P}}^{r}_{G}=(((K_{j},\Omega_{j}),({\mathcal{D}}_{j},d_{j}))_{j\in J},{\mathcal{B}})$. In view of what was said in the previous paragraph, to conclude the proof of Theorem 10.1 we need to show that the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ is modelled on ${\mathcal{P}}^{r}_{G}$. To do this, we will explicitly describe some ${\mathcal{P}}^{r}_{G}$-atlas ${\mathbf{A}}^{a}$ for $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$. Recall first that we have two kinds of constituent spaces in the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$. The constituent spaces of the first kind are (the copies of) the cores $C_{z}$ of the V-trees ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{z})$ of extended Whitehead graphs $\widetilde{W}_{z}$ (these graphs are described in Section 10, as part of the structure of ${\mathcal{R}}_{G}$, and in Section 9). By Lemma 9.6(1), each such core $C_{z}$ (and hence each constituent space of $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ which is its copy) is canonically homeomorphic to the space with peripherals $(X_{z},{\mathcal{D}}_{z})$ associated to the corresponding rigid cluster factor $(G(z),{\mathcal{H}}_{z})$ of $G$. The corresponding rigid cluster vertex $z$ of $T^{r}_{G}$ is then in the same $G$-orbit as precisely one of the vertices $v_{j}$ mentioned in Example 11.10, and therefore there is $g\in G$ which maps $z$ to $v_{j}$. This element $g$, in its action on the boundary $\partial G$, maps the space with peripherals $\partial G(z)=X_{z}$ onto the analogous space with peripherals $\partial G(v_{j})=K_{j}$, and we identify $C_{z}$ (via $\partial G(z)$ and $\partial G(v_{j})$) with the corresponding $K_{j}$, using this map (and this identification essentially does not depend on a choice of $g$ as above, since any two such identifications differ by a decoration-preserving homeomorphism). Thus, for each constituent space of the first kind we get in this way its identification with an appropriate space $K_{j}$, and we take it as the corresponding element of the atlas ${\mathbf{A}}^{a}$. The constituent spaces of the second kind in the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ are those spaces $K_{R}$ which correspond to the arm-subtrees $R\subset T_{*}$. By the discussion in the previous subsection, each such $K_{R}$ is naturally a densely decorated punctured circle, with types in the set $U_{y}$, where $y\in Y$ is the Y-label associated to $R$. Recalling that $y$ is a $G$-orbit in the set of flexible vertices in $T^{r}_{G}$, we note that the decorated space with peripherals $K_{j}$ with this index $j\in J$ for which $v_{j}\in y$, is also a densely decorated punctured circle. Recalling further that $U_{y}$ is the set of $G$-orbits of the edges of $T^{r}_{G}$ which are adjacent to any fixed vertex $v\in y$, we deduce (by referring to Lemma 11.4(2)(3)) that $K_{R}$ is then isomorphic to $K_{j}$. Moreover, by referring to Lemma 11.4(2)(3), we can (and will) identify $K_{R}$ with the above mentioned $K_{j}$, in an essentially unique way, via a decoration-preserving homeomorphism which sends each peripheral of any type $\varepsilon\in U_{y}$ in $K_{R}$ to a peripheral of type $\tau(\partial G(u^{\prime}))$ in $K_{j}$, where $u^{\prime}$ is any vertex adjacent to $v_{j}$ such that the edge $[u^{\prime},v_{j}]$ belongs to the $G$-orbit $\varepsilon$. Thus, for each constituent space of the second kind in the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$, this gives us an identification with an appropriate space $K_{j}$ of ${\mathcal{P}}^{r}_{G}$, and we take this identification as the corresponding element of the atlas ${\mathbf{A}}^{a}$.
It remains to show that the above described family ${\mathbf{A}}^{a}$ of identifications satisfies conditions (1)-(4) of Definition 11.8 (as applied to the connecting system ${\mathcal{P}}^{r}_{G}$). Fix the notation $(T^{a},\{K^{a}_{\tilde{t}}\},\{\Sigma^{a}_{\tilde{u}}\},\{\varphi^{a}_{\tilde{e}}=\varphi^{a}_{[\tilde{u},\tilde{t}]}\})$ for the data in $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$. We pass to checking that condition (1) of the Definition 11.8 holds. Let $\Sigma^{a}_{\tilde{u}}$ be some peripheral space in the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$. We need to show that the set of types (induced by the above identifications) of the corresponding peripheral subspaces $\varphi^{a}_{[\tilde{u},\tilde{t}]}(\Sigma^{a}_{\tilde{u}})$, where $\tilde{t}$ runs through all vertices adjacent to $\tilde{u}$ in the tree $T^{a}$, coincides with some single block of $\mathcal{B}$. To do this, we consider first the case when $\Sigma^{a}_{\tilde{u}}$ is a core-peripheral (as explained in the previous subsection). Denote by $\tilde{v}$ the (unique) core vertex of $T^{a}$ adjacent to $\tilde{u}$, where by a core vertex we mean here a vertex for which the constituent space $K^{a}_{\tilde{v}}$ is a copy of some core $C_{z}=\partial G(z)$. View then the peripheral $P=\varphi^{a}_{[\tilde{u},\tilde{v}]}(\Sigma^{a}_{\tilde{u}})$ of $K^{a}_{\tilde{v}}$ as a peripheral in the space $\partial G(z)$ (viewed as a space with peripherals associated to the rigid cluster factor $G(z)$), and denote by $u$ the white vertex of $T^{r}_{G}$ for which $P=\partial G(u)$. Denoting by $[u]\in A$ the $G$-orbit of the vertex $u$ (where $A$ is as in Example 11.10), we claim that the set of types of the peripherals of the form $\varphi^{a}_{[\tilde{t},\tilde{u}]}(\Sigma^{a}_{\tilde{u}})$ coincides with the block $B_{[u]}$ of $\mathcal{B}$. To see this, note that if we interpret the types of the peripherals in ${\mathcal{P}}^{r}_{G}$ as $G$-orbits of edges, as in Observation 11.11, then the above described identifications from the atlas ${\mathbf{A}}^{a}$ simply preserve those types. Moreover, since the peripherals in the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ correspond to the C-peripherals of the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$, we may view the observations from Facts 13.2 and 13.3, after appropriate reinterpretation, as concerning the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$. Our claim follows then fairly directly from Fact 13.2(0). We pass now to considering the remaining case when $\Sigma^{a}_{\tilde{u}}$ is a wedge-peripheral, corresponding to some wedge core. This wedge core is contained in a V-tree of copies of precisely one $\theta$-graph $\Gamma_{z}$ from the connecting system ${\mathcal{R}}_{G}$, and the vertex $z$ is white in $T^{r}_{G}$ and not adjacent to any rigid cluster vertex. We denote by $[z]\in A$ the $G$-orbit of the vertex $z$, and we observe (in a way similar as in the previous case) that, in view of Fact 13.3(0), the set of types of the peripherals of the form $\varphi^{a}_{[\tilde{t},\tilde{u}]}(\Sigma^{a}_{\tilde{u}})$ coincides with the block $B_{[z]}$ of $\mathcal{B}$. This completes the verification of condition (1).
In order to verify the remaining conditions (2)-(4) of Definition 11.8, we again interpret Facts 13.2 and 13.3 as concerning the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$, as in the previous paragraph. In view of this, condition (2) of Definition 11.8 follows directly from Facts 13.2(1)(2A) and 13.3(1)(2A), condition (3) follows from Facts 13.2(1) and 13.3(1), while condition (4) is a consequence of Facts 13.2(1)(2B) and 13.3(1)(2B). We omit further details, thus completing the proof of the fact that the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ is modelled on the connecting system ${\mathcal{P}}^{r}_{G}$. In view of the comments at the beginning of this section and at the beginning of this subsection, this completes also the whole proof of Theorem 10.1.
13.6. Proofs of the missing parts of Theorems A and B
We complete the proofs of Theorems A and B of the introduction. To do this, we only need to justify the implications (3$\Rightarrow$(2) in Theorem A and (3$\Rightarrow$(2) in Theorem B.
Proof of the implication (3$\Rightarrow$(2) in Theorem A.
It obviously follows from Theorem 10.1 that, assuming condition (3) in Theorem A, the Gromov boundary $\partial G$ is homeomorphic to a regular tree of graphs. It is thus sufficient to show that the graphs $\Gamma_{z}$ appearing in the graphical connecting system ${\mathcal{R}}_{G}$ are all 2-connected. To do this, note that if $\Gamma_{z}$ is the extended Whitehead graph of a rigid cluster factor $G(z)$ of $G$ (i.e. if $z$ is a rigid cluster vertex of $T^{r}_{G}$), then $\Gamma_{z}$ is 2-connected by Fact 9.5. On the other hand, if $z$ is a white vertex of $T^{r}_{G}$ not adjacent to a rigid cluster vertex, $\Gamma_{z}$ is a $\theta$-graph with $k(z)$ edges, where $k(z)$ is the degree of the vertex $z$ in the tree $T^{r}_{G}$. Since $k(z)\geq 3$, this $\theta$-graph $\Gamma_{z}$ is thick and 2-connected. It follows that $\partial G$ is indeed homeomorphic to a regular tree of 2-connected graphs, as asserted in condition (2) of Theorem A, which completes the proof. ∎
Proof of the part (3$\Rightarrow$(2) in Theorem B.
Suppose that condition (3) of Theorem B holds, i.e., $G$ has no rigid factor in its Bowditch JSJ splitting. Obviously, $G$ has then no rigid cluster factor in its reduced Bowditch JSJ splitting. As a consequence, all graphs $\Gamma_{z}$ in the graphical connecting system ${\mathcal{R}}_{G}$ are some thick $\theta$-graphs. It follows then from Theorem 10.1 that $\partial G$ is a regular tree of $\theta$-graphs, as asserted in condition (2) of Theorem B, which completes the proof. ∎
14. Examples
We start this section with the explanations concerning the examples provided in the introduction (in Figures 1–4). In the later part of the section, in Section 14.2, we discuss some Coxeter group examples.
14.1. Examples from the introduction.
In order to show that the examples from the introduction are correct, we need to justify that the described splittings of the corresponding groups $G_{i}$ are indeed their Bowditch (canonical) JSJ splittings. To do this, we will need several auxilliary results.
The following result characterizes the Bowditch JSJ splittings of 1-ended hyperbolic groups in terms of the features of their associated graph of groups decompositions (see Theorem 3.22 in [DT19], where a more general case of relatively hyperbolic groups is addressed). Recall that a splitting $\mathcal{G}$ of some group $G$, as a graph of groups, is nontrivial if no vertex group of $\mathcal{G}$ coincides with $G$. A splitting of $G$ is relative to a family $P$ of subgroups of $G$ if each subgroup from $P$ conjugates in $G$ into some vertex group of $\mathcal{G}$.
Theorem 14.1 (F. Dahmani, N. Touikan [DT19]).
Let $G$ be a 1-ended hyperbolic group which is not cocompact Fuchsian. The canonical JSJ splitting (i.e. the Bowditch JSJ splitting) of $G$ is the unique splitting over 2-ended groups that satisfies the following condition:
The underlying graph $X$ of the associated graph of groups $\mathcal{G}=(\{G_{v}\},\{G_{e}\},\{\varphi_{e}\})$ is bipartite, with black and white vertices, so that
  1. (1)
    the white vertex groups $G_{u}$ are maximal 2-ended subgroups of $G$;
  2. (2)
    the black vertex groups $G_{t}$ are non-elementary (i.e. infinite and not 2-ended); moreover, the images of the incident edge groups in each such $G_{t}$ are pairwise distinct, and denoting by $P_{t}$ the family of all such images, the pair $(G_{t},P_{t})$ falls into one of the following two categories
    1. (a)
      $(G_{t},P_{t})$ is an abstract flexible factor (as in Definition 8.7),
    2. (b)
      $(G_{t},P_{t})$ is algebraically rigid (i.e. there is no nontrivial splitting of $G_{t}$ relative to $P_{t}$ along finite or 2-ended subgroups) and $(G_{t},P_{t})$ is not an abstract flexible factor; moreover, we require in this case that the images in $G_{t}$ of any two incident edge groups are not conjugated in $G_{t}$ into the same 2-ended subgroup of $G_{t}$;
  3. (3)
    for any white vertex $u$ of degree 1 in $X$, the incident edge group $G_{e}$ does not coincide with $G_{u}$ (i.e. it is a proper subgroup of $G_{u}$); moreover, if the neighbouring vertex of $u$ in $X$ carries a flexible factor, then the index of $G_{e}$ in $G_{u}$ is at least 3;
  4. (4)
    for any white vertex $u$ of degree 2 in $X$, if both neighbours of $u$ carry flexible factors then at least one of the incident edge groups $G_{e}$ is a proper subgroup of $G_{u}$.
The next result, which seems to be well known, allows us to recognize the rigid factors of the groups $G_{i}$ from the introduction. For completeness, we include a sketch of a proof of this result, since we haven’t found an appropriate reference in the literature.
Proposition 14.2.
Let $S$ be a compact surface, possibly with boundary, equipped with a hyperbolic metric (so that the boundary is totally geodesic), and let $L$ be a finite collection of closed geodesics contained in the interior of $S$. Suppose that the family $L$ is filling, i.e., any connected component obtained by cutting $S$ along the union of $L$ is either a disk or an annulus, one of whose boundary components coincides with some boundary component of $S$. Let $G=\pi_{1}S$, and let $P$ be the family of (the conjugacy classes of) the cyclic subgroups of $G$ corresponding to the curves in $L$ and to the boundary components of $S$. Then $(G,P)$ is algebraically rigid, i.e., $G$ has no nontrivial splitting relative to $P$ over a finite or 2-ended subgroup.
In order to justify the above proposition, we will make use of the following lemma, which is essentially a restatement of Theorem III.2.6 in [MS84] (see also Lemmas 2.2 and 2.3 in [DS99]). Recall that a splitting of a group $G$ as a graph of groups is minimal if the corresponding action of $G$ on the associated Bass-Serre tree $T$ is minimal, i.e., there is no $G$-invariant proper subtree $T_{0}\subset T$, $T_{0}\neq T$. Recall also that a simple closed curve on a compact surface is essential if it is homotopically non-trivial and not parallel to a boundary component.
Lemma 14.3.
Let $G=\pi_{1}S$ be the fundamental group of a compact connected hyperbolic surface $S$ (we allow that $S$ has nonempty totally geodesic boundary), and let $Q$ be the family of the conjugacy classes of cyclic subgroups of $G$ corresponding to the boundary components of $S$. If $G$ splits non-trivially and minimally relative to $Q$ as a graph of groups $\mathcal{G}$ with cyclic edge groups, then there is a collection $\Lambda$ of pairwise disjoint essential simple closed curves contained in the interior of $S$ such that every edge group of $\mathcal{G}$ is conjugate to the $\pi_{1}$-image of some component of $\Lambda$ (in particular, every edge group is infinite cyclic), and every vertex group is conjugate to the $\pi_{1}$-image of some connected component of the complement $S\setminus\bigcup\Lambda$. Moreover, the components of $\Lambda$ can be realized as geodesics, except for the families of mutually parallel components, which can be realized as arbitrarily close to a geodesic.
Remark 14.4.
Note that the above lemma implies in particular a rather well known but useful fact that the fundamental group of a compact hyperbolic surface is freely indecomposable relative to its boundary subgroups, i.e., that the pair $(G,Q)$ as in the lemma is freely indecomposable.
Proof of Proposition 14.2:.
Suppose, a contrario, that $G$ admits a notrivial splitting $\mathcal{G}$ (as a graph of groups) relative to $P$ over finite or 2-ended subgroups. We can assume, without loss of generality, that this splitting is minimal. Since $G$ is torsion-free, this splitting is over infinite cyclic or trivial subgroups. Let $Q$ denote the set of the conjugacy classes of those cyclic subgroups of $G$ which correspond to the boundary components of $S$. Since we have $Q\subset P$, $\mathcal{G}$ is obviously a splitting relative to $Q$. By Lemma 14.3, we can realize this splitting geometrically, as induced by some collection $\Lambda$ of pairwise disjoint essential simple closed curves on $S$. This shows that the splitting is actually over infinite cyclic subgroups only (compare Remark 14.4). Furthermore, we may assume that the curves of $\Lambda$ are either geodesic, or very close to a geodesic. Since $\mathcal{G}$ is a relative splitting (with respect to $P$), an easy geometric observation shows that one can slightly perturb $\Lambda$ so that each curve $C$ of $\Lambda$ is contained in a single connected component of the complement $S\setminus\bigcup L$. By the fact that $L$ is a filling collection, each connected component of $S\setminus\bigcup L$ is either a disk or an annulus adjacent to a boundary component of $S$. Obviously, being an essential simple closed curve, $C$ cannot be contained in a disk component since it is homotopically nontrivial in $S$, but it also cannot be contained in an annulus component since it is not parallel to a boundary component. It follows that the family $\Lambda$ is empty. However, if this is the case, the induced splitting is obviously trivial, which contradicts our assumption that the splitting was nontrivial. This completes the proof. ∎
We are now ready to discuss and explain the examples from the introduction. Note first that for each graph of groups $\mathcal{G}_{i}$, $i=1,2,3,4$, the underlying graph $X_{i}$ carries a natural bipartite structure, with white vertices corresponding to the curves of the gluings, and with black vertices corresponding to the surface pieces ($F_{j}$ and $H_{j}$) of $M_{i}$. It is also easy to realize that each of the white vertex groups is a maximal cyclic (and hence also maximal 2-ended) subgroup of the corresponding torsion free group $G_{i}$. On the other hand, each of the black vertex groups is the fundamental group of a hyperbolic surface with nonempty boundary, so it is a non abelian free group, and hence it is non-elementary. Moreover, each of the black vertex groups of the form $\pi_{1}(F_{j})$ is easily seen to be an abstract flexible factor (when equipped with the family of images of the incident edge groups). We will show that any black vertex group of the form $G_{t}=\pi_{1}(H_{j})$, equipped with the corresponding family $P_{t}$ of images of incident edge groups, satisfies condition (2)(b) of Theorem 14.1. The fact that the pair $(G_{t},P_{t})$ as above is algebraically rigid follows directly from Proposition 14.2. To see that $(G_{t},P_{t})$ is not an abstract flexible factor, note that any flexible abstract factor $(G,P)$ fails to be still flexible after deleting from $P$ some of the subgroups. On the other hand, if we delete from $P_{t}$ those subgroups of $G_{t}=\pi_{1}H_{j}$ which correspond to the gluing curves contained in the interior of $H_{j}$, the resulting new abstract factor $(G_{t},Q_{t})$ will be obviously flexible. It remains to check the last assertion of condition (2)(b), namely that no two incident edge groups from $P_{t}$ conjugate into the same 2-ended subgroup of $G_{t}$. This follows however from the fact that these subgroups (as well as any pair of their conjugates), when viewed as groups of isometries of a hyperbolic plane, are generated by loxodromic isometries having distinct axes, so their intersection is trivial.
Since the verification of the remaining conditions (3) and (4) of Theorem 14.1 is straightforward, we omit its details, and we conclude that the splittings of the groups $G_{i}$ related to the graph of groups expressions $\mathcal{G}_{i}$ are indeed the Bowditch JSJ splittings.
We now pass to a more detailed description of the Gromov boundary in one of the above examples, namely the Gromov boundary $\partial G_{2}$. We will do this by explicitly describing the graphical connecting system $\mathcal{R}_{G_{2}}$, as introduced in Section 10, and as appearing in the statement of Theorem 10.1. This will do the job, since this graphical connecting system uniquely determines the corresponding regular tree of graphs $\mathcal{X}(\mathcal{R}_{G_{2}})$, and by Theorem 10.1 we have $\partial G_{2}\cong\mathcal{X}(\mathcal{R}_{G_{2}})$.
Observe that the set $Z$ of orbit representatives (as in Section 10) consists of a single vertex, $z_{0}$, which corresponds to the rigid cluster factor $\pi_{1}H_{0}$ in the splitting $\mathcal{G}_{2}$. This is so since, as it is easy to realize, the Bass-Serre tree $T^{r}_{G_{2}}$ corresponding to $\mathcal{G}_{2}$ contains no white vertices which are not adjacent to a rigid cluster vertex. The graph $\Gamma$ in $\mathcal{R}_{G_{2}}$ coincides then with the graph $\Gamma_{z_{0}}$, and the latter is isomorphic to the extended Whitehead graph $\widetilde{W}$ of the corresponding rigid cluster factor $\pi_{1}H_{0}$. It is also not hard to realize that the degree in $T^{r}_{G_{2}}$ of any (white) vertex adjacent to $z_{0}$ equals 2, and hence the extended Whitehead graph $\widetilde{W}$ coincides with the ordinary Whitehead graph $W$ of this factor.
To get the form of the Whitehead graph $W$, observe that for an appropriately chosen freely generating set $\{a,b\}$ in $\pi_{1}H_{0}$, the peripheral subgroups of $\pi_{1}H_{0}$ are generated by the elements $a$, $b$ and $aba^{-1}b^{-1}$, respectively. As a consequence, the Whitehead graph $W$ has the form of a square with both diagonals (i.e. $W=K(4)$), where the four boundary edges correspond to the generator $aba^{-1}b^{-1}$. The peripheral subgroups are obviously orientable, and if we choose their orientations corresponding to the above choices of the generators, the order induced orientations of the edges of $W$ are arranged so that the edges on the boundary of the square are oriented consistently with one of their cyclic order. The associated V-involution $\bf a$ is described by transpositions of the pairs of opposite vertices in the square, and the corresponding maps $\alpha_{v}:{\rm Lk}_{v}\to{\rm Lk}_{a(v)}$ are given by the assignments of points corresponding to the diagonals to the relevant points corresponding to the diagonals, and by the assignments of points corresponding to the remaining edges in the unique possible way respecting the orientations.
Let $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3},\varepsilon_{4}$ be the edges of $W$ on the boundary of the square, and let $\delta_{1},\delta_{2}$ be the diagonals, with the above mentioned orientations. Then the set $\mathcal{A}$ of E-connections is easily seen to have the following form:
$\mathcal{A}=\{(\delta_{1},\delta_{1})\}\cup\{(\delta_{2},\delta_{2})\}\cup(\{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3},\varepsilon_{4}\}\times\{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3},\varepsilon_{4}\}).$
By the assertion of Theorem 10.1, the Gromov boundary $\partial G_{2}$ is homemorphic to the regular tree of graphs $K(4)$ associated to the above described graphical connecting system $\mathcal{R}_{G_{2}}$. However, due to the large symmetry of the graph $K(4)$, any two tree systems of punctured graphs $K(4)$ are in fact isomorphic, and consequently their limits are homeomorphic, so we can just say that $\partial G_{2}$ is homeomorphic to the (unique) tree of graphs $K(4)$.
This finishes our discussion of the examples from the introduction.
14.2. More examples - right angled Coxeter groups with 1-dimensional nerves.
In this subsection we illustrate the results of the present paper by means of various examples of (hyperbolic) Coxeter groups, or more precisely right angled Coxeter groups (which we shortly call RACG’s). We assume that the reader is familiar with basic terminology and facts related to Coxeter groups, and we refer to the book [Dav08] for an introduction.
In an earlier paper [Ś21] the second author of the present paper has shown that for a large class of 1-ended hyperbolic RACG’s their Gromov boundaries are trees of graphs. This class of RACG’s can be described in terms of their defining graphs (or nerves) $N$ as follows:
$N$ is a triangulation of a topological graph $X$ such that
  • $X$ is connected and has no cut vertex, and it is distinct from the circle $S^{1}$;
  • each (essential) edge of $X$, when viewed as a subcomplex of $N$, consists of at least 3 edges of $N$.
We denote the class of graphs $N$ satisfying the above two conditions by $\mathcal{C}_{JS}$, and the corresponding RACG’s associated to such graphs by $W_{N}$. The argument in [Ś21] for showing that Gromov boundaries $\partial W_{N}$ of such groups are trees of graphs is different than the one given in the present paper. The trees of graphs appearing as boundaries $\partial W_{N}$ for $N\in\mathcal{C}_{JS}$ belong to a subclass consisting of so called reflection trees of graphs, as mentioned in Example 6.5 in this paper. Strictly speaking, it is shown in [Ś21] that if $N\in\mathcal{C}_{JS}$ is a triangulation of a topological graph $X$ then $\partial W_{N}\cong\mathcal{X}^{r}(X)$ (the reflection tree of graphs $X$).
A wider class of examples has been studied in the paper [DT17] by Pallavi Dani and Anne Thomas. That paper contains a complete description of the Bowditch JSJ splittings for a family of 1-ended hyperbolic RACG’s whose defining graphs $N$ satisfy the following set of conditions (which appear in [DT17] under the name Standing Assumptions 1.1):
  • $N$ has no embedded cycles consisting of less than 5 edges,
  • $N$ is connected and has no cut vertex and no cut edge,
  • $N$ is not a cycle.
We denote the class of graphs $N$ satisfying the above conditions by $\mathcal{C}_{DT}$. The class is strictly larger than $\mathcal{C}_{JS}$. In the light of the results in [DT17], and using Theorems A and B of the present paper (or rather their more precise version given as Theorem 10.1), one can indicate many more hyperbolic RACG’s $W_{N}$ with the boundary $\partial W_{N}$ recognized as an explicitly given tree of graphs (see Examples 14.5 and 14.6 below, Figures 5 and 6). One can also indicate examples with boundaries $\partial W_{N}$ of topological dimension 1 and not homeomorphic to any tree of graphs, even the ones for which all rigid factors are virtually free (see e.g. an example presented in Figure 7 below).
We now describe and discuss the promised examples from the class $\mathcal{C}_{DT}$. In all of them, as presented in Figures 5–7, we indicate their graph of groups decompositions related to the Bowditch JSJ splittings. These decompositions are obtained by applying the results of [DT17], and we do not provide the details, referring the interested readers directly to that paper. The underlying graphs of the indicated graph of groups decompositions in all of these examples are the quotients of the corresponding Bass-Serre trees by the actions of the groups, so one can easily deduce various properties of the corresponding Bowditch JSJ splittings (as well as of their reduced variants).
Example 14.5.
Let $G=W_{N}$ be the RACG for the defining graph $N$ as presented in the left part of Figure 5. The right part of the same figure presents the graph of groups decomposition $\mathcal{G}$ of $G$ related to the Bowditch JSJ splitting of this group. Vertices represented by circles are flexible, i.e., of type (v1), while those represented by little squares are of type (v2). The symbols inside circles and squares, as well as those which accompany the edges, are the generators of the corresponding special subgroups of $W_{N}$ appearing as vertex groups and edge groups of $\mathcal{G}$, respectively.
Figure 5.
There are no rigid vertices in $\mathcal{G}$, so there are also no rigid vertices in the corresponding Bass-Serre tree $T_{G}$. It follows then from Theorem B that the boundary $\partial G$ is a regular tree of $\theta$-graphs. Note that the defining graph $N$ clearly does not belong to the class $\mathcal{C}_{JS}$, so this examples goes beyond the class studied in an earlier mentioned paper [Ś21]. (Obviously, it is easy to show many more examples of this sort, and another one, whose JSJ splitting contains a rigid factor, is described below in Example 14.6, see the group $G_{1}$ presented in Figure 6.) To get the explicit form of the regular tree of $\theta$-graphs homeomorphic to $\partial G$, i.e., to describe the corresponding graphical connecting system $\mathcal{R}_{G}$ as in Section 10, we need a closer analysis.
One deduces from the form of the graph of groups $\mathcal{G}$ that preimages in $T_{G}$ of the vertices $\langle A,B,8\rangle$ and $\langle A,C,7\rangle$ have degree 4, so that for any such preimage vertex $v$ there are precisely two edges issuing from $v$ that are mapped to each edge issuing from $\langle A,B,8\rangle$ or $\langle A,C,7\rangle$ (this is indicated in Figure 5 by little symbols “×$\times 2$”). Preimages in $T_{G}$ of the vertex $\langle B,C\rangle$ have degree 3. Each of the 2-ended groups at vertices of type (v2) is non-orientable, because it is either an infinite dihedral group, or the product of an infinite dihedral group with a group of order 2. It follows that all flexible factors of the Bowditch JSJ splitting of $G$ are also non-orientable.
According to our exposition in Section 10, the graphical connecting system $\mathcal{R}_{G}=(\Gamma,{\bf a},\mathcal{A})$ is described as follows. First, $\Gamma$ is the disjoint union of two copies of the $\theta$-graph $\Xi_{4}$ (corresponding to vertices $\langle A,B,8\rangle$ and $\langle A,C,7\rangle$), which we denote $\Gamma_{2}$ and $\Gamma_{3}$, respectively, and one copy of the $\theta$-graph $\Xi_{3}$ (corresponding to the vertex $\langle B,C\rangle$), which we denote $\Gamma_{1}$. The $V$-involution ${\bf a}$ of $\mathcal{R}_{G}$ is just the union of the standard $V$-involutions (as described in Example 7.11) for the $\theta$-graphs $\Gamma_{1}$, $\Gamma_{2}$ and $\Gamma_{3}$.
To describe the set $\mathcal{A}$ of $E$-connections for $\Gamma$, denote by $a,b,c$ the edges of $\Gamma_{1}$, by $d,e,f,g$ the edges of $\Gamma_{2}$, and by $h,i,j,k$ the edges of $\Gamma_{3}$, and for each of those edges fix one of its orientations (so that these letters actually denote oriented edges). According to the description in Section 10, for each $\Gamma_{i}$ we fix a bijection between the set of edges of $\Gamma_{i}$ and the set of vertices in $T_{G}$ adjacent to the vertex $z\in Z$ representing this component. We have indicated our choices of these bijections by assigning letters $a,\dots,k$ to flexible vertices of the graph in the right part of Figure 5. Any set $Y$ of representatives of flexible vertices in $T_{G}$ can be canonically identified with the set of vertices represented by circles in Figure 5. It follows that the sets $E_{y}:y\in Y$ coincide with the sets $\{a\}$, $\{b\}$, $\{c,d,e,h,i\}$, $\{f,g\}$ and $\{j,k\}$. Due to non-orientability of all flexible factors, the set of $E$-connections has then the form
$\mathcal{A}=\bigcup_{y\in Y}E_{y}^{\pm}\times E_{y}^{\pm},$
with $E_{y}:y\in Y$ as mentioned above.
Example 14.6.
Let $G_{1}=W_{N_{1}}$ and $G_{2}=W_{N_{2}}$ be the 1-ended hyperbolic RACG’s for the defining graphs $N_{i}$ as presented in the left parts of Figures 6 and 7, respectively. The central parts of these figures present the graph of groups decompositions $\mathcal{G}_{i}$ of these groups related to their Bowditch JSJ splittings. These decompositions have been obtained by referring to the results in [DT17]. Both underlying graphs $X_{i}$ of $\mathcal{G}_{i}$ contain a single vertex related to a rigid factor, namely the vertex represented by a hexagonal symbol, which corresponds to the special subgroup $H=\langle A,B,C,D\rangle<G_{i}$ in each of the two cases. It is also not hard to deduce from the form of $\mathcal{G}_{i}$’s that the corresponding rigid vertices in the Bass-Serre trees $T_{G_{i}}$ are not isolated, i.e they have some rigid neighbours at combinatorial distance 2. Thus, in both cases we have nontrivial rigid clusters, and we have rigid cluster factors which contain rigid factors as proper subgroups.
Figure 6.
Figure 7.
Consider the graph of groups decompositions $\mathcal{G}^{r}_{i}$ of the groups $G_{i}$ related to their reduced Bowditch JSJ splittings. It is not hard to realize that they have forms as presented in the right parts of Figures 6 and 7. The rigid cluster factors correspond to the vertices indicated by the symbols of the bold hexagons. The rigid cluster factors in the group $G_{1}$ are the conjugates of the special subgroup $H_{1}=\langle A,B,C,D,3\rangle$, and it is not hard to see that they are virtually free. It follows then from Theorem A that the Gromov boundary $\partial G_{1}$ is a regular tree of 2-connected graphs. The explicit form of this tree of graphs, i.e., the corresponding graphical connecting system $\mathcal{R}_{G_{1}}$, can be extracted from the data in $\mathcal{G}^{r}_{1}$, and the least obvious, though still starightforward part of this extraction is the determination of the Whitehead graph associated to the rigid cluster factor $H_{1}$. We skip the details.
On the other hand, the rigid cluster factors in the group $G_{2}$ are the conjugates of the special subgroup $H_{2}=\langle A,B,C,D,5\rangle$ (see Figure 7). Since the defining graph of this subgroup is a 5-cycle, it is a cocompact Fuchsian group (it can be realized e.g. as the group generated by reflections in the sides of a right angled hyperbolic pentagon). The Gromov boundary of this subgroup is a circle, so the subgroup is not virtually free, even though the rigid factors from which it is formed are virtually free. It follows then from Theorem A that the Gromov boundary $\partial G_{2}$ is not homeomorphic to any tree of graphs.
15. Final remarks and comments
Remark 15.1.
It is worth noting that the observations and results from sections 9-13 above (concerning an explicit expression of the Gromov boundary $\partial G$ as a regular tree of graphs) are still valid in more general situations than those corresponding to Bowditch JSJ or reduced Bowditch JSJ splittings. We now state a few such generalizations.
  1. (1)
    Let ${\mathcal{G}}=(X,\{G_{t}\}\sqcup\{G_{u}\},\{G_{e}\})$ be a graph of groups such that:
    1. (a)
      the underlying graph $X$ of $\mathcal{G}$ is bipartite, and its vertex set is accordingly bipartitioned into black vertices $t$ and white vertices $u$;
    2. (b)
      the white vertex groups $G_{u}$ of $\mathcal{G}$ are 2-ended, and the edge groups $G_{e}$ of $\mathcal{G}$ are 2-ended as well;
    3. (c)
[…]
    Under these assumptions, the fundamental group $G:=\pi_{1}({\mathcal{G}})$ is hyperbolic (by Bestvina-Feighn combination theorem), and the vertex groups of $\mathcal{G}$ are quasi-convex in $G$ (e.g. by Theorem 1.2 in [Kapo]). Consider the action of $G$ on the Bass-Serre tree $T_{\mathcal{G}}$ of $\mathcal{G}$. The vertex set of $T_{\mathcal{G}}$ is then equipped with the induced bipartition into black and white vertices. By our assumptions on $\mathcal{G}$, no two distinct white vertices of $T_{\mathcal{G}}$ are fixed (for the above action) by an infinite subgroup of $G$, and hence the action (as well as $\mathcal{G}$ itself) is acylindrical. If we associate to the splitting $\mathcal{G}$ of $G$ the tree system $\Theta^{\mathcal{G}}(G)$, in the analogous way as in Section 3 for Bowditch and reduced Bowditch splittings, the arguments as in the proof of Theorem 3.2 (based on the discussion in Section 2.6.2 of [CaMa]) still yield that $\partial G\cong\lim\Theta^{\mathcal{G}}(G)$. In particular, this holds even if $G$ is not 1-ended.
  2. (2)
[…]
  3. (3)
    Finally, in addition to the assumptions of (1) and (2), suppose that each black vertex factor $G_{t}$ is weakly rigid, which means that $G_{t}$ has no nontrivial splitting along finite subgroups relative to its incident edge groups $G_{e}$. (This assumption is automatically fulfilled by the factors $G_{t}$ which are flexible, so it is sufficient to demand that all non-flexible factors of $\mathcal{G}$ are weakly rigid.) Suppose also that the corresponding action of $G$ on the associated Bass-Serre tree is minimal, which is equivalent to saying that for each white vertex $u$ of $X$, either the degree of $u$ in $X$ is at least 2, or this degree is 1 and the image in $G_{u}$ of the unique incident edge group $G_{e}$ is a proper subgroup. Then $G$ is 1-ended (e.g. by Corollary 1.5 in [Toui]). Note that under the above minimality assumption for the $G$-action on the Bass-Serre tree, the $\theta$-graphs appearing in $\mathcal{R}_{\mathcal{G}}$ have the forms $\Xi_{k}$ with $k\geq 2$, so they are all 2-connected. Moreover, if we choose the Whitehead graphs $W_{t}$ mentioned above in part (2) to be minimal, then they are 2-connected as well (by the argument as in the proof of Proposition 9.3). (Such a choice is then analogous to the one declared in our previous considerations in Section 8, as explained in Remark 8.24.) Under such a choice of the Whitehead graphs (and of their extended variants), the above mentioned equality $\partial G\cong\lim\Theta[\mathcal{R}_{\mathcal{G}}]$ expresses the boundary $\partial G$ as an explicit regular tree of 2-connected graphs.
Example 15.2.
We present an example which demonstrates that the discussion of Remark 15.1 allows us to easily deduce that boundaries of various hyperbolic groups have an explicit form of a regular tree of graphs, without any need to first understand their canonical JSJ decompositions.
Consider the graph of groups $\mathcal{G}$ whose underlying graph $X$ is the barycentric subdivision of a single edge, so that the original vertices $t,s$ of this edge are black, and the barycenter vertex $u$ is white. Let $G_{t}$ be the free group $F_{2}$ of rank 2, with its standard generators denoted $x,y$, and let $G_{s}$ be another free group viewed as the fundamental group of a compact hyperbolic surface $S$ with boundary $\partial S$ consisting of a single component. Let $G_{u}$ be the infinite cyclic group, with fixed generator $a$, and let the edge groups $G_{[u,t]}$, $G_{[u,s]}$ coincide with $G_{u}$. Let the incidence homomorphism $G_{[u,s]}\to G_{s}$ map the generator $a$ of $G_{[u,s]}$ to some generator of the subgroup $\pi_{1}\partial S<\pi_{1}S=G_{s}$. Finally, let the incidence homomorphism $G_{[u,t]}\to G_{t}$ map the generator $a$ of $G_{[u,t]}$ to the element of $G_{t}=F_{2}$ represented by some nontrivial cyclically reduced word $w=w(x,y)$ which is not a proper power of any other word. The fundamental group $G=\pi_{1}{\mathcal{G}}$ of this graph of groups, which we will focus on, can be then also viewed as the fundamental group of the space obtained by gluing $S$ along its boundary $\partial S$, via an appropriate map corresponding to the word $w$, to the wedge of two circles.
Note that any graph of groups $\mathcal{G}$ as above satisfies conditions (a)-(c) of part (1) of Remark 15.1, so that $G=\pi_{1}{\mathcal{G}}$ is hyperbolic. The abstract factor $G_{s}$ is easily seen to be flexible, so if we view the other factor $G_{t}$ as non-flexible (which we can do even if it is actually flexible), it will be isolated in $\mathcal{G}$. Moreover, if the word $w$ involves both generators $x,y$, the Whitehead graph $W_{t}$ for the basis $B=\{x,y\}$ is essential, and the assumtions of part (2) of Remark 15.1 are fulfilled. The extended Whitehead graph $\widetilde{W}_{t}$ coincides then with $W_{t}$, and the corresponding graphical connecting system ${\mathcal{R}}_{\mathcal{G}}$ involves no $\theta$-graphs (because the unique white vertex $u$ of $X$ is adjacent to a non-flexible vertex $t$). By the arguments as in the proof of Theorem 10.1 (given in Section 13) we get that the Gromov boundary $\partial G$ is then a regular tree of the (copies of the) Whitehead graphs $W_{t}$.
Finally, if the word $w$ as above is Whitehead-minimal (in the sense of the Whitehead algorithm, as described e.g. in Section 3 of [CaMac]), then the factor $G_{t}$ is weakly rigid, the group $G$ is 1-ended, and the Whitehead graph $W_{t}$ is 2-connected, so that $\partial G$ is a regular tree of 2-connected graphs $W_{t}$.
Remark 15.3.
After completing the main part of this paper, we have realized that groups $G$ satisfying condition (3) of Theorem A (and condition (3) of Theorem B) are actually residually finite, and hence virtually torsion free. This has been pointed to us by Marco Linton, and the justification goes via virtual specialness of all such $G$, guaranteed by the fact that they obviously have quasiconvex virtual hierarchies $\mathcal{QVH}$ (compare Theorem 13.1 in [Wise] and Theorem 1.3 in [HagW]). As a consequence, each group $G$ as above contains a torsion free subgroup $G^{\prime}$ of finite index. The restricted action of $G^{\prime}$ on the Bass-Serre tree $T^{r}_{G}$ represents then the reduced Bowditch JSJ splitting of $G^{\prime}$, which we denote by ${\mathcal{G}}^{\prime}$. (The latter follows either by referring to Dahmani-Touikan characterization of the canonical JSJ splittings, as recalled in Theorem 14.1 of the present paper, or by using the Bowditch approach to canonical JSJ splittings, as presented in [Bow98], where the splitting is expressed in terms of configurations of local cut points in the Gromov boundary, which is the same for $G$ and $G^{\prime}$.) The factors of the reduced Bowditch JSJ splitting ${\mathcal{G}}^{\prime}$ of $G^{\prime}$ are obviously either free or infinite cyclic. Applying our approach from Sections 6–13 above to this splitting (for proving the implication (3$\Rightarrow$(1) in Theorem A) we would avoid some of the technicalities (e.g. the idea of the Whitehead graph for a virtually free factor, or many details related to potential non-orientability of some peripheral subgroups in the factors). Since obviously $\partial G=\partial G^{\prime}$, this would lead to a slightly simpler proof of the implications (3$\Rightarrow$(1) in Theorems A and B. Some further simplifications could be obtained by passing (via residual finiteness of $G$) to an even smaller finite index subgroup $G^{\prime\prime}<G^{\prime}<G$, for which additionally all flexible factors of the corresponding induced splitting ${\mathcal{G}}^{\prime\prime}$ would be orientable (in the sense of Definition 8.10). However, since the passages from $G$ to $G^{\prime}$ (or to $G^{\prime\prime}$) and from the reduced Bowditch JSJ splitting $\mathcal{G}$ of $G$ to the corresponding splitting ${\mathcal{G}}^{\prime}$ of $G^{\prime}$, or to ${\mathcal{G}}^{\prime\prime}$ of $G^{\prime\prime}$, (and, as a consequence, the passages from the connecting systems $\mathcal{R}_{G}$ to $\mathcal{R}_{G^{\prime}}$ or to $\mathcal{R}_{G^{\prime\prime}}$) do not seem amenable to easy explicit calculation, we have decided to leave this part of the paper (and the current form of Theorem 10.1 and its proof) in the current form. We believe that the possibility of directly expressing $\partial G$ as an explicit regular tree of graphs (in the form $\partial G\cong\mathcal{X}(\mathcal{R}_{G})$, as in Theorem 10.1 and Remark 15.1) justifies our decision.
Remark 15.4.
As we have noted in Remark 15.3, we can always pass from a group $G$ satisfying condition (3) of Theorem A to its finite index subgroup $G^{\prime\prime}$ which is torsion free and for which all flexible factors of the corresponding reduced Bowditch JSJ splitting ${\mathcal{G}}^{\prime\prime}$ are orientable. This allows us, by referring to Theorem 10.1, to express the Gromov boundary $\partial G=\partial G^{\prime\prime}$ of any such $G$ as a regular tree of graphs $\mathcal{X}(\mathcal{R}_{G^{\prime\prime}})$ of a more restricted form than the corresponding tree of graphs $\mathcal{X}(\mathcal{R}_{G})$. More precisely, in the corresponding graphical connecting systems $\mathcal{R}_{G^{\prime\prime}}$ the set $\mathcal{A}$ of $E$-connections has a simpler form, and corresponds to connected sum operations of the involved graphs which always respect the distinguished orientations $e^{\#}$ of the edges in the graphs (such orientations are in this case assigned to all edges of the involved graphs, see Section 10). This observation suggests the following question.
Question 15.5.
Does there exist a regular tree of graphs which cannot be realized as the Gromov boundary of a hyperbolic group?
Remark 15.6.
We briefly comment on how the ideas and results of this paper contribute to the problem of classifying (up to homeomorphism) Gromov boundaries of topological dimension 1.
  1. (1)
    Recall that on the opposite extreme to trees of graphs among 1-dimensional Gromov boundaries $\partial G$ we have those boundaries corresponding to 1-ended groups $G$ that do not split over 2-ended subgroups. In these cases the boundary has no local cut points (cf. [Bow98]), and it is known to be homeomorphic either to the Sierpiński curve or to the Menger curve (cf. [KK00]). We call the corresponding groups $G$ carpet groups and Menger groups, respectively.
  2. (2)
    Observe also that a 1-ended hyperbolic group $G$ which is not cocompact Fuchsian has 1-dimensional Gromov boundary $\partial G$ if and only if the boundaries $\partial G_{v}$ of all rigid factors $G_{v}$ (in the Bowditch JSJ splitting of $G$) satisfy $\dim(\partial G_{v})\leq 1$. To justify this, note that in view of Theorem 3.2 (which allows us to express $\partial G$ as the limit of the tree system $\Theta(G)$ composed of boundaries of the factors in the Bowditch JSJ splitting of $G$), and in view of the fact that peripheral spaces in the tree system $\Theta(G)$ are all doubletons, this observation follows fairly directly from Proposition 2.D.5 in [Ś20].
  3. (3)
    In view of the above observation (2), Theorem A of the present paper shows that trees of graphs are the simplest 1-dimensional Gromov boundaries, and that there is a large variety of 1-dimensional Gromov boundaries that are not trees of graphs. It is in particular worth observing that the latter Gromov boundaries include, among others, a vast class of examples in which the corresponding groups $G$ have no carpet factors or Menger factors (or factors containing such groups as subgroups) in their JSJ splittings. For instance, it is possible that (as in the example $G_{4}$ in the introduction) a group has all rigid factors virtually free, but some of its rigid cluster factors are not virtually free. Furthermore, there are obviously groups $G$ with some rigid factors having boundaries of topological dimension 1 and not coinciding with a carpet group ar a Menger group (or a group containing such a group as subgroup). It is possible e.g. that a rigid factor $G_{v}$ of $G$ is itself a group with $\partial G_{v}$ homeomorphic to some tree of graphs. This shows that the zoo of 1-dimensional connected Gromov boundaries $\partial G$ is quite rich and consists of quite complex species and we believe it deserves further study from the topological perspective.